A Bibliography of Publications in IEEE Annals of the History of Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 August 2025 Version 1.116

Title word cross-reference

$14.95 [Smi03]. $24.95 [Ber03d]. $44.95 [Mor03]. $8 [Gra21b]. $8.50 [Smi78]. $95.00 [Win03]. π [RG02, She12].

. [Ber03d].

0 [RAG^{+}04]. 0-19-514025-7 [RAG^{+}04]. 0-262-03303-8 [RAG^{+}04]. 0-262-06224-0 [Ber03d]. 0-262-08305-1 [RAG^{+}04]. 0-7382-0561-3 [RAG^{+}04]. 0-801-86895-5 [RAG^{+}04]. 0-8018-6809-2 [Mor03]. 0-929306-27-9 [RAG^{+}04]. 0-930405-85-4 [Cer03]. 000-year-old-computer [Mar10a]. 01 [Sta07]. 085-296-7926 [Win03].

1 [Cor15c, GR02, GH18b, Ich06, Kap07, Kos03b, RAG^{+}04, Sac07, Spi15a]. 1-2-3 [Kap07, Sac07]. 1-4033-1517-5 [Smi03]. 10NES [O’D09]. 11 [Kis95]. 1620 [Spi05]. 1870s [Tym96]. 1880s [Hei11]. 18th [Cro03a, Cro03b]. 18th-Century [Cro03a, Cro03b]. 1920s [Cor17, Cor25a]. 1930s [Cor24b, Hei11]. 1940s [Asp00, Cor06a]. 1950s [Cla10, Cor06a, DMD18, Hai09, Kos03a, Tin10]. 1958 [Cob23]. 1958-1985 [Vog17]. 1960s [Cor02, Don10, Gup07, Hai02, Hai09, Hen05, Joh98, Joh02, Nof25, Pee09, Smi16b, Tin10, Tym96]. 1970s [Cor15b, Cor15a, Gre95, HIIT05, Pee09, RP12, Vid20, VR22, WMB22]. 1976-1985 [Ber06a]. 1978 [Hus97]. 1980s [Cor15a, Cor17, Cor24b, Cor25a, Gra21b, Hod19, Ñañ22, OARHC10, RP12, Sta15]. 1985-1995 [Ber06b]. 1990 [Wei93d]. 1990s [Cor15b, Gra21b, HII^{+}05, MR22]. 19th [Sha23]. 1st [Smi03].

2 [KD21, McJ17b, Roe16b, Swa05a, Wil92a]. 2.0.1 [Kid00h]. 20 [Mur15]. 2000 [Bru09, Ros04]. 2000s [Smi16b]. 2002 [RAG^{+}04]. 2016 [Ano15a]. 2017 [Ano17a, Sid20]. 2018 [Yam19a, Yam19b]. 2019 [Wal19]. 2020 [Ano20p, Ano20-28, Ano20-38]. 2022 [Ano21a, AH23]. 2024 [ABL^{+}24, Cor24a]. 2025 [BMFB25]. 20th [Kid96a, EK24]. 210 [Run08]. 21st [Ano20n, Ano22-59, Ano22-61, Ano23-49, Ano23-50, Ano22-60]. 28a [Wil00c, Bro00].

3 [Kap07, RAG^{+}04, Sac07]. 30 [Smi07]. 30-Million-Rule [Len22]. 33 [Per99]. 370 [Wad12]. 3rd [Kra22].

4004 [Asp97]. 440 [JMSV10, JMV10, Sie10].

5 [RAG^{+}04]. 50-Year [Nel22]. 50th [Wal19].

60 [DA14]. 6000 [Var94]. 650 [Gil07].

7 [RAG^{+}04]. 70 [Sta03].

8 [RAG^{+}04]. 80 [de 74]. 8086 [Maz10]. 80s [Der20]. 850 [OARHC10].

9 [RAG^{+}04]. 940 [KH21b]. 978 [GH18b]. 978-0-262-03398-5 [GH18b]. 978-1-4214-1868-1 [GH18b].

A. [Kid96b]. AAAS [Wal19]. ABA [McK95]. Abacus [Che25]. Abbate [Rus13c]. ABC [Swe05]. Absence [Cam09]. Abstract [GR22]. Abstraction [VT21]. Academy [JB07]. Access [Bac09]. Accomplishments [HII^{+}05]. Account [Kos03b, McK11]. Accreditation [AH23]. Accurate [Smi99a]. Achieve [Ano17b, Ano17c]. Achievements [Rei23b].

ACM [PP21, Shr24, Ano19a, Ano21b, Ber09]. ACM/IEEE [Ano19a, Ano21b]. Acoustic [ES05]. Acoustics [Swe05]. Acquire [Cor19a, Cor19b, Cor20]. Acquisition [Bau06]. Across [RCP^{+}13]. ACS [Ran15, Rob16]. act [BW00, Wil00b]. Action [GH18b]. Activist [Nyi11]. activity [CK92c, CK98]. Ad [Ano18c, Ano18d, Ano18p, Ano18o, Ano18t, Ano18v, Ano18u, Ano18-27]. Ada [Lee00a, Spi03, Too96]. Adabas [Mag09]. ADAPSO [Hai04b, Hai05b]. Adder [HT12]. Adding [FVW25, Kid01d, Roe08, Kid00a, Kis98]. Addressing [Day21]. Adelle [Asp01a]. administration [Gri98]. Administrative [Ano18m, Ano18n]. Adobe [WG19]. Adopter [Gal11]. Ads [Ben13]. advantage [Asp00]. Advent [McJ20, Mor20, Yos13a]. Advert [Ano22a]. Advertisement [Ano13b, Ano13c, Ano14d, Ano14f, Ano14g, Ano14k, Ano15b, Ano15c, Ano15e, Ano15f, Ano15g, Ano15l, Ano15m, Ano15a, Ano15n, Ano15o, Ano15q, Ano15p, Ano15v, Ano16a, Ano16b, Ano16h, Ano16i, Ano16m, Ano17f, Ano17m, Ano17u, Ano17w, Ano17x, Ano17y, Ano17z, Ano13g, Ano13f, Ano14c]. Advertising [Hic10]. Advocate [Smi11]. AEG [JMSV10, JMV10]. AEG-Telefunken [JMSV10, JMV10]. aerospace [Joh96]. Affair [van19]. AFIPS [Gro92]. AfroNet [McG21]. After [BMFB25, BS11, GM96, GS09, LH95, Lon08]. AG [Mag09, Lei17]. Again [Mir17]. Against [GS07]. Age [MT06, Mor03, Nof23, Bow96c, Kid99a]. Agenda [Cor04]. agendas [Owe96]. Ages [Kid99e, Smi03]. Agricultural [Gri00a]. Ahead [Mar12]. AI [Ano22b, Ano24a, Cop23, Gar21, Kel22, PPM^{+}23, Poi25, Sch23, SST22, Ste22]. Aided [Cro18, Gra24, Lla24, ORB^{+}92, Ric24]. Aids [Gri01b, Mil11, Pol95, CF92]. Aiken [Coh98, Cop04b, Cro00b, Lee00c]. Aiken-Dahlgren [Lee00c]. Air [Dor94]. airborne [Bro99b]. Airline [Ekl94, RAG^{+}04]. Airy [CK92a]. Akera [RAG^{+}04]. al [Cop23]. Alamos [HPR14a, Lew17]. Alan [Bar07, Hsu20, Sau93]. Albert [Ano12a]. Alberta [SW96, Smi07]. Alcoa [Ges08]. Aldrich [Ano00b]. Alfred [Nyi11]. Algol [Alb14, AD14, Bra13, Dur14, DA14, MK14, Nof10]. algorithm [RG02]. Algorithmic [DA14, GG92, TYD21b]. Algorithms [Dry18, RG02, Fit23]. All-Important [Nof23]. Allan [BH99, Wil00c, Wil00d]. Allen [Sim98]. Allocation [VT21]. Almanac} [Cro03b]. alone [Bis98]. Alper [Sid20]. Alternate [Wil98a]. Alternative [Alp15]. Alto [Spr18b]. Alvar [Fai24]. Always [SB20]. Alwin [Tod93]. Amateur [Got14]. Amazon [Léc21]. Ambiguous [Lee97, Sjö11]. Amendment [AG19]. America [Alb10, Cor25a, Mag09, McK22a, da 15b]. American [Ano00b, Ken22, AW94, Bur02, Cor96, Cor07, Dia12, Don10, Hai10a, Hei97, Hei11, Mai16, Mor19, Rio00, Sch17, Uss96, Vog17]. Americanization [Alb10]. Amiga [Rus13c, Sza18]. among [Kid99b]. Amsterdam [AD14]. Amy [Fit23]. Anacom [Asp93a, Ges08]. Analog [Asp93a, Cly93, Coh15, Col14, Hem14a, Kra22, Sma93, Bow96c, Joh96, Per99, Puc96, Tym96]. Analysis [And22, Neu09]. Analyst [Wil03, Too96]. Analytical [Bro00, Wil00c, Bro98]. analyzer [Hol96, Tym96]. analyzers [Hai03a]. Ancestor [Har13]. Ancient [GB13, de 59]. Anderson [Gui99]. Andrew [CH15]. Andries [Mac98]. Anecdotal [Smi01b]. Anecdotes [Ano97a, Ano02b, Ano10a, Ber03f, Fit02a, Fit02b, Fit03a, Fit03b, FAP^{+}03, Fit03c, FRB04, Gri01a, Kap97, LM06, MS09, Pat06, Rob05a, Rob04a, Rob04b, Rob04c, Rob05b, Rob05c, Rob05d, Rob05e, Roe09, Tom92a, Tom92b, Tom92c, Tom93a, Tom93b, Tom93c, Tom94, Tom95b, Tom95a, Tom95c, Tom99, Tom00a, Tom00c, Tre01]. Animated [Was13]. Annals [Ano03a, Ano06, Ano07, Ano19f, Ano20f, Ano20g, Ano21h, Ano22g, Ano22h, Ano22i, Ano23e, Ano23f, Ano23g, Ano24h, Ano24i, Ano24j, Ano98b, Ano19b, Ano20-27, Cer02, Gal04, Ano20-51, Ano21-52, Ano22-68, Ano23-58, Ano24-61]. Anne [Nik17a]. Anne-Louise [Nik17a]. Anniversary [Kid96a, Wal19, Win96]. Annual [Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05, Ano06, Ano07, Ano09a, Ano09b, Ano11a, Ano13a, Dod00, Wal19]. Answer [Ber14, Cor04]. Anthropomorphosis [Sta20b]. Antikythera [Ash76, Bal76, Bro77, Dra76a, Nas75, Nor77, Ole77, Smi78, Bro86, Mor84, Spi08, de 74]. Antiquity [Smi78]. Antivirus [ZS23]. anyone [Kis99b]. Aotearoa [CM25]. Apollo [Kid99d]. App [Ano24-32, Ano24-33, SH19, Ano24-34]. Application [Cas06, Cor97]. Applications [Ano97b, GR02, Ray02, Sch06a, SST22, Whe92b, Cor96, Ano22-31, Ano19h, Ano20b, Ano20j, Ano22j, Ano23h, Ano24k, Ano24l]. Applicon [Ric24]. Applied [CD14, Sch23]. Appreciation [Mor84]. Apprehension [Tha17]. Approach [Lat13]. Appropriating [Alb10, Don10]. Appropriation [BLH12, Yos13a]. Appropriations [Ken22]. April [Cor24a]. Arc [Ges07b]. Archeologist [Ber03d]. Architect [And07b, Nik17b]. Architects [Kid99e, MK20]. Architectural [Hal18, Cer97]. Architecture [Ano20h, Ano20i, MA96, Paw22, VT21, Fal98, Roj97, Ano24s]. Architectures [Lav93]. Archive [ORB^{+}92, Sha00]. Archives [BK01, Fei10, ORB^{+}92].

Archway [Laz24]. Area [Ges06b, Ges09a]. Argentina [Car15]. Arithmetic [Roe15, Roe16b, Wil97a]. Arming [AW94]. ARPA [O’N95]. ARPANET [FC15, O’N95, Wal19, VR22, WMB22, FC16, Kir99, Luk11, WI14]. Array [Smi11]. Art [Col14, Gar21, McD10a, OB08]. Arthur [Asp07, Asp21, Mis07a, Wei92e]. article [Ano98b]. Articles [Ano20l, Ano21l, Ano21m, Ano21k, Ano22m, Ano22u, Ano22k, Ano22l, Ano24m]. Artifacts [Ens07, HC20, Yam18]. Artificial [Ano21-46, Ano21-47, Ano22p, Ano22-63, Ano23-52, Ano23-53, BG22, Har22b, Kan21, Kli11, PS23, Shi19, dPRG21, vH23, Ada96, Kan21]. artillery [Per99]. Artist [Gri22c]. Arts [SWH18]. Asia [Cho16, RCP^{+}13]. Asian [NM09]. Ask [Nov16]. Asociación [FHD17]. Aspects [DMS18]. Aspray [VM18]. Assembling [Ñañ22]. assembly [Hol97]. Assertions [Hoa03]. Assessing [CG20]. Assisted [BH95]. Assistive [Alp15]. Associates [Moo03]. Association [CCK00]. associative [Ber03f]. Assurance [CK94]. Astonishing [RC23]. Astronomical [Roe12, SB20]. ATT [Sim25]. Atanasoff [Gri00a, Lee95d, Wil01b]. Atex [Dra18, Sey18a]. Atlas [How99]. Atsushi [RAG^{+}04]. Attempt [Bau06]. Attention [Cor22, Fra23a]. Audio [Gri22a]. Augmentative [Alp15]. Augusta [Lee00a]. Australia [BT23a, Mac03]. Australian [Ain23, Tat13]. Author [Ano13a, Ano18a, Ano24b, Ano25a, Ano00a]. Authoritarian [Per20]. Authoritarianism [da 15a]. Authoritative [RAG^{+}04]. Authority [Gal11]. Authors [Ano92c]. autobiography [Hal00]. Autodesk [HG24]. Automata [Kli11]. Automated [Ekl94]. Automatic [Hus97, Jon04, VT21, Kis95]. Automatics [dPRG21]. Automating [Mar02]. Automation [BB02, BGM02, Hai06, McC02, Sla15, Yos23, De 95, Mac95, Ken22]. Autoscritcher [CF92]. Availability [EHP22]. Award [Ano16h, Ano21a, Ano22-69, Ano15a, Ano17a, Ano17l, Ano19a, Ano21b, Ano21a, Ano23z]. Awards [Ano16a, Ano17d, Ano19o, Ano19p, Ano20r, Ano20s, Ano21-37, Ano21-38, Ano23n, Ano23-41, Ano24-60, Ano25-27, Bow98b, Ano17y, Ano22-51]. Away [Est22].

B [Fit23, VM24]. B-Splines [VM24]. B.C. [Bro77, de 74, Bal76, Dra76a, Nas75, Nor77, Ole77, Smi78]. B5000 [GS03, GS09]. Babbage [Ano00b, Ano17a, Ano21a, Asp07, CK01b, Gra24, Lee95e, Mis07a, Nor01, Sei95, Wil00c, Yos21, Ber92, Bro98, Bro00, BW00, CK94, CK00, CK01a, Dod00, FF03, GG92, Gri10, Hym92, LaV24, Rob93, Roe09, Roj21, Swa05a, Wil00a, Wil00b, Wil98b]. Babel [Nof25]. Baby [Cop11b, Bur05]. Bachman [Hai11]. Back [Fel11, Mul03, van19, Tym96]. Bag [Cor13]. Baggy [DAT25]. Balances [Vid20]. Balls [Was13]. Baltimore [GH18b]. Banking [FM93, Mar12, MF93, Tol21, McK95]. Banks [Neu04]. Barcelona [FH14]. Barclays [Mar12]. Bardeen [Sei03]. Base [Hai09]. Based [Nov11, Ste03, CMM95, Est00]. Bassett [Mor03]. Battalion [Hea01a]. Battle [Lus21]. Battles [Mul20, Kid00e]. Bauer [Léc21]. BBC [Jon04]. BBN [Bar05, Ber05, Cas06, Feu06, FC16, Hea05, Lev05, Mak06, PB06, Sch06a, Swe05, Wei06]. BC [Ash76]. Be [Ano16i, Ano17k, Gal04, Tes18]. beaten [Joh99]. Beatrice [Cam03b]. Beauchamp [Win03]. Beautiful [CCK00, Gon24b]. Became [Kat10, Pau14]. Beckmann [OB08]. Before [AC16, BMFB25, Fra04, Ger04, GS03, LaV24, Lig03, Mor20, Pol97, Roe16a, RAG^{+}04, Tak96]. began [Joh98]. Beginning [BGM02, Car15, Dar12, FBK06, Kid00d]. Beginnings [Cro93, LML92, Mur09, Neb98]. Behavioral [Swe05]. Behind [Cam03a].

Being [Bur23]. Belarus [Sto99]. Bell [RAG^{+}04, Sim25, Irv01, Spi15b, Tra24]. Bells [OK07]. Benchmark [Kap07]. Benson [Dia16]. bent [Cos96]. Beranek [WN05, WN06]. Berkeley [Lon04, Yat97]. Berlin [Kir20]. Bernard [Fra23b, HW03b, Ake08, Hai04b]. Bernie [Hai04b, Ano01b]. Berry [BS96]. BESK [Pet05]. Best [Fev13, Ano17o, Ano17p, Cor24b]. Bets [HPR14a]. Better [Ano17w, Ano17x, Ano17v]. Between [Cor21a, Gre05, Kel22, RAG^{+}04, SR14, VHCSCC15, DMD18, DeN15a, De 95, Yat97, Ger04]. Beyond [CP23, Sta19]. Bibliographic [Ano97b]. Bibliography [Ano92e]. Big [Ano16m, Ano19k, Ano19l, Ano20k, Ano23j, DAT25, Shr24, Joh99, Med08, Ano25g]. Bigelow [Wal18]. Bill [Eck07b]. Billion [Gra21b]. Biographies [Ada05, Ada06, AD07, And09, Ano92c, Ano97c, Ano98a, Ber06c, DSW92, DH08, For97, HW03a, Hai03a, HW03b, Hai03b, Hai04a, Hai04b, Hai04c, Hai05a, Hai05b, Hai05c, LB06, Lee99a, Lee00b, Lee01d, Lee01c, Lee01a, Lee01b, Lee02, LH03, Pol03, RAG^{+}04, Van09, Wei92e, Wei92b, Wei92c, Wei92d, Wei93d, Wei93a, Wei93b, Wei93c, Wei94b, Wei94c, Wei94d, Wei94e, Wei95c, Wei95a, Wei95b, Wei96a, Wil92b]. Biography [Asp21, Gio93, Gri11a]. Biological [Cob23]. Biology [Sau93, Ste22]. Biomedical [Nov11, Rus13a]. Bion [Kid97b]. Birth [CG94, Lip15, McQ09, Urq24, Kid99a, Win96]. Birthday [Lee93d]. Bitnet [GC00]. Black [McG21, Cou95, Nor23, McG21]. Blackett [And07b]. Blaise [Kis98]. Blanch [Gri97]. Bletchley [Ano97e, Wei94a, Ano22-67, LH95, ORB^{+}92, Sal05]. Blomkamp [RCP^{+}13]. Blue [Joh99, Med08]. Board [Ano09c, Ano09d, Ano15e, Ano15f, Ano19s, Ano15d, Ano19-34, Ano19-35, Ano21-34, Ano21-35, Ano21-36, Ano22-47, Ano22-48, Ano22-49, Ano22-50]. Boat [Ges07a]. Bob [Car23]. Body [Ano97d, Lea23]. Boeblingen [Gan04a, Gan04b, Roj04a, End04]. Boeing [Kas24a]. Bohn [Wal20]. Bolt [WN05, WN06]. Bombes [EMW02, LBA00]. Bond [McJ20]. Bone [Roe17]. Bonner [Moo03]. Book [Ake00, Ano97b, Ano97e, Ano97d, Ano97f, Bal76, Bar03, BS11, Ber03d, Bro77, Cam22, CK97, Cer03, Dra76a, EA17, Fie03, Ger04, Gui99, Kid99a, Lee97, Lee99b, Mis99, Mor03, Nas75, Nor77, Ole77, Pos00, Rio00, Roj04a, Roj04b, Rus16d, Rus17, Sei03, Sid20, Smi99a, Smi02, Smi03, Smi78, Smi99b, Spi03, Spr18a, Tym99, Wal00, Wei94a, Wil00a, Wil01a, Win03, Yos22, Rus15c, Rus16a, Lip15]. Books [Smi03, Ano14e, CB19, GH18b, KW19, Rus13b, Rus14b, Rus14a, Rus15d, VM18]. Bootstrapping [Bar03]. Boroughs [Ges09b]. Both [Fev13]. Bottomless [Med08]. Boucas [Per20]. Bouncing [Was13]. Boundaries [Jes07]. Boys [Pet05]. Brain [Ano00b, Bow96c, Wil00a]. Brake [Hic08]. Branch [BBW18, BBW19, Cor17]. Brazil [Per20, RC15, da 15a]. Brazilian [CdLF^{+}22]. Breaking [GM96, LH95]. Bresenham [RG02]. Brief [DAGVR^{+}15, Léc21, Rob19, Tak96, Tri01, Wir08, Kid00h]. Brinch [LH03]. Britain [Cop23, Mac03]. British [Bow96c, CCK00, Hic10, Urq24]. Brno [SWH18]. Broadcasts [Jon04]. Brogramming [Hic13]. Bromley [Wil00c, Wil00d]. Brooks [Hsu23]. Bruderer [Kra22]. BT [Per99]. BTM [Yos13a]. Buffer [Sho01]. bug [Kid98, Ekl92]. Build [Ano00b, Roj04a, Wil00a]. Build-Up [Roj04a]. Building [ACCH23, Ake98, DGG^{+}08, ERS03, Hai03a, Wil01a, Lin16]. Buildup [Gan04a]. Built [Hea03, Luk11]. Bundling [Pug02]. Bureau [Gur96, Hai04b, Kir98, vdE94]. Bureaucratic [Hey08a, Hey08b]. Burroughs [Bra13, GS03, GS07, GS09, Kos03a, Mar12, Yos13a]. Bush [Ano00b, Kid96f, OK07, Puc96, Rio00]. Business [All18, Cam22, Cor06a, Cor07, Gra13, Har22a, Kid99e, Ste22, Lan00, Old95, Uss96]. Busman [Ges06a]. Byron [Too96]. BYTE [Kid96a].

C [Kit03, Shr24, Yam19a, Yat97, Nof23]. ca [Ash76, Bal76, Bro77, Dra76a, Nas75, Nor77, Ole77, Smi78, de 74, Ano24r, Kru09]. Cable [Lig03]. CAD [Bal24, Ber24, Bis98, Bro23, EK24, Eva24, Gra24, GK24, Kas24a, Kas24b, Laz24]. CAD/CAM [Bal24, Kas24b]. CAD/CAM/CAE [Kas24a]. CAD/MCAE [EK24]. CAE [Kas24a]. CAL [MR23]. Calculating [GHV05, Jon04]. Calculation [Alo04, Smi12]. Calculations [Swe98, vdE92, Pol97]. Calculator [Kis01, Roe15, Roe16b, Kis95, Lee00c]. Calculators [Ano92d, Fla98, Hen10, Kis01, Pre96, RP12, San13, Swa95a, Swa95b, Swa95c]. Calculus [Car15]. Calcumeter [Otn98]. Calendar [Ash76, Bal76, Bro77, Dra76a, Nas75, Nor77, Ole77, Smi99a, Smi78, Smi00, de 74]. Calendrical [Swe98]. California [CK12]. Call [Ano15b, Ano15c, Ano16h, Ano17d, Ano18b, Ano19o, Ano19p, Ano20l, Ano20n, Ano20m, Ano20o, Ano20-32, Ano20-33, Ano20-34, Ano21l, Ano21m, Ano21n, Ano21a, Ano21-28, Ano21-29, Ano21-30, Ano21k, Ano22b, Ano22m, Ano22k, Ano22l, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22t, Ano22s, Ano22-30, Ano22-35, Ano22-36, Ano23n, Ano23r, Ano23o, Ano23p, Ano23q, Ano23-27, Ano24t, Ano24s, Ano24u, Ano24v, Ano24w, Ano24-36, Ano24-37, Ano24m, Ano24o, Ano25l]. Caller [Ano16a]. Calls [Ano22u]. Calvin [Wei95c, Wil01b].

CAM [Kas24a, Bal24, Kas24b]. Cambridge [GH18b, Sid20, Ben93a, CK92b, CK92c, CK98, Cro92, Eva24, Ges07b, Nee92, PHL92]. Came [Gal04, Tes18]. Campbell [RAG^{+}04]. Campbell-Kelly [RAG^{+}04]. Campuses [Ran14]. Can [Cor22]. Canada [Cam03b, Dor94, DAT25, Ges07a]. Canadian [Var94]. Canyon [Cou95]. CAP [Nee92]. Capability [Lin16]. capacity [Uss96]. Capitalism [Fid18, Sch17]. Capitalisms [KC20, KD21]. Capture [GA22]. Capturing [BT23a]. Card [Kis92b, Kis05, vdE94, Hei97, Kis97]. Cards [Wil02]. Career [Ano01b, Ano16j, Ano20t, Ano20-40, Ano23-28, Ano23-29, Ano23-30, Ano24r, Ano24-38, Ano24-39, Ano24-40, Ano25m, Ano25n, Nel22, Pri24, Ano17b, Ano17c]. Careers [McK22b]. Carl [Bar97, Ber06c]. Carlo [HPR14a]. Carr [Lee01d]. Carries [Roe17]. Carrying [Cor13]. Carteron [MK10]. Case [Ayl10a, BB02, Cro99, CdLF^{+}22, Kid23, Lei17, Lan00, Per99, Veh99]. Cash [BLR11]. Cash-Dispensing [BLR11]. Cashless [Smi16b]. Casinos [Tur11]. Caught [De 95, BW00, Wil00b]. CBI [Ano02e, BK01, Yos01, Yos22]. CBI/Tomash [Yos01]. CBI’s [Yos23]. CDC [Kos07, Neu09]. ceiling [Goy96]. Celebrates [Woo95]. Celebrating [Shr24, Win96]. Celebrities [Gri11a]. Census [GA22, LM94, Kis97]. Census-Taking [LM94]. Center [Ano16i, Ano17k, Ano23-29, Ano23-30, Gri07a, Kid00a, Ano23-28, Ano24r, Ano24-38, Ano24-39, Ano24-40, Ano25m, Ano25n, Fei10, Mag20, Min07b, ORB^{+}92]. Centers [Wil02]. Central [vdE94]. Centre [Wil94b]. Centro [dÁ24]. Century [Ano00b, Ano22-59, Ano22-61, Ano23-49, Ano23-50, Kid19b, MPDT12, Sha23, Mar10a, Rio00, Ano22-60, Cor25b, Cro03a, Cro03b, EK24]. century-long [Mar10a]. CEO [Moo03]. Ceremony [AH23, Yam19a]. Certifies [Yam18]. Ceruzzi [RCP^{+}13]. CFP [Ano25g, Ano25i, Ano25h, Ano25y, Ano25z]. CGA [Ano19q, Ano24x]. Chains [Bur24]. Challenge [Lee97]. Challenges [Del08, MR22]. Chalmers [Joh96]. Champaign [BB02]. Champion [Nik17a]. Chang [Pat24]. Change [BLH12, Dra18, Gio93, Sei97, Var94]. Changed [Hic10, Mis07c, Pat14]. Changes [KH21a, KH21b]. Changing [Nor96, Rus13c, Vog17, Owe96]. Chaotic [Abb10]. Chapter [CK92a]. Character [Par16a]. Characteristics [Wie10]. Charisma [Der20]. Charles [Ano00b, Ano17a, Ano21a, Eck07b, Ano99a, Asp07, Bro98, CK94, CK00, CK01a, CK01b, Dod00, Gra24, GG92, Gri10, Hai11, Mis07a, Nor01, Roj21, Sei95, Swa05a, Wal18, Wil00a, Wil98b, Yos21]. Chatbots [LHS^{+}25]. ChatGPT’s [RC23]. Checks [Vid20]. Chemists [HW17]. Cheque [De 95]. Chicago [Kid01d]. Chief [And07b, Ano14h, Ano22-62]. Child [Tha17]. Children [Ran14]. Chile [AG12, Med08]. Chilean [Med05]. China [Wu15, ZS23]. Chinese [Ano92d, JB07]. Chip [Maz07]. CHM [Bro23, Spi17]. Choice [Ano24b, Ano25a]. Choose [Ano16j]. Chris [Shr24]. Christine [Ber03d]. Chromium [Hai01]. Chromium-Plated [Hai01]. chronology [Cor98, Sto99]. Cincom [Nie09]. cipher [CF92, Kid00k]. Circuits [Gre13]. Circulation [Ake04]. Cis [Hic19]. Cis-tem [Hic19]. CiSE [Ano25j]. Cities [Rei23a]. Citizens [Per20]. Claims [Lee92e, Wei96b]. Clamor [Day16]. Claritas [McK22a]. Clark [Sha23]. Clarke [Lee01d]. Class [Est22, Wil01a]. Classification [Tha17, Kid00j]. Classified [Poi25]. Classrooms [Ran14]. Clearing [De 95]. Clement [Rob93, Wil92b]. Cleve [Hai08].

Clinics [Lus21]. Clippinger [Lee98a]. Cloned [Ñañ22]. Clones [Sta15]. Closed [Lla24]. Clothes [Hic10]. Cloud [Ano14g]. Club [Pet17]. Clustering [McK22a]. Coast [ERS03]. Coaxial [Roe17]. Code [Buc20, HPR14b, Par16a, Pri21, Rus06, Fra23b]. Codebreaker [And07a]. Codebreakers [Wei94a]. Codebreaking [Hai18, Smi01a]. Codes [GM96, LH95]. Codevelopment [DF16]. coevolution [Bar03]. Coffee [Cor21a]. Cognitive [Bur23]. Cogwheel [Ano00b, Wil00a]. Cohen [HW03b]. Cold [Don10, Tat10, Tin10]. collaboration [Est00]. Collect [Shu06]. Collected [Kid96b]. Collecting [GJ12, Gra21a]. Collection [Wil01a, SWH18]. Collective [Med05]. Collectives [Gri22c]. collector [Bro99a]. Colombian [Ñañ22]. Colombian-Cloned [Ñañ22]. Colonization [Est22]. Colorant [Hem14a]. Colossus [Ano97e, Cop04a, HP18, Hai18, Lav06, Sal05, Wil18]. Column [Sei95]. Com [Fei11]. Combining [Ekl94]. Come [Sta20b]. Comes [Day16, Sub14]. COMIC [Hem14a]. Coming [Joh06]. Command [Asp99, Sta20b]. Comments [Ben93a, Ben93b, Ber92, Buc95a, Buc95b, BS96, Car92, Dum94, Ekl92, GM96, Gro92, Gro93, Hal92, Hym92, Jac93, Kis92b, Kis92a, Rob93, Ros92, Sha00, Tod93, Tro93, Tro95, WF92, Wil00b]. Commercial [Bal24, Cor96, Dow09, Sta14, Cor97]. Commercialization [BH09, Kel22, Ste22]. Commercializing [BG22]. Committee [Ano18m, Ano18n, CCK00, Gri01b]. Commodity [DK23]. Commodore [Rus13c]. Common [All18, McK95]. Communication [BMFB25, Bar07, Kid99e, Sid20]. Communications [Ekl94]. Communities [Ran14]. Community [DGG^{+}08, PPM^{+}23, Est00, Nor96]. Companies [Hai05b, Mor03, Oba22]. Company [EK24, JMSV10, Joh02, Law10, RAG^{+}04, Klu99]. Comparative [CK97]. Comparison [Bab97]. Compatibility [Sum07]. Compatible [Tak05, Lee92f]. Compelling [Nof25]. competence [CMM95]. Competing [Abb10]. competition [CMM95]. competitive [Asp00]. Compiler [End13, McJ17a]. Compiling [Hol94, Wad12]. Complete [Ano00a, Smi01a]. Complex [Dow09, Low09, Nik17b]. compromise [Uss96]. COMPSAC [Ano20p, Ano18c]. Computation [Den15b, Der20, EF11, Ens13, Gro93, RG02, Sei00, Gri01b, Pol95, Wil94b]. Computational [Fer22, HW17, Sha23, SK11]. Computer [ACCH23, Ain23, Alo04, And09, Ano92d, Ano97b, Ano00b, Ano09e, Ano15m, Ano16h, Ano16i, Ano16g, Ano16j, Ano17n, Ano17l, Ano17m, Ano17k, Ano18k, Ano18j, Ano19g, Ano19h, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-34, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-35, Ano20y, Ano20z, Ano20h, Ano20i, Ano20j, Ano20-29, Ano20-35, Ano20-30, Ano20-32, Ano20-36, Ano20-33, Ano20-31, Ano20-34, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21-28, Ano21-31, Ano21y, Ano21-29, Ano21-34, Ano21-37, Ano21-32, Ano21z, Ano21-30, Ano21-38, Ano21-35, Ano21-33, Ano21-27, Ano21-36, Ano22n, Ano22o, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22j, Ano22-31, Ano22-32, Ano22-35, Ano22-37, Ano22-41, Ano22-47, Ano22-33, Ano22-36, Ano22-38, Ano22-42, Ano22-48, Ano22-46, Ano22-39]. Computer [Ano22-43, Ano22-49, Ano22-44, Ano22-34, Ano22-51, Ano22-45, Ano22-40, Ano22-50, Ano23o, Ano23p, Ano23q, Ano23y, Ano23h, Ano23-33, Ano23-37, Ano23-28, Ano23-31, Ano23-34, Ano23-38, Ano23-35, Ano23-29, Ano23-32, Ano23-39, Ano23-27, Ano23-41, Ano23-36, Ano23-40, Ano23-30, Ano24s, Ano24u, Ano24v, Ano24k, Ano24l, Ano24-41, Ano24-44, Ano24-42, Ano24-38, Ano24-45, Ano24-43, Ano24-39, Ano24-36, Ano24-35, Ano24-37, Ano24-40, Ano25d, AH23, Ash76, Asp01a, AG21, Bal76, Bar07, BLH12, Bau07, BS11, Ber14, Bow98b, Bra22, Bro77, BH95, Cam22, CK97, CKGS08, Cam03b, Car15, Car20, Cer25, CK08, CRB24, Cop11a, Cop11b, Cor24a, Cor01b, Cor25b, Cou95, Cro93, Cro18, Dav17, Day23, DMD18, De 21, Don10, Dra76a, DF16, ERS03, EF11, Ens01, Est00, Est02, Fed21, FVW25]. Computer [FBK06, Fra13, Ges07a, Gil07, GH93, GG46, Gra24, Gri03a, Gri11a, Gru05, Gup07, Hai10a, HO23, Hem14a, Hem16, Hem19, Hey08a, Hey08b, How99, Hum94, Hun12, Ich06, Jac16, Jes13, JMSV10, JMV10, Jor09, Ken22, Kid97a, Kid99e, Kli99, Kne21, Kru94, Lap11, LS08, Lav93, Lav22a, Lee97, LS00, Lla24, Lon08, Low09, Maz08, McG12, Med05, Mir17, Mis16, Mol14, Moo01, MK20, Ñañ22, Nas75, Neu05, Nof23, Nor77, Not20, NK98, Nyi11, ORB^{+}92, Oba22, OB08, Ole77, OARHC10, Par07, Pet05, Pet94, Pet17, Pet20, PA96, Ral04, Ran94, Ray02, Rei23a, RR04, Ric24, Roj04b, Roj21, Rop07, SS09, Sch08, Sch23, Sie10, Sla15, Smi78, Spi08, Spr18a, Sta13, Sta15, Sta20b, Ste03, SW95, Sub12, Tak96, Tat13, Twe93, Van10b, Ver11]. Computer [War15, Was13, Wei96b, Wie10, Wil01b, Wil03, WD98, Wyv98, Yos13a, Yos13b, Yos15a, Yos15b, Yos16a, ZR06, da 15a, de 59, Aga01, Ake00, Ano99b, Bro99a, Bro99b, Cer97, Cor96, Cro00b, DFM^{+}99, Eck98, Est96, Fal98, Gas99, Gre95, Kat97, Kid98, Kid99c, Kid00c, Kid00e, KS97, Kit08b, Klu99, Lan00, Lee95f, MP97, Mar10b, Mar10a, Old95, Per99, Roj98, Sha00, Tom00b, Wil00a, Wil92b, Wri98, Yat97, de 74, Ano18l, Ano19-36, ERS03, Gio93, GG96, Gra24, GR02, GH18b, Hus97, Kid99d, Min07a, Min07b, Spi21, Spi23a, Spi24a, WH17, Woo95, Smi02, Ano20a, Ano20b, Ano20-50, Ano22t]. Computer-Aided [Cro18, Gra24, Lla24, ORB^{+}92, Ric24]. Computer-Assisted [BH95]. Computer-Based [Ste03]. Computer-Generated [Fra13]. Computerization [Lap11, Paj08, Joh99, Nor98]. Computerized [Dor94, Fra23a, Lus21, SRG02]. Computerizing [Sey18b, VHCSCC15]. Computers [Asp13, Ayl10a, Ber06a, Cas06, Cer25, CP23, Cly93, Cor08a, Ens07, Fel11, Gal11, Gon24a, Got14, GS01, GS04, GS09, Irv01, KKW09, Lon04, Moo03, Neu04, Nis23, Nov16, ORB^{+}92, Pro99, Roj04b, Smi04, Sto99, Swe05, TZ99, Tin11, Tom96, VN11, Van95, Was13, Wes17, Yam19b, Coh98, Cor98, DMPS99, GS98, Gri96, Hal00, Har99, Joh96, Kid96g, Puc96, Sei00, WD98, GS03, OB08, Ano20m, Ano24o]. Computervision [Laz24, VM24]. Computing [Abb03, Ake04, Alb10, Alo07, Alp15, AG12, And07a, And07b, And22, Ano22-30, Asp93a, AW94, Asp07, Ben93a, BT23a, BT23b, Bru17, Bul15, CK92b, CK01b, CK02b, Cam13, CD14, Cer01, Cer03, Cer14, CVRM15, Cho16, Cla10, Coh15, Col14, CH15, Cor04, Cor06a, Cro92, Cro99, Cro03a, Cro03b, CG94, Dav17, DT21, Dia12, DAGVR^{+}15, DGG^{+}08, Dra76b, DMS18, Dun21, Dur10, Ekl94, Ens04b, Ens18a, Ens18b, Fei22, FH14, FHD17, Ger04, GB13, Gra21a, Gra22, Gra02b, Hai10a, Hal18, HC20, Hen05, HIIT05, HII^{+}05, Hic10, Hic13, Hoc22, HP24, HmWK23, Jes07, JB07, Jon19, KB21, KC20, KD21, Kid96c, Kid19b, Kid23, KBSW11, Lap09, Lav22b, LBA00, Lew16, Lew17, Lun18, Mac03, Med05, MPDT12, Mis07a, Mis07b, Mis07c, Mul03, Nan09]. Computing [Nel22, Nel17, Nik17a, Nik17b, Nor23, Nov11, Par07, PM12, PN16, Raj15, Ran14, Raw24, RAG^{+}04, Rus13a, Rus13c, Sch06a, Sch06b, Sjö11, Sla08, Sma93, SW96, Smi03, Smi07, Spr16, Sta20b, Sum07, Tin10, Tol21, Tra24, Tri01, TMS14, Vog17, WN05, WN06, Wei08, Wu21, Yoo05, Yos21, da 15b, dPRG21, Ano24-54, Asp00, Bar97, Bar03, BH99, Cor97, FAP^{+}03, Goy96, Gri98, Gri00a, Gur96, ICKDL99, Lee96, MH11, Nor96, Owe96, Sze99, Uss96, Veh99, Win96, Ano03a, Ano06, Ano07, Ano13b, Car23, CdLF^{+}22, ORB^{+}92, RAG^{+}04, RCP^{+}13, Kra22, Ano19c, Ano19d, Ano19e, Ano19f, Ano19i, Ano19m, Ano19n, Ano20d, Ano20c, Ano20f, Ano20g, Ano21c, Ano21d, Ano21h, Ano22c, Ano22g, Ano22h, Ano22i, Ano23a, Ano23e, Ano23f, Ano23g, Ano23k, Ano23l]. Computing [Ano23m, Ano24d, Ano24e, Ano24h, Ano24i, Ano24j, Ano24m, Ano25b, Ano25c, Ano25h, Ano24c]. ComputingEdge [Ano21o, Ano20e, Ano21e, Ano21f, Ano21g, Ano22d, Ano22e, Ano22f, Ano23b, Ano23c, Ano23d, Ano24f, Ano24g]. Comrie [Cro99]. COMTRAC [Has16]. Concept [BK01, Gri12, Hai02, HPR14c, Sjö11]. Conception [McK11, Pat09]. Concepts [Roy11, Bar97]. Conceptualization [Geo08]. Conceptualizations [Pet17]. concern [Kat97]. Concise [RCP^{+}13]. Conclusion [Gra02b]. conclusions [Kis98]. Condition [Ano97f]. Conference [Ano02e, Ano20n, Ano22p, Kli11, ORB^{+}92, PP21]. Conferences [Rei08, Shr24]. Conferencing [Sub12]. Configuring [Sou19]. Congress [Ano19-37]. Connect [Ano19r]. Connected [Ano19-47, Ano23-44, Ano24-34, Ano24-52, Ano24-53, Ano15q]. Conner [Sav23]. Consciousness [Sub20]. Consensus [Rus06]. Consent [Jon19]. consequences [Kid00j, Nor98]. Conservation [FVW25]. Consolidation [AG12, Ber06b]. Consortium [Tol21]. constituted [Asp97]. Construction [Kid97b, Swa05a, Vag24, Fal98]. Constructions [Ada96]. Consumer [Spr16]. Consumer-Led [Spr16]. Consumption [Yoo05]. contemporary [Bar97]. Content [Nof23, Kat97]. Contents [Ano13i, Ano14i, Ano15r, Ano15s, Ano15t, Ano15u, Ano16n, Ano16o, Ano16p, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano18z, Ano18w, Ano18x, Ano18y, Ano19-49, Ano19-50, Ano19-51, Ano19-52, Ano20-46, Ano20-47, Ano20-48, Ano20-49, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano22-64, Ano22-65, Ano22-66, Ano23-54, Ano23-55, Ano23-56, Ano23-57, Ano24-55, Ano24-56, Ano24-57, Ano24-58, Ano25w, Ano25x, Ano14j]. Contested [Dia16]. Context [Jor09, Mal02, Roe15, Var94, Gri00a]. Contexts [Dow04]. Contextualizing [Hai10a]. Continuum [FVW25]. Contract [Lin16]. Contributions [Asp01a, BBC^{+}09a, BBC^{+}09b, ERS03, End04, Cer03, Gur96]. Contributor [Ber09]. Control [Ayl10a, Bar05, Ger04, Got23, Has16, RAG^{+}04, Sjö11, Sta20b, SW95, Asp99, Hal00, Per99, Wor95]. controlled [Kis95]. Controversy [Par16a, Tin11, Joh99]. Convention [SWH18]. Conventional [Cor20]. Conversational [Wie10]. Conway [Spi23b]. Cooper [Hsu20]. Coordinating [Mis99]. Copies [Tin11]. Coping [ES05, Mor19]. Copper [MPDT12]. Copy [O’D09]. Corbató [Ano12b, Spi15a, Spi16a]. Core [Ano15v, Cou95]. Cores [Car92]. Corinna [Ken22]. Corner [Lew14]. Corporate [Cor21a]. Corporation [Bru09, Cou95, ERS03, Lee95f, McK22a, Pag08, Row12, Wor95]. Corporations [Lus21, Cor96]. Correction [Cor20, Hem24a, KH21a, Ros92, Tro95]. Cortada [RCP^{+}13]. Corydon [McK22b]. Costa [VHCSCC15]. costs [Gre95]. could [Kis99b]. Council [Lap11, Mar02]. Counterinsurgency [Fei22]. Counters [Roe17]. Countess [Lee00a]. Couple [Asp01a]. Courses [AKM08]. Courts [Lus21]. Cover [Ano13d, Ano15h, Ano15i, Ano15j, Ano15k, Ano16c, Ano16d, Ano16e, Ano17g, Ano17h, Ano17i, Ano17j, Ano18i, Ano18f, Ano18g, Ano18h, Ano19t, Ano19u, Ano19v, Ano19w, Ano20u, Ano20v, Ano20w, Ano20x, Ano21p, Ano21r, Ano21s, Ano22v, Ano22w, Ano22x, Ano22y, Ano23u, Ano23v, Ano23w, Ano23x, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano25r, Ano25s, Ano14a, Ano14b, Ano21q]. Covered [Ano20-35, Ano20-36, Ano21-31, Ano21-32, Ano21-33, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano24-42, Ano24-43]. CPL [Har13]. CPU [Maz07]. Cracking [Hod19]. Craft [Tha17]. Cray [Ano17y]. Craze [HVB04]. created [Rus16a, RAG^{+}04]. Creating [Cor15c, Joh02, Len22, Moo03, PA96, Kid00a]. Creation [FF03, Gra07, Lei17, Nof10, Too10]. Creative [Sou19]. Creativity [Ber05, Sha23]. Creator [Hai08]. Crews [McJ20]. Crime [Par07]. Cripping [Wu21]. Crisis [Kir20, Pay14]. Crispin [GH18b]. Critical [Cro18, Geo08]. Criticism [Gar21]. Critique [Rus12b]. Crossing [Cer97]. Crossroads [Mir17]. Crown [RAG^{+}04]. Cruelty [Pla16]. Crunching [CK07b]. cryptanalysis [CF92]. Crypto [Doy01]. Cryptographic [Rei23b]. Cryptography [DF16, Fid18]. Cryptologia [Kid99e]. Cryptonomicon [Kid00b]. Crystal [Kid99a]. Crystallography [SK11]. Crystals [Bur24]. CS [Ano17a, Ano19a, Ano21b, Ano21a, Ano14k, Ano18e, Ano18d, Ano19s, Ano23s, Ano24y, Ano24z, Ano25m, Ano25n, Ano25l, Ano25o, Ano25p, Ano25k]. CSIRAC [Ain23]. CTSS [Lee92f, RL92, Van12]. Cubicle [KB21]. Cultural [Vid23]. Culture [Alb14, Asp15, Ber05, Cor21a, MK14, TYD21a, TYD21b, Yos21]. Cultures [Fer22, Poi25]. Cunningham [Ano11f]. Cups [Cor21a]. Current [GS07, Goy96]. Curriculum [Gup07, ICKDL99]. Curtain [Cam03a, Lew16, Gaw19]. Customer [Cor24b]. Cutting [Ano21-46, Ano21-47, Ano22-63, Ano23-52, Ano23-53]. CVPR [Ano20q]. Cyber [Sub20]. Cybernetic [Aum11]. Cybernetics [Kli11, Ger04, RAG^{+}04]. Cybersecurity [Ano14f, Ano15n, Ano15o, DF16]. Cyberspace [Cer03, Fid17, Kru07]. Cyberwar [Hun12]. Cycle [Kne17]. Cycorp [Len22]. Czechoslovakia [Dur14]. Czechoslovakian [Dur10, Kis99a].

D [Ano17l, Ash76, Kis95]. D&I [Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano23-31, Ano23-32, Ano24-41]. Dahlgren [Lee00c]. Dam [Mac98]. Dame [Spi18]. Dana [Nik17b]. Dark [Par07]. DARPA [Nor96]. Dartmouth [Kli11]. Darwin [Got23, Kid99b]. Dassault [Ber24]. Data [Ack15, Ano16m, Ano20n, Ano21-45, Ano23-51, Bac09, Cor25a, DAT25, Fer22, GA22, Hai09, HVB04, HP96, Kid96d, Kos07, Lea23, MR22, PB06, WP94, Wor95, Bir00, Kis99a, Ano19k, Ano19l, Ano20k, Ano23j, Ano25g]. Data-Flow [WP94]. Databank [Tol21]. Database [BH09, GB09, Gra12, Gra13, Gri12, Hai11, Kru09, WC12]. Databasing [Fei22]. Datacom [Ste09]. Datacom/DB [Ste09]. Datamation [Cer02, Hea04]. DataPort [Ano22-52, Ano24-46, Ano24-47]. DATAR [Var94]. Date [Gil07]. David [Shr24, Kid97a, McK22b, RAG^{+}04, Smi99a]. Davis [POD23]. Dawn [RC15, Smi16a]. day [Neb98, Alo04, Hus97]. Days [Ber05, Bul15, Cor15b, Hea04, Kee04, Kee05, Lav22b, Ric24, Gri94]. DB [Ste09]. DB2 [Boh13, HS13]. DBMS [Bac09]. DDoS [BYO^{+}22, Gri22c]. De-Brogramming [Hic13]. Dealer [Was13]. Death [Lip15]. Debate [Ben93a, Ben93b, Ber92, Buc95a, Buc95b, BS96, Car92, Dum94, Ekl92, GM96, Gro92, Gro93, Hal92, Hym92, Jac93, Kis92b, Kis92a, Rob93, Ros92, Tod93, Tro93, Tro95, WF92].

Debated [Day16]. Debates [Lon08, Sha00, Wil00b]. Deborah [Spi14]. Debugging [CK92a]. decline [CCK00]. Decode [EF11]. Decoding [Mar10b, Mar10a]. Defense [FVW25]. Defining [Hey08a, Hey08b, Nof23]. DEHOMAG [Kis95]. Delay [Ben93b, Hal92, Kis92a, WF92]. Demise [Gra07]. Democracy [FHD17, da 15a]. Democratic [Cor12]. Demographic [And22]. Demographics [Cor01b, Fra23a]. Demography [MG22]. Demos [Was13]. Denial [BYO^{+}22]. Denning [Wal12]. department [Lee95f, Gio93, LS00, Min07a]. Departments [Ano20-27, Kid99e, MS09]. dependencies [DV18]. Derek [Bal76, Bro77, Dra76a, Mor84, Nas75, Ole77, Smi78, Nor77]. Desch [Jac93, LBA00]. Description [Lav06]. Deserve [Ano15l]. Design [And09, DMD18, DeN15a, Ens18a, Fre07, Gra24, Hal18, Hin18, KB21, Lin16, Lla24, Ric24, Ste03, Yos23, HTAY97, Roh99, ORB^{+}92]. Designed [Kir20]. Designer [Hai18]. Designing [Lla24]. Designs [Nor23]. Desk [Ano04c, Ast24a, Ast24b, Ast25a, Ast25b, Ber01c, Ber02a, Ber02b, Ber02c, Ber03c, Ber03e, Con19, Con20a, Con20b, Dia19a, Dia19b, Dia21, Ens15a, Ens15b, Ens16a, Ens16b, Ens17a, Ens17b, Ens17d, Ens17c, Gri04b, Gri04c, Gri04d, Gri05a, Gri05b, Gri05c, Gri05d, Gri06a, Gri06b, Gri06c, Gri06d, Gri07b, Gri07c, Gri07d, Hei12a, Hei12b, Hei13a, Hei13b, Hei13c, Hei13d, Hei14a, Hei14b, Hem20, Hem21, Hem22a, Hem22b, Hem22c, Hem23a, Hem23b, Hem24b, Hie12a, Hie12b, Sub23a, Sub23b, WE14, Yos08a, Yos08b, Yos08c, Yos08d, Yos09a, Yos09b, Yos09c, Yos09d, Yos10a, Yos10b, Yos10c, Yos10d, Yos11a, Yos11b, Yos11c, Yos11d, Hem24a, Mal02]. Desks [HA24]. Desktop [BBW18, BBW19, GH18a, GH19, GH20, Spr18a, WH17]. Desolla [Mor84]. Details [Roe16b]. Detection [ES05, Yos16b]. Determination [She12]. Determine [Smi99a]. Developed [Kas24b]. Developing [Fit03b, Kas24a, McK95, Min07a, OARHC10, Sac07, Van10b, Bar97]. Development [Ain23, BLR11, Boh13, BH95, Con15, ERS03, EM94, FM93, Gan04b, Has16, Hum94, JMV10, Jon19, Jor09, Kis05, Kne17, LR12, Lon08, Maz07, McD10a, MB09, Ñañ22, Par15, Par08, Sat11, Sie10, Spr16, Vid20, CMM95, DFM^{+}99, O’N95]. Developments [End13, Gup07, Nee92, Ano99b, Fal98]. Device [Dra76b]. Devices [Bru17, Kid19b]. Dial [HT12]. Dialectic [Gal11]. Diamond [Lat13]. Diana [Cer03]. Did [Cor04, FAP^{+}03]. Died [Har22a]. Difference [BS96, Sha97, LaV24, Rob05b, Roe09, Swa05a]. differential [Hai03a]. Diffusion [RCP^{+}13]. Digital [Ano13c, Bor08, Bul15, Coh15, Col14, ES05, Fra18, Fra23a, Gon24a, Irv01, Kra22, McJ20, Mor03, Mul03, PB22, Roe17, RCP^{+}13, San11, Smi01b, Smi16a, Sub13, TMS14, Bro99b, BH99, Cor96, Eck98, Gri96, Hal00, Har99, Per99, Rus16a, Tym96]. Digitization [Tur11]. Digitizing [Rus13a]. Dijkstra [Pay14]. Dionysius [Fra23b]. Direct [Bac09, Fra23a]. Direct-Access [Bac09]. directed [RCP^{+}13]. Directions [Mis07b]. Director [PM12, Sei95]. Dirty [Pet94]. Disability [Alp15, Sid20]. Disappeared [PB22]. Discipline [McD10a]. Disciplines [Del08]. Disclaimer [GM96]. Discourse [Mis16]. discover [Mar10a]. Discoveries [Neb98]. Discovery [Bru17, Win96]. Discussion [Wil18]. Discussions [Kid15]. Dispensing [BLR11]. Displacement [San13]. Display [MB09]. Disputes [Gon24a]. Dissecting [Lea23]. Dissent [Lew17]. Dissertations [Fra13]. Distinctiveness [Aum11]. Distinguished [Ano15a]. Distributed [AH23, BYO^{+}22, RAG^{+}04, Sch06a]. Distributing [Mul20]. Distribution [Sta14]. Disturbance [Gri22c]. Diversity [Nik17a]. divide [Cer97]. DIY [Tin11]. DLL [DV18]. Do [Wes17]. Doctor [Nov16, ZS23]. Doctoral [Fra13]. Documentation [Wil02, Asp99]. Documenting [Bro23, Gal11]. Documents [Bru17]. Doddeson [Kid99a]. Does [Gar21]. Doing [Cor15b]. Domain [Fei11, BV19]. Domains [Asp15]. Domestic [Ver11]. Donald [Smi11]. Donates [ORB^{+}92]. doomed [Lee96]. Door [Mul03]. Doorsteps [Pos22]. Doran [Car23]. Dorothy [Ano00b, Cro00a]. Dot [Fei11]. Douglas [Bar97, Bar03]. Dov [Kat18]. down [Del08]. DPMA [ORB^{+}92]. Draft [vN93]. drawing [RG02]. Dream [Kid97a, Spi03]. dreams [Tom00b]. Driven [Hen10, Roe08, Lan00]. Driving [Jor09]. DRTE [Pet94]. DS [HN13]. Dummer [Gre13]. Duncan [Smi99a]. During [Ano92d, Fei22, Gaw19, Nof25, DMD18]. Dutch [De 95, Van95, Ver11, vdE94]. DVDs [O’D09]. Dwarfs [Lee97]. Dylan [Vid23]. Dynamics [Hei11]. Dynasty [Ano92d].

Earliest [Ber05, Neb98, Har99]. Early [AL12, AG12, AC16, AG21, Boh13, BL12, Bru17, Bul15, CK92a, Cer25, Cha12, Chi14, Cop23, Cow94, End04, End13, Fal98, Fra13, Fra18, Fre07, Gal11, Ges07a, Hai10a, Hem14b, Hem16, Hen10, HIIT05, HP96, Irv01, Jes13, Joh96, Jon03, Jor09, Kee04, Kee05, Kid96d, Kid15, Kid19b, Kir99, Kos07, Lav22b, MA96, Med08, Neu04, Nov11, OB08, OARHC10, Paj08, Pee09, Phi12, Poi25, Pri24, Ric24, RP12, Roe08, Run08, SB98, Sho01, Smi07, Smi12, SST22, Spe98, Sta14, Sta20b, VN11, WC12, Yat97, Yos15b, Yos16b, dÁ24, Asp00, CK92c, CK98, Cos96, DMPS99, Gri94, Gur96, Puc96, RG02, Van95]. East [Cho16]. Easter [Alo04]. Eastern [Hen05, HIIT05, HII^{+}05]. Easy [Maz24]. Ebony [Nel17]. EC [Van10b]. Eckert [Ano19a, Ano21b, Eck96, Lee95c].

Eckert-Mauchly [Ano21b]. Ecologies [Ake04]. Economic [CKGS08, Cor97, Day16, Jac16, Sch23, Kid00f]. Economy [Fra23a, Smi16b]. ECREA [BMFB25]. Ed [Kid97a, Kra22, RCP^{+}13]. Edge [Ano21-46, Ano21-47, Ano22-63, Ano23-52, Ano23-53, DF16, Tri01, Tym99, Ano19c, Ano19d, Ano19e, Ano20d, Ano20a, Ano20c, Ano24d, Ano24e, Ano25b, Ano25c]. Editing [Tes18]. Edition [Too10]. Editions [Ral04]. Editor [Ano04c, Ano22-58, EMW02, Gri02, Gri04e, Hem24a, Hem24b, Yos08c, Yos09c, Alb14, Gra12, Abb03, Ano09e, Ano14h, Ano20-45, Ano21-43, Ast24a, Ast24b, Ast25a, Ast25b, Ber01c, Ber02a, Ber02b, Ber02c, Ber03c, Ber03e, Bös08, Con19, Con20a, Con20b, Dia19a, Dia19b, Dia21, Ens15a, Ens15b, Ens16a, Ens16b, Ens17a, Ens17b, Ens17d, Ens17c, Gra13, Gra02b, Gri04b, Gri04c, Gri04d, Gri05a, Gri05b, Gri05c, Gri05d, Gri06a, Gri06b, Gri06c, Gri06d, Gri07b, Gri07c, Gri07d, Hei12a, Hei12b, Hei13a, Hei13b, Hei13c, Hei13d, Hei14a, Hei14b, Hem20, Hem21, Hem22a, Hem22b, Hem22c, Hem23a, Hem23b, Hie12a, Hie12b, Nyc96, Sei96, Sub23a, Sub23b, WE14, Wal21, Yos08a, Yos08b, Yos08d, Yos09a, Yos09b, Yos09d, Yos10a, Yos10b, Yos10c, Yos10d, Yos11a, Yos11b, Yos11c, Yos11d]. Editorial [Ano09c, Ano09d, Ano15e, Ano15f, Buc95b, Ano15d]. Editors [Gaw19, PPM^{+}23, Ano22-62, CG06, CG07, GB09, GR02, HA24, Mis07b, Yos15a, da 15b]. Edmund [Lon04, Yat97]. Eds [Gui99, RAG^{+}04]. EDSAC [CK92c, CK98, Whe92b, Whe92a, Wil92a, Wil97a]. Education [Ano16a, Ano17d, Ano20r, Ano20s, AKM08, Cor15a, Day24, Jac16, Pro99]. Educational [Feu06, Nis23, Tat13]. Educator [Nik17a]. EDVAC [Rei97, Wil93b, vN93]. Edward [Gri13, Gui99, McC22, Out03]. Edwards [Cro03a]. Edwin [Asp93a]. Effects [And22]. Efficiency [KBSW11]. Effort [Dur14].

Efforts [HIIT05]. EIC [Wil96b]. Eight [Ral04]. El’brus [WD98]. ELEA [Mor20]. Electric [Coh15, Lee95f, LS00, Lee97, Old95]. Electrical [KBSW11, Bab97, Tym96]. Electronic [Ano97d, Hai01, Kid99e, MT06, Mir17, Neb98, Sma93, Sta14, Van12, Vid20, ZR06, Bir00, Lee00c, Sha00, GG46, GG96, Gri22c]. Electronics [CK08, Mil11, Ñañ22, Neb98, Ric24, SL12, Pee09]. Elisabeth [Lee01d]. ELIZA [LHS^{+}25]. Eloge [Wei92e, Wei93d]. elusive [Kid98]. Elysium [RCP^{+}13]. Email [Par08, Spi16b, Spi16c]. Embedded [MK14]. Embodied [Con15, Ens09]. Embracing [Dur14]. Emergence [Cro92, Ano02e, Cer97, Gri96]. Emerging [Ano25q]. Encounters [Bul21, DMD18]. Encyclopedia [RAG^{+}04, Roj04b, Ral04]. End [Fid17]. Endless [Ano00b, Rio00]. energy [Est00]. Engelbart [Bar97, Bar03, OK07]. Engine [Bro00, BS96, LaV24, Swa05a, Wil00c, Bro98, LaV24, Rob05b, Roe09]. Engineer [Ano00b, Gon20, Pag08, Rio00, Wil92b]. Engineering [Ano18m, Ano18n, BLH12, Ber05, Bre25, Cor24b, DT21, HPR14b, Jes13, Lin16, Mah04, McD10a, Sat11, Sch08, Wir08, Ari00, Bab97, DFM^{+}99, Fal98, Mey99, Sha97, Win96, Ano21c, Ano21d, Ano22c, Ano23a, Ano24c]. Engines [Kid99c, CK01a]. England [Cro03a, Cro03b, Dav17]. ENIAC [Ano00b, GG46, Ake00, Cor06a, Fri94, Fri96, GG96, Gri96, GH18b, HPR14a, HPR14b, MA96, McT98, Neu06, Pol97, Rop07, She12, Wil06, ZR06, BF96]. Enigma [CH15, CF92, Kid96e]. Enough [All19]. entered [Asp00]. enters [Old95]. enthusiasm [Bow96c]. Entrepreneurial [Oba22]. entrepreneurs [Bro99a]. entry [Asp00]. Environment [Ens13, Hea05]. Epic [Smi99a]. Episodes [Roe17]. Epistemic [Poi25]. Equipment [Cor25a].

Era [CH15, Dar12, FBK06, Wor95, Gri98]. ERMA [FM93, Hea01a, MF93]. ERMETH [Neu05]. Errata [Ano23t]. Erwin [Asp01a, Asp01b, Yos13b]. Essays [Fer22, Sei97]. Establishing [Mol06]. Establishment [Gre13]. Estrin [Spi14]. Eternal [Fid17]. Ethical [Ens07]. ETHW [Mag20]. Eulogy [Dod00, CK00]. EUNET [PB22]. Europe [RCP^{+}13, End13, Nof10, PS23]. European [Alb10, ORB^{+}92, Paj08, PB22, Sch08]. Evaluation [Fei22, GK02]. Events [Abb10, Ano01c, Ano02e, Ano02c, Ano02d, Ano04b, Ano20-27, Asp01c, Bau07, Bra22, Cam04, CE02, CE03a, CE03b, CE03c, CEG^{+}03, CRM^{+}04, FC00a, FC00b, FC01a, FC01b, Gri99, Gri04a, Kit05a, Kit05b, Kit05c, Kit05d, Kit06a, Kit06b, Kit06c, Kit06d, Kit07, KSFY07, KYG^{+}07, KFS^{+}07, KMA^{+}08, KHS^{+}08, Kit08a, Kit09a, Kit09b, Kit09c, Kit09d, Kit10a, Kit10b, Kit10c, Kit10d, Kit11a, Kit11b, Kit11c, Kit11d, Kit12c, Kit12a, Kit12b, Kit13a, Kit13b, Kit13c, Kit13d, Kit14b, Kit14c, Kit14a, Kit14d, Kit15a, Kit15b, Kit15c, Kit15d, Kit16a, Kit16b, Kit16c, Kit16d, WS17, YS12, Cor98, Sto99]. Ever’ [Cam13]. Everyday [Asp13]. Everyone [CG06, CG07, Gra21a]. Evidence [Nov11, Sha00]. Evidence-Based [Nov11]. Evolution [BBW19, CK97, CK07b, Cor07, Gra02b, GK02, Kel22, Lei08, MK20, Nan09, Rob24, War15, Kid99b, Kid99c, Pos98]. Evolving [Ano20t]. Ewing [Smi99a]. examples [Joh96]. Excavating [Bur24]. Excursion [Ges07a]. execution [CMM95]. Exercise [Bro10]. Exhibition [Sza18, Gon20, Yam19b]. Exist [Gar21]. Expansion [AG12, Cla10, Per20, Sch08]. Experience [Fra13, De 95, HTAY97, MR22]. Experiences [Mac03, Mil06, Wyv98, Kir99, Mey99]. Experiment [Gon24b]. Experimental [Dav17]. Experiments [Hem16]. Expert [BG22, Gre05, Har22a, Yos16b]. Expertise [Sou19]. Experts [Ano15v, Lew16]. Explanations [Cor20]. Exploitation [Tha17]. exploration [Mac95]. Exploring [MA96, Sei00]. Explosion [Gra13]. External [Wie10]. Extra [Ano02e]. Eyes [Cor06b, Lew16]. Eyewitness [McK11].

F [Est02, Lee98a, Van09]. Fabrications [RC23]. Faces [Gru05]. Facilities [AW94]. Facit [Pet05, San13]. FACT [Hea03]. Factory [Lin16, vdE94]. Facts [Bru17]. Failed [Har13]. Failing [Cer14]. Failure [CO12, Cro18, Ich06, Mai16, Neu09, Ste06, DV18]. Fair [Sim25]. Fairchild [Law10, LR12, Maz08]. Fake [HC20]. Fall [Dur10, Gaw19, Gri01b, Hou21, Oba22, Rub06, Spi21, Tak05, Lee95f]. Familiar [Roy11]. Family [Sch17, KS97, Lee00a, WD98]. Fantasy [Cro18]. Far [Mar12, Hen05, HIIT05, HII^{+}05]. farm [Hal00]. FastLane [GH18b]. Fate [Wil93b]. Father [Sub12]. Featured [Ano20-50]. features [HTAY97]. Featuring [Bul21]. federal [Gre95]. Federation [Bra22, Tat10, Not20]. Feedback [Ger04, RAG^{+}04]. Feigenbaum [Gri13, McC22]. Fellows [Yos23]. Fellowship [Yos01]. Female [Cam03b]. Fermat [Cor08a]. Fernando [Ano12b, Spi15a, Spi16a]. Fernbach [Ano17y]. Ferranti [Twe93]. Ferrite [Car92]. Ferut [Hum94, Wil94b]. Feynman [Sei00]. Field [McG12, SW95, Mey99]. Fields [Ens01]. Fifties [DMD18]. fiftieth [Win96]. Figure [Wu15]. film [Rus15c]. Final [Ber03e, Ekl92]. Finance [Der20]. Finding [Gra20a, Mah04]. findings [Kis98]. Finis [Sav23]. Finland [Paj08, SS09, SR14, Veh99]. Finn [Ber03d]. Finnish [VN11]. fire [Kid99a, Per99, Kne21]. firing [Pol97]. Firms [Hai05b]. First [Ain23, Ano00b, Bac09, Cam22, Car20, CP23, Dav17, Doy01, Dra76b, DA14, EF11, Gal04, Gil07, GS04, HN13, HO23, Ich06, Kid99e, LHS^{+}25, Neu05, Roe09, WN05, WN06, Wil18, ZR06, vN93, Ake00, Bis98, Bro99b, FAP^{+}03, Gil97, Joh99, Kis95, Lan00, LBA00, Mar10b, Rei97, Wei00, Wil00a, Wil92b, Too10]. First-Generation [Ain23]. Fit [Ano16j, Ano17w, Ano17x, Ano17v, Ano17b, Ano17c]. FITS [SB20]. Five [Jon04, McK22b]. Fixed [Est02]. Fixed-Plus-Variable [Est02]. Fixes [Hod19]. flaws [Ano16f]. Flood [RCP^{+}13]. Flow [WP94]. Flowcharting [Kid19a]. Flowers [Hai18]. Flows [Fra18]. Floyd [Hai04c]. Fluor [Laz24]. Flyable [Bro99b]. Focus [Ano15g, Ano16b, Ano17f, Ano17e]. Follow [Wes17]. Followed [Wes17]. Following [Bor08]. Font [Big20a, Big20b]. fonts [Wri98]. Footsteps [Lav06]. Forbe [Lee99b]. Forbes [Kid99e]. Forces [Jor09]. Ford [Bur02, Cly93]. Forensic [Tha17]. forget [Lee96]. Forgotten [CH15]. Fork [NL16]. Form [McD10a]. formal [MP97]. Format [Hos24]. Formative [Gra12]. former [Ber03f]. Forming [GK24, Min07b]. Formula [Ram97]. Fortran [Bac98, McJ17a]. Forward [Cor07]. Fostering [Uss96]. Foundation [GH18a, Roj04a, Bur02, Nor01]. Founders [Joh02]. Founding [Ber05, ERS03, JB07, Pag08, WG19]. Four [Ral04, Rus13b]. Fourth [Raw14, AG19]. Fowler [GHV05]. FP [Var94]. FP-6000 [Var94]. Fragments [Roe09]. Frame [Sho01, Mur19]. FrameMaker [Mur19]. Framework [EHP22]. France [BBC^{+}09a, BBC^{+}09b, PM12, MK14]. Frank [Hai04b]. Franz [KK22]. fraternity [Kid00a]. Frederick [RAG^{+}04]. French [Fra23b, LaV24]. Friedland [Gri22b]. Friedman [Buc95b]. Friendliness [Nor23]. Frohman [Kat18]. Front [Ano13d, Ano14b, Ano15h, Ano15i, Ano15j, Ano15k, Ano16c, Ano16d, Ano16e, Ano17g, Ano17h, Ano17i, Ano17j, Ano18i, Ano18f, Ano18g, Ano18h, Ano19t, Ano19u, Ano19v, Ano19w, Ano20u, Ano20v, Ano20w, Ano20x, Ano21p, Ano21r, Ano21s, Ano22v, Ano22w, Ano22x, Ano22y, Ano23u, Ano23v, Ano23w, Ano23x, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano25r, Ano25s, Ano14a, Ano21q]. Frontier [Ano97d, Ano00b, Sou19, Bir00, Rio00]. Fuhgedaboudit [Ges09b]. Full [HIIT05]. Fund [Vid20, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano23-31, Ano23-32, Ano24-41, Wei96d]. Furber [Gri14]. further [LS00]. Future [Hai06, Lee92h, Pay14, Rob24, Tym99, Rus13c]. Futures [Cor24a]. Fuze [Lat13].

G [Kid97a]. Galileo [Dra76b]. Galler [Ake08, Ano01b]. Game [Jor09, Nyi11]. Games [Ano09e, Ver11]. Gaming [SS09]. Ganzhorn [Roj04a]. Gaude [BBC^{+}09a, BBC^{+}09b]. GDR [Cor12, Don10, Sch23]. Gears [Smi78, de 74, Nor77, Ash76, Bal76, Bro77, Dra76a, Nas75, Ole77]. geeks [Rus16a]. Gehani [RAG^{+}04]. Gender [Abb03, Rus13c, Sch17, Ste03, Veh99, Ada96]. General [Ano19b, Ano19q, Ano19-38, Ano19-39, Ano19-44, Ano19-48, BG93a, ERS03, Gar21, Sma93, Kru94, Lee95f, Lee97, LS00, Old95]. General-Purpose [ERS03, Sma93]. Generated [Fra13]. Generation [Ain23, GS01, GS04, GS09, Raw14, Yos01]. Generations [Swa95c]. Genesis [Hol94, RAG^{+}04, Run08]. Genie [KH21a, KH21b, Kid99e]. genius [Sei03]. geniuses [Rus16a]. Geoghegan [Fra23b]. Geologist [Smi11]. Gerard [Wei96f]. German [Ken22, PPM^{+}23, Bul21, Car92, Cor12, CF92, Lei08, Sie10]. Germany [Aum11, LM94, Sch17, Cer25]. Geromat [Kir20]. Gertie [Pet94]. Gertrude [Gri97].

Get [Ano15l, Ano20y, Ano20z, Ano21t, Ano21u, Ano21v, Ano21w, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano23y, Ano24t, Ano24-31, FAP^{+}03]. Gets [Bro10, Joh99]. Getting [Hea02]. Giant [AC16, CVRM15, Sat11]. Giants [Mor19]. Giving [Sid20]. Glaser [Out03]. glass [Goy96]. Global [Ano97f, Cor04, Ens13, Ens18b, Hou21, Lis24, Shi19, Kid99b]. globalization [Kid00f]. Gloria [Nel22]. Glory [Hea04]. Go [Cor04, Der20, Ano14k, Gas99]. Go-Go [Der20]. goals [Ano17b, Ano17c]. going [Owe96]. Gold [Est22]. Goldstein [Hai04b]. Gopher [Fra04]. Gordon [Spi15b]. Got [Hai09, Ano16f]. Gottschalk [Dia16]. Governance [Par16b]. government [Gre95, Uss96, Hun12, War15]. Governments [Kid19b]. Graduate [Mil06]. Graphics [Ano19h, Ano20b, Ano20j, Ano22j, Ano22-31, Ano23h, Ano24k, Ano24l, BBB^{+}98, CHGW17, Kru94, LS08, Mol14, Sho01, Spr18a, Wyv98, Fal98, Ano19g]. Great [Bro10, Gri03b, SR14]. Greatest [Kid99e, Lee99b]. Greece [DMS18, Ges11a]. Greek [de 59]. Greeks [Ash76, Bal76, Dra76a, Nas75, Nor77, Ole77, Smi78, de 74, Bro77]. Green [Fai24]. Grew [Har22a]. Grips [Joh06]. Grosch [Wei96b]. Ground [Hea02]. Grounding [TMS14]. Group [Hea05, Rus16a, Bro23]. Groups [SR14]. Growing [WG19]. Growth [HS13]. Grupo [BLH12]. Guatelli [HC20]. Guest [Alb14, Abb03, Ano09e, Bös08, CG06, CG07, GB09, Gra12, Gra13, Gra02b, GR02, Mis07b, Nyc96, Sei96, Yos15a, da 15b]. Guichard [Nik17a]. Guidance [Kid99d]. Guide [Ano97b].

H [Cer03, Wal11]. Haanstra [Lee01c]. Hacked [PB22]. hackers [Rus16a]. Hacking [Hic19, Vag24]. Haeff [CH15, CHGW17]. Haigh [GH18b, RCP^{+}13]. HAL [Kid97a]. Half [Ano19b, Ano19q, Ano19-39, Mac03, MPDT12]. Halloween [Fit02b]. Hamlet [Fei22]. Hammer [Ber06c]. Hamming [Lee98b]. Hands [Abb01]. Hannibal [Cly93]. Hanratty [Bal24]. Hansen [LH03]. Happened [Deu13]. Happenings [Bau07, Bow94a, Bow94b, Bow95, Bow96a, Bow96b, ORB^{+}92, O’N92a, O’N92b, O’N93a, O’N93b, O’N93c, O’N94, Spi17, Wei92a]. Hard [Mah08]. hardcover [Ber03d]. Harder [Asp93a]. Hardman [Nel22]. Hardware [Ano20-37, Day23, GS04, Lis24]. Hardware-Oriented [Ano20-37]. Harlan [Ano17l]. Harley [SRG02]. Harold [Hai18]. Harry [Pol03]. Hartley [ORB^{+}92]. Hartree [Gon24a]. Harvard [RCP^{+}13, Kos07, Paw22]. Hashing [Ram97]. Hasselblad [San11]. HCI [Mor20]. Health [Dav17]. Hearing [Mil11]. Heavens [Cro03b, Mar10b, Mar10a]. Heaviside [Bab97]. Hedgehog [RAG^{+}04]. Heights [Ano15v]. Heinrich [Van10a]. Heinz [Woi25]. Held [Yam19b]. Helen [Cam03b]. Hell [DV18]. Helmut [Wei98]. Helped [CRB24, Dra18]. her [Lee00a]. Herbert [Kra22, Hey08a, Hey08b, Van09]. Here [Ano14c, Sta20b, Rus13c]. Heritage [PB93, Yam18]. Hermeneutics [Ack15]. HES [Fei22]. Hidden [Cor24b]. Hiding [GB13, Lap09]. High [Ges09a, KB21, Nof23, VR22, WMB22, De 95, Gil97, WD98, Sou19]. high-level [Gil97]. high-speed [WD98]. High-Tech [KB21]. Higinbotham [Nyi11]. Hiltz [Sub13]. Hinds [RAG^{+}04]. Hiroshi [CK08]. Historian [Ber09, Cer01, Smi11]. Historians [Cer14, Cor16, Cor22, Cor25a, Ens04a]. Historical [Gri11a, Hol94, KH21a, KH21b, Kat97, KBSW11, Mal02, MA96, Roe15, Smi01b, Sub20, Swa05b, De 95, Goy96, Mac95, Sha97]. Historicizing [Wu15]. Histories [BT23b, Cor22, Hoc22, HP24, MH11, PN16, RCP^{+}13]. Historiographic [Geo08].

Historiography [FC16, Pet20]. History [Abb03, Ake04, Alb10, Alp15, AG12, Ano97b, Ano99b, Ano03a, Ano06, Ano07, Ano09e, Ano19f, Ano20-27, Ano20f, Ano20g, Ano21h, Ano22g, Ano22h, Ano22i, Ano23e, Ano23f, Ano23g, Ano24h, Ano24i, Ano24j, AKM08, Asp01a, Asp07, AC16, BMFB25, BT23a, Bös08, Bre25, BL12, Bro18, Bro23, BYO^{+}22, Bru17, BK01, CK92a, CK01b, CK07a, CKGS08, CK14, Cer02, Cer03, Cer14, CVRM15, Cha12, Cho16, Con15, Cop11a, Cop11b, CHGW17, Cor02, Cor15b, Cor15c, Cor16, Cor18, Cow94, Dia12, DAGVR^{+}15, Dow09, DMS18, Dun21, Ens04b, Ens09, Ens18a, Ens18b, Eps13, EK24, Eva24, Fai24, Fev13, FM93, Fra18, Ges07a, GR22, GB13, Got23, GB09, GJ12, Gra20a, Gra20b, Gra21a, Gra22, Gra24, Gri03b, Gri07a, Gro93, Gui99, HS13]. History [Has16, HUA23, HC20, Hen05, HIIT05, HII^{+}05, Hic13, HP96, HG24, Hsu20, Hsu23, ICKDL99, Jes07, Jes13, Joh20, Kat18, Kid96d, Kru09, KK22, Lap09, LS08, Lea23, Lev05, Low09, Lun18, Mag20, Mah04, Mah08, McD10a, McG12, MR23, MF93, Med05, Mil11, Mis07a, Mis07b, Mit08, NM09, Nor01, NL16, ORB^{+}92, Phi12, Pla16, Pol95, Pri21, PP21, Raj15, Ral04, Ran14, Raw14, Ray02, Rob19, Roe17, RAG^{+}04, Row12, Sch06b, Shu06, SAV24, Sjö11, Sla08, SW96, Smi78, Spi18, Spi21, Spi23a, Spi24a, Ste09, Tak96, Tri01, TMS14, Tym99, WH17, Wei08, WP94, Wil98a, WT03, Win03, Wir08, Wri98, Wu21, Yoo05, Yos01, Yos15b, Yos16b, ZS23, da 15b, Ada96, Ano02e, Bac98, Cor98, DFM^{+}99, GC00, Kid99d, Kid00h, KS97, Lee96, Smi00, SD98, Sze99]. history [Tym96, Zwe99, Léc21, Roj04b, RCP^{+}13]. historyofscience.com [Cer03]. Hobby [Sta15]. Hobbyists [SS09]. Hodgkin [Ano00b, Cro00a]. Hoelzer [Wei98]. Holiday [Ges06a]. Hollerith [Kis92b, Kis05]. Home [Was13]. Homebrew [Pet17]. Honeywell [KS97].

Hong [Smi16b]. Honor [Bow98a]. Honoring [Ano22-30, Asp01a]. Hook [Cer03]. Hooligan [ZS23]. hopes [De 95]. Hopkins [GH18b, Mor03, RAG^{+}04]. HOPL [PP21]. Hopper [Pay14]. Hospital [VT21]. Host [Fei11, Ano20-28]. Hot [MT06]. House [Ano13b, Ano13c, Ano15b, Ano15c, Ano15e, Ano15f, Ano15g, Ano15l, Ano15m, Ano15a, Ano15n, Ano15o, Ano15q, Ano15p, Ano15v, Ano16a, Ano16b, Ano16h, Ano16i, Ano16m, Ano17f, Ano17m, Ano17u, Ano17y, Ano18d, Ano18o, Ano18t, Ano18u, Ano18-27, Hea03, Ano17w, Ano17x, Ano17z, SWH18]. Howard [Coh98, Cro00b, Lee00c]. Hughes [Smi99b]. Human [Ano97f, Ger04, Gru05, Hic08, MK20, Pet20, RAG^{+}04, Sou19, Tom96, Tom00b]. Human-Computer [Gru05]. Humanity [Smi99a]. Humor [Bau07]. Humphrey [Ano23z]. Hungary [Sze99]. Hush [Tra19]. Hush-A-Phone [Tra19]. Hustle [Oba22]. HWMS [Gri22a].

I. [HW03b]. IBM [Roj04a, Sch17, Ake98, BBC^{+}09a, BBC^{+}09b, BJ03, Bro99a, BH95, CRB24, Cor13, Cor14, Cor15a, Cor17, Cor18, Cor21a, Cor24b, Cor25a, End04, End13, Gan04a, Gan04b, Gil07, Gra02a, HS13, Hal18, HN13, Hen05, HIIT05, HII^{+}05, Med08, Mey99, Per20, Rob16, Sch08, Spe98, Spi05, SD98, WC12, Yos13a, Hou21]. IC [CO12, LR12]. ICA [BLH12]. Icon [Hin18, Hou21]. ICs [Phi12, SL12]. ICT [HUA23]. Ida [Tom00b]. Idea [RAG^{+}04]. Identity [Sub14]. Ideology [Kan21, Tym96]. IDES [Yos16b]. IDIIOM [Bis98]. IDMS [Kru09]. IDS [Bac09]. IEE [Win03]. IEEE [Ano17a, Ano20n, Ano21a, Ano97e, Ano98b, Ano03a, Ano06, Ano07, Ano14c, Ano15m, Ano16h, Ano16i, Ano16g, Ano16j, Ano17n, Ano17l, Ano17m, Ano17k, Ano18k, Ano18j, Ano18l, Ano18m, Ano18n, Ano19a, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-34, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-35, Ano19-36, Ano19-37, Ano20n, Ano20-29, Ano20-35, Ano20-30, Ano20-32, Ano20-36, Ano20-33, Ano20-31, Ano20-34, Ano20-37, Ano20-38, Ano21b, Ano21x, Ano21-28, Ano21-31, Ano21y, Ano21-29, Ano21-34, Ano21-37, Ano21-32, Ano21z, Ano21-30, Ano21-38, Ano21-35, Ano21-33, Ano21-27, Ano21-36, Ano22n, Ano22o, Ano22p, Ano22-31, Ano22-32, Ano22-35, Ano22-37, Ano22-41, Ano22-47, Ano22-33, Ano22-36, Ano22-38, Ano22-42, Ano22-48, Ano22-46, Ano22-39, Ano22-43, Ano22-49, Ano22-44, Ano22-34]. IEEE [Ano22-51, Ano22-45, Ano22-40, Ano22-50, Ano22-52, Ano22-53, Ano23o, Ano23p, Ano23q, Ano23-33, Ano23-37, Ano23-28, Ano23-31, Ano23-34, Ano23-38, Ano23-35, Ano23-29, Ano23-32, Ano23-39, Ano23-27, Ano23-41, Ano23-36, Ano23-40, Ano23-30, Ano23-42, Ano23-43, Ano24u, Ano24v, Ano24-34, Ano24-32, Ano24-33, Ano24-41, Ano24-44, Ano24-42, Ano24-38, Ano24-45, Ano24-43, Ano24-39, Ano24-36, Ano24-35, Ano24-37, Ano24-40, Ano24-46, Ano24-47, Asp92a, Asp92b, Asp92c, Asp93b, Asp93c, Asp93d, Asp93e, Bow98b, Bow98a, Mag20, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano20m, Ano20y, Ano20z, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21t, Ano21u, Ano21v, Ano21w, Ano21h, Ano21i, Ano21j, Ano22m, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22g, Ano22h]. IEEE [Ano22i, Ano22j, Ano23y, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24t, Ano24s, Ano24-31, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano25d]. IEEE-CS [Ano17a]. II [Bac98, Car92, Lon08, Smi01a, TYD21b]. III [Bac98, Kir20, Lee00c, Lee01d]. Illegal [Tin11]. Illinois [BB02, Mil06]. Illustrators [Kid00c]. Image [Kru07, Nof25, Vog17, Kir98]. Images [Tha17]. Imagines [Est22]. Imaging [San11]. Imagining [Mal02, Pet17]. IMP [VR22, WI14, WMB22]. Impact [EK24, Lee92i, Wei08]. Implementation [AD14, Sch23]. Implementing [HPR14b]. Implications [Hun12, KBSW11]. Important [Nof23, Cor98]. impossibility [Kid00g]. Impossible [TYD21a, TYD21b]. Improvement [Kid01d]. IMS [McG09, Pat09]. Inc. [DW20, WG19]. Incidents [Vag24]. Inconsistent [Gri10]. Inconspicuous [Sat11]. Incorporating [AKM08]. Increases [And22]. Incredible [Kid99e]. Independence [Yos13a]. Index [Ano92a, Ano92b, Ano93, Ano94, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05, Ano06, Ano07, Ano09a, Ano09b, Ano11a, Ano13a]. India [Raj15, Sub14, Sub20, Fer22]. Indian [Day24, Sin09]. Indianapolis [Ges12a]. induced [Ano99e]. Industrial [AL12, Smi16b, Var94, Van95]. industries [Yat97]. Industry [Ake98, Bro23, CK97, CKGS08, CK12, Car93, Cor06b, Cor21b, Don10, Dra18, EK24, GJ02, Gra03, GJ12, Gra20a, Gra22, GK24, Hai04b, Hai10a, Hem14a, Joh02, Joh20, Kas24b, Lee97, Lei08, Lei17, MPDT12, Pet05, PA96, SS09, Sey18b, Tak96, VN11, Vog17, Yos13a, Yos13b, Yos15b, Fal98, Joh96, Joh98, Kid00e, Kid00f, Pro99, Van95]. Industry-Developed [Kas24b]. Inequality [Sid20]. Influence [Cor06a, PHL92, Roh99]. Influencers’ [Bri06]. Influences [Cro93]. Info [Ano18d, Ano24y, Ano24z, Ano25o, Ano25p]. Informática [FHD17]. Informatics [AKM08, Bau06, Bös08, FH14]. Information [Ano18e, Ano18k, Ano20n, Ano22-46, Ano23-37, Ano23-38, Ano23-39, Ano24-44, Ano24-45, Asp13, Asp15, Car93, CM25, Cor08b, Cor12, Cor15c, Cor19a, Cor19b, Cor20, Cor25b, Cro18, Day24, Fei10, Fra18, Fra23b, Geo08, GB13, Gri12, GK02, Hai09, Kid96f, Kid99e, Mal02, ORB^{+}92, RCP^{+}13, Sip13, Tat10, Tym99, Yos01, Yos21, Ano18a, Asp99, CMM95, Kid99a, McG09]. information-based [CMM95]. Informix [Sip13]. Infotech [Day23]. Infrastructure [FC16, GA22, Gra22, Par15, Sch06a, SB20]. ing [Bro10]. Ingres [Row12]. Initial [Hen05, HIIT05]. Initiates [HW17]. Initiative [Dav17]. Initiatives [Rei23a]. Innovation [Lew14, dÁ24, Lan00, Uss96]. Innovative [Has16]. innovators [Rus16a]. Inscriptive [Dun21]. Insights [Cor25a]. Instead [VT21]. Institute [JB07, Sei95, Yos21, Klu99, Asp07, CK01b, Car15, Gra24, Mis07a, Nor01, WD98]. Institutional [Ake04, Gio93]. Institutionalizing [Hai01]. Institutions [McG12]. Instruction [Cro18, Kid15]. Instructional [BH95]. Instruments [Kid97b, Phi12]. insurance [Yat97]. Integrated [Bac09]. Integration [Ano20n, Log98]. Integrator [GG46, GG96]. Intel [Asp97, Maz07, MS09, Maz10, Sta07]. Intelligence [Ano21-46, Ano21-47, Ano22-63, Ano23-52, Ano23-53, BG22, Har22b, PS23, Shi19, dPRG21, vH23, Ada96, Kid99b, Ano22p, Kli11, Kan21]. Intelligent [Kel22, Tym96]. Intelligentsia [TYD21a]. intentioned [Fit02b]. Interaction [Gru05, MK20, Pet20]. interactions [Yat97]. Interactive [Hem16]. Interdisciplinary [Yos21]. Interest [Bro23]. Interests [Vid20]. Interface [Ano19r, Bak17, MK20, Wil00c, Bro00]. Interfaces [Mor20, Yos23]. Interleaf [DW20]. International [Ano20n, Ano20-37, CK97, Fra18, Mis99, Paj08, Sch06b, Sch08, Par07, Tat10]. Internet [Ano00b, Ano19i, GH18b, McG21, Abb10, Cer06, DeN15a, Dow09, Kid00h, Kir99, Kru07, Lig03, McK11, Par08, Par16b, RC15, Rus06, VHCSCC15, Wal00, Wu15]. Interpretation [FC15, GR22]. intersection [Est96]. Interview [Ano01b, Ano12b, Ano20-39, BG93b, Gri13, Gri22b, Mac98, McJ20, Pat24, Sav23, Spi23b, SB23, Wal11, Wal18, Wal20, Yos24, Ano12a, Spi14]. Interviews [Gri14, LR92, LH95, RL92]. Introduction [Abb03, Ano09e, Bös08, BG93a, CK92b, CG06, CG07, GB09, GR02, KC20, Lee92g, Mis07b, MK20, Nyc96, Sei96, Smi07, VN11, Wil00c, Wil01b, Alb14, Gra12, Gra13, Per99, RCP^{+}13, Yos15a, da 15b]. Introductory [VR22]. Intrusion [Yos16b]. Intrusion-Detection [Yos16b]. Inventing [Ano00b, Ari00, Wal00]. Invention [Rob24, Asp97, Kid00a, Kid00i, Puc96]. Inventions [Neb98]. Inventor [Gal11]. Investigating [CRH^{+}21]. INWG [Day16, McK11]. IP [Yam18]. IPSJ [SWH18, Yam18]. IPTO [Kit03]. IPUMS [MR22]. IRI [Ano20n]. Irma [Gri03a]. Iron [Gaw19]. Irresistible [Mer06]. Isaacson [Rus16a]. ISBN [Ber03d, Cer03, GH18b, Mor03, RAG^{+}04, Smi03, Win03]. Island [Ges10b]. Israel [Ges10a]. Issue [Ano20o, Ano21n, Ano22q, Ano22r, Ano22t, Ano22s, Ano24w, Ber94, Ber00, Ber01c, Ber01a, Ber01b, Ber02a, Ber03c, Ber03a, Ber03b, Gal04, Lee92a, Lee92b, Lee92c, Lee92d, Lee93a, Lee93b, Lee93c, Lee94a, Lee94b, Lee95a, MK20, Wil93a, Wil94a, Wil96a, Wil97b, Lee95b]. Issue/From [Ber01c, Ber02a, Ber03c]. issues [Cer97]. Italian [CG20, Hen10, HT12]. Italy [AG21, DMPS99]. IV [PP21, Pos98]. Ivan [SB23].

J [Bal24, Cro99, GH18b, Hai05c, RAG^{+}04, Rus13a, Rus13c, Shr24, Wal12]. J. [Eck96, Kit03, Lee95c, Yos24]. Jack [Yos24, Eck07a]. James [RCP^{+}13, Gui99, RR23]. Japan [Che25, CK08, CO12, HII^{+}05, Sat11]. Japanese [Tak96]. Jean [Ber09, MK10, SWH18]. Jeffrey [GH18b, Lee99b, RAG^{+}04]. Jeremy [Cer03]. Jersey [Ges06a]. Jevons [BH99]. Jewel [RAG^{+}04]. JFEAC [Bar24]. JHUP [Yos23]. Joan [Lee01d]. Job [Ano15g, Ano16b, Ano17f, Ano17e, Ano17o, Ano17p, Gra21a]. Jobs [Ano19s, Ano19-34, Ano19-35, Ano21-34, Ano21-35, Ano21-36, Ano22-47, Ano22-48, Ano22-49, Ano22-50]. Joe [Bro18, Jac93]. John [Ano99f, Ano98a, Ano99c, CD14, Cos96, DSW92, Gri00a, Gri03a, Hai10b, Lee95d, Lee01d, Lee01c, Pri24, Sei03, Sha23, Wil01b]. Johns [GH18b, Mor03, RAG^{+}04]. Johnson [Shr24]. Join [NL16]. Jonathan [Ano99d]. Joseph [LBA00, Wil92b]. Journal [Ano20y, Ano20z, Ano21t, Ano21u, Ano21v, Ano21w, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano23y, Ano25d, Pol95, Sha00, SD98]. Journey [DK23, Kid99d, Sin09]. JSTOR [Sha00]. July [Hus97]. Jumping [Dow04]. Junction [Yoo05]. just [Van95]. Justice [McG21].

Kangxi [Ano92d]. Kanji [HIIT05]. Kare [Hin18]. Karl [Roj04a]. Kasik [Shr24]. Katapayadi [Ram97]. Kay [Bar07]. Keep [Ano20-40]. Keller [HO23]. Kelly [RAG^{+}04]. Ken [Ano17y, Win03]. Kennedy [Ano17y]. Kenneth [HO23]. Kerberos [Sal21]. Key [CG20, Hei11, Hen10, Par15, Roe08]. Key-Driven [Hen10, Roe08]. Key-Set [Hei11]. Keystone [McC02]. KI [PPM^{+}23]. KI-Rundbrief [PPM^{+}23]. Kidwell [Ano00b]. Kiesler [RAG^{+}04]. Kilburn [And09, BG93b]. Kilby [Eck07a]. Killer [Ben13, SH19]. King [Lee97]. Kingdom [Eva24, Kir99, Wyv98]. Kings [Lee95e]. Kissinger [Lap11]. Kling [Hai03b]. Knowing [Sei97]. Knowledge [Ake04, Cob23, Poi25, Ano17z, Puc96]. Known [Bru17, EF11]. Knox [Mor03]. Knuth [CK14]. Kong [Smi16b]. Konrad [CRB24, Gil97, Roj97, RDGH05, Wei96c, Cer25]. Korean [Par15, Par16a].

L [Asp93a, Nor23]. L. [Cro99]. Lab

[Har22b]. Labor

[Sch17, Sim25, Sti24, VT21]. Laboratories [Roj04a, Irv01]. Laboratory [Spe98, BBC^{+}09a, BBC^{+}09b, Bur02, CdLF^{+}22, Gan04a, Gan04b, How99, Lat13, LBA00, Moo01, PHL92]. Labs [Mor03, RAG^{+}04, Sim25, Tra24]. Lady [Too96, Gon24a]. Lagos [Oba22]. Landscape [McD10b]. Lane [Del08]. Language [All18, Ber09, DA14, End13, KKW09, NM09, Nan09, Nof10, Raw14, Sin09, Smi11, Wad12, Wei06, Gil97, McK95]. Languages [All18, Ber94, Den15b, Hen05, HIIT05, HII^{+}05, Nof25, Pau14, PP21, Smi04, WP94, Roh99]. Lansdown [Ano99f]. Lapel [Cor21a]. Laptop [Lin16]. Large [JMSV10, MR22, Nik17b, Wie10, Joh99, KS97, Log98]. Large-Scale [JMSV10, MR22, KS97]. Largest [Ano97f]. Larry [Ano98b, Wal20]. Last [And22, Wil98b]. Late [RP12, Sch23]. Later [Nee92]. Latin [Cor25a, Sha23, da 15b]. Launch [CO12]. Lautenberg [Hai04b]. Law [Mol06]. Lawrence [Asp21, Hai05c, Kit12c]. Lay [Gre05]. Laying [GH18a]. Lead [HII^{+}05, MT06]. Leadership [Ano17n, Asp07]. Leading [Ano15v, Hea05]. Leap [CVRM15]. Learn [Cor22]. Learning [Cro19, Kel22, Pla16, Sla15, Sta20a]. least [Gas99]. Lebedev [CG94]. Led [Spr16]. Lee [Wei92e, Wei96a]. Legacy [HO23, Kid97a, Law10, Roj97]. Legal [Hod19, Kid00e, Mul20]. Legend [Law10]. Legends [CP23]. Legibility [All18]. Lehmers [Cor08a]. Leibniz [Kis99b]. LEO [Ari00, Cam03a, Kid99e, PHL92, Kis99a]. Leonardo [Gon20, dPRG21]. Less [DMD18]. Lesson [Nel17]. Lessons [Cor21a, Ens04a, FM93, Mai16, MF93, Ste06, Lee96]. Let [Ano23-44, Ano24-34]. Letter [Ber92, EMW02]. Letters [Ano19-36, Gri02, Gri04e, Ano18l, Ano20h, Ano20i]. Level [Fei11, Gil97]. Levinson [Tym99]. LFK [Ano02f]. LGP [Smi07]. LGP-30 [Smi07]. Liability [Mul20]. Libraries [GR02, Ray02, Wil02]. Library [BB02, Bur02, Cer03, Gra02b, Mal02, Mar02, McC02, SRG02, Smi03, Asp99, Ano14k, BGM02, Wil01a]. License [Sal20]. Licklider [Kit03]. Lies [Fit23]. Life [Ake08, Ano98b, Ano00b, Asp01b, Asp13, BJ03, CRH^{+}21, Kne17, McC22, Moo03, Neu06, Par15, RAG^{+}04, Rus13a, Ste22, Sti24, ZR06, Cos96, Cro00a, Sei03, Yat97]. Lifetime [Mac03]. Light [Smi16a]. Lighthill [van19]. Lillian [Kid99a]. Limits [EM94, Sei00, Cor20]. LINC [Kit08b]. Line [Hal92, Lee92g, WF92, RG02]. line-drawing [RG02]. Lines [Ben93b, Kis92a, Dor94]. Link [GA22, McQ09]. Link-State [McQ09]. Links [Cor21a]. Lions [Ano99c]. Lisa [Kid00i]. LISP [McJ17b, KK22]. listings [Ano00a]. Listserv [GC00]. Literature [vH23]. Lithuania [TZ99]. Little [OK07, Kid00i]. Lives [CK94, McG21, Fer22]. Living [Ano22-67, Cro03a, Tri01, Wei96b]. Liz [McJ20]. Local [CVRM15, De 21, Ens18b, Ges06b, Ges09a]. Locality [AD14]. Locating [Ber92, Kis97, LM94]. Lock [Gre95]. Lock-in [Gre95]. LOCOMAT [Roe12]. LOCS [Ano18o]. Logic [Bul21, De 21, DT21, Dun21, Pri21]. Logical [DMD18]. Logics [TYD21a, TYD21b]. Logistical [Hoc22, HP24]. Logistics [Sti24]. Logos [Was13]. London [Yos24]. Long [Wei08, Mar10a, Ges10b]. Long-Term [Wei08]. Longitudinal [GA22]. Look [Rop07, Tha17, LS00]. Looking [Ano17o, Ano17p, Cor06b, Mit08]. LOSC [Ano18p]. Losing [Ver11]. Lost [Hea01a]. Lotus [Kap07, Sac07]. Louis [Out03, Kid00a]. Louise [Nik17a]. Lovelace [FF03, Gon24a, Lee00a, Spi03, Too96]. Lowther [Lee01d]. Ltd [Tol21]. Ltd. [Cro99]. Luc [Léc21]. Ludgate [CRH^{+}21, CRB24, RC23]. Lynn [Spi23b].

M [Cer03, Kid96b, Smi03]. MA [Sid20]. MAC [LFS^{+}92, LR92, Bur23]. Machin [RG02]. Machine [Abb01, Ano97f, Ger04, GHV05, Got23, Gri03b, Hei11, Kir20, LBA00, MA96, Pla16, Roe08, RAG^{+}04, SR14, Sha23, Sou19, Ste22, Tur11, Wad12, Aga01, CF92, Kid00a, Kid00k, Kis98, Lee00c, McK95, Smi02]. Machines [Bul21, CVRM15, Den01, Den15b, Dic13, Hai18, Jon04, Ken22, Kid01d, Sei97, vdE94, Kid99b, Tym96]. Macintosh [SH19]. Maddida [ERS03]. Made [Ano92d, Cor97, Lee95e]. Magazine [Nel17]. Magazines [Ano13c]. Magic [McK22a]. Magnavox [MS09]. Magnetic [Wei00]. Maher [Rus13c]. Mahoney [RCP^{+}13, Asp14, CK13]. Mail [Fra23a, Van12]. main [Sto99]. Mainframe [Lap09, Gre95]. Mainframes [Tak05, Kid99c]. Maintenance [Cor24b, VR22, WMB22, DV18]. Major [Ano16h, Ano19o, Ano19p, Ano23n, Yos21]. make [Roj98]. Maker [HC20]. Makers [FVW25]. Makes [Mah08, Sum07]. Making [GA22, GH18b, Hos24, Kid00f, LR12, Sjö11, Sta03, dÁ24, Aga01, Bro99a, Kid00g, Smi02]. Malta [Alo07]. man [Gas99]. man-in-space [Gas99]. Manage [Ano21-45, Ano23-51]. Management [Ano18m, Ano18n, BH09, GB09, Gra12, Gra13, Kos07, Kru09, McG09, Sip13]. Managing [Boh13, GH18b]. Manchester [Ibb99, And07b, Bur05, Cop11a, Cop11b, Cro93, Gio93, Hai03a, Lav93, Lav22b, Lav22a, PB93, SB98, Twe93]. Manuals [Cor24b]. Manufacturing [Lin16, MF93]. Mapping [Kru07, Smi00]. Maps [FC15, FC16]. MARC [McC02]. March [Yos16b]. Mark [GH18b, Lee00c, Pos98, SB98]. Marked [Wal19]. Market [da 15a]. Marketing [Ben13, Woh06]. Markets [Sch08]. Marquand [BH99]. Mars [Est22]. Marshall [Léc21]. Martin [POD23, RAG^{+}04]. Mary [Shr24, Cro03a, HO23]. Maryland [Min07a, Min07b]. Mass [Ken22]. Masthead [Ano13e, Ano16k, Ano16l, Ano17q, Ano17r, Ano18q, Ano18r, Ano18s, Ano19-43, Ano19-41, Ano19-42, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano22-54, Ano22-55, Ano22-56, Ano22-57, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24-48, Ano24-49, Ano24-50, Ano24-51, Ano25t, Ano25u]. Mataré [Van09]. Matching [Gil07]. Material [Cor21a]. Materiality [Cer06, Ens13]. Materials [EHP22, Gra24]. Math [Gri98]. Mathematical [CCK00, Gri01b, Hai08, Hai10b, Kid97b, Kid99e, PHL92, Roe12, Puc96, Gri97, Pol95]. Mathematician [CD14]. Mathematics [DT21, Lon08, MP97, Puc96]. Matilde [VHCSCC15]. Matlab [Hai08]. Matter [McG21]. Matters [Ano14d]. Mauchly [Ano19a, Ano21b, Cos96]. Maurice [Lee93d]. Max [And07a, Lav22b]. MCAE [EK24]. MCC [Len22]. McClean [Ano98a]. McIntyre [Smi11]. MCM [Sta03, Sta17]. MCM/70 [Sta03]. Me [BJ03, VM24]. Meaning [Ano00b, Gon24a, Jac16, Pos00]. Meaningless [Nof23]. Meanings [CO12, Ver11]. Measuring [Sla15]. Mechanical [Ben93b, Che25, Cly93, Dra76b, Hal92, Kis92a, Roe15, Roe16b, San13, WF92]. Mechanics [WD98]. Mechanism [Ash76, Bal76, Bro86, Bro77, Dra76a, Mor84, Nas75, Nor77, Ole77, Smi78, Spi08, de 74]. Mechanization [Cob23]. Medal [Bow98a]. Media [Bri06]. Medical [Cas06]. Medicine [Nov11]. Medium [Bar07]. Meeting [ABL^{+}24, PM12, WH17, Wal19]. Meets [Cor08a, Bro99a]. Member [Ano98b]. Membership [Ano13f, Ano14d, Ano17w, Ano17x, Ano17v, Ano18t, Ano18v, Coo02, Ano17z]. Memex [Kid96f]. Memoir [Smi04, Bir00, Est00, Log98, Mey99, Pos98]. Memoirs [Cor13, Goe02b, Goe02a, Tri01]. Memorial [Wei96d]. Memoriam [Hai04c]. Memory [Cop11a, CHGW17, Del08, Med05, Pow19, Ber03f, Eck98]. Mercury [Gas99]. Merger [Bau06, GS07]. Mertia [Fer22]. Merwin [Ano15a, Ano17n, Ano21-44]. Meryl [Sid20]. MESM [FBK06]. Message [Bri06]. Messages [Bur05]. Messaging [Hem14b, Van12]. Metallurgy [MPDT12]. Metaphor [Hal18]. metaphysician [Too96]. Methods [Bur05, Lon04, SK11]. Metrics [Sla15]. Mexican [dÁ24]. Mexico [BLH12, CVRM15, DAGVR^{+}15, OARHC10]. MF7114 [Sta10]. Michael [Cer03, Nor23, Ano20-39, Asp14, CK13, Der20, Kid99a, RCP^{+}13]. Microchip [Léc21, RAG^{+}04, Kid00g]. microchips [Kid00f]. Microcircuit [SL12]. Microcircuitry [AL12, BL12]. Microcircuits [Lat13]. Microcomputer [Sta03, Sub14]. Microcomputers [Fev13, Tin11]. Microhistory [Gal11]. Micrologic [LR12]. Microprocessor [Sta10, Sta13, Asp97]. microprocessors [Kid99c]. Microsoft [Ano97d, DV18]. Middle [Smi03]. MIL [Sta10]. Milestones [Lav22a, Kra22]. Military [McD10a, MP97]. Milken [Der20]. Million [Len22]. Mills [Ano17l]. Mina [Ano98c]. Mind [CH05, Hey08a, Hey08b, Kid99c]. Mindell [RAG^{+}04]. Miniaturization [Mil11]. Minicomputer [Fre07]. Minimalism [DMD18]. Mining [MPDT12]. Minitel [Mai16]. Miracle [HPR14b]. Misa [GH18b]. Missing [GA22]. Mobile [Alp15, Ano14g, Sid20]. Model [Dar12, FVW25, HC20]. Model-Makers [FVW25]. Modeless [Tes18]. Models [Cob23, Kne17]. Modern [CD14, DMS18, GB13, GH18b, HPR14b, HUA23, Lon04, Mir17, Ram97, RAG^{+}04, Aga01, Smi02, Win96, Che25]. Modernism [Nof23]. Modular [Roe17]. Molecular [SL12]. Moler [Hai08]. Molnar [Ano99a]. Monte [HPR14a, Rei23b]. Montreal [Ges07a]. Mooers [Wei95c, Wil01b]. moon [Kid99d]. Moore [Mol06, Win96]. morphogenesis [Kid96b]. Morris [Pat24]. Morse [Owe96]. Mortimer [Gar21]. MOS [Mor03]. most [Cor98]. Mother [LHS^{+}25]. motion [PM12]. Motives [Bur05]. Motors [Kru94]. Mourice [Ano98a]. Movement [Sta15]. MRI [Bru09]. MRP [Pee09]. MU5 [Ibb99, Roh99]. Multi [Ano15v]. Multi-Core [Ano15v]. Multimedia [Ano97d, Ano23s]. Multiplying [McP95]. Mulvin [Vid23]. Murray [Lee01d, Sub12]. Museum [Ano92d, Car23, AH23, Gra24, Spi21, Spi23a, Spi24a, WH17]. Museums [HC20]. MuseumsForum [Woi25]. My [Bar24, Gri94, Lew14, Moo03]. myComputer [Ano13g]. myCS [Ano17u, Ano17s, Ano17t]. mystery [Mar10b]. Myth [Mir17, Paw22, Tra19].

NA-Net [DGG^{+}08]. NAE [Yos23]. Name [BV19, Lee00a]. Names [Fei11]. Narain [RAG^{+}04]. nation [Coh98]. National [Fid18, Ñañ22, ORB^{+}92, Paj08, Par16a, Sub14, Dav17, Gur96, Kir98, Lap11]. Natural [Tym99, Wei06]. Natural-Language [Wei06]. Nature [Nof23]. Nautical [Cro03b]. Naval [LBA00, Moo01]. Navy [Doy01]. Nazi [LM94]. NCC [Hus97]. NCR [LBA00]. near [Kid00g]. Neat [Poi25]. Nebeker [RAG^{+}04]. NEC [Yam19a]. necessity [Sha97]. Need [Ano20t, Cor04, Kid00l]. needed [Coh98]. negotiated [Fit03b]. Negotiating [Sta13]. Neil [Wei96d]. Neill [RCP^{+}13]. Nelson [VM18]. Nema [Kid00k].

Nerds [Kid00h]. Net [Nor98, DGG^{+}08]. Netherlands [vdE92]. Netlib [DGG^{+}08]. Netnews [Bel25]. Nets [Gui99]. Network [Ano02f, Cro19, CH05, Hem16, Mai16, Sch06a, Vag24, Est00, Fei10, Gri22c]. network-based [Est00]. Networked [PM12]. Networking [Ano18u, Cer03, Day16, Fie03, Ges06b, PB06, Rei23a]. Networks [Ayl10b, BMFB25, DF16, Ges09a, HP96, Kid96d, PB22, RP12, Rus12b, Gui99]. Neumann [Gil97, GH93, Gri03a]. Neural [Gui99]. Never [All19, Tat13]. Newell [Cly93, Sim98]. Newman [And07a, Lav22b, WN05, WN06]. Newspapers [MT06]. Nicolas [Kid97b]. Nigeria [Oba22]. Niklaus [Spi24b]. Nimbi [Jor09]. Nineteenth [Kid19b]. Nixdorf [Woi25]. No [Cor15a, Swa05a, Ano97e, Win03]. Noble [Har13]. NOL [Wil01b]. NOMAD [Raw14]. Nominations [Ano16a, Ano16h, Ano17d, Ano17y, Ano19o, Ano19p, Ano20r, Ano20s, Ano21a, Ano22b, Ano22-30, Ano23n]. Nominees [Ano17d]. non [Gil97]. nonrole [Kis97]. Nonuniform [VM24]. Noon [Sou19]. Norberg [Asp07, Asp21, Mis07a]. Norman [Cer03]. Norris [Eck07b, Wor95]. North [Mag09]. Northern [SR14, CK12]. Norway [Lun18]. Norwegian [Bro99a]. Notch [Hea05]. Note [Buc95b, VR22]. Notes [ABL^{+}24, Ano20-45, Ano21-43, Ano22-58, BMFB25, Bro86, NL16, Wal21, War15, FF03]. Novel [Hem19, Kid19b]. November [Rus13a, Yam19b]. NSA [Mis16]. NSF [AW94, Yos22, Yos23]. NTT [Sat11]. Nuclear [HPR14a]. Number [CK07b, Cor08a, vdE94, Coh98]. numbers [CCK00]. Numeracy [Che25]. Numerical [GG46, GG96].

O [Gro92]. Obituaries

[Hai03b, Lee01d, Lee01c, Wei96e, Wei96f]. Obituary [Ano98c, HW03b, Kit12c, Lee00d, Mor84, Out03, Pol03, POD23, RR23, Spi24b, Wei96c, Ano99a, Ano99c]. Objection [Gon24a]. Objects [Cor22]. Obsolescence [Cam09]. obviousness [Owe96]. occupation [Zwe99]. Oceania [BT23b, HUA23]. October [Fid17]. Odyssey [Har22b, MS09]. Oedipus [Lav06]. off [Hea02]. Offer [Bau06]. Office [Hai06, Hei11, KB21, Ste03, Alo07]. Offices [Cor17]. Official [All19]. Old [Rop07, Mar10a]. Olden [Cor15b]. Oldest [Wei96b]. Olivetti [Mor20]. Olivier [Léc21]. Once [Mer06, Mir17, RP12, SB20, Wes17]. Once-Irresistible [Mer06]. One [Ano17z, LHS^{+}25, Rei23a]. Online [Doy01, Mai16]. Only [Hic10, HW17]. Ontologies [Dia16]. Open [Ano20y, Ano20z, Ano20-38, Ano20-40, Ano21t, Ano21u, Ano21v, Ano21w, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano23y, Ano25d, Ens04a]. Opening [Car23]. Openness [HW17]. OpenType [Hos24]. Operating [KH21a, KH21b, HTAY97, Rei97]. Operation [Sim25, ZR06]. Operations [Urq24, Wie10]. Operators [Hic10]. Opportunities [Ano20t, Mag20, Goy96]. Opportunity [Mer06]. Options [Ano17w, Ano17x, Ano17v, Ano20-40, Fit03b]. Oracle [Ben13, Men13, Pre12]. Oral [BT23a, Bro18, Fai24, Gui99, HG24, Hsu20, Hsu23, Kat18, Spi18]. Orange [Lip15]. orbits [Gas99]. Ordinary [Bur24]. Ordnance [Lat13, Moo01]. Organization [Smi16b, Van95]. Organizational [BB02]. Organizations [Cor19a, Cor19b, Cor20]. Oriented [Ano20-37, Cam13]. Origin [Alo04, Bac09, Bel25, Den01, Fei11, Kru94, Rei08, Sal20, Sal21, Sha00, Yos15b]. Original [HT12, McJ17a]. Originators [Cop04a]. Origins [Ber06a, BV19, Cer03, CP23, Cop04a, Day24, Est02, FH14, Gri22c, Hai06, Hal18, Jes13, Pug02, Ran94, Tat10, War18, Wil93b, Asp99, Bar03, BH99, Cro99, Fit03c]. Ornstein [Smi03]. oscillations [Ano99e]. OSI [Rus06]. Oslo [Hol96]. Other [Del08, Den01, Gri01b, Kid96f, Pol95]. Ottawa [Ges07a]. Our [Pos22, Bab97]. Outcomes [McG12]. Outer [Ges09b]. Overcoming [Ano99e]. Overview [BL12, LM94, Smi01b]. Owens [Ano98b]. Ownership [Dia12]. Oxford [RAG^{+}04, RCP^{+}13, Ges07b].

P [Smi99b, Yos24]. Padua [ACCH23]. Paid [Nel22]. Paint [Smi01b]. Pamela [RAG^{+}04]. Paper [CRB24]. Papers [Ano15b, Ano15c, Ano18b, Ano20n, Ano20m, Ano20-32, Ano20-33, Ano20-34, Ano21-28, Ano21-29, Ano21-30, Ano22n, Ano22o, Ano22p, Ano22-35, Ano22-36, Ano23o, Ano23p, Ano23q, Ano23-27, Ano24t, Ano24s, Ano24u, Ano24v, Ano24-36, Ano24-37, Ano24o, Ano25l, Lus21, NK98]. Paradigm [HPR14b]. Parametron [Yam19b]. PARC [Raw24]. Park [Ano97e, Ano22-67, LH95, ORB^{+}92, Sal05, Wei94a]. Parker [DSW92]. Part [BBC^{+}09a, BBC^{+}09b, BBW18, BBW19, Big20a, Big20b, Cop11a, Cop11b, Cor16, Cor18, Cor19a, Cor19b, Cor20, Goe02a, GH19, GH20, Hen05, HIIT05, HII^{+}05, Hey08a, Hey08b, Kee04, Kee05, Pre12, Roe15, Spi16a, Spi16b, Spi16c, WN05, WN06, Yos16a, Cor15c, Goe02b, GR02, KD21, Roe16b, Spi15a, TYD21a, TYD21b]. Participants [Ano92c, Cor21b]. Participation [Ano23r, Rus13c]. Partisans [Coo02]. Parts [Ñañ22]. Pascal [Kis98]. Past [Ber14, Bur05, Cor09, Cor21b, EHP22, Gra03, Nan09]. Pasts [Cor24a]. Pasts/Computer [Cor24a]. Patches [ABL^{+}24]. Patent [CRB24, McT98]. Patents [Con15]. Path [Lei17]. Patrick [And07b, Bal24]. Patterns [Cor08b]. Paul [RCP^{+}13, Tym99]. Pay [Cor22]. PC [CG06, CG07, Mer06, Sta07, Sum07]. PEACESAT [HUA23]. Pearcey [Ain23].

Peering [Lew16]. Peggy [Ano00b]. Pelkey [RR23]. Penetration [Hun12]. Pennsylvania [Win96]. People [Cro93, Ens04b, Kid15, RAG^{+}04, Smi04, Sta20b, Cor98]. Percy [CRH^{+}21, CRB24, RC23]. Peresus [RAG^{+}04]. perfection [CK01a]. perpetual [Goy96]. Persecution [LM94]. Person [Kir20]. Personal [Ber06a, Cam03a, End13, Gal11, Gra02a, Hoa03, Hum02, Jac16, Kat10, Kee04, Kee05, Nor23, Oba22, OARHC10, Pet17, Phi12, Roj04a, Sta17, Sta20b, Bar03, Bir00, Hai03a, Kli99, Log98, Mey99, Sha00]. Perspective [Asp92a, Asp92b, Asp92c, Asp93b, Asp93c, Asp93d, Asp93e, CK12, Hoa03, Hum02, Nor01, PPM^{+}23, Spi08, Joh99]. Perspectives [Ano09e, CKGS08, De 21, Kid96c, NM09, PS23]. Persuasion [Jac16]. Pervasive [Ano24m]. Peter [Gri22b, Wal12]. Ph.D. [HO23]. Phenomenon [And07b, Cro93, Eva24]. Phil [Owe96]. Philadelphia [Hea01b]. Philco [Kos03a, Ros04]. Philips [Pee09]. Philosophers [Hem19]. Phone [Tra19]. Phonetic [Hen05]. Photolithographic [Lat13]. Physicist [And07b]. picture [PM12]. Pictures [RCP^{+}13]. Piece [CK02a, Cor01a, Cor09, Dic13, DA14, Edw01, Gri00b, Lap09]. Pillars [CK02b]. pilot [Ano99e]. pilot-induced [Ano99e]. Pinkerton [Ano98a]. Pins [Cor21a]. Pioneer [And07a, And09, Bal24, Cam03b, CD14, CH15, Goe02b, Goe02a, Hai08, Hai10b, Hai11, Nik17b, Nyi11, SL12, SK11, Spi15a, Spi16a, Spi16b, Spi16c, Sub13, Twe93, WH17, Yos13b, dPRG21, Cro00b, Kis99a, Hus97]. Pioneering [Bir00, CK08, Ste06, SW95, Hal00]. Pioneers [CM25, RAG^{+}04, Kid00d]. Piracy [Lis24]. Pisa [CG20]. Pit [Med08]. Pittsburgh [Asp99, Ges08]. Pivot [Day23]. PKI [Par15]. Plain [Lap09]. Plains [Ges09a]. Planar [LR12]. Plankalkül [Gil97]. Planned [GH93]. Planning [Gri11b]. Plans [Bro00, Wil00c]. Plated [Hai01]. Platform [Spr18b, Bis98]. Platforms [Sum07]. Play [Rus16d]. Plug [Tak05]. Plug-Compatible [Tak05]. Plus [Est02]. Pocket [RP12]. Podcast [Ano22-59, Ano22-60, Ano22-61, Ano23-49, Ano23-50]. Point [Cor15a, FVW25]. Policy [Bak17, Lon08, Sub14, Van10b, War15, Kid00f]. Polish [Gaw19]. Political [Ich06, McD10b, Mul20, Wei08, Joh99]. Politics [Don10, Kan21, Nor23, Per20, Tin10]. Pollak [Kis99a]. Pong [Low09]. Poor [Kid00i]. portrait [Cro00b, Spi03]. Portrayal [Tra24]. Portrayals [Vog17]. possible [Cor97]. Postal [De 95]. Postcolonial [PN16]. Postel [Ano99d]. PostScript [War18]. Postwar [CK08, CO12, Kid23, Sch17]. Potential [Ano24-59, Nel17, Ano25k]. Power [And22, Coh15, Ens04b, Lew17, Sla08]. Pp [Smi78, Ano97e, GH18b, Mor03, RCP^{+}13, Smi03, Win03]. pp. [Cer03, RCP^{+}13]. Practical [Kid15]. Practice [Ayl10b, Dun21]. Practices [Cor07, Cor08b]. Precision [WD98]. preconditions [Cor97]. Presence [Cor25a]. Present [EHP22, Nan09, Neb98]. Preserve [Shu06]. Preserving [Cor09, Cor21b, Gra03, Joh20]. President [Ber09]. Presper [Eck96, Lee95c]. Press [Ber03d, GH18b, Mor03, RAG^{+}04, RCP^{+}13, Sid20, Win03]. Price [Bro77, Nor77, Smi78, Ash76, Bal76, Dra76a, Mor84, Nas75, Ole77]. Priestley [GH18b]. Primes [Bul15]. Principal [Kid97b]. Printed [Kid15]. Privacy [Ano19-45, Ano22-59, Ano22-60, Ano22-61, Ano23-49, Ano23-50, Ano19j, Ano21i, Ano21j, Ano23i, Ano24t, Ano24-31, Ano24n, Ano24p, Ano24q, Ano25e, Ano25f]. Private [Hym92]. Privatizing [Abb10]. Prize [Yam19a, Ano98b]. Probability [Dry18].

Problem [Fri94, Gar21, Nel17, Wu15, Fit02b]. Problems [HS14]. Process [Ayl10a, SW95, Woh06]. Process-Control [Ayl10a]. Processes [Cob23]. Processing [Ber06a, Ber06b, CG06, Cor25a, ES05, Hai06, MR22, Mak06, NM09, Sin09, Urq24, Yos01, Bir00, Kir98, Kis99a, Tat10]. produced [RCP^{+}13]. Product [Boh13, Cor25b, Gan04b, Hai02, Hen05, McG09, Ano02e, Pos98]. Production [FC15, Ken22, O’D09, Van95]. productivity [Kid96g, Ken22]. Products [GJ02, Yos15b]. Profession [Zwe99]. Profession/occupation [Zwe99]. Professional [Ano19-40, Ano21k]. Professionalism [Ens01]. Professionalization [FHD17]. Profit [Moo03]. Program [Day21, HPR14c, Hem19, McG09, Rop07, WI14, Cer97, Gri96]. Programmability [HP18]. Programmable [RP12]. Programme [Dav17]. Programmer [Vog17, Wil00c, Bro00]. Programming [Alb14, All18, Ber94, Ber09, Bul21, CK92c, CK98, CK07b, De 21, Got23, Gri11b, Kos03a, Kos03b, Nof10, Nof25, Pau14, Pay14, Pri21, PP21, Ran94, Whe92a, Gil97, Roh99]. Programs [Dun21, Hun12, Jon03, Lus21, Pau14, Roj21, SB98]. Progression [Smi16b]. Project [CG20, Dur10, KH21a, McJ17b, Moo01, MK14, OARHC10, PHL92, Rob16, Spi05, Wil01b, Ibb99, Joh99, Bur23, Gas99, Gri97, Gri98, KH21b, Lav22b, LFS^{+}92, LR92, Ran15, Roe12, Wil18, Yos22]. Projects [CdLF^{+}22, Paj08, Gri98]. Prolegomena [Ray02]. Proliferation [Ber06b, Nof25]. Prolog [Lee92h]. Prometheus [PM12, Smi99b, Kid99e]. Promoting [Ayl10a, Kid23]. proof [Mac95]. Proposals [Ano20o, Ano21n, Ano22q, Ano22r, Ano22t, Ano22s, Ano24w, Cor25b]. Prosaic [Ayl10a]. Prosopography [Nel17]. Protection [O’D09]. Proto [Sta07].

Proto-PC [Sta07]. Prototype

[Roe09, SB98, Spr18a]. Prototypes [Sta20a]. Provision [AW94]. Proxies [Vid23]. Prudential [Yat97]. Psychological [Lus21]. Public [Par15, Wil18]. Public-Key [Par15]. Publications [Ano22-62, Ano23-40, CB19, GH18b, KW19, Sey19, Smi78, VM18]. Publics [Vid20]. Published [Ano20y, Ano20z, Ano21t, Ano21u, Ano21v, Ano21w, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano23y, Ano24t, Ano24-31]. Publishers [Ral04]. Publishing [BBW18, BBW19, GH18a, GH19, GH20, Pag08, RAG^{+}04, Sey18b, Spr18a, Spr18b, WH17]. Punched [Kis92b, Kis05, Wil02, vdE94, Hei97, Kis97]. Punched-Card [vdE94]. Purchasing [Lew17]. Purdue [RR04]. Purloined [Lap09]. Purpose [ERS03, Sma93]. Pushing [Jes07]. PVC [Ano19-44].

Quadrupole [CO12]. Quantum

[Ano20-38, Ano22-53, Ano23-42, Ano23-43]. Quarter [Ano19-38, Ano19-48]. Queens [Buc20]. Queries [Ben93a, Ben93b, Ber92, Buc95a, Buc95b, BS96, Car92, Dum94, Ekl92, GM96, Gro92, Gro93, Hal92, Hym92, Jac93, Kis92b, Kis92a, Rob93, Ros92, Sha00, Tod93, Tro93, Tro95, WF92, Wil00b]. query [Fit02b]. Quest [Ano00b, Cer25, Wil00a]. Question [Ayl10b, Day21, Ens01]. Quetelet [Ber92]. Quevedo [Gon20, dPRG21]. Quing [Ano92d].

R [Ano20-39, Cer03, GH18b, Hai10b, Kit03, PM12, Shr24]. R&D [Bar05, Day23, Hea05]. Race [Nel17, Tom96]. Racial [McG21]. Radar [Gre13]. Radical [And07b]. Radimsky [Nik17a]. RAID [Kat10]. Rail [Urq24]. Rainbow [Ano22-59, Ano22-60, Ano22-61, Ano23-49, Ano23-50]. RAND [Mis16, GS98, GS01, GS04, Sch08]. Random [MB09]. Ranking [Rob19]. Rare [Bru17]. Raster [MB09]. Rational [VM24].

Rationality [TYD21b]. Raúl [Cer25, Roj04b]. Raymond [Spi16b, Spi16c]. Raymund [Mis99]. Raytheon [Kos07]. RCA [AL12]. RDBMS [CK12, HN13]. Reach [Lis24]. Readings [Bul21]. Ready [WJ05]. Real [Cer14, Got23, Har22b, Mar12, Gas99]. Real-Time [Got23, Mar12, Gas99]. Reality [DK23, Kid97a]. Realizing [BK01]. Really [Pla16]. Reanimated [LHS^{+}25]. Reasoning [Jon03]. Rebuilding [Sal05]. Reciprocals [Maz24]. Reckoners [WJ05]. Recoding [Rus13c]. Recognition [Ano15l]. Recollection [Gra02a, Kat10, Kee04, Kee05]. Recollections [Bau07, Joh02, Kap07, Ros04, Rub06, Sac07, Kid00d]. Recomputing [Roe12]. Reconfigurable [Est02]. Reconsidering [HPR14c]. Reconstruction [RDGH05]. Reconstructions [Swa05b]. recording [Wei00]. Records [Cor09, ORB^{+}92, Rop07]. Recovering [DAT25, Fed21, Med05, TYD21a]. Recovery [Sta19]. Red [Sta15]. Rediscovered [HT12]. Rees [Ano98c]. References [Ano92e]. Reflecting [van19]. reflection [Hai03a]. Reflections [Gra02b]. Regeneration [Gio93]. Regnecentralen [Klu99]. Reign [Ano92d, Dry18]. Reimagining [Alp15]. Reinterpreting [TYD21b]. Reinvention [Hou21]. Related [Ano20-27]. Relational [Dar12, Gra12, Gra13, Gri12, WC12]. Relay [Yam19b]. Relevance [Cob23, MA96]. Reliability [VR22, WMB22]. reluctant [Gri98]. Remaking [GH18b]. Remember [Wil06]. Remembering [Ano02f, CK13, Hai06]. remembrances [BBB^{+}98]. Remington [Sch08]. Reminiscence [Cam03a]. Reminiscences [Ber03f, Ran15]. Remote [Rei23a]. repeat [Lee96]. Replicated [HC20]. Replicating [Bur05]. Report [Dod00, Kid96a, PP21, vN93]. Representation [Cob23, FC16, Poi25].

Reprint [CK01b]. Republic [Cor12]. Repurposing [Hic08]. Requirements [Hen05]. Rescuing [Kid99e, Smi99b]. Research [AL12, Ano21-45, Ano23-51, Cor04, Cro92, ERS03, EHP22, Har22b, Mor03, ORB^{+}92, Spe98, Ber03f, Est00, Klu99]. Researchers [Poi25]. Researching [Cor02]. Reservation [Dor94, Ekl94]. Reservations [RAG^{+}04]. Reserve [da 15a]. ReserVec [Dor94]. Reservisor [Ekl94]. Resilience [HmWK23]. Resistance [HmWK23, Sla08]. Resolution [Ano95]. Resources [Ano16j, Mag20, Mar02]. Response [All19]. Responses [ABL^{+}24]. rest [Lee95e, Hos24]. Restoration [Spi05, Too10]. Restoring [LHS^{+}25]. Restructuring [Par16b, Kid00f]. Results [Rob19]. Resurrection [WI14]. Retail [VN11]. Rethinking [Cor19a, Cor19b, Sla08, Cor20]. Retrieval [GK02, Mal02]. Retrospect [Twe93, Veh99]. Retrospection [End13]. Retrospective [Spr16, WI14]. Reunion [ORB^{+}92, LBA00]. Reuse [Ano20n]. Rev. [Ano97e]. Revealed [HW17]. Review [Ake00, Ash76, Bal76, Bar03, Ber03d, Bro77, Cer03, Cer25, Cor98, Dra76a, EMW02, Fer22, Fie03, Fit23, Fra23b, Ger04, Gui99, Hou21, KH21b, Kan21, Ken22, Kid97a, Kid97b, Kid99a, Kne21, Kra22, Léc21, Lee99b, McG21, Mis99, Mor03, Nas75, Nor77, Nor23, Ole77, Pos00, Rio00, Roj04a, Roj04b, Sei03, Shr24, Sid20, Smi99a, Smi02, Smi03, Smi78, Smi99b, Spi03, Spi24a, Sza18, Tym99, Vid23, Wal00, Wei94a, Wil00a, Win03, PM12, Rus16a, Rus16d, Rus17, KH21a, Roj04a]. reviewed [Ano14e, CB19, GH18b, KW19, Rus13b, Rus14b, Rus14a, Rus15d, VM18]. Reviews [AC17, Ake98, Ald02, Ann12, Ano97b, Ano97e, Ano97d, Ano97f, Ano00b, Ano03b, Ano04d, Ano08a, Ano08b, Ano09f, Ano09g, Ano10b, Ano11b, Ano11c, Ano11d, Ano11e, Ano14e, BBS06, CK97, CK01a, Cer92a, CFSL92, Cer92b, Cer92c, Cer93, CB19, Con17, Dow09, EA17, GH18b, Hey09, HC06, Kid93a, Kid93b, Kid93c, Kid94a, Kid94b, Kid94c, Kid94d, Kid95, Kid96a, Kid96c, Kid96d, Kid96e, Kid99e, Kid00d, Kid00b, Kid00c, Kid00e, Kid00f, Kid00h, Kid00g, Kid00i, Kid00k, Kid00j, Kid00l, Kid01b, Kid01c, Kid01a, Kid02, KW19, LA12, Lee97, Neb98, Nor98, PM12, Roj03a, Roj03b, RAG^{+}04, Roj04c, Roj04d, Roj05a, Roj05b, Roj05c, Roj05d, Roj06a, Roj06b, Roj07, RSCF07, Rus12a, Rus13a, Rus13b, Rus13c, RCP^{+}13, Rus14b, Rus14a, Rus15d, Rus15c, Rus15a, Rus15b, Rus16d, Rus16b, Rus16c, Rus17, Smi01a]. Reviews [Swe98, VM18, Wei94a, WFE07]. Revised [Cop11a, Cop11b]. revisited [Old95]. Revolution [ES05, Hai01, Léc21, Pay14, RAG^{+}04, Sla08, Tym99, Rus16a]. REXX [Cow94]. rhetoric [Joh99, Owe96]. Rhodes [Tom00b]. Rica [VHCSCC15]. Rice [Hai10b]. Richard [Ano15a, Ano17n, Ano21-44, Lee98a, Lee98b]. right [Ano17b, Ano17c, O’D09]. Ring [Chi14, Nee92]. Ringing’ [OK07]. Riordan [Kid99a]. Ripples [Bor08]. Rise [Dur10, Fra23a, Gri01b, Hou21, Mor03, Oba22, Rub06, Tak05, CCK00, Lee95f, LS00]. Risk [Sla15, Ste06]. Rivalry [Lew17, SR14]. Rob [Hai03b]. Robert [Ano99f, Hai04c]. Roberto [HC20]. Rocappi [Sey18b]. Rock [Ano14f, Ano14g, Ano15n, Ano15o, Ano15p, Ano16m]. Rodney [Hsu23]. Roger [Wei96a]. Rojas [Cer25]. Role [Cor07, Kid19b, Sat11, O’N95, Puc96]. roles [Goy96]. Rolf [Ano13h]. Romania [Ano99b]. Rombauer [Gri03a]. root [RG02]. Roots [Mol14, Nov11]. Rope [GH18b]. Rosenfeld [Gui99]. Ross [Mor03]. Rosseland [Hol96]. ‘Rough [Rus06]. Routing [McQ09]. Roxanne [Sub13]. Royal [Gre13, Lav22b, Pee09]. Rule [Kid15, Len22]. Rules [Wes17]. Rundbrief [PPM^{+}23]. Running [Rus06]. Ruptures [da 15a]. Rush [Est22]. Russia [Kli99, Pro99]. Russian [TYD21a, TYD21b].

S [Ano98c, CK13, Gri03a, Ros04]. S-2000 [Ros04]. S. [CG94]. Saab [Per99]. Sabre [CMM95, Hea02]. Sala [Rei23b]. sale [LS00]. Sales [Cor15a]. Salesman [Cor13, Cor14, Bro99a]. Salton [Wei96f]. Sammet [SWH18, Ber09]. Samuel [Wei92e]. Sand [Bur24]. Sandeep [Fer22]. Sandhill [RAG^{+}04]. SAP [Lei08]. SAPO [Dur10]. Sara [RAG^{+}04]. Saturation [Cor15a]. sausage [Lee95e]. Saved [SH19]. Scale [Ayl10b, Hei11, JMSV10, MR22, KS97, Log98]. Scales [WT03]. Scarcely [Bru17]. Scene [Hod19]. Schafer [PM12]. Schickard [Kis01]. Schlombs [Ken22]. Schlumberger [SST22]. Schmidt [Mis99]. Schoenberg [Hai05c]. Scholars [Yos01]. Scholarship [Ano21-44, Day21, Ano17n]. School [Win96, Cor15a, FH14]. Schools [Nis23]. Schwilgué [Roe16a]. Science [Ano20n, Ano21c, Ano21d, Ano22c, Ano23a, Ano24c, Aum11, AG21, Car15, Cro92, De 21, Gio93, GH18b, Gup07, HO23, Hem19, HC20, JB07, Kel22, Lon08, Min07a, Min07b, Nof23, NK98, Sch23, SW96, Smi78, Spi08, Asp99, DFM^{+}99, Est96, Kat97, Kli99, Pro99, Sei03, Ral04, Yam19b]. Sciences [RR04, Ste22, Swe05]. Scientific [AW94, DGG^{+}08, Gro93, HW17, McC22, Pri24, Est00, Kis99a, Cro99]. Scientist [Nyi11]. Scientists [AW94, LH95]. Scope [Hei11]. Scott [PM12]. Screen [CHGW17]. Scripts [Sin09]. Scritchers [Dum94]. Scruffy [Poi25]. SDC [Hem14b, Mis16]. SDRC [EK24]. SDS [KH21a, KH21b]. SDS-940 [KH21a, KH21b]. SEAC [Kir98]. Sean [Asp14, RCP^{+}13]. Search [Ano15g, Ano16b, Ano17f, Ano17e, Bur02, Gra20b, Jon03, McJ17a, Rob19, Mar10a]. Searching [SRG02]. Second [Neu06]. Secondary [AKM08]. secrecy [Kid96f]. Secret [Wil18]. Secrets [HW17, Mar10a]. Sector [Day24, Tol21, WT03]. Security [Ano19j, Ano19-45, Ano20-37, Ano21i, Ano21j, Ano22-59, Ano22-60, Ano22-61, Ano23i, Ano23-49, Ano23-50, Ano24n, DeN15a, Fid18, Lap11, Mis16, Sla15, Sub20, Vag24, War15, Yos15a, Yos15b, Yos16a, MP97]. Seduction [Bak17]. Seek [Ano22-62, Ano23-40]. Seeking [Ano14h, VR22, WMB22]. seen [Kis99b]. Selected [NK98]. Selection [Lew17]. Seller [Cor24b]. Selling [Mer06, Was13]. semiconductor [Kid00f, Law10, LR12]. Semiconductores [dÁ24]. Seminal [CG20]. Seminar [Mit08]. Seminars [Sey19]. sequence [Kis95]. sequence-controlled [Kis95]. Sequences [Roe15, Roe16b]. Serbian [MPDT12]. Series [CK01b, ORB^{+}92, Win03, Yos22, Yos23]. Service [Ano15a, Ano21-37, Ano21-38, Ano22-51, Ano23-41, Ano24-60, Ano25-27, BYO^{+}22, Cro99, Hai02, Hai04b, Dav17, De 95]. Services [Day23, Gra02a]. session [Hus97]. Set [Hei11, Mal02]. Seven [Lee97, Gas99]. Severo [Smi03]. Seybold [Sey19]. Seymour [Ano17y]. Shape [Bro10]. shaped [Kid00e]. Shaping [Ake98, Hei97]. Share [Ano21-45, Ano23-51]. Shared [Fev13, Gal11]. Sharing [CKGS08, Hai05b, Hem14b, LHS^{+}25, Lee92f, Lee92e, Lew14, Ros92, Run08, Spi15a, Spi16a, Wie10]. Shift [San11]. Shinkansen [Has16]. Shirley [Spi18]. Short [Got23, Smi04, Sze99]. Shorts [DAT25]. Should [Cor22, Shu06]. shuttle [Ano99e]. Side [Par07]. Sidney [Ano17y]. SIGCIS [ABL^{+}24]. SIGGRAPH [Shr24]. Sight [Lap09]. Sightings [Ano01c, Ano02e, Ano02c, Ano02d, Ano04b, Ano20-27, Asp01c, Cam04, CE02, CE03a, CE03b, CE03c, CEG^{+}03, CRM^{+}04, FC00b, FC01a, FC01b, Gri99, Gri04a, Kit05a, Kit05b, Kit05c, Kit05d, Kit06a, Kit06b, Kit06c, Kit06d, Kit07, KSFY07, KYG^{+}07, KFS^{+}07, KMA^{+}08, KHS^{+}08, Kit08a, Kit09a, Kit09b, Kit09c, Kit09d, Kit10a, Kit10b, Kit10c, Kit10d, Kit11a, Kit11b, Kit11c, Kit11d, Kit12c, Kit12a, Kit12b, Kit13a, Kit13b, Kit13c, Kit13d, Kit14b, Kit14c, Kit14a, Kit14d, Kit15a, Kit15b, Kit15c, Kit15d, Kit16a, Kit16b, Kit16c, Kit16d, WS17, YS12, FC00a]. Sign [Mor20]. Signal [ES05]. Signetics [Sti24]. SIGPLAN [PP21]. Silicon [Ber03d, Bur24, RAG^{+}04, dÁ24, Est22]. SIM8 [Sta07]. SIM8-01 [Sta07]. SimH [Sup15]. Simon [Hey08a, Hey08b]. Simple [HS14]. SIMULA [Hol94]. Simulating [ZR06]. Simulation [Bak17, Rei08]. Simulations [HPR14a]. simulator [Per99]. Sing [Ano97d]. Sinhala [Nan09]. Sisson [Wei96a]. Sixty [Kne17]. Size [Kis92b]. Skår [Ano13h]. skepticism [Bow96c]. Skills [Ano20t]. Slashdot [Cer02]. Slice [Lun18]. Slicing [Tol21]. Slide [Kid15]. Slot [Tur11]. Slovakia [DFM^{+}99]. ‘Slowly [OK07]. Small [CVRM15]. Smart [RAG^{+}04]. Smithsonian [Dod00]. Social [Ano18u, Cor22, EM94, Ens04b, Lee92i, Par15, Ran14, RP12, Wei08, GC00]. Socialist [Lis24, Sch23]. Societal [Sch23]. SOCIETY [Ano22-38, Ano17n, Ano17l, Ano18m, Ano18n, Ano19-34, Ano19-35, Ano21-34, Ano21-35, Ano21-36, Ano22-37, Ano22-47, Ano22-48, Ano22-39, Ano22-49, Ano22-51, Ano22-40, Ano22-50, Ano23-28, Ano23-31, Ano23-32, Ano24-41, Ano24-38, Ano24-39, Ano24-40, Asp15, Kid99e, Lav22b, Ano15m, Ano16h, Ano16i, Ano16g, Ano16j, Ano17m, Ano17k, Ano18k, Ano18j, Ano18l, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-36, Ano20-29, Ano20-35, Ano20-30, Ano20-32, Ano20-36, Ano20-33, Ano20-31, Ano20-34, Ano21x, Ano21-28, Ano21-31, Ano21y, Ano21-29, Ano21-37, Ano21-32, Ano21z, Ano21-30, Ano21-38, Ano21-33, Ano21-27, Ano22n, Ano22o, Ano22-32, Ano22-35, Ano22-41, Ano22-33, Ano22-36, Ano22-42, Ano22-46, Ano22-43, Ano22-44, Ano22-34, Ano22-45, Ano23o, Ano23p]. Society [Ano23q, Ano23-33, Ano23-37, Ano23-34, Ano23-38, Ano23-35, Ano23-29, Ano23-39, Ano23-27, Ano23-41, Ano23-36, Ano23-40, Ano23-30, Ano24u, Ano24v, Ano24-44, Ano24-42, Ano24-45, Ano24-43, Ano24-36, Ano24-35, Ano24-37, Bow98b, Woo95, Ano20y, Ano20z, Ano21t, Ano21u, Ano21v, Ano21w, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano23y, Ano25d]. sociological [Mac95]. Sociologist [Sub13]. Sociotechnical [Par15]. Sociotechnological [EHP22]. Soft [Tym99]. Software [Ano25i, Ano25v, Bal24, Ber06a, Ber06b, CK97, CK07a, CG06, CG07, Con15, Cor02, Dia16, End04, Ens09, Fed21, Goe02b, Goe02a, Gra02a, GJ02, Gra03, GJ12, Gra20a, Gra20b, GS04, Hai02, Hai05b, Hai08, Hai09, Hai10b, Hai11, HW17, Hod19, Hum94, Hum02, Joh02, Joh20, Kap07, Kas24b, Kee04, Kee05, Kne17, Lei08, Lei17, Mah04, Mah08, McD10a, McG21, Mer06, Pos22, Pug02, Sat11, Sch06a, Shu06, Sie10, Sta14, Sta17, Sta19, Wir08, Woh06, Yos15b, ZS23, Ano02e, Bab97, DV18, Joh98, Kid00d, Pos98, Sha97, Bau06, Bro23, Gra21b, Lei17, Mag09, Pag08, Ano22m]. Solicited [Ano17y]. Solid [Gre13, Har99]. solid-state [Har99]. Solla [Bal76, Bro77, Dra76a, Nas75, Nor77, Ole77, Smi78]. Solutions [Hen05, HIIT05]. Solver [Fri94, Gar21]. solving [Mar10b]. Some [Ber03e, OK07]. Sonic [RAG^{+}04]. Soroban [Che25]. Sorting [Kid00j]. Source [Ens04a]. Sources [Nel17, Wil98a, Van95]. Soviet [Ber03f, CP23, CG94, FBK06, Ich06, Lew16, Sta15]. Sovietization [Dur10]. SP Subscribe hhalf [Ano19-46]. Space [Dow04, Low09, VT21, Ano99e, Gas99]. Spain [FHD17]. Special [Ano20o, Ano21n, Ano22q, Ano22r, Ano22t, Ano22s, Ano24w, Ber94, Bro23, MK20, Woi25]. Specified [OB08]. Spectrum [CK14]. Speech [Mak06]. Speed [EM94, WD98]. Sperry [GS98, GS01, GS04, GS07]. Spies [Fit23]. Spitting [Vog17]. Splines [VM24]. Splitting [Sha97]. Sponsoring [Yos01]. Spread [Cor08b, Kid19b]. Spreadsheet [CK07b, Kap07]. Spreadsheets [CG07, Der20]. Spring [Shi19]. SQL [Cha12, Deu13, HN13, Wad12]. SQL/DS [HN13]. square [RG02]. square-root [RG02]. SRI [Par07]. St. [Kid00a]. Stages [CG20, Uss96]. Stalking [Kid98]. stand [Bis98]. stand-alone [Bis98]. Standard [Deu13, Hos24]. Standardization [Lis24, Mis99]. Standards [Hos24, McD10a, Par16b, Rus06, Rus12b, Sum07, Gur96, Hus97, Kir98]. Standing [Vid23]. Stanford [Bra13, Gil07]. Stanley [BH99]. Starr [Sub13]. Stars [Ano14f, Ano14g, Ano15n, Ano15o, Ano15p, Ano16m]. Start [Ano92d, GJ02, Mor03, Gri98, Kir98]. Start-up [Mor03]. Startup [Ano15p]. State [Cla10, Lis24, McQ09, Pri24, Har99]. State-Socialist [Lis24]. States [Rus13a, Cor97, Jes13, Kid19b, Kid23, Lla24, Mir17, Tra19, Tro93, Tro95, Vag24]. Stating [McG12]. Statistical [Nik17b, Smi12]. Statistics [Alo07, vdE94]. Stay [Ano15q, Ano19-47, Ano21-46, Ano21-47, Ano22-63, Ano23-44, Ano23-52, Ano23-53, Ano24-34, Ano24-52, Ano24-53]. steel [Van95]. Step [CVRM15]. Stephanie [Spi18]. stepped [Kis99b]. Stepping [Rob24]. Sterling [Bau06, Gra21b]. Steve [Gri14]. Still [Ben93b]. Stockholm [Ges12b]. Storage [Hai09, Kat10]. Store [Bac09]. Stored [HPR14c, Pow19, Rop07, Cer97]. Stored-Program [HPR14c, Rop07]. Stories [Cor14, Kid99e, Ste03, Lee99b]. Stork [Kid97a]. Story [Bel25, Ber24, Cer14, Dia16, Fel11, Gra21b, Hea01b, Kid99e, Lei08, Lor19, Men13, Pre12, Sup15, Wei94a, Lee95e, McK95, Smi01a].

Strano [Gri22c]. Strategies [Ano15v]. Strategy [JMSV10, Lew17]. Strela [Ich06]. Strela-1 [Ich06]. Stretch [Bro10]. Stretch-ing [Bro10]. Stripped [Roe17]. Structure [CK97, Est02, JMV10, Nof23, Roe15]. Structures [BS11]. Struggle [GS07, Smi99a]. Student [Ano17n, Ano18v, Ano21-44, Cam13, Mil06]. Student-Oriented [Cam13]. Studies [Kli11, Mis99, Est96]. Studio [OB08]. Study [BB02, CK97, Hem19, Hol94, Par07, Lan00, Lee96]. Studying [Cor07, Cor15c, Cor16, Cor18]. Style [AD14]. Styles [Alb14]. subcommittee [McK95]. Subject [Ano13a, Ano00a]. Submissions [Ano20-38]. Submit [Ano19-53]. Success [Ber24, Gra21b, Ich06]. Supercomputer [Lew17]. Supercomputers [EM94]. superconductive [Ber03f]. SuperPaint [Sho01]. Superscritcher [CF92]. Supply [Bur24, Kid99e]. Support [Cla10]. Surveillance [DeN15a]. Survey [Geo08, Eck98]. Survive [GS07]. Susan [Hin18]. Susanne [Mis99]. Sustainability [HmWK23]. Sustainable [Ano19m, Ano19n, Ano23k, Ano23l, Ano23m, Ano24-54]. Sutherland [SB23]. Svein [Hol96]. SW [Ano19-48]. SWAC [Cor08a, HTAY97, Hus97]. Sweden [Ges12b, Joh96, LS08, Pet05]. Swedish [Joh96, Joh99]. Swiss [Kid00k, Neu04, Neu05]. Switch [Rob24]. Switches [Pau14]. Switching [Tra24, Gre95]. Switzerland [Bru17]. Sybase [Eps13]. Sydis [HVB04]. Symantec [AC16]. Symbol [Maz08]. Symposia [Yos22]. Symposium [Ano20-37, Cob23]. synthesis [Sha97]. Synthesizers [Col14]. System [ABL^{+}24, BH95, Dor94, Doy01, Fei22, FM93, Gri12, GK02, Has16, Hem14b, KH21a, KH21b, Kat10, Kis05, Len22, MR23, MF93, Nie09, Run08, Sho01, Gas99, Rei97, BV19, Bru09, DAT25, Ekl94, Kru09, Lee92f, McG09, Mor20, Urq24, Wad12]. System/370 [Wad12]. Systèmes [Ber24]. Systems [Ayl10b, Bal24, Bar05, BH09, BG22, Bru09, Coh15, Got23, GB09, Gra12, Gra13, Har22a, Hum94, Kas24a, Kel22, Kos07, LHS^{+}25, Nik17b, Pee09, Smi01b, SD98, Tol21, WC12, WG19, Whe92a, Wie10, Wil03, Yos16b, Ari00, Eck98, Hei97, KS97].

T&W [Laz24]. T&W/Versacad [Laz24]. Table [Ano13i, Ano14i, Ano14j, Ano15r, Ano15s, Ano15t, Ano15u, Ano16n, Ano16o, Ano16p, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano18z, Ano18w, Ano18x, Ano18y, Ano19-49, Ano19-50, Ano19-51, Ano19-52, Ano20-46, Ano20-47, Ano20-48, Ano20-49, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano22-64, Ano22-65, Ano22-66, Ano23-54, Ano23-55, Ano23-56, Ano23-57, Ano24-55, Ano24-56, Ano24-57, Ano24-58, Ano25w, Ano25x, Pol97]. Tables [CCK00, Fei11, Gri97, Gri98, Gri01b, Pol95, Roe12]. Tabulating [Cro03b]. Tabulation [LM94]. Tabulator [Hai01, Kis95]. Tactics [ABL^{+}24]. Taiwan [Lin16, Tin11]. Taiwanese [Tin10]. Take [Ano14k, Ano15v]. Takeover [Bau06]. Taking [HII^{+}05, LM94, MT06]. Tale [Pag08, Hol97]. Talking [Gui99]. Tape [CK92a]. Tarui [CO12]. Taube [Gar21]. tax [Bro99a]. TBD [Ano19-53]. TC [Ano25y]. Teaching [Cro92, Kir20]. Tech [Ano17o, Ano17p, Ano25q, Ges09a, KB21]. TechIgnite [Ano16q]. Technical [Hod19, Par08, Sei97, McK95]. Technique [Ram97]. Techno [Sim25]. Techno-Utopia [Sim25]. Technological [BB02, Hol94, Ich06, Mol14, Var94, Bow96c, Joh99]. Technologies [Ake98, CK08, Cor12, Cor15c, Cor19a, Cor19b, Cor20, Cor25b, HmWK23, MB09]. Technology [Ano18m, Ano18n, Ano20-50, BLR11, BLH12, Bre25, Car93, CM25, Cor06b, Cor08b, Day24, Don10, Feu06, GA22, Ges09b, Hal18, JMV10, Ken22, Kid99e, Lee99b, Lev05, LM94, McD10b, Mis99, Mor03, RAG^{+}04, RCP^{+}13, Sub14, Win03, Yam18, DFM^{+}99, Hei97, Joh99, Kis97, MP97, Per99, Pos00, Wri98, Ano00b, JB07, Joh96, Mur19, WD98, Kid00i]. Técnicos [FHD17]. Tecnología [dÁ24]. TECO [Mur09]. Telecommunications [Ano97f, BBC^{+}09a, BBC^{+}09b, Cer03, Pau14, PM12, Mis99]. Telefunken [JMSV10, JMV10]. Telegraphy [Win03]. Telematics [Fra18, Urq24]. Telenet [Vag24]. Telephone [Irv01]. Teletext [Ste06]. Teletext/Videotex [Ste06]. Television [ORB^{+}92]. Telidon [DAT25]. Tell [Cer14]. Telnet [Fit03b]. tem [Hic19]. Templates [Kid19a]. TEMS [Ano18m, Ano18n]. Ten [Ano24a]. Tenet [Run08]. TENEX [Mur15]. Tension [DeN15a]. Tensions [Sch23]. Term [Lee92e, Wei08, Sha00]. Terminal [HVB04]. Ternary [GHV05]. Territory [Roy11]. Test [Gon24b]. Testing [Lus21]. Texas [Phi12]. Text [Day21, Gaw19, Hos24, Van12]. texts [Kat97]. Textual [Cor01b]. Thai [KKW09]. Theater [Gri22c]. Theatre [Gri22a]. Their [Cor20, Cor25b, Est96]. Them [RP12]. Theoretical [AG21]. Theory [Cor08a, Fra23b, Gri03b, Kid99e, Fra23b]. There [Fra04, LaV24, Lig03, Roe16a, Cor15a]. These [Cor22]. things [Kid00j]. Think [CK02a, Cor01a, Cor09, Dic13, DA14, Edw01, Gri00b, Lap09]. Thinker [GG92]. Thinking [Che25, Lon04]. Third [GS01, GS09, Lee00c]. Third-Generation [GS01, GS09]. Thomas [GH18b, RCP^{+}13, GHV05, GH18b, Hai18, Smi99b]. Thompson [Bro18]. Those [Lee96]. Thought [Cob23, Fel11, Gon24b]. Thoughts [BS11, Ber03e, Sum07]. Three [Bul21, Gru05, Kir20, Kos07, RG02]. Three-Person [Kir20]. Thwart [CRB24].

Tidal [vdE92]. Tillitt [SRG02]. Time [Bar24, CKGS08, Dow04, Fev13, Got23, Hai05b, Hem14b, LHS^{+}25, Lee92f, Lee92e, Lee92g, Lew14, Mar12, Ros92, Run08, Spi15a, Spi16a, Tol21, Wie10, Gas99, Smi00, Van95]. Time-Shared [Fev13]. Time-Sharing [CKGS08, Hai05b, Hem14b, LHS^{+}25, Lee92f, Lee92e, Lew14, Ros92, Run08, Spi15a, Spi16a, Wie10]. Times [Pat14, Bab97]. Timesharing [MR23]. timetable [Cor98]. Tinkering [Tin11]. Titan [Nee92]. Title [Wei96b]. Today [Cor09]. Token [Chi14]. Tokyo [Yam19b]. Tom [And09, BG93b]. Tomash [Asp01a, Asp01b, Wil01a, Yos01, Yos13b]. Tomlinson [Spi16b, Spi16c]. Tommy [Hai18]. Tomorrow [Cer02, Spi03]. Too [Mar12]. Tool [ZS23]. Tooling [Cor16]. Tools [Spr18a]. Toolset [Got23]. Top [Ano20-50, Fei11, Hea05]. Top-Level [Fei11]. Top-Notch [Hea05]. Topics [AY11]. Topologist [And07a]. TOPS [Mur15]. TOPS-20 [Mur15]. Torchi [Roe16a]. Toronto [BGM02, Ges07a, Gri94, Hum94, Wil94b]. Torres [Gon20, dPRG21]. Total [Urq24, Nie09]. Tower [Nof25]. TR [JMSV10, JMV10, Sie10]. Traces [Mit08]. Tracing [CP23]. Tractable [Jon03]. Trade [Ñañ22]. TRADIC [Bro99b]. Traffic [Has16]. Traffic-Control [Has16]. Tragedies [Ano00b, Ake00]. Training [Day24]. Trans [Dor94]. Trans-Canada [Dor94]. Transac [Ros04]. Transactions [Ano19k, Ano19l, Ano19m, Ano19n, Ano20m, Ano20k, Ano23j, Ano23k, Ano23l, Ano23m, Ano24t, Ano24-31, Ano24o, Ano24p, Ano24q, Ano25h, Ano25e, Ano25f]. Transfer [Lev05]. Transfers [Vid20]. Transformation [Per99, Owe96]. Transformations [Bul15]. Transforming [Bar07, Car93]. Transistor [CO12, GS98]. transistorized [Bro99b]. Transition [FH14, SB20]. Transitions [Coh15]. Transnational [Ñañ22].

Transparent [Nor23]. Trapped [Nor98]. Trenches [Smi03]. Trends [Ano20-50, KBSW11]. Trevor [Ain23]. Trials [Gil07]. Trip [Del08]. TriStar [RCP^{+}13]. Triumphs [Ano00b, Ake00]. trouble [Kid96g]. True [HC20, Sei03, Smi99a]. trust [MP97, Ano20-37]. TSUSC [Ano18-27, Ano25z]. Tube [CH15, CHGW17]. tubes [Log98]. Turing [Aga01, Bul21, Cop23, Day21, Gon24a, Gon24b, Hic08, Kid96b, OARHC10, Sau93, Smi02, Tro93, Tro95]. Turing-850 [OARHC10]. Turn [Yos21]. Turoff [Sub12]. Tutoring [Kel22]. Twentieth [Cor25b, Kid19b]. twig [Cos96]. Twin [CK02b]. Two [CB19, GH18b, Hem16, Hen10, KW19, Lav22a, LH95]. Tymshare [KH21a, KH21b]. Typography [McJ20].

U.S [RCP^{+}13, MP97]. U.S. [Bow96c, GA22, HO23]. Ubiquitous [Raw24]. Ubisco [Neu09]. UCLA [Est02, Kid23]. UK [BLR11, Cla10, Gre13, ORB^{+}92]. Ulery [Nik17b]. Ulm [Kir20]. Ultra [Kid96f]. unanticipated [Nor98]. Unbundling [Gra02a, Hum02, Ano02e]. Uncertainty [Hod19]. underlying [Bar97]. Understand [Cor01b, Kid00l]. Understanding [EHP22, Mis07c, Wei06]. Unfair [Cop04b]. Unfamiliar [Roy11]. Unfolds [Cor15c, Cor16]. Unicode [Hos24]. Unifying [Aum11]. Unigraphics [SAV24]. Union [Ber03f, FBK06]. Unisys [GS07]. United [Kid19b, Tra19, Cor97, Eva24, Jes13, Kid23, Kir99, Lla24, Mir17, Rus13a, Tro93, Tro95, Vag24, Wyv98]. Univ [RAG^{+}04]. Univac [Hea01b, Joh06, Kos03b]. Universal [Gon20, MK14, Aga01, Roj98, Smi02]. Universality [AD14]. Universities [AW94, Asp00]. University [Cla10, CdLF^{+}22, GH18b, Joh96, Mor03, RAG^{+}04, RCP^{+}13, Yam19b, ACCH23, Asp99, Ben93a, Bra13, BB02, BGM02, CK92b, CK92c, CK98, Cam13, Gio93, Hum94, Ibb99, Kos07, Min07a, Min07b, RR04, SW96, Smi07, VHCSCC15, Wil94b, Win96]. Unix [Sip13, Too10]. Unknown [Bru17]. Unlimited [Ano17z]. Unlock [Ano24-59, Ano25k]. Unraveling [Nof10]. Unsolicited [Bau06]. Update [ABL^{+}24, Spi21, Spi23a]. Upon [RP12]. Urbana [BB02]. Urbana-Champaign [BB02]. USA [Sid20]. Usability [CK07b, Kid96g]. Use [Alo04, Cor01b, EM94, Kis01, Neu04, Sch23, Ste03, VN11, Ver11, Wil02, Wu15]. Useful [Kid15, Sjö11]. Usefulness [Kid96g]. User [Bur23, SR14, Wu15, Lan00, Nor23]. user-driven [Lan00]. User-Friendliness [Nor23]. Users [Paj08]. Uses [Cor25b, Kid97b, Wil93b, WT03]. Using [Cor01b, EF11, Kas24a]. USPNet [HUA23]. UTEC [Wil94b]. Utility [Mir17]. Utopia [Sim25].

V [CH15, Est02, Lee93d, Lee95d, PM12, Dia16, Tra19]. vacuum [Log98]. Valentim [Per20]. Validity [BB02]. Valley [Ber03d, Est22, dÁ24]. Vancouver [Ges11b]. Vand [SK11]. Vandiver [Cor08a]. Vannevar [Ano00b, Kid96f, Puc96, Rio00]. Variable [Est02]. vendors [Gre95]. Ventura [Lor19]. Venture [Har13]. verb [Gri96]. verification [MP97]. Versacad [Laz24]. Verse [Sha23]. Versus [Poi25, AD14]. very [Log98]. Vezza [Ano12a]. via [Day21]. Victim [LM94]. victims [Kis97]. Video [Col14]. Videogames [Low09, MB09]. Videotex [Ste06]. Vietnam [Fei22]. View [Cer01, Phi12, Smi03, Joh98, Kli99]. Vignettes [LS08]. Vincent [Wil01b]. Vintage [Bra22, Not20]. Virtual [AG19, Ano20q, Den01, DK23, ZR06]. Virtualities [Den01]. VisiCalc [Gra07]. Vision [DK23, Gre13, Kit03, Raw24].

Visions [Abb10]. Visit [Tro93, Tro95]. Vladimir [SK11]. Voder [Sim25]. Voice [HVB04, Sid20]. Voice/Data [HVB04]. Voices [AY11]. Vol [PM12, Ano97e, RAG^{+}04]. Volume [Ano03a, Ano05, Ano06, Ano07]. Volunteer [Ano21-37, Ano21-38, Ano22-51, Ano23-41, Ano24-60, Ano25-27]. vs [Tom96].

W [Hai04c, Hai11, Lee01c, RCP^{+}13, Rus16a]. W. [BH99]. Wada [CK08]. Walden [McK22b]. Walking [VT21]. Wall [AG19]. Walther [Tod93]. Wang [Wei93d]. War [Bur24, Car92, Cor14, Day16, Lon08, Rus06, Smi01a, Don10, Fei22, Tat10, Tin10]. Ward [Ano11f]. Ware [Wal11]. Warhol [Sza18]. Wars [Big20a, Big20b, SR14]. Was [Asp00, AC16, Bis98, BW00, BS96, DA14, Fra04, LaV24, Lig03, Luk11, Roe16a, Rus13c, Tat13, Wil00b, Ano14c, Lee95e, Lee00a, McK22a]. ‘Wat [Cam13]. Watch [Ano15v, Ano22b, Ano24a, Wil03]. Waterloo [Cam13]. Watson [BJ03]. WAVES [LBA00]. Way [Aum11, Cor07, Hai10a, Kis95]. Ways [Jon03, McP95]. Wealth [Cor25b]. weapons [Pol97]. Web [Ano02e, BMFB25, Fra04]. Weber [Lee01d]. Week [Ano20-38, Ano23-42, Ano23-43, Ano22-53]. Welfare [Sch17]. Welke [Kit12c]. Welker [Van10a]. Well [Nel22, Fit02b]. well-intentioned [Fit02b]. Wenzel [Kis99a]. Were [Cor17, Got14, OK07]. Werle [Mis99]. Wesley [Lee98b]. West [Aum11, ERS03, PPM^{+}23]. Western [Hus97, Lew16]. Westinghouse [Asp93a, SL12]. What1468371Whom [CdLF^{+}22]. wheel [Kis99b]. Where [Owe96, Ano16j]. wherever [Ano14k]. Whirlwind [Fed21]. White [Roe16a]. Whiteness [Kan21]. Whitton [Shr24]. Who [BS96, Dic13, Lee96]. Wiki [Bre25]. Wilder [Léc21]. Wilkes [Lee93d]. Will [HW17]. William [Buc95b, Cly93, Eck07b, Nyi11]. Williams [Cer03, Ano20-39]. Willis [Wal11]. Win [Bro10]. Windows [DV18]. wins [Ano98b]. Winter [Shi19, Rei08]. Wipro [Day23]. Wirth [Spi24b]. Wiseman [Wei96d]. Wishes [Lee93d]. within [Kru94, Lei08]. without [CK07b, Zwe99]. Woman [Kos03b, Nel22, Vog17]. Women [Abb03, Ano22-30, BF96, Est96, Goy96, Gur96, Hic10, Rus13c, Vog17, Fri96]. Womersley [CD14, Pri24]. Word [Alo04, Ber06a, Ber06b, CG06, Ekl92, Hai06, Wil98b]. Words [Woh06]. WordStar [Rub06]. Work [Ake08, Asp01b, Chi14, CRH^{+}21, Dun21, Raw24, Ste03, SW95, Vid23, Gri98, Kis97, RAG^{+}04]. Worked [Cor17]. Workers [Kid99e, Wes17]. Working [Ano22-67, Est22, Hea03, Pri24, Roe16b, VT21]. works [Kid96b]. Workshop [BMFB25]. World [Ano97f, Ano00b, Ano15v, Ano19-37, Cor08b, Dra18, Har22b, Hos24, Kid99e, LHS^{+}25, Lew14, Lla24, Per20, Rei23a, Sim25, Wil01a, Ake00, Fie03, Mar10b, Mis07c, Car92, GH18b, Lon08, Smi01a]. World-Class [Wil01a]. Worlds [Fev13]. Worsley [Cam03b]. WP [Woh06]. Write [Ano20-51, Ano21-52, Ano22-68, Ano23-58, Ano24-61, Dic13]. Writing [Dun21, Pla16]. Wyatts [Ano22-69].

Xerox’ [Got23, Raw24, Spr18b]. xxiv [Win03].

Y2K [Mul20, Kid00l]. Yasuo [CO12]. Year [Ano97d, Ber03d, Nel22, Woi25, Mar10a, Smi99a]. Years [AL12, AG12, And22, AG21, BS11, Cer01, ES05, GM96, Gra12, Kne17, LH95, Med08, Spe98, WC12, WN05, WN06, Woo95, Yam18, dÁ24, Wei00]. York [Sim25, Smi78, Ges09b, Ges10b]. Yost [GH18b]. Young [Lee99b]. ‘Yours [Kid01d]. Youth [Gri10].

Z1 [Roj97]. Z3 [Roj97, Roj98, RDGH05]. Z80 [Lis24]. Zambia [Mac03]. Zealand [Car20, CM25, Nis23, Tol21]. Zegart [Fit23]. Zilog [Lis24]. Zurich [Spe98]. Zuse [Cer25, CRB24, Gil97, Roj97, Roj98, RDGH05, Wei96c]. Zygmont [RAG^{+}04].

References

Abbate:2001:HM

[Abb01]  Janet Abbate.  Hands on

the machine. IEEE Annals of the History of Computing, 23(3):88, 87, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3088abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3088.
pdf.

Abbate:2003:GEI

[Abb03] Janet Abbate. Guest Editor’s introduction: Women and gender in the history of computing. IEEE Annals of the History of Computing, 25(4):4–8, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL a4004.pdf;

http://csdl.computer.org/
dl/mags/an/2003/04/a4004.
htm.

Abbate:2010:PIC

[Abb10] Janet Abbate. Privatizing the Internet: Competing visions and chaotic events, 1987–1995. IEEE Annals of the History of Computing, 32(1):10–22, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ames:2024:NSM

[ABL^{+}24] Morgan Ames, Ekaterina Babintseva, Xiaochang Li, Colette Perold, and Gili Vidan. Notes from the 2024 SIGCIS meeting: System update: Patches, tactics, responses. IEEE Annals of the History of Computing, 46(3): 94–95, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aspray:2016:BIW

[AC16] William Aspray and James W. Cortada. Before it was a giant: The early history of Symantec, 1982–1999. IEEE Annals of the History of Computing, 38 (4):26–41, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Abbate:2017:R

[AC17] Janet Abbate and Elinor Carmi. Reviews. IEEE Annals of the History of Computing, 39(4):9–11, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Agosti:2023:BCU

[ACCH23] Maristella Agosti, Alberto Cammozzo, Francesco Contin, and Silvio Hénin. Building a computer at the University of Padua, 1958–1961. IEEE Annals of the History of Computing, 45(1):77–85, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Acker:2015:THD

[Ack15]  Amelia Acker.  Toward a

hermeneutics of data. IEEE Annals of the History of Computing, 37(3):70–75, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030070-abs.html.

Anderson:2007:B

[AD07] David Anderson and Janet Delve. Biographies. IEEE Annals of the History of Computing, 29(4):90–102, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alberts:2014:UVL

[AD14] G. Alberts and E. G. Daylight. Universality versus locality: The Amsterdam style of Algol implementation. IEEE Annals of the History of Computing, 36(4):52–63, October/ December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Adam:1996:CGH

[Ada96] A. Adam. Constructions of gender in the history of artificial intelligence. IEEE Annals of the History of Computing, 18(3):47–53, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11096/00511944.pdf.

Adamson:2005:B

[Ada05] Michael R. Adamson. Biographies. IEEE Annals of the History of Computing, 27(3): 98–102, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Adamson:2006:B

[Ada06] Michael R. Adamson. Biographies. IEEE Annals of the History of Computing, 28(1):99– 103, January/March 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alvarez:2012:HCC

[AG12] Juan Alvarez and Claudio Gutierrez. History of computing in Chile, 1961–1982: Early years, consolidation, and expansion. IEEE Annals of the History of Computing, 34(3): 22–33, July/September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Andoh:2019:VWF

[AG19] Chris Andoh and Andreas Godderis. The virtual wall of the Fourth Amendment.

IEEE Annals of the History of Computing, 41(1):47–50, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://ieeexplore.
ieee.org/document/8667940/

.

Ausiello:2021:TCS

[AG21] Giorgio Ausiello and Giorgio Gambosi. Theoretical computer science in Italy: The early years. IEEE Annals of the History of Computing, 43(3): 44–56, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Agar:2001:TUM

[Aga01] Jon Agar. Turing and the universal machine: the making of the modern computer. Revolutions in science. Icon, Cambridge, UK, 2001. ISBN 1-84046-250-7. iv + 153 pp. LCCN QA76.2.T87 A43 2001.

Asahi:2023:DCM

[AH23] Hiroharu Asahi and Hiromichi Hashizume. 2022 Distributed Computer Museum accreditation ceremony. IEEE Annals of the History of Computing, 45(2):70–71, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ainsworth:2023:TPD

[Ain23] Barbara Ainsworth. Trevor Pearcey and the development of CSIRAC: an Australian firstgeneration computer. IEEE Annals of the History of Computing, 45(4):53–63, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Akera:1998:BIS

[Ake98] A. Akera. Building IBM: Shaping an industry and its technologies [reviews]. IEEE Annals of the History of Computing, 20(1):81–82, January/March 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646212.pdf.

Akera:2000:ETT

[Ake00] A. Akera. ENIAC: the triumphs and tragedies of the world’s first computer [book review]. IEEE Annals of the History of Computing, 22 (4):65–66, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887993.pdf.

Akera:2004:CKI

[Ake04] Atsushi Akera. The circulation of knowledge, institutional ecologies, and the history of computing. IEEE Annals of the History of Computing, 26 (3):88, 86–87, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/03/a3088.
htm;   http://csdl.computer.
org/dl/mags/an/2004/03/a3088.
pdf.

Akera:2008:LWB

[Ake08] Atshushi Akera. The life and work of Bernard A. Galler (1928–2006). IEEE Annals of the History of Computing, 30 (1):4–14, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Antonitsch:2008:IHS

[AKM08] Peter K. Antonitsch, Robert Kleinhagauer, and Peter Micheuz. Incorporating history in secondary education informatics courses. IEEE Annals of the History of Computing, 30(3): 56–63, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ahrons:2012:IRM

[AL12] Richard W. Ahrons and David Laws. Industrial research in microcircuitry at RCA: The early years, 1953–1963. IEEE Annals of the History of Computing, 34(1):60–73, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alberts:2010:AAA

[Alb10] Gerard Alberts. Appropriating America: Americanization in the history of European computing. IEEE Annals of the History of Computing, 32(2): 4–7, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alberts:2014:ACP

[Alb14] Gerard Alberts. Algol culture and programming styles [guest editor’s introduction]. IEEE Annals of the History of Computing, 36(4):2–5, October/December 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aldrich:2002:R

[Ald02]  Peggy Aldrich.  Reviews.

IEEE Annals of the History of Computing, 24(2):94–99, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2094.pdf;   http://www.
computer.org/annals/an2002/
a2094abs.htm.

Allen:2018:CLB

[All18] Ben Allen. Common language: Business programming languages and the legibility of programming. IEEE Annals of the History of Computing, 40(2):17–31, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8415737/.

Allen:2019:ORN

[All19]  Kera Allen.  “The official

response is never enough”. IEEE Annals of the History of Computing, 41(1):34–46, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://ieeexplore.
ieee.org/document/8633419/

.

Aloisio:2004:CED

[Alo04] Mario Aloisio. The calculation of Easter Day, and the origin and use of the word computer. IEEE Annals of the History of Computing, 26(3):42–49, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/03/a3042.
htm;   http://csdl.computer.
org/dl/mags/an/2004/03/a3042.
pdf.

Aloisio:2007:CMS

[Alo07] Mario Aloisio. Computing at the Malta Statistics Office, 1947–1970. IEEE Annals of the History of Computing, 29(3): 49–61, July/September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alper:2015:AAA

[Alp15] M. Alper. Augmentative, alternative, and assistive: Reimagining the history of mobile computing and disability.

IEEE Annals of the History of Computing, 37(1):96, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anderson:2007:MNT

[And07a] David Anderson. Max Newman: Topologist, codebreaker, and pioneer of computing. IEEE Annals of the History of Computing, 29(3):76–81, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anderson:2007:PBP

[And07b]  David Anderson.  Patrick

Blackett: Physicist, radical, and chief architect of the Manchester computing phenomenon. IEEE Annals of the History of Computing, 29(3): 82–85, July/September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anderson:2009:BTK

[And09] David P. Anderson. Biographies: Tom Kilburn: a pioneer of computer design. IEEE Annals of the History of Computing, 31(2):82–86, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anderson:2022:EIC

[And22] Barbara A. Anderson. The effects of increases in computing power on demographic analysis over the last 50 years.

IEEE Annals of the History of Computing, 44(4):67–70, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Annals:2012:R

[Ann12] Annals. Reviews. IEEE Annals of the History of Computing, 34(1):99–101, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:1992:Ia

[Ano92a] Anonymous. 1991 index. IEEE Annals of the History of Computing, 14(4):78–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4078.pdf.

Anonymous:1992:Ib

[Ano92b] Anonymous. 1992 index. IEEE Annals of the History of Computing, 14(4):86–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4086.pdf.

Anonymous:1992:BPA

[Ano92c]  Anonymous.  Biographies

of participants and authors. IEEE Annals of the History of Computing, 14(2):6– ??, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a2006.pdf.

Anonymous:1992:HSC

[Ano92d] Anonymous. How to start a computer museum: Chinese calculators made during the Kangxi reign in the Quing dynasty. IEEE Annals of the History of Computing, 14(4):59–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4059.pdf.

Anonymous:1992:RB

[Ano92e] Anonymous. References and bibliography. IEEE Annals of the History of Computing, 14(1):52–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1052.pdf.

Anonymous:1993:I

[Ano93] Anonymous. 1993 index. IEEE Annals of the History of Computing, 15(4):89–??, October/ December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4089.pdf.

Anonymous:1994:I

[Ano94]  Anonymous.  1994 index.

IEEE Annals of the History of Computing, 16(4):95– ??, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4095.pdf.

Anonymous:1995:R

[Ano95] Anonymous. Resolution. IEEE Annals of the History of Computing, 17(3):10, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397058.pdf.

Anonymous:1997:A

[Ano97a]  Anonymous.  Anecdotes.

IEEE Annals of the History of Computing, 19(1): 60–64, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560740.pdf.

Anonymous:1997:BGH

[Ano97b] Anonymous. A bibliographic guide to the history of computer applications, 1950–1990 [book reviews]. IEEE Annals of the History of Computing, 19(1):80, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:1997:B

[Ano97c]  Anonymous.  Biographies.

IEEE Annals of the History of Computing, 19(1): 65–73, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560742.pdf.

Anonymous:1997:SBE

[Ano97d] Anonymous. I sing the body electronic: a year with Microsoft on the multimedia frontier [book reviews]. IEEE Annals of the History of Computing, 19(1):80, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560757.pdf.

Anonymous:1997:CBP

[Ano97e] Anonymous. The Colossus of Bletchley Park, IEEE Rev., vol. 41, no. 2, pp. 55–59 [book reviews]. IEEE Annals of the History of Computing, 19(1):80, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560758.pdf.

Anonymous:1997:WLM

[Ano97f]  Anonymous.  The world’s

largest machine: Global telecommunications and the human condition [book reviews]. IEEE Annals of the History of Computing, 19 (1):80, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560755.pdf.

Anonymous:1998:BJM

[Ano98a]  Anonymous.  Biographies,

John Mourice McClean Pinkerton. IEEE Annals of the History of Computing, 20(3): 69–71, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707576.pdf.

Anonymous:1998:LOW

[Ano98b] Anonymous. Larry Owens wins IEEE Life Member Prize for his Annals article on MIT. IEEE Annals of the History of Computing, 20 (3):76, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707578.pdf.

Anonymous:1998:OMR

[Ano98c] Anonymous. Obituary Mina S. Rees. IEEE Annals of the History of Computing, 20(1): 65–66, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646216.pdf.

Anonymous:1999:CMO

[Ano99a] Anonymous. Charles E. Molnar [obituary]. IEEE Annals of the History of Computing, 21(3):67–69, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778987.pdf.

Anonymous:1999:HCD

[Ano99b] Anonymous. History of computer developments in Romania. IEEE Annals of the History of Computing, 21(3): 58–60, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778983.pdf.

Anonymous:1999:JLO

[Ano99c]  Anonymous.  John Lions

[obituary]. IEEE Annals of the History of Computing, 21 (3):67, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778986.pdf.

Anonymous:1999:JP

[Ano99d] Anonymous. Jonathan Postel. IEEE Annals of the History of Computing, 21(3): 69–70, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778988.pdf.

Anonymous:1999:OPI

[Ano99e] Anonymous. Overcoming pilotinduced oscillations in the space shuttle. IEEE Annals of the History of Computing, 21 (4):69–70, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801535.pdf.

Anonymous:1999:RJL

[Ano99f]  Anonymous.  Robert John

Lansdown. IEEE Annals of the History of Computing, 21(3): 66–67, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778985.pdf.

Anonymous:2000:AIC

[Ano00a] Anonymous. 2000 annual index: Complete author and subject listings. IEEE Annals of the History of Computing, 22(4):80–??, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4080.pdf.

Anonymous:2000:RPA

[Ano00b] Anonymous. Reviews: Peggy Aldrich Kidwell: The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer: Endless Frontier: Vannevar Bush, Engineer of the American Century: Dorothy Hodgkin: a Life: ENIAC: The Triumphs and Tragedies of the World’s First Computer: Inventing the Internet: Meaning in Technology. IEEE Annals of the History of Computing, 22(4):62–63, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4062.pdf.

Anonymous:2001:AI

[Ano01a] Anonymous. Annual index. IEEE Annals of the History of Computing, 23(4):107– 108, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

computer.org/annals/an2001/
a4107abs.htm;   http://dlib.
computer.org/an/books/an2001/
pdf/a4107.pdf.

Anonymous:2001:CIB

[Ano01b] Anonymous. A career interview with Bernie Galler. IEEE Annals of the History of Computing, 23(1):22–33, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1022.pdf;   http://www.

computer.org/annals/an2001/
a1022abs.htm.

Anonymous:2001:ES

[Ano01c]  Anonymous.  Events and

sightings. IEEE Annals of the History of Computing, 23(3):74–77, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3074abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3074.
pdf.

Anonymous:2002:AI

[Ano02a] Anonymous. 2002 annual index. IEEE Annals of the History of Computing, 24 (4):60–61, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4060.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4060.
pdf.

Anonymous:2002:A

[Ano02b] Anonymous. Anecdotes. IEEE Annals of the History of Computing, 24(1):80–89, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1080abs.htm;
http://dlib.computer.org/
an/books/an2002/pdf/a1080.
pdf.

Anonymous:2002:ESa

[Ano02c] Anonymous. Events and sightings. IEEE Annals of the History of Computing, 24(2):100– 103, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/articles/a2-2002/
eands.htm;   http://dlib.
computer.org/an/books/an2002/
pdf/a2100.pdf;   http://
www.computer.org/annals/
an2002/a2100abs.htm.

Anonymous:2002:ESb

[Ano02d]  Anonymous.  Events and

sightings. IEEE Annals of the History of Computing, 24(3):82–84, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3082.pdf.

Anonymous:2002:ESW

[Ano02e] Anonymous. Events and sightings: Web extra: CBI conference — “Unbundling history: The emergence of the software product”. IEEE Annals of the History of Computing, 24(1):90–94, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1090abs.htm;
 http://dlib.computer.org/
an/books/an2002/pdf/a1090.
pdf.

Anonymous:2002:RLN

[Ano02f]  Anonymous.  Remembering

the LFK network. IEEE Annals of the History of Computing, 24(3):79–81, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3079.pdf.

Anonymous:2003:IAH

[Ano03a] Anonymous. IEEE Annals of the History of Computing, 2003 annual index, volume 25. IEEE Annals of the History of Computing, 25(4):97–103, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL a4097.pdf;

http://csdl.computer.org/
dl/mags/an/2003/04/a4097.
htm.

Anonymous:2003:R

[Ano03b] Anonymous. Reviews. IEEE Annals of the History of Computing, 25(1):70–71, January/March 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/01/a1070.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1070.pdf;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1070abs.htm.

Anonymous:2004:AI

[Ano04a] Anonymous. 2004 annual index. IEEE Annals of the History of Computing, 26(4):92– 100, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/04/a4092.pdf;
 http://csdl.computer.org/
dl/mags/an/2004/04/a4092.
htm.

Anonymous:2004:ES

[Ano04b]  Anonymous.  Events and

sightings. IEEE Annals of the History of Computing, 26(2):84–85, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/02/a2084.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2084.
pdf.

Anonymous:2004:ED

[Ano04c] Anonymous. From the editor’s desk. IEEE Annals of the History of Computing, 26(3):2–3, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2004/03/
a3002.pdf;   http://csdl.
computer.org/dl/mags/an/
2004/03/a3002.htm.

Anonymous:2004:R

[Ano04d] Anonymous. Reviews. IEEE Annals of the History of Computing, 26(2):90–93, April/ June 2004. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/02/a2090.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2090.
pdf.

Anonymous:2005:AIV

[Ano05] Anonymous. 2005 annual index, volume 27. IEEE Annals of the History of Computing, 27(4):97–101, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2005/04/
a4097.pdf.

Anonymous:2006:IAH

[Ano06] Anonymous. IEEE Annals of the History of Computing 2006 annual index, volume 28. IEEE Annals of the History of Computing, 28(4):104–109, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2006/04/
a4104.pdf.

Anonymous:2007:AII

[Ano07]  Anonymous.  2007 annual

index, IEEE Annals of the History of Computing, volume 29. IEEE Annals of the History of Computing, 29(4):103–109, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2007/04/
                                        man2007040103.pdf.

Anonymous:2008:Ra

[Ano08a] Anonymous. Reviews. IEEE Annals of the History of Computing, 30(3):82–83, July/ September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2008:Rb

[Ano08b] Anonymous. Reviews. IEEE Annals of the History of Computing, 30(4):104–105, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:AIa

[Ano09a] Anonymous. Annual index. IEEE Annals of the History of Computing, 31(1):0, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:AIb

[Ano09b] Anonymous. Annual index. IEEE Annals of the History of Computing, 31(4):0, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:EB

[Ano09c] Anonymous. Editorial board. IEEE Annals of the History of Computing, 31(1):3, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:EEB

[Ano09d] Anonymous. Editorial: Editorial board. IEEE Annals of the History of Computing, 31(3): 3, July/September 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:GEI

[Ano09e] Anonymous. Guest Editor’s introduction: Perspectives on the history of computer games. IEEE Annals of the History of Computing, 31(3):4, July/ September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:Ra

[Ano09f] Anonymous. Reviews. IEEE Annals of the History of Computing, 31(3):74–76, July/ September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2009:Rb

[Ano09g] Anonymous. Reviews. IEEE Annals of the History of Computing, 31(4):116–117, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2010:A

[Ano10a] Anonymous. Anecdotes. IEEE Annals of the History of Computing, 32(2):80–83, April/ June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2010:R

[Ano10b] Anonymous. Reviews. IEEE Annals of the History of Computing, 32(2):70–71, April/ June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:AI

[Ano11a] Anonymous. Annual index. IEEE Annals of the History of Computing, 33(4):0, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:Ra

[Ano11b] Anonymous. Reviews. IEEE Annals of the History of Computing, 33(1):83–85, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:Rb

[Ano11c] Anonymous. Reviews. IEEE Annals of the History of Computing, 33(2):102–104, April/ June 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:Rc

[Ano11d] Anonymous. Reviews. IEEE Annals of the History of Computing, 33(3):80–81, July/ September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:Rd

[Ano11e] Anonymous. Reviews. IEEE Annals of the History of Computing, 33(4):81–83, October/ December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2011:WC

[Ano11f] Anonymous. Ward Cunningham. IEEE Annals of the History of Computing, 33(4):62– 67, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2012:AVI

[Ano12a] Anonymous. Albert Vezza [interview]. IEEE Annals of the History of Computing, 34(2): 80–85, April/June 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2012:IFC

[Ano12b] Anonymous. Interview: Fernando Corbató. IEEE Annals of the History of Computing, 34(1):83–87, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:AAS

[Ano13a] Anonymous. 2013 annual author and subject index. IEEE Annals of the History of Computing, 35(4):90–93, October/ December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:CNH

[Ano13b] Anonymous. Computing Now house advertisement. IEEE Annals of the History of Computing, 35(4):c3, October/ December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:DMH

[Ano13c]  Anonymous.  Digital mag-

azines [house advertisement]. IEEE Annals of the History of Computing, 35(4):c4, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:FC

[Ano13d]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 35(4):c1, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:M

[Ano13e]  Anonymous.  [Masthead].

IEEE Annals of the History of Computing, 35(4):4, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:MAb

[Ano13f] Anonymous. Membership [advertisement]. IEEE Annals of the History of Computing, 35(4):55, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:MAa

[Ano13g] Anonymous. myComputer [advertisement]. IEEE Annals of the History of Computing, 35(4):31, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:RS

[Ano13h] Anonymous. Rolf Skår. IEEE Annals of the History of Computing, 35(1):72–77, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2013:TC

[Ano13i] Anonymous. Table of contents. IEEE Annals of the History of Computing, 35(4):c2, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:FCa

[Ano14a] Anonymous. [Front cover].

IEEE Annals of the History of Computing, 36(2):c1, April/ June 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:FCb

[Ano14b]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 36(3):c1, July/ September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:IWH

[Ano14c] Anonymous. IEEE was here [advertisement]. IEEE Annals of the History of Computing, 36(4):c3, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:MMA

[Ano14d] Anonymous. Membership matters [advertisement]. IEEE Annals of the History of Computing, 36(1):c3, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:RBR

[Ano14e] Anonymous. Reviews [3 books reviewed]. IEEE Annals of the History of Computing, 36 (4):87–91, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:RSC

[Ano14f] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Annals of the History of Computing, 36(3):c3, July/ September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:RSM

[Ano14g] Anonymous. Rock stars of mobile cloud [advertisement]. IEEE Annals of the History of Computing, 36(1):c4, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:SEC

[Ano14h] Anonymous. Seeking Editor in Chief. IEEE Annals of the History of Computing, 36(3): c4, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:TCa

[Ano14i] Anonymous. Table of contents. IEEE Annals of the History of Computing, 36(1): c2, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:TCb

[Ano14j] Anonymous. Table of contents. IEEE Annals of the History of Computing, 36(3): c2, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2014:TCL

[Ano14k] Anonymous. Take the CS Library wherever you go! [advertisement]. IEEE Annals of the History of Computing, 36(4): c4, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:RMD

[Ano15a] Anonymous. 2016 Richard E. Merwin Distinguished Service Award house advertisement. IEEE Annals of the History of Computing, 37(3):55, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030055.pdf.

Anonymous:2015:CPHa

[Ano15b] Anonymous. Call for papers house advertisement. IEEE Annals of the History of Computing, 37(1):c3, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:CPHb

[Ano15c] Anonymous. Call for papers house advertisement. IEEE Annals of the History of Computing, 37(4):c3, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:EB

[Ano15d] Anonymous. Editorial Board. IEEE Annals of the History of Computing, 37(2):5, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020005.pdf.

Anonymous:2015:EBHa

[Ano15e] Anonymous. Editorial board house advertisement. IEEE Annals of the History of Computing, 37(1):3, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:EBHb

[Ano15f] Anonymous. Editorial board house advertisement. IEEE Annals of the History of Computing, 37(4):13, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:FYJ

[Ano15g] Anonymous. Focus on your job search house advertisement. IEEE Annals of the History of Computing, 37(2): c4, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man20150200c4.pdf.

Anonymous:2015:FCa

[Ano15h]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 37(1):c1, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:FCb

[Ano15i]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 37(2):c1, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man20150200c1.pdf.

Anonymous:2015:FCc

[Ano15j]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 37(3):c1, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man20150300c1.pdf.

Anonymous:2015:FCd

[Ano15k] Anonymous. Front cover. IEEE Annals of the History of Computing, 37(4):c1, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:GRY

[Ano15l] Anonymous. Get the recognition you deserve house advertisement. IEEE Annals of the History of Computing, 37 (4):97, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:ICS

[Ano15m] Anonymous. IEEE Computer Society house advertisement. IEEE Annals of the History of Computing, 37(3):67, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030067.pdf.

Anonymous:2015:RSCa

[Ano15n] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Annals of the History of Computing, 37(2): c3, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man20150200c3.pdf.

Anonymous:2015:RSCb

[Ano15o] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Annals of the History of Computing, 37(3):c3, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man20150300c3.pdf.

Anonymous:2015:SRS

[Ano15p]  Anonymous.  Startup rock

stars house advertisement. IEEE Annals of the History of Computing, 37(1):c4, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:SCH

[Ano15q] Anonymous. Stay connected house advertisement. IEEE Annals of the History of Computing, 37(3):c4, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/an/2015/03/
                                        man20150300c4.pdf.

Anonymous:2015:TCa

[Ano15r] Anonymous. Table of contents. IEEE Annals of the History of Computing, 37(1): c2, January/March 2015. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:TCb

[Ano15s] Anonymous. Table of contents. IEEE Annals of the History of Computing, 37(2): c2, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/an/2015/02/
                                        man20150200c2.pdf.

Anonymous:2015:TCc

[Ano15t] Anonymous. Table of contents. IEEE Annals of the History of Computing, 37(3):c2, c1, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/an/2015/03/
                                        man20150300c2.pdf.

Anonymous:2015:TCd

[Ano15u] Anonymous. Table of contents. IEEE Annals of the History of Computing, 37(4):c2–1, ???? 2015. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2015:WWL

[Ano15v] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. IEEE Annals of the History of Computing, 37(4):c4, ???? 2015. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:CNE

[Ano16a] Anonymous. Caller for nominations education awards nominations house advertisement. IEEE Annals of the History of Computing, 38(3):c3, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:FYJ

[Ano16b] Anonymous. Focus on your job search house advertisement. IEEE Annals of the History of Computing, 38(3):c4, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:FCa

[Ano16c] Anonymous. Front cover. IEEE Annals of the History of Computing, 38(2):c1, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:FCb

[Ano16d] Anonymous. Front cover. IEEE Annals of the History of Computing, 38(3):c1, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:FCc

[Ano16e] Anonymous. Front cover. IEEE Annals of the History of Computing, 38(4):c1, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:GF

[Ano16f] Anonymous. got flaws? IEEE Annals of the History of Computing, 38(4):c4, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:ICSc

[Ano16g] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 38(4):9, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:ICSa

[Ano16h] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. IEEE Annals of the History of Computing, 38 (2):c3, April/June 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:ICSb

[Ano16i] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Annals of the History of Computing, 38(3):c2, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:ICSd

[Ano16j] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. IEEE Annals of the History of Computing, 38(4):c3, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:Ma

[Ano16k] Anonymous. Masthead. IEEE Annals of the History of Computing, 38(2):5, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:Mb

[Ano16l] Anonymous. Masthead. IEEE Annals of the History of Computing, 38(3):3, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:RSB

[Ano16m] Anonymous. Rock stars of big data house advertisement. IEEE Annals of the History of Computing, 38(2):c4, April/ June 2016. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:TCa

[Ano16n] Anonymous. Table of contents. IEEE Annals of the History of Computing, 38(2):c2–1, April/ June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:TCb

[Ano16o] Anonymous. Table of contents. IEEE Annals of the History of Computing, 38(3):1–2, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:TCc

[Ano16p] Anonymous. Table of contents. IEEE Annals of the History of Computing, 38(4):1–2, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2016:T

[Ano16q] Anonymous. TechIgnite. IEEE Annals of the History of Computing, 38(4):c2, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2017:ICC

[Ano17a]  Anonymous.  2017 IEEE-

CS Charles Babbage Award. IEEE Annals of the History of Computing, 39(3):8, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030008.pdf.

Anonymous:2017:AYCa

[Ano17b] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Annals of the History of Computing, 39(2):??, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:AYCb

[Ano17c] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Annals of the History of Computing, 39(3):72, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030072.pdf.

Anonymous:2017:CNE

[Ano17d] Anonymous. Call for nominees education awards nominations. IEEE Annals of the History of Computing, 39(3):3, July/September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030003.pdf.

Anonymous:2017:FYJb

[Ano17e] Anonymous. Focus on your job search. IEEE Annals of the History of Computing, 39 (2):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:FYJa

[Ano17f] Anonymous. Focus on your job search house advertisement. IEEE Annals of the History of Computing, 39(1):c4, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/01/
                                        man20170100c4.pdf.

Anonymous:2017:FCa

[Ano17g]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 39(1):c1, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/01/
                                        man20170100c1.pdf.

Anonymous:2017:FCb

[Ano17h] Anonymous. Front cover. IEEE Annals of the History of Computing, 39(2):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:FCc

[Ano17i]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 39(3):c1, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man20170300c1.pdf.

Anonymous:2017:FCd

[Ano17j] Anonymous. Front cover. IEEE Annals of the History of Computing, 39(4):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:ICSd

[Ano17k] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Annals of the History of Computing, 39(4):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:ICSb

[Ano17l] Anonymous. IEEE Computer Society Harlan D. Mills Award. IEEE Annals of the History of Computing, 39(3):68, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030068.pdf.

Anonymous:2017:ICSc

[Ano17m] Anonymous. IEEE Computer Society house advertisement. IEEE Annals of the History of Computing, 39(3):c4, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.

org/csdl/mags/an/2017/03/
man20170300c4.pdf.

Anonymous:2017:ICSa

[Ano17n] Anonymous. IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. IEEE Annals of the History of Computing, 39(1):c3, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man20170100c3.pdf.

Anonymous:2017:LBTa

[Ano17o]  Anonymous.  Looking for

the BEST tech job for you? IEEE Annals of the History of Computing, 39(3):c2, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man20170300c2.pdf.

Anonymous:2017:LBTb

[Ano17p] Anonymous. Looking for the BEST tech job for you? IEEE Annals of the History of Computing, 39(4):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:Ma

[Ano17q] Anonymous. Masthead. IEEE Annals of the History of Computing, 39(1):3, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010003.pdf.

Anonymous:2017:Mb

[Ano17r] Anonymous. Masthead. IEEE Annals of the History of Computing, 39(2):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:Mc

[Ano17s] Anonymous. myCS. IEEE Annals of the History of Computing, 39(3):24, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030024.pdf.

Anonymous:2017:Md

[Ano17t] Anonymous. myCS. IEEE Annals of the History of Computing, 39(4):??, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:MHA

[Ano17u] Anonymous. myCS house advertisement. IEEE Annals of the History of Computing, 39 (2):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:NMOc

[Ano17v] Anonymous. New membership options for a better fit.

IEEE Annals of the History of Computing, 39(3):c3, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man20170300c3.pdf.

Anonymous:2017:NMOa

[Ano17w]  Anonymous.  New mem-

bership options for a better fit. House advertisement. IEEE Annals of the History of Computing, 39(1):c2, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man20170100c2.pdf.

Anonymous:2017:NMOb

[Ano17x] Anonymous. New membership options for a better fit. House advertisement. IEEE Annals of the History of Computing, 39(2):??, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:NSS

[Ano17y]  Anonymous.  Nominations

are solicited for the Seymour Cray, Sidney Fernbach, & Ken Kennedy Awards house advertisement. IEEE Annals of the History of Computing, 39 (2):??, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:OMU

[Ano17z] Anonymous. One membership.

Unlimited knowledge. House advertisement. IEEE Annals of the History of Computing, 39(2):??, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:TCa

[Ano17-27] Anonymous. Table of contents.

IEEE Annals of the History of Computing, 39(1):1–2, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010001.pdf.

Anonymous:2017:TCb

[Ano17-28] Anonymous. Table of contents. IEEE Annals of the History of Computing, 39(2):1–2, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2017:TCc

[Ano17-29] Anonymous. Table of contents. IEEE Annals of the History of Computing, 39(3):1–2, July/September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030001.pdf.

Anonymous:2017:TCd

[Ano17-30] Anonymous. Table of contents. IEEE Annals of the History of Computing, 39(4):1–2, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:2018:AI

[Ano18a] Anonymous. Author information. IEEE Annals of the History of Computing, 40(4):C2, October/December 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:CP

[Ano18b]  Anonymous.  Call for pa-

pers. IEEE Annals of the History of Computing, 40(3): c2, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:CA

[Ano18c] Anonymous. Compsac ad. IEEE Annals of the History of Computing, 40(3):37, July/ September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:CIH

[Ano18d] Anonymous. CS info house ad. IEEE Annals of the History of Computing, 40(3):94, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://ieeexplore.
                                        ieee.org/document/8509564/

.

Anonymous:2018:CI

[Ano18e]  Anonymous.  CS informa-

tion. IEEE Annals of the History of Computing, 40(2):c3, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8415732/.

Anonymous:2018:FCa

[Ano18f]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 40(1):c1, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man20180100c1.pdf.

Anonymous:2018:FCb

[Ano18g] Anonymous. Front cover. IEEE Annals of the History of Computing, 40(2):c1, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:FCc

[Ano18h]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 40(3):c1, July/ September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:FC

[Ano18i]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 40(4):C1, October/December 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:ICSb

[Ano18j] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 40(4): C3, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:ICSa

[Ano18k]  Anonymous.  IEEE Com-

puter Society information. IEEE Annals of the History of Computing, 40(1):18, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010018.pdf.

Anonymous:2018:ILC

[Ano18l] Anonymous. IEEE letters of the Computer Society. IEEE Annals of the History of Computing, 40(4):C4, October/ December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:ITEa

[Ano18m] Anonymous. IEEE Technology and Engineering Management Society TEMS Administrative Committee. IEEE Annals of the History of Computing, 40 (4):3, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:ITEb

[Ano18n] Anonymous. IEEE Technology and Engineering Management Society TEMS Administrative Committee. IEEE Annals of the History of Computing, 40 (4):4, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:LHA

[Ano18o]  Anonymous.  LOCS house

ad. IEEE Annals of the History of Computing, 40(3):c3, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:LA

[Ano18p] Anonymous. LOSC ad. IEEE Annals of the History of Computing, 40(2):c4, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:Ma

[Ano18q] Anonymous. Masthead. IEEE Annals of the History of Computing, 40(1):5, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010005.pdf.

Anonymous:2018:Mb

[Ano18r] Anonymous. Masthead. IEEE Annals of the History of Computing, 40(2):5, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8415736/.

Anonymous:2018:Mc

[Ano18s] Anonymous. Masthead. IEEE Annals of the History of Computing, 40(3):7, July/ September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://ieeexplore.
                                        ieee.org/document/8509539/

.

Anonymous:2018:MHA

[Ano18t]  Anonymous.  Membership

house ad. IEEE Annals of the History of Computing, 40(3): c4, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:SNH

[Ano18u] Anonymous. Social networking house ad. IEEE Annals of the History of Computing, 40 (3):104, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:SMA

[Ano18v] Anonymous. Student membership ad. IEEE Annals of the History of Computing, 40 (2):c2, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:TCa

[Ano18w] Anonymous. Table of contents.

IEEE Annals of the History of Computing, 40(1):1–2, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010001.pdf.

Anonymous:2018:TCb

[Ano18x] Anonymous. Table of contents. IEEE Annals of the History of Computing, 40(2):1–2, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:TCc

[Ano18y] Anonymous. Table of contents. IEEE Annals of the History of Computing, 40(3):1– 2, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:TC

[Ano18z] Anonymous. Table of contents. IEEE Annals of the History of Computing, 40(4):1–2, October/December 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2018:THA

[Ano18-27] Anonymous. TSUSC house ad. IEEE Annals of the History of Computing, 40(3):77, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:AIC

[Ano19a]  Anonymous.  ACM/IEEE

CS Eckert–Mauchly Award. IEEE Annals of the History of Computing, 41(1):C2, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:AGH

[Ano19b] Anonymous. Annals general half. IEEE Annals of the History of Computing, 41(2):63, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CEa

[Ano19c] Anonymous. Computing Edge. IEEE Annals of the History of Computing, 41(2):C4, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CEb

[Ano19d] Anonymous. Computing Edge. IEEE Annals of the History of Computing, 41(3):C4, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CEc

[Ano19e] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 41(4):C4, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:IAH

[Ano19f] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 41(1):3, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://ieeexplore.
ieee.org/document/8667928/

.

Anonymous:2019:ICGa

[Ano19g] Anonymous. IEEE Computer Graphics. IEEE Annals of the History of Computing, 41(1): 57, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICGb

[Ano19h]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 41(3):88, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:IIC

[Ano19i] Anonymous. IEEE Internet Computing. IEEE Annals of the History of Computing, 41(4):124, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ISP

[Ano19j] Anonymous. IEEE Security & Privacy. IEEE Annals of the History of Computing, 41(1): 46, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ITBa

[Ano19k] Anonymous. IEEE Transactions on Big Data. IEEE Annals of the History of Computing, 41(1):6, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ITBb

[Ano19l] Anonymous. IEEE Transactions on Big Data. IEEE Annals of the History of Computing, 41(1):54, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ITSa

[Ano19m] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Annals of the History of Computing, 41(1):50, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ITSb

[Ano19n] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Annals of the History of Computing, 41(2):C3, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CMAa

[Ano19o] Anonymous. Call for 2019 major awards nominations. IEEE Annals of the History of Computing, 41(3):C2, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CMAb

[Ano19p] Anonymous. Call for 2019 major awards nominations. IEEE Annals of the History of Computing, 41(4):C2, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CGH

[Ano19q] Anonymous. CGA general half.

IEEE Annals of the History of Computing, 41(2):47, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CI

[Ano19r] Anonymous. Connect on interface. IEEE Annals of the History of Computing, 41(1): 67, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:CJB

[Ano19s] Anonymous. CS jobs board. IEEE Annals of the History of Computing, 41(2):71, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:FCa

[Ano19t]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 41(1):C1, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:FCb

[Ano19u] Anonymous. Front cover. IEEE Annals of the History of Computing, 41(2):C1, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:FCc

[Ano19v] Anonymous. Front cover. IEEE Annals of the History of Computing, 41(3):C1, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:FCd

[Ano19w]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 41(4):C1, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSa

[Ano19x] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(1): 4, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSb

[Ano19y] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(1): C3, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSc

[Ano19z] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(2):4, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSd

[Ano19-27] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(3):4, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSf

[Ano19-28] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4):4, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSg

[Ano19-29] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4): 60, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSh

[Ano19-30] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4): 70, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSi

[Ano19-31] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4): 96, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSj

[Ano19-32] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4): 118, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSk

[Ano19-33] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 41(4): 123, October/December 2019.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSe

[Ano19-34] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 41(3):C3, July 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ICSl

[Ano19-35] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 41(4):C3, October/ December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:ILC

[Ano19-36] Anonymous. IEEE letters of the Computer Society. IEEE Annals of the History of Computing, 41(1):68, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:IWC

[Ano19-37] Anonymous. IEEE World Congress. IEEE Annals of the History of Computing, 41 (2):C2, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:GQ

[Ano19-38] Anonymous. IS general quarter. IEEE Annals of the History of Computing, 41(2):66, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:IGH

[Ano19-39] Anonymous. IT general half. IEEE Annals of the History of Computing, 41(2):41, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:IP

[Ano19-40] Anonymous. IT professional. IEEE Annals of the History of Computing, 41(1):63, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:Ma

[Ano19-41] Anonymous. Masthead. IEEE Annals of the History of Computing, 41(2):3, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:Mb

[Ano19-42] Anonymous. Masthead. IEEE Annals of the History of Computing, 41(3):3, July 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:M

[Ano19-43] Anonymous. Masthead. IEEE Annals of the History of Computing, 41(4):3, October/ December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:PG

[Ano19-44] Anonymous. PVC general.

IEEE Annals of the History of Computing, 41(2):28, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:SP

[Ano19-45] Anonymous. Security & privacy. IEEE Annals of the History of Computing, 41(1): 66, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:S

[Ano19-46] Anonymous. SP Subscribe hhalf.

IEEE Annals of the History of Computing, 41(2):21, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:SC

[Ano19-47] Anonymous. Stay connected.

IEEE Annals of the History of Computing, 41(1):C4, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:SGQ

[Ano19-48] Anonymous. SW general quarter. IEEE Annals of the History of Computing, 41(2):66, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:TCa

[Ano19-49] Anonymous. Table of contents. IEEE Annals of the History of Computing, 41(1):1– 2, January/March 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:TCb

[Ano19-50] Anonymous. Table of contents. IEEE Annals of the History of Computing, 41(2):1–2, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:TCc

[Ano19-51] Anonymous. Table of contents. IEEE Annals of the History of Computing, 41(3):1–2, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:TCd

[Ano19-52] Anonymous. Table of contents. IEEE Annals of the History of Computing, 41(4):1–2, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2019:TS

[Ano19-53] Anonymous. TBD submit. IEEE Annals of the History of Computing, 41(2):72, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CEb

[Ano20a] Anonymous. Computer Edge.

IEEE Annals of the History of Computing, 42(3):C4, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CGA

[Ano20b] Anonymous. Computer Graphics and Applications. IEEE Annals of the History of Computing, 42(3):10, July/September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Cb

[Ano20c] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 42(1):75, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CEa

[Ano20d] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 42(2):C3, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Cc

[Ano20e] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 42(4):C4, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:IAH

[Ano20f] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 42(1):80, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:IAT

[Ano20g] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 42(2):54, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICAa

[Ano20h]  Anonymous.  IEEE Com-

puter Architecture Letters. IEEE Annals of the History of Computing, 42(1):C3, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICAb

[Ano20i]  Anonymous.  IEEE Com-

puter Architecture: Letters. IEEE Annals of the History of Computing, 42(3):99, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICG

[Ano20j]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 42(4):126, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ITB

[Ano20k] Anonymous. IEEE Transactions on Big Data. IEEE Annals of the History of Computing, 42(4):127, October/ December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CAb

[Ano20l] Anonymous. Call for articles. IEEE Annals of the History of Computing, 42(3):91, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CPIb

[Ano20m] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Annals of the History of Computing, 42(4): 98, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CPIa

[Ano20n]  Anonymous.  Call for pa-

pers: IEEE 21st International Conference on Information Reuse and Integration for Data Science (IEEE IRI 2020). IEEE Annals of the History of Computing, 42(1):78, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CSI

[Ano20o] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 42(4):99, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Caa

[Ano20p] Anonymous. COMPSAC 2020. IEEE Annals of the History of Computing, 42(1):C2, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:CV

[Ano20q] Anonymous. CVPR: Virtual. IEEE Annals of the History of Computing, 42(4):C2, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:EANa

[Ano20r] Anonymous. Education awards nominations. IEEE Annals of the History of Computing, 42 (1):C4, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:EANb

[Ano20s] Anonymous. Education awards nominations. IEEE Annals of the History of Computing, 42 (3):100, July/September 2020.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ECO

[Ano20t]  Anonymous.  Evolving ca-

reer opportunities need your skills. IEEE Annals of the History of Computing, 42(3): 4, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:FCa

[Ano20u]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 42(1):C1, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:FCb

[Ano20v] Anonymous. Front cover. IEEE Annals of the History of Computing, 42(2):C1, April/June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:FCc

[Ano20w]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 42(3):C1, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:FCd

[Ano20x]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 42(4):C1, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:GPNa

[Ano20y] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 42(3):C3, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:GPNb

[Ano20z] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 42(4):C3, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:HES

[Ano20-27] Anonymous. History of the events and sightings and related Annals departments. IEEE Annals of the History of Computing, 42(3):92–98, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:H

[Ano20-28] Anonymous. HOST 2020. IEEE Annals of the History of Computing, 42(2):C4, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSa

[Ano20-29] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 42(1): 4, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSc

[Ano20-30] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 42(2): 4, April/June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSg

[Ano20-31] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 42(4):4, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSd

[Ano20-32] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 42(2):41, April/June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSf

[Ano20-33] Anonymous. IEEE Computer Society call for papers. IEEE Annals of the History of Computing, 42(3):25, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSh

[Ano20-34] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 42(4):23, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSb

[Ano20-35] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 42(2):C2, April/June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:ICSe

[Ano20-36] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 42(3): C2, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:IIS

[Ano20-37] Anonymous. IEEE International Symposium on Hardware-Oriented Security and Trust. IEEE Annals of the History of Computing, 42(1): 77, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:IQW

[Ano20-38] Anonymous. IEEE Quantum Week 2020 is open for submissions. IEEE Annals of the History of Computing, 42(1): 79, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:IMR

[Ano20-39] Anonymous. Interview with Michael R. Williams. IEEE Annals of the History of Computing, 42(2):72–86, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:KYC

[Ano20-40] Anonymous. Keep your career options open. IEEE Annals of the History of Computing, 42(1):76, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Ma

[Ano20-41] Anonymous. Masthead. IEEE Annals of the History of Computing, 42(1):3, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Mb

[Ano20-42] Anonymous. Masthead. IEEE Annals of the History of Computing, 42(2):3, April/June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Mc

[Ano20-43] Anonymous. Masthead. IEEE Annals of the History of Computing, 42(3):3, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:Md

[Ano20-44] Anonymous. Masthead. IEEE Annals of the History of Computing, 42(4):3, October/ December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:NE

[Ano20-45] Anonymous. Notes from the Editor. IEEE Annals of the History of Computing, 42(4): 124–126, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:TCa

[Ano20-46] Anonymous. Table of contents. IEEE Annals of the History of Computing, 42(1):1– 2, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:TCb

[Ano20-47] Anonymous. Table of contents. IEEE Annals of the History of Computing, 42(2):1–2, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:TCc

[Ano20-48] Anonymous. Table of contents. IEEE Annals of the History of Computing, 42(3):1– 2, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:TCd

[Ano20-49] Anonymous. Table of contents. IEEE Annals of the History of Computing, 42(4):1–2, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:TTT

[Ano20-50] Anonymous. Top technology trends for 2020 featured in Computer. IEEE Annals of the History of Computing, 42(1): 80, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2020:WA

[Ano20-51] Anonymous. Write for Annals. IEEE Annals of the History of Computing, 42(4):128, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICC

[Ano21a]  Anonymous.  2022 IEEE

CS: Charles Babbage Award: Call for award nominations. IEEE Annals of the History of Computing, 43(3):5, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:AIC

[Ano21b] Anonymous. ACM/IEEE CS Eckert-Mauchly Award. IEEE Annals of the History of Computing, 43(2):41, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CSEa

[Ano21c] Anonymous. Computing in Science & Engineering. IEEE Annals of the History of Computing, 43(1):99, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CSEb

[Ano21d] Anonymous. Computing in Science & Engineering. IEEE Annals of the History of Computing, 43(2):29, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Cb

[Ano21e] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 43(1):C2, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Cd

[Ano21f] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 43(3):C2, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Ce

[Ano21g] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 43(4):C2, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:IAH

[Ano21h] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 43(3):88, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ISPa

[Ano21i] Anonymous. IEEE Security & Privacy. IEEE Annals of the History of Computing, 43(1): 100, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ISPb

[Ano21j] Anonymous. IEEE Security & Privacy. IEEE Annals of the History of Computing, 43 (2):47, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:IPC

[Ano21k] Anonymous. IT Professional: Call for articles. IEEE Annals of the History of Computing, 43(4):56, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CAa

[Ano21l] Anonymous. Call for articles. IEEE Annals of the History of Computing, 43(1): 5, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CAb

[Ano21m] Anonymous. Call for articles. IEEE Annals of the History of Computing, 43(3):21, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CSI

[Ano21n] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 43(4):112, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Cc

[Ano21o] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 43(2):C2, April/ June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:FCa

[Ano21p]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 43(1):C1, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:FCb

[Ano21q] Anonymous. [Front cover].

IEEE Annals of the History of Computing, 43(2):C1, April/ June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:FCc

[Ano21r]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 43(3):C1, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:FCd

[Ano21s]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 43(4):C1, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:GPIa

[Ano21t] Anonymous. Get published in the IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 43(2):C4, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:GPIb

[Ano21u] Anonymous. Get published in the IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 43(3):C4, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:GPNa

[Ano21v] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 43(1):C3, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:GPNb

[Ano21w] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 43(4):C3, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSa

[Ano21x] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 43(1): 4, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSd

[Ano21y] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 43(2): 4, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSi

[Ano21z] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 43(3): 4, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSn

[Ano21-27] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 43(4):4, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSb

[Ano21-28] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 43(1): 96, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSe

[Ano21-29] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 43(2):17, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSj

[Ano21-30] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 43(3): 67, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSc

[Ano21-31] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 43(1): C4, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSh

[Ano21-32] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 43(2):C3, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSm

[Ano21-33] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 43(3): C3, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSf nals of the History of Computing, 43(2):50, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSl

[Ano21-35] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 43(3):87, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSo

[Ano21-36] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 43(4):C4, October/ December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSg

[Ano21-37] Anonymous. IEEE Computer Society volunteer service awards. IEEE Annals of the History of Computing, 43(2): 51, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:ICSk

[Ano21-38] Anonymous. IEEE Computer Society: Volunteer service awards. IEEE Annals of the History of Computing, 43 (3):86, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Ma

[Ano21-39] Anonymous. Masthead. IEEE Annals of the History of Computing, 43(1):3, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Mb

[Ano21-40] Anonymous. [Masthead]. IEEE Annals of the History of Computing, 43(2):3, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Mc

[Ano21-41] Anonymous. Masthead. IEEE Annals of the History of Computing, 43(3):3, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:Md

[Ano21-42] Anonymous. Masthead. IEEE Annals of the History of Computing, 43(4):3, October/ December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:NE

[Ano21-43] Anonymous. Notes from the Editor. IEEE Annals of the History of Computing, 43(2): 48–49, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:RMS

[Ano21-44] Anonymous. Richard E.

Merwin student scholarship. IEEE Annals of the History of Computing, 43(3):80, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:SMY

[Ano21-45] Anonymous. Share and manage your research data. IEEE Annals of the History of Computing, 43(3):85, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:SCEa

[Ano21-46] Anonymous. Stay on the cutting edge of artificial intelligence. IEEE Annals of the History of Computing, 43(1): 94, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:SCEb

[Ano21-47] Anonymous. Stay on the cutting edge of artificial intelligence. IEEE Annals of the History of Computing, 43(2): 49, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:TCa

[Ano21-48] Anonymous. Table of contents. IEEE Annals of the History of Computing, 43(1):1– 2, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:TCb

[Ano21-49] Anonymous. Table of contents. IEEE Annals of the History of Computing, 43(2):1–2, April/ June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:TCc

[Ano21-50] Anonymous. Table of contents. IEEE Annals of the History of Computing, 43(3):1– 2, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:TCd

[Ano21-51] Anonymous. Table of contents. IEEE Annals of the History of Computing, 43(4):1–2, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:WA

[Ano21-52] Anonymous. Write for Annals. IEEE Annals of the History of Computing, 43(2):52, April/ June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:A

[Ano22a] Anonymous. Advert. IEEE Annals of the History of Computing, 44(1):136, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:AWC

[Ano22b] Anonymous. AI’s 10 to watch:

Call for nominations. IEEE Annals of the History of Computing, 44(2):80, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CSA

[Ano22c] Anonymous. Computing in Science & Engineering. IEEE Annals of the History of Computing, 44(1):72, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Ca

[Ano22d] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 44(1):C2, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Cc

[Ano22e] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 44(2):C2, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Cd

[Ano22f] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 44(3):95, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:IAHa

[Ano22g] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 44(1):138, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:IAHb

[Ano22h] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 44(2):83, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:IAHc

[Ano22i] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 44(4):90, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICGa

[Ano22j]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 44(1): 122, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CAb

[Ano22k] Anonymous. Call for articles.

IEEE Annals of the History of Computing, 44(2):7, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CAc

[Ano22l] Anonymous. Call for articles. IEEE Annals of the History of Computing, 44(4):5, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CAI

[Ano22m] Anonymous. Call for articles for IEEE Software. IEEE Annals of the History of Computing, 44(3):94, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CPIa

[Ano22n]  Anonymous.  Call for pa-

pers: IEEE Computer Society. IEEE Annals of the History of Computing, 44(3):69, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CPIb

[Ano22o]  Anonymous.  Call for pa-

pers: IEEE Computer Society. IEEE Annals of the History of Computing, 44(4):56, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CPIc

[Ano22p] Anonymous. Call for papers: IEEE Conference on Artificial Intelligence. IEEE Annals of the History of Computing, 44(4):87, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CSIa

[Ano22q] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 44(1):87, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CSIb

[Ano22r] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 44 (2):81, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CSId

[Ano22s] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 44(4):89, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CSIc

[Ano22t] Anonymous. Call for special issue proposals for Computer. IEEE Annals of the History of Computing, 44(3): 99, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:CAa

[Ano22u] Anonymous. Calls for articles. IEEE Annals of the History of Computing, 44(1): 7, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:FCa

[Ano22v]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 44(1):C1, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:FCb

[Ano22w] Anonymous. Front cover. IEEE Annals of the History of Computing, 44(2):C1, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:FCc

[Ano22x]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 44(3):C1, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:FCd

[Ano22y]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 44(4):C1, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:GPNa

[Ano22z] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 44(1):C3, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:GPNb

[Ano22-27] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 44(2):C3, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:GPNc

[Ano22-28] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 44(3):C3, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:GPNd

[Ano22-29] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 44(4):C3, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:HWC

[Ano22-30] Anonymous. Honoring women in computing: Call for nominations. IEEE Annals of the History of Computing, 44(3):100, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICGb

[Ano22-31] Anonymous. IEEE Computer Graphics and Applications. IEEE Annals of the History of Computing, 44(4):83, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSa

[Ano22-32] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 44(1): 4, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSf

[Ano22-33] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 44(2): 4, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSp

[Ano22-34] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 44(4):4, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSb

[Ano22-35] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 44(1): 108, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSg

[Ano22-36] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 44(2):43, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSc

[Ano22-37] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 44(1):137, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSh

[Ano22-38] Anonymous. IEEE Computer SOCIETY D&I Fund. IEEE Annals of the History of Computing, 44(2):82, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSl

[Ano22-39] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 44(3):97, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSs

[Ano22-40] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 44(4):92, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSd

[Ano22-41] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 44(1): 140, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSi

[Ano22-42] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 44(2):84, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSm

[Ano22-43] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 44(3): 98, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSo

[Ano22-44] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 44(4):C2, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSr

[Ano22-45] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 44(4):91, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSk

[Ano22-46] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 44(3):4, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSe nals of the History of Computing, 44(1):C4, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSj

[Ano22-48] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 44(2):C4, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSn

[Ano22-49] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 44(3):C4, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSt

[Ano22-50] Anonymous. IEEE Computer Society Jobs Board. IEEE Annals of the History of Computing, 44(4):C4, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ICSq

[Ano22-51] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE Annals of the History of Computing, 44(4): 88, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:ID

[Ano22-52] Anonymous. IEEE DataPort. IEEE Annals of the History of Computing, 44(4):66, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:IQW

[Ano22-53] Anonymous. IEEE Quantum Week. IEEE Annals of the History of Computing, 44(3): C2, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Ma

[Ano22-54] Anonymous. Masthead. IEEE Annals of the History of Computing, 44(1):3, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Mb

[Ano22-55] Anonymous. Masthead. IEEE Annals of the History of Computing, 44(2):3, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Mc

[Ano22-56] Anonymous. Masthead. IEEE Annals of the History of Computing, 44(3):3, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:Md

[Ano22-57] Anonymous. Masthead. IEEE Annals of the History of Computing, 44(4):3, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:NE

[Ano22-58] Anonymous. Notes from the editor. IEEE Annals of the History of Computing, 44(1): 130, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:RCSa

[Ano22-59] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Annals of the History of Computing, 44 (2):77, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:RCSb

[Ano22-60] Anonymous. Over the rainbow:

21st Century security & privacy podcast. IEEE Annals of the History of Computing, 44 (3):79, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:RCSc

[Ano22-61] Anonymous. Over the rainbow:

21st century security & privacy podcast. IEEE Annals of the History of Computing, 44(4):

43, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:PSE

[Ano22-62] Anonymous. Publications seek 2023 Editors in Chief. IEEE Annals of the History of Computing, 44(1):139, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:SCE

[Ano22-63] Anonymous. Stay on the cutting edge of artificial intelligence. IEEE Annals of the History of Computing, 44(1): 19, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:TCa

[Ano22-64] Anonymous. Table of contents. IEEE Annals of the History of Computing, 44(2):1–2, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:TCb

[Ano22-65] Anonymous. Table of contents. IEEE Annals of the History of Computing, 44(3):1– 2, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:TCc

[Ano22-66] Anonymous. Table of contents. IEEE Annals of the History of Computing, 44(4):1–2, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:WLB

[Ano22-67] Anonymous. Working and living at Bletchley Park. IEEE Annals of the History of Computing, 44(2):65–68, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:WAb

[Ano22-68] Anonymous. Write for Annals. IEEE Annals of the History of Computing, 44(3): 96, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2022:WAa

[Ano22-69] Anonymous. Wyatts award.

IEEE Annals of the History of Computing, 44(2):75, April/ June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CSE

[Ano23a] Anonymous. Computing in Science & Engineering. IEEE Annals of the History of Computing, 45(3):107, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Ca

[Ano23b] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 45(1):C4, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Cb

[Ano23c] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 45(3):C4, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Cc

[Ano23d] Anonymous. ComputingEdge. IEEE Annals of the History of Computing, 45(4):C4, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:IAHa

[Ano23e] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 45(1):96, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:IAHb

[Ano23f] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 45(2):79, April/ June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:IAHc

[Ano23g] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 45(4):108, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICG

[Ano23h]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 45(4):38, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ISP

[Ano23i] Anonymous. IEEE Security & Privacy. IEEE Annals of the History of Computing, 45(4): 107, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ITB

[Ano23j] Anonymous. IEEE Transactions on Big Data. IEEE Annals of the History of Computing, 45(4):10, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ITSa

[Ano23k] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Annals of the History of Computing, 45(1):85, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ITSb

[Ano23l] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Annals of the History of Computing, 45(2): 5, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ITSc

[Ano23m] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Annals of the History of Computing, 45(3):104, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CMA

[Ano23n] Anonymous. Call for 2023 major awards nominations. IEEE Annals of the History of Computing, 45(2):3, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CPIa

[Ano23o]  Anonymous.  Call for pa-

pers: IEEE Computer Society. IEEE Annals of the History of Computing, 45(1):30, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CPIb

[Ano23p] Anonymous. Call for papers:

IEEE Computer Society. IEEE Annals of the History of Computing, 45(2):28, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CPIc

[Ano23q]  Anonymous.  Call for pa-

pers: IEEE Computer Society. IEEE Annals of the History of Computing, 45(3): 5, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CP

[Ano23r] Anonymous. Call for participation. IEEE Annals of the History of Computing, 45(1): 95, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:CM

[Ano23s] Anonymous. CS multimedia.

IEEE Annals of the History of Computing, 45(2):61, April/ June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:E

[Ano23t] Anonymous. Errata. IEEE Annals of the History of Computing, 45(4):107, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:FCa

[Ano23u]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 45(1):C1, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:FCb

[Ano23v] Anonymous. Front cover. IEEE Annals of the History of Computing, 45(2):C1, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:FCc

[Ano23w]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 45(3):C1, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:FCd

[Ano23x]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 45(4):C1, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:GPN

[Ano23y] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 45(2):62, April/ June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:HA

[Ano23z]  Anonymous.  Humphrey

Award. IEEE Annals of the History of Computing, 45(1): 94, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSk

[Ano23-27] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 45(4):52, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSc

[Ano23-28] Anonymous. IEEE Computer Society Career Center. IEEE Annals of the History of Computing, 45(1):C3, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSh

[Ano23-29] Anonymous. IEEE Computer Society career center. IEEE Annals of the History of Computing, 45(3):C3, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSo

[Ano23-30] Anonymous. IEEE Computer Society career center. IEEE Annals of the History of Computing, 45(4):C3, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSd

[Ano23-31] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 45(2):69, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSi

[Ano23-32] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 45(4):C2, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSa

[Ano23-33] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 45(1): C2, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSe

[Ano23-34] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 45(2):78, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSg

[Ano23-35] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 45(3): 65, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSm

[Ano23-36] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 45(4):80, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSb

[Ano23-37] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 45(1):3, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSf

[Ano23-38] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 45(3):3, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSj

[Ano23-39] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 45(4):3, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSn

[Ano23-40] Anonymous. IEEE Computer Society: Publications seek. IEEE Annals of the History of Computing, 45(4):81, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:ICSl

[Ano23-41] Anonymous. IEEE Computer Society: Volunteer service awards. IEEE Annals of the History of Computing, 45(4):64, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:IQWa

[Ano23-42] Anonymous. IEEE quantum week. IEEE Annals of the History of Computing, 45(2): C2, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:IQWb

[Ano23-43] Anonymous. IEEE quantum week. IEEE Annals of the History of Computing, 45(3): C2, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:LSC

[Ano23-44] Anonymous. Let’s stay connected. IEEE Annals of the History of Computing, 45(1): 5, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Ma

[Ano23-45] Anonymous. Masthead. IEEE Annals of the History of Computing, 45(1):4, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Mb

[Ano23-46] Anonymous. Masthead. IEEE Annals of the History of Computing, 45(2):4, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Mc

[Ano23-47] Anonymous. Masthead. IEEE Annals of the History of Computing, 45(3):4, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:Md

[Ano23-48] Anonymous. Masthead. IEEE Annals of the History of Computing, 45(4):4, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:RCSa

[Ano23-49] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Annals of the History of Computing, 45 (3):102, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:RCSb

[Ano23-50] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Annals of the History of Computing, 45(4): 91, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:SMY

[Ano23-51] Anonymous. Share and manage your research data. IEEE Annals of the History of Computing, 45(2):80, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:SCEa

[Ano23-52] Anonymous. Stay on the cutting edge of artificial intelligence. IEEE Annals of the History of Computing, 45(1): 87, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:SCEbb

[Ano23-53] Anonymous. Stay on the cutting edge of artificial intelligence. IEEE Annals of the History of Computing, 45(4):5, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:TCa

[Ano23-54] Anonymous. Table of contents. IEEE Annals of the History of Computing, 45(1):1– 2, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:TCb

[Ano23-55] Anonymous. Table of contents.

IEEE Annals of the History of Computing, 45(2):1–2, April/ June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:TCc

[Ano23-56] Anonymous. Table of contents. IEEE Annals of the History of Computing, 45(3):1– 2, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:TCd

[Ano23-57] Anonymous. Table of contents. IEEE Annals of the History of Computing, 45(4):1–2, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2023:WA

[Ano23-58] Anonymous. Write for Annals. IEEE Annals of the History of Computing, 45(3):108, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ATW

[Ano24a] Anonymous. AI’s ten to watch. IEEE Annals of the History of Computing, 46(2):91, April/ June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:AC

[Ano24b] Anonymous. Author choice. IEEE Annals of the History of Computing, 46(4):C2, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CSE

[Ano24c] Anonymous. Computing in Science & Engineering. IEEE Annals of the History of Computing, 46(4):147, October/ December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CEa

[Ano24d] Anonymous. Computing Edge. IEEE Annals of the History of Computing, 46(3):C4, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CEb

[Ano24e] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 46(4):C4, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Ca

[Ano24f] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 46(1):C4, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Cb

[Ano24g] Anonymous. ComputingEdge.

IEEE Annals of the History of Computing, 46(2):C4, April/ June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAHa

[Ano24h] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 46(1):80, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAHb

[Ano24i] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 46(2):92, April/ June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAHc

[Ano24j] Anonymous. IEEE Annals of the History of Computing. IEEE Annals of the History of Computing, 46(4):152, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICGa

[Ano24k]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 46(2): 71, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICGb

[Ano24l]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Annals of the History of Computing, 46(4):83, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IPC

[Ano24m] Anonymous. IEEE Pervasive Computing: Call for articles. IEEE Annals of the History of Computing, 46(4):6, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ISP

[Ano24n] Anonymous. IEEE Security and Privacy. IEEE Annals of the History of Computing, 46(4):37, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ITC

[Ano24o] Anonymous. IEEE Transactions on Computers: Call for papers. IEEE Annals of the History of Computing, 46(4): 149, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ITP

[Ano24p] Anonymous. IEEE Transactions on Privacy. IEEE Annals of the History of Computing, 46 (4):5, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:TP

[Ano24q] Anonymous. Transactions on Privacy. IEEE Annals of the History of Computing, 46(3): C2, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CCC

[Ano24r] Anonymous. CA Career Center. IEEE Annals of the History of Computing, 46(3):C3, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CPIa

[Ano24s] Anonymous. Call for papers: IEEE Computer Architecture. IEEE Annals of the History of Computing, 46(1):19, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CPG

[Ano24t]  Anonymous.  Call for pa-

pers: Get published in the new IEEE Transactions on Privacy. IEEE Annals of the History of Computing, 46(1): 73, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CPIb

[Ano24u]  Anonymous.  Call for pa-

pers: IEEE Computer Society. IEEE Annals of the History of Computing, 46(1):65, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CPIc

[Ano24v] Anonymous. Call for papers: IEEE Computer Society. IEEE Annals of the History of Computing, 46(2):12, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CSI

[Ano24w] Anonymous. Call for special issue proposals. IEEE Annals of the History of Computing, 46 (2):88, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Cc

[Ano24x] Anonymous. CGA. IEEE Annals of the History of Computing, 46(3):49, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CIa

[Ano24y] Anonymous. CS info. IEEE Annals of the History of Computing, 46(3):3, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:CIb

[Ano24z] Anonymous. CS info. IEEE Annals of the History of Computing, 46(4):3, October/ December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:FCa

[Ano24-27] Anonymous. Front cover.

IEEE Annals of the History of Computing, 46(1):C1, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:FCb

[Ano24-28] Anonymous. Front cover. IEEE Annals of the History of Computing, 46(2):C1, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:FCc

[Ano24-29] Anonymous. Front cover. IEEE Annals of the History of Computing, 46(3):C1, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:FCd

[Ano24-30] Anonymous. Front cover. IEEE Annals of the History of Computing, 46(4):C1, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:GPN

[Ano24-31] Anonymous. Get published in the new IEEE Transactions on Privacy. IEEE Annals of the History of Computing, 46(2): C2, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAa

[Ano24-32] Anonymous. IEEE app. IEEE Annals of the History of Computing, 46(3):93, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAb

[Ano24-33] Anonymous. IEEE app. IEEE Annals of the History of Computing, 46(4):123, October/ December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IAL

[Ano24-34] Anonymous. The IEEE APP: Let’s stay connected . . .. IEEE Annals of the History of Computing, 46(1):79, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSi

[Ano24-35] Anonymous. IEEE Computer Society. IEEE Annals of the History of Computing, 46(3): 35, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSh

[Ano24-36] Anonymous. IEEE Computer Society call for papers. IEEE Annals of the History of Computing, 46(3):6, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSj

[Ano24-37] Anonymous. IEEE Computer Society: Call for papers. IEEE Annals of the History of Computing, 46(4):69, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSd

[Ano24-38] Anonymous. IEEE Computer Society Career Center.

IEEE Annals of the History of Computing, 46(1):C3, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSg

[Ano24-39] Anonymous. IEEE Computer Society Career Center. IEEE Annals of the History of Computing, 46(2):C3, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSk

[Ano24-40] Anonymous. IEEE Computer Society Career Center. IEEE Annals of the History of Computing, 46(4):C3, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSa

[Ano24-41] Anonymous. IEEE Computer Society D&I Fund. IEEE Annals of the History of Computing, 46(1):C2, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSc

[Ano24-42] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 46(1): 61, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSf

[Ano24-43] Anonymous. IEEE Computer Society has you covered! IEEE Annals of the History of Computing, 46(2):37, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSb

[Ano24-44] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 46(1):3, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:ICSe

[Ano24-45] Anonymous. IEEE Computer Society information. IEEE Annals of the History of Computing, 46(2):3, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IDa

[Ano24-46] Anonymous. IEEE Dataport.

IEEE Annals of the History of Computing, 46(3):85, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:IDb

[Ano24-47] Anonymous. IEEE Dataport.

IEEE Annals of the History of Computing, 46(4):113, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Ma

[Ano24-48] Anonymous. Masthead. IEEE Annals of the History of Computing, 46(1):4, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Mb

[Ano24-49] Anonymous. Masthead. IEEE Annals of the History of Computing, 46(2):4, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Mc

[Ano24-50] Anonymous. Masthead. IEEE Annals of the History of Computing, 46(3):4, July/ September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:Md

[Ano24-51] Anonymous. Masthead. IEEE Annals of the History of Computing, 46(4):4, October/ December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:SCa

[Ano24-52] Anonymous. Stay connected. IEEE Annals of the History of Computing, 46(1):59, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:SCc

[Ano24-53] Anonymous. Stay connected.

IEEE Annals of the History of Computing, 46(4):136, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:SCb

[Ano24-54] Anonymous. Sustainable computing. IEEE Annals of the History of Computing, 46(3): 80, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:TCa

[Ano24-55] Anonymous. Table of contents. IEEE Annals of the History of Computing, 46(1):1– 2, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:TCb

[Ano24-56] Anonymous. Table of contents.

IEEE Annals of the History of Computing, 46(2):1–2, April/ June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:TCc

[Ano24-57] Anonymous. Table of contents. IEEE Annals of the History of Computing, 46(3):1– 2, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:TCd

[Ano24-58] Anonymous. Table of contents. IEEE Annals of the History of Computing, 46(4):1–2, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:UYP

[Ano24-59] Anonymous. Unlock your potential. IEEE Annals of the History of Computing, 46(4):9, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:VSA

[Ano24-60] Anonymous. Volunteer service awards. IEEE Annals of the History of Computing, 46 (2):65, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2024:WA

[Ano24-61] Anonymous. Write for Annals. IEEE Annals of the History of Computing, 46(3): 96, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:AC

[Ano25a] Anonymous. Author choice. IEEE Annals of the History of Computing, 47(1):C2, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CEa

[Ano25b] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 47(1):C4, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CEb

[Ano25c] Anonymous. Computing Edge.

IEEE Annals of the History of Computing, 47(2):C4, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:IOJ

[Ano25d] Anonymous. IEEE Open Journal of the Computer Society. IEEE Annals of the History of Computing, 47(2):6, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TPa

[Ano25e] Anonymous. Transactions on Privacy. IEEE Annals of the History of Computing, 47(1): 5, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TPb

[Ano25f] Anonymous. Transactions on Privacy. IEEE Annals of the History of Computing, 47(2): 59, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CBD

[Ano25g] Anonymous. CFP Big Data. IEEE Annals of the History of Computing, 47(2):63, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CTC

[Ano25h] Anonymous. CFP Transactions on Computing. IEEE Annals of the History of Computing, 47 (2):61, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CS

[Ano25i] Anonymous. CFP software. IEEE Annals of the History of Computing, 47(2):32, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:C

[Ano25j] Anonymous. CiSE. IEEE Annals of the History of Computing, 47(1):71, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CUY

[Ano25k] Anonymous. CS — unlock your potential. IEEE Annals of the History of Computing, 47 (2):19, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CCP

[Ano25l] Anonymous. CS call for papers.

IEEE Annals of the History of Computing, 47(2):43, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CCCa

[Ano25m] Anonymous. CS Career Center. IEEE Annals of the History of Computing, 47(1):C3, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CCCb

[Ano25n] Anonymous. CS Career Center.

IEEE Annals of the History of Computing, 47(2):C3, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CIa

[Ano25o] Anonymous. CS info. IEEE Annals of the History of Computing, 47(1):3, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:CIb

[Ano25p] Anonymous. CS info. IEEE Annals of the History of Computing, 47(2):3, April/June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:ET

[Ano25q] Anonymous. Emerging tech. IEEE Annals of the History of Computing, 47(2):C2, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:FCa

[Ano25r]  Anonymous.  Front cover.

IEEE Annals of the History of Computing, 47(1):C1, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:FCb

[Ano25s] Anonymous. Front cover. IEEE Annals of the History of Computing, 47(2):C1, April/June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:Ma

[Ano25t] Anonymous. Masthead. IEEE Annals of the History of Computing, 47(1):4, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:Mb

[Ano25u] Anonymous. Masthead. IEEE Annals of the History of Computing, 47(2):4, April/June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:S

[Ano25v] Anonymous. Software. IEEE Annals of the History of Computing, 47(1):80, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TCa

[Ano25w] Anonymous. Table of contents. IEEE Annals of the History of Computing, 47(1):1– 2, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TCd

[Ano25x] Anonymous. Table of contents.

IEEE Annals of the History of Computing, 47(2):1–2, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TCb

[Ano25y] Anonymous. TC CFP. IEEE Annals of the History of Computing, 47(1):35, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:TCc

[Ano25z] Anonymous. TSUSC CFP.

IEEE Annals of the History of Computing, 47(1):49, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2025:VSA

[Ano25-27] Anonymous. Volunteer service awards. IEEE Annals of the History of Computing, 47 (2):67, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aris:2000:ISE

[Ari00] J. Aris. Inventing systems engineering [LEO]. IEEE Annals of the History of Computing, 22(3):4–15, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18655/00859522.pdf.

Ashfaque:1976:RBG

[Ash76] S. M. Ashfaque. Review: Gears from Greeks — Antikythera Mechanism — Calendar Computer from ca 80 BC — by D. D. Price. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):175–176, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aspray:1992:IPa

[Asp92a] William Aspray. From IEEE’s perspective. IEEE Annals of the History of Computing, 14(2):5–??, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/

an/books/an1992/pdf/a2005.
pdf.

Aspray:1992:IPb

[Asp92b]  William Aspray.  From

IEEE’s perspective. IEEE Annals of the History of Computing, 14(3):5–??, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3005.pdf.

Aspray:1992:IPc

[Asp92c] William Aspray. From IEEE’s perspective. IEEE Annals of the History of Computing, 14(4):5–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4005.pdf.

Aspray:1993:EHA

[Asp93a] William Aspray. Edwin L.

Harder and the Anacom: Analog computing at Westinghouse. IEEE Annals of the History of Computing, 15(2):35– 52, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2035.pdf;   http://www.
computer.org/annals/an1993/
a2035abs.htm.

Aspray:1993:IPa

[Asp93b]  William Aspray.  From

IEEE’s perspective. IEEE Annals of the History of Computing, 15(1):5–??, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a1005.pdf.

Aspray:1993:IPb

[Asp93c] William Aspray. From IEEE’s perspective. IEEE Annals of the History of Computing, 15(2):5–??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://dlib.computer.org/
                                        an/books/an1993/pdf/a2005.
                                        pdf.

Aspray:1993:IPc

[Asp93d]  William Aspray.  From

IEEE’s perspective. IEEE Annals of the History of Computing, 15(3):4–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a3004.pdf.

Aspray:1993:IPd

[Asp93e] William Aspray. From IEEE’s perspective. IEEE Annals of the History of Computing, 15(4):4–??, October/ December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4004.pdf.

Aspray:1997:IMW

[Asp97] W. Aspray. The Intel 4004 microprocessor: what constituted invention? IEEE Annals of the History of Computing, 19(3): 4–15, July/September 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13172/00601727.pdf.

Aspray:1999:CCD

[Asp99] W. Aspray. Command and control, documentation, and library science: the origins of information science at the University of Pittsburgh. IEEE Annals of the History of Computing, 21(4): 4–20, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801528.pdf.

Aspray:2000:WEE

[Asp00] W. Aspray. Was early entry a competitive advantage? US universities that entered computing in the 1940s. IEEE Annals of the History of Computing, 22(3): 42–87, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18655/00859525.pdf.

Aspray:2001:AET

[Asp01a] William Aspray. Adelle and Erwin Tomash: Honoring a couple’s contributions to computer history. IEEE Annals of the History of Computing, 23(4):2–3, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2001/pdf/
                                        a4002.pdf;   http://www.
                                        computer.org/annals/an2001/
                                        a4002abs.htm.

Aspray:2001:ETH

[Asp01b]  William Aspray.  Erwin

Tomash: His life and work. IEEE Annals of the History of Computing, 23(4):4–11, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2001/a4004abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2001/pdf/a4004.
                                        pdf.

Aspray:2001:ES

[Asp01c]  William Aspray.  Events

and sightings. IEEE Annals of the History of Computing, 23(4):105–106, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/

annals/an2001/a4105abs.htm;
http://computer.org/annals/
articles/a4-2001/eands.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a4105.
pdf.

Aspray:2007:LCH

[Asp07] William Aspray. Leadership in computing history: Arthur Norberg and the Charles Babbage Institute. IEEE Annals of the History of Computing, 29 (4):16–26, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aspray:2013:CIE

[Asp13] William Aspray. Computers, information, and everyday life. IEEE Annals of the History of Computing, 35(4):96, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aspray:2014:MSM

[Asp14] W. Aspray. Michael Sean Mahoney (1939–2008). IEEE Annals of the History of Computing, 36(3):70–79, July/ September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aspray:2015:ISD

[Asp15]  William Aspray.  Informa-

tion society, domains, and culture. IEEE Annals of the History of Computing, 37(2):2– 4, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020002-abs.html.

Aspray:2021:BAL

[Asp21]  William Aspray.  Biogra-

phy: Arthur Lawrence Norberg (1938–2021). IEEE Annals of the History of Computing, 43 (4):70–76, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Astarte:2024:EDa

[Ast24a] Troy Astarte. From the Editor’s desk. IEEE Annals of the History of Computing, 46 (3):5–6, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Astarte:2024:EDb

[Ast24b] Troy Astarte. From the Editor’s desk. IEEE Annals of the History of Computing, 46 (4):6, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Astarte:2025:EDa

[Ast25a] Troy Astarte. From the Editor’s desk. IEEE Annals of the History of Computing, 47(1): 6, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Astarte:2025:EDb

[Ast25b] Troy Astarte. From the Editor’s desk. IEEE Annals of the History of Computing, 47 (2):5, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aumann:2011:DUS

[Aum11] Philipp Aumann. The distinctiveness of a unifying science: Cybernetics’ way to West Germany. IEEE Annals of the History of Computing, 33(4):17– 27, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aspray:1994:AAS

[AW94] William Aspray and Bernard O.

Williams. Arming American scientists: NSF and the provision of scientific computing facilities for universities, 1950– 1973. IEEE Annals of the History of Computing, 16(4): 60–74, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4060.pdf;   http://www.
computer.org/annals/an1994/
a4060abs.htm.

Aspray:2011:NVN

[AY11] William Aspray and Jeffrey R.

Yost. New voices, new topics. IEEE Annals of the History of Computing, 33(2):4–8, April/ June 2011. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic).

Aylen:2010:PPC

[Ayl10a]  Jonathan Aylen.  Promot-

ing the prosaic: The case for process-control computers. IEEE Annals of the History of Computing, 32(3):94–96, July/ September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Aylen:2010:QSN

[Ayl10b] Jonathan Aylen. A question of scale: Networks, systems, and practice. IEEE Annals of the History of Computing, 32 (4):94–96, October/December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Baber:1997:CEE

[Bab97] R. L. Baber. Comparison of electrical “engineering” of Heaviside’s times and software “engineering” of our times. IEEE Annals of the History of Computing, 19(4): 5–17, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/13659/00627895.pdf.

Backus:1998:HFI

[Bac98] J. Backus. The history of Fortran I, II, and III. IEEE Annals of the History of Computing, 20 (4):68–78, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728232.pdf.

Bachman:2009:OID

[Bac09] Charles W. Bachman. The origin of the integrated data store (IDS): The first directaccess DBMS. IEEE Annals of the History of Computing, 31 (4):42–54, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Baker:2017:SBU

[Bak17]  Kevin T. Baker.  Sim-

ulation and ‘seduction’ at the policy interface. IEEE Annals of the History of Computing, 39(3):6–7, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030006-abs.html.

Ballin:1976:BRB

[Bal76] Theodore N. Ballin. Book review: Gears from the Greeks, The Antikythera Mechanism: A Calendar Computer from ca. 80 B.C. by Derek de Solla Price. The Classical World, 70(3):202, November 1976. ISSN 0009-8418 (print), 1558-9234 (electronic).

URL http://www.jstor.org/
stable/4348621.

Baltes:2024:PJH

[Bal24] Cheryl Baltes. Patrick J. Hanratty: Pioneer in commercial CAD/CAM software systems. IEEE Annals of the History of Computing, 46(4):50–57, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Barnes:1997:DCE

[Bar97] S. B. Barnes. Douglas Carl Engelbart: developing the underlying concepts for contemporary computing. IEEE Annals of the History of Computing, 19(3):16–26, July/September 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/13172/00601730.pdf.

Barnes:2003:BDE

[Bar03]  S. B. Barnes.  Bootstrap-

ping: Douglas Engelbart, coevolution, and the origins of personal computing [book review]. IEEE Annals of the History of Computing, 25(3):84– 85, July/September 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Baron:2005:CSR

[Bar05] Sheldon Baron. Control systems R&D at BBN. IEEE Annals of the History of Computing, 27(2):52–64, April/ June 2005. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2052.
htm;   http://csdl.computer.
org/dl/mags/an/2005/02/a2052.
pdf.

Barnes:2007:AKT

[Bar07] Susan B. Barnes. Alan Kay:

Transforming the computer into a communication medium. IEEE Annals of the History of Computing, 29(2):18–30, April/June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Barr:2024:MTJ

[Bar24] Michael Barr. My time with the JFEAC. IEEE Annals of the History of Computing, 46(2): 66–71, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bauer:2006:IAS

[Bau06]  Walter F. Bauer.  Infor-

matics acquisition by Sterling Software: Unsolicited offer, takeover attempt, and merger. IEEE Annals of the History of Computing, 28(3):32–40, July/ September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bauer:2007:CRE

[Bau07] Walter F. Bauer. Computer recollections: Events, humor, and happenings. IEEE Annals of the History of Computing, 29(1):85–89, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bregman:2002:LAU

[BB02] Alvan Bregman and Robert H. Burger. Library automation at the University of Illinois at Urbana-Champaign, 1965– 2000: a case study of technological and organizational validity. IEEE Annals of the History of Computing, 24(2):71– 85, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2071.pdf;   http://www.
computer.org/annals/an2002/
a2071abs.htm.

Bloomenthal:1998:GR

[BBB^{+}98] J. Bloomenthal, R. E. Barnhill, B. A. Barsky, P. Bezier, R. Forrest, N. Max, D. M. Palyka, D. F. Rogers, H. Rushmeier, A. R. Smith, R. Stock, N. M. Thalmann, and D. Thalmann. Graphics remembrances. IEEE Annals of the History of Computing, 20 (2):35–51, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667296.pdf.

Bastian:2009:IFGa

[BBC^{+}09a] Michel Bastian, Marc Boisseau, Robert Cohendet, Alain Croisier, Claude Galand, Etienne Gorog, Philippe Hernandez, Michel Humbert, Cuong Ngo Mai, Pierre Secondo, and Robert F. Steen. IBM France La Gaude Laboratory contributions to telecommunications: Part 1. IEEE Annals of the History of Computing, 31(2):4–17, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bastian:2009:IFGb

[BBC^{+}09b] Michel Bastian, Marc Boisseau, Robert Cohendet, Alain Croisier, Claude Galand, Etienne Gorog, Philippe Hernandez, Michel Humbert, Cuong Ngo Mai, Pierre Secondo, and Robert F. Steen. IBM France La Gaude Laboratory contributions to telecommunications: Part 2. IEEE Annals of the History of Computing, 31(2):18–30, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bowen:2006:R

[BBS06] Jonathan P. Bowen, Tim Bergin, and Christopher H. Sterling. Reviews. IEEE Annals of the History of Computing, 28(2):77–80, April/ June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/02/a2077.pdf.

Beeton:2018:TBD

[BBW18] Barbara Beeton, Karl Berry, and David Walden. TEX: A branch of desktop publishing, part 1. IEEE Annals of the History of Computing, 40(3):78–93, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8509554/.

Beeton:2019:TBD

[BBW19] Barbara Beeton, Karl Berry, and David Walden. TEX: A branch in desktop publishing evolution, part 2. IEEE Annals of the History of Computing, 41(2):29–41, April 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bellovin:2025:NOS

[Bel25] Steven M. Bellovin. Netnews: The origin story. IEEE Annals of the History of Computing, 47 (1):7–21, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bennett:1993:CQDa

[Ben93a] J. M. Bennett. Comments, queries, and debate: Computing at the University of Cambridge. IEEE Annals of the History of Computing, 15(2):6– 7, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2006.pdf.

Bennett:1993:CQDb

[Ben93b] J. M. Bennett. Comments, queries, and debate: Still more on mechanical delay lines. IEEE Annals of the History of Computing, 15(2): 7, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2006.pdf.

Bennett:2013:OMK

[Ben13] Rick Bennett. Oracle marketing: Killer ads. IEEE Annals of the History of Computing, 35(2):24–30, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Berg:1992:CQD

[Ber92] Herman Berg. Comments, queries, and debate: On locating the Babbage–Quetelet letter. IEEE Annals of the History of Computing, 14(1):7–9, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1006.pdf.

Bergin:1994:SIP

[Ber94] Thomas J. Bergin. Special issue on programming languages. IEEE Annals of the History of Computing, 16(4): 4–12, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4004.pdf.

Bergin:2000:AI

[Ber00] Tim Bergin. About this issue. IEEE Annals of the History of Computing, 22(4):2–??, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4002.pdf.

Bergin:2001:AIa

[Ber01a] Tim Bergin. About this issue. IEEE Annals of the History of Computing, 23(1):2–??, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1002.pdf.

Bergin:2001:AIb

[Ber01b] Tim Bergin. About this issue. IEEE Annals of the History of Computing, 23(3):2, July/September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3002abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a3002.
pdf.

Bergin:2001:AIE

[Ber01c] Tim Bergin. About this issue/from the Editor’s desk. IEEE Annals of the History of Computing, 23(2):2– ??, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2002.pdf.

Bergin:2002:AIE

[Ber02a] Tim Bergin. About this issue/from the Editor’s desk. IEEE Annals of the History of Computing, 24(4):2, October/December 2002. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2002/04/a4002.htm;   http:
//csdl.computer.org/dl/mags/
an/2002/04/a4002.pdf.

Bergin:2002:EDa

[Ber02b] Tim Bergin. From the Editor’s desk. IEEE Annals of the History of Computing, 24(1):2, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1002abs.htm;
http://dlib.computer.org/
an/books/an2002/pdf/a1002.
pdf.

Bergin:2002:EDb

[Ber02c] Tim Bergin. From the Editor’s desk. IEEE Annals of the History of Computing, 24(2): 2, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2002.pdf;   http://www.
computer.org/annals/an2002/
a2002abs.htm.

Bergin:2003:AIa

[Ber03a] Tim Bergin. About this issue. IEEE Annals of the History of Computing, 25(2):2, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2002.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2002.
pdf.

Bergin:2003:AIb

[Ber03b]  Tim Bergin.  About this

issue. IEEE Annals of the History of Computing, 25(3):2–3, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3002.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3002.
pdf.

Bergin:2003:AIE

[Ber03c] Tim Bergin. About this issue/from the Editor’s desk. IEEE Annals of the History of Computing, 25(1):2– 3, January/March 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/01/a1002.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1002.pdf.

Bergin:2003:BRC

[Ber03d] Tim Bergin. Book review:

Christine A. Finn, An Archeologist’s Year in Silicon Valley, MIT Press, 2001, ISBN 0-262-06224-0, $24.95 (hardcover). IEEE Annals of the History of Computing, 25(4): 96, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bergin:2003:EDS

[Ber03e] Tim Bergin. From the Editor’s desk: Some final thoughts. IEEE Annals of the History of Computing, 25(4):2–3, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL a4002.pdf;

http://csdl.computer.org/
dl/mags/an/2003/04/a4002.
htm.

Berkovich:2003:ARS

[Ber03f] S. Berkovich. Anecdotes: Reminiscences of superconductive associative memory research in the former Soviet Union. IEEE Annals of the History of Computing, 25(1):72–75, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Beranek:2005:BED

[Ber05] Leo L. Beranek. BBN’s earliest days: Founding a culture of engineering creativity. IEEE Annals of the History of Computing, 27(2):6–14, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2006.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/02/a2006.
                                        pdf.

Bergin:2006:OWP

[Ber06a] Thomas J. (Tim) Bergin. The origins of word processing software for personal computers: 1976-1985. IEEE Annals of the History of Computing, 28 (4):32–47, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bergin:2006:PCW

[Ber06b] Thomas J. (Tim) Bergin. The proliferation and consolidation of word processing software: 1985-1995. IEEE Annals of the History of Computing, 28 (4):48–63, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bergin:2006:BCH

[Ber06c] Tim Bergin. Biographies: Carl Hammer (1914–2004). IEEE Annals of the History of Computing, 28(2):81–86, April/ June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/02/a2081.pdf.

Bergin:2009:JSP

[Ber09] Thomas J. (Tim) Bergin. Jean Sammet: Programming language contributor and historian, and ACM President. IEEE Annals of the History of Computing, 31(1):76–85, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bergin:2014:WCA

[Ber14] T. J. Bergin. What is a computer? An answer from the past. IEEE Annals of the History of Computing, 36(3):80– 84, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bernard:2024:DSC

[Ber24] Francis Bernard. Dassault Systèmes: a CAD success story. IEEE Annals of the History of Computing, 46(4):124–135, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Barkley-Fitz:1996:WE

[BF96] W. Barkley-Fitz. The women of Eniac. IEEE Annals of the History of Computing, 18 (3):13–23, Fall 1996. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bowker:1993:GI

[BG93a] Geof Bowker and Richard Giordano. General introduction. IEEE Annals of the History of Computing, 15(3):6–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3006.pdf.

Bowker:1993:ITK

[BG93b] Geof Bowker and Richard Giordano. Interview with Tom Kilburn. IEEE Annals of the History of Computing, 15(3):17–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3017.pdf;   http://www.
computer.org/annals/an1993/
a3017abs.htm.

Brock:2022:ESC

[BG22] David C. Brock and Burton Grad. Expert systems: Commercializing artificial intelligence. IEEE Annals of the History of Computing, 44(1):5– 7, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bregzis:2002:BAU

[BGM02] Ritvars Bregzis, Calvin Gotlieb, and Carole Moore. The beginning of automation in the University of Toronto Library, 1963–1972. IEEE Annals of the History of Computing, 24(2):50–70, April/ June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an2002/pdf/a2050.
pdf;   http://www.computer.
org/annals/an2002/a2050abs.
htm.

Buck:1995:DIC

[BH95] George Buck and Steve Hunka.

Development of the IBM 1500 computer-assisted instructional system. IEEE Annals of the History of Computing, 17(1): 19–31, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Buck:1999:WSJ

[BH99] G. H. Buck and S. M. Hunka.

W. Stanley Jevons, Allan Marquand, and the origins of digital computing. IEEE Annals of the History of Computing, 21 (4):21–27, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801529.pdf.

Bergin:2009:CDM

[BH09] Thomas J. Bergin and Thomas Haigh. The commercialization of database management systems, 1969–1983. IEEE Annals of the History of Computing, 31 (4):26–41, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bigelow:2020:FWPa

[Big20a] Charles Bigelow. The font wars, part 1. IEEE Annals of the History of Computing, 42 (1):7–24, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bigelow:2020:FWPb

[Big20b] Charles Bigelow. The font wars, part 2. IEEE Annals of the History of Computing, 42(1):25–40, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Birkenstock:2000:PFE

[Bir00] J. W. Birkenstock. Pioneering: on the frontier of electronic data processing, a personal memoir. IEEE Annals of the History of Computing, 22 (1):4–47, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17665/00815465.pdf.

Bissell:1998:WIF

[Bis98] D. Bissell. Was the IDI-IOM the first stand-alone CAD platform? IEEE Annals of the History of Computing, 20(2):14–19, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667292.pdf.

Black:2003:WML

[BJ03] Don Black and Walter Dickson Jones. Watson and me: a life at IBM. IEEE Annals of the History of Computing, 25(3):4–18, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a3004abs.htm;   http:
//computer.org/annals/an2003/
extras/a3004.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/03/a3004.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/03/a3004.pdf.

Bruemmer:2001:RCH

[BK01] Bruce H. Bruemmer and Elisabeth Kaplan. Realizing the concept: a history of the CBI archives. IEEE Annals of the History of Computing, 23(4):29–38, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a4029.pdf;   http://www.
computer.org/annals/an2001/
a4029abs.htm.

Brock:2012:EHM

[BL12] David C. Brock and David A.

Laws. The early history of microcircuitry: An overview. IEEE Annals of the History of Computing, 34(1):7–19, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Batiz-Lazo:2012:ECA

[BLH12] Bernardo Batiz-Lazo and Thomas Haigh. Engineering change: The appropriation of computer technology at Grupo ICA in Mexico (1965–1971). IEEE Annals of the History of Computing, 34(2):20–33, April/June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Batiz-Lazo:2011:DCD

[BLR11] Bernardo Batiz-Lazo and Robert J. K. Reid. The development of cash-dispensing technology in the UK. IEEE Annals of the History of Computing, 33(3): 32–45, July/September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Barcella:2025:NEC

[BMFB25] Deborah Barcella, Petra Mazzoni, Martin Fomasi, and Gabriele Balbi. Notes from the 2025 ECREA Communication History Workshop: Communication Networks Before and After the Web. IEEE Annals of the History of Computing, 47(2):64–66, April/June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bohl:2013:PMD

[Boh13] Marilyn Bohl. Product managing DB2’s early development. IEEE Annals of the History of Computing, 35(2): 67–68, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Borg:2008:FDR

[Bor08] Kevin L. Borg. Following digital ripples. IEEE Annals of the History of Computing, 30(2): 84, 82, April/June 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Boszormenyi:2008:GEI

[Bös08] Laszlo Böszörmenyi. Guest Editor’s introduction: History of informatics. IEEE Annals of the History of Computing, 30 (3):4–7, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bowker:1994:Ha

[Bow94a] Geoffrey C. Bowker. Happenings. IEEE Annals of the History of Computing, 16(3): 57–??, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a3057.pdf.

Bowker:1994:Hb

[Bow94b] Geoffrey C. Bowker. Happenings. IEEE Annals of the History of Computing, 16(4): 75–??, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a4075.pdf.

Bowker:1995:H

[Bow95] Geoffrey Bowker. Happenings. IEEE Annals of the History of Computing, 17(1): 54–61, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bowker:1996:Ha

[Bow96a] G. Bowker. Happenings. IEEE Annals of the History of Computing, 18(1):56, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476562.pdf.

Bowker:1996:Hb

[Bow96b]  G. Bowker.  Happenings.

IEEE Annals of the History of Computing, 18(4): 67, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11673/00539921.pdf.

Bowles:1996:UTE

[Bow96c] M. D. Bowles. U.S. technological enthusiasm and British technological skepticism in the age of the analog brain. IEEE Annals of the History of Computing, 18(4):5– 15, October/December 1996.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539911.pdf.

Bowker:1998:IMH

[Bow98a] G. C. Bowker. 1998 IEEE Medal of Honor. IEEE Annals of the History of Computing, 20(2):80, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667313.pdf.

Bowker:1998:ICS

[Bow98b] G. C. Bowker. IEEE Computer Society awards. IEEE Annals of the History of Computing, 20(2):79–80, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667312.pdf.

Braden:2013:BAS

[Bra13] Robert Braden. Burroughs Algol at Stanford University, 1960–1963. IEEE Annals of the History of Computing, 35 (4):69–73, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Brace:2022:VCF

[Bra22] Jeffrey Brace. Vintage computer federation events. IEEE Annals of the History of Computing, 44(1):129–130, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Brewer:2025:ETH

[Bre25] Nathan Brewer. The engineering and technology history wiki. IEEE Annals of the History of Computing, 47(2): 62–63, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bride:2006:MMI

[Bri06] Ed Bride. The media are the message: ’The Influencers’. IEEE Annals of the History of Computing, 28(4):74–79, October/December 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Brookes:1977:BRB

[Bro77] A. M. P. Brookes. Book review: Gears from the Greeks, by Derek de Solla Price. The Antikythera Mechanism: A Calendar Computer from ca. 80 B.C. The Classical Review, 27 (1):94–95, ???? 1977. CO-DEN ???? ISSN 0009-840X (print), 1464-3561 (electronic).

                                        URL http://www.jstor.org/
                                        stable/710979.

Bromley:1986:NAM

[Bro86] Allan G. Bromley. Notes on the Antikythera mechanism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29 (1):5–27, March 1986. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bromley:1998:CBA

[Bro98]  A. G. Bromley.  Charles

Babbage’s Analytical Engine, 1838. IEEE Annals of the History of Computing, 20(4):29– 45, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728228.pdf.

Brosveet:1999:ISM

[Bro99a] J. Brosveet. IBM salesman meets Norwegian tax collector: computer entrepreneurs in the making. IEEE Annals of the History of Computing, 21(2):5–13, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16485/00761790.pdf.

Brown:1999:FTF

[Bro99b] L. C. Brown. Flyable TRADIC:

the first airborne transistorized digital computer. IEEE Annals of the History of Computing, 21 (4):55–61, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801533.pdf.

Bromley:2000:BAE

[Bro00]  Allan G. Bromley.  Bab-

bage’s Analytical Engine Plans 28 and 28a. The programmer’s interface. IEEE Annals of the History of Computing, 22(4):5–19, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2000/pdf/
                                        a4005.pdf;   http://ieeexplore.
                                        ieee.org/iel5/85/19196/00887986.
                                        pdf;   http://www.computer.
                                        org/annals/an2000/a4005abs.
                                        htm.

Brooks:2010:SIG

[Bro10] Frederick P. Brooks, Jr. Stretch-ing is great exercise — it gets you in shape to win. IEEE Annals of the History of Computing, 32(1):4–9, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Brock:2018:OHJ

[Bro18] David C. Brock. Oral history of Joe Thompson. IEEE Annals of the History of Computing, 40(2):6–16, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8415731/.

Brock:2023:DCH

[Bro23] David Brock. Documenting CAD history with CHM’s Software Industry Special Interest Group. IEEE Annals of the History of Computing, 45(4): 84, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Brueck:2009:SMS

[Bru09] Robert L. Brueck. System 2000: The MRI Systems Corporation. IEEE Annals of the History of Computing, 31 (4):76–86, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bruderer:2017:EHC

[Bru17] Herbert Bruderer. Early history of computing in Switzerland: Discovery of rare devices, unknown documents, and scarcely known facts. IEEE Annals of the History of Computing, 39(1):65–72, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010065-abs.html.

Burns:1996:CQD

[BS96] Coral Burns and Natalie Simpson. Comments, queries, and debate: Another difference engine, or who was Berry? IEEE Annals of the History of Computing, 18(4): 65, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11673/00539920.pdf.

Bell:2011:BCS

[BS11] Gordon Bell and Daniel P. Siewiorek. The book Computer Structures: Thoughts after 40 years. IEEE Annals of the History of Computing, 33(2): 89–95, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Boell:2023:COH

[BT23a] Sebastian K. Boell and Peter Thorne. Capturing an oral history of computing in Australia. IEEE Annals of the History of Computing, 45(4):82–83, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Boell:2023:HCO

[BT23b] Sebastian K. Boell and Janet M. Toland. Histories of computing in Oceania. IEEE Annals of the History of Computing, 45(4): 6–10, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Buchholz:1995:CQDa

[Buc95a] Werner Buchholz. Comments, queries, and debate. IEEE Annals of the History of Computing, 17(1):4–5, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Buchholz:1995:CQDb

[Buc95b] Werner Buchholz. Comments, queries, and debate: Editorial note: William Friedman. IEEE Annals of the History of Computing, 17(3):6, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397054.pdf.

Buckholtz:2020:QC

[Buc20]  E. Buckholtz.  Queens of

code. IEEE Annals of the History of Computing, 42(2): 55–62, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bullynck:2015:CPT

[Bul15] Maarten Bullynck. Computing primes (1929–1949): Transformations in the early days of digital computing. IEEE Annals of the History of Computing, 37(3):44–54, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030044-abs.html.

Bullynck:2021:GEL

[Bul21] Maarten Bullynck. German encounters of logic and programming (1948–1958), featuring three readings of Turing machines. IEEE Annals of the History of Computing, 43 (4):10–26, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Burke:2002:FFS

[Bur02] Colin Burke. The Ford Foundation’s search for an American Library Laboratory. IEEE Annals of the History of Computing, 24(3):56–74, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a3056.pdf;   http://www.
                                        computer.org/annals/an2002/
                                        a3056abs.htm.

Burton:2005:RMB

[Bur05] Christopher P. Burton. Replicating the Manchester Baby: Motives, methods, and messages from the past. IEEE Annals of the History of Computing, 27(3):44–60, July/ September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Burke:2023:CBU

[Bur23] Caitlin Burke. The cognitive being as the user at Project Mac. IEEE Annals of the History of Computing, 45(2): 6–14, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Burrington:2024:WCO

[Bur24] Ingrid Burrington. From war crystals to ordinary sand: Excavating silicon supply chains.

IEEE Annals of the History of Computing, 46(2):13–23, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bonastre:2019:ODN

[BV19] O. M. Bonastre and A. Veà.

Origins of the Domain Name System. IEEE Annals of the History of Computing, 41(2): 48–60, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Buchholz:2000:WBC

[BW00] W. Buchholz and M. Wilkes.

Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4):69– 70, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887996.pdf.

Brooks:2022:DDS

[BYO^{+}22] Richard R. Brooks, Lu Yu, Ilker Ozcelik, Jon Oakley, and Nathan Tusing. Distributed denial of service (DDoS): a history. IEEE Annals of the History of Computing, 44(2):44– 54, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Caminer:2003:BCL

[Cam03a] David Tresman Caminer. Behind the curtain at LEO: a personal reminiscence. IEEE Annals of the History of Computing, 25(2):3–13, April/June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2003.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2003.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/02/a2003abs.
htm.

Campbell:2003:BHW

[Cam03b] Scott M. Campbell. Beatrice Helen Worsley: Canada’s female computer pioneer. IEEE Annals of the History of Computing, 25(4):51–62, October/ December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4051abs.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4051.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4051.pdf.

Campbell:2004:ES

[Cam04] Dan Campbell. Events & sightings. IEEE Annals of the History of Computing, 26(3):60–64, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/03/a3060.
htm;   http://csdl.computer.
org/dl/mags/an/2004/03/a3060.
pdf.

Campbell:2009:AO

[Cam09] Scott M. Campbell. On the absence of obsolescence. IEEE Annals of the History of Computing, 31(4):118–120, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell:2013:WES

[Cam13] Scott Campbell. ‘Wat For Ever’ student-oriented computing at the University of Waterloo. IEEE Annals of the History of Computing, 35(1):11– 22, January/March 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Caminer:2022:NBA

[Cam22] Hilary Caminer. A new book about the first business computer. IEEE Annals of the History of Computing, 44(4):86, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Carter:1992:CQD

[Car92] Robert Donald Carter. Comments, queries, and debate: German ferrite cores from World War II. IEEE Annals of the History of Computing, 14(3):8, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3006.pdf.

Carlson:1993:TIT

[Car93] Walter M. Carlson. Transforming an industry through information technology. IEEE Annals of the History of Computing, 15(1):39–43, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a1039.pdf;   http://www.
                                        computer.org/annals/an1993/
                                        a1039abs.htm.

Carnota:2015:BCS

[Car15] Raul Jorge Carnota. The beginning of computer science in Argentina and the Calculus Institute, 1957–1970. IEEE Annals of the History of Computing, 37(4):40–52, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Carpenter:2020:FCN

[Car20] B. E. Carpenter. The first computer in New Zealand. IEEE Annals of the History of Computing, 42(2):33–41, April/ June 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Carpenter:2023:OBD

[Car23] Brian E. Carpenter. Opening of the Bob Doran Museum of Computing. IEEE Annals of the History of Computing, 45(4):82, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Castleman:2006:MAC

[Cas06] Paul Castleman. Medical application of computers at BBN. IEEE Annals of the History of Computing, 28(1):6–16, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Collopy:2019:RRP

[CB19] P. S. Collopy and M. J.

Baranay. Reviews of recent publications [two books reviewed]. IEEE Annals of the History of Computing, 41(2): 67–70, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Croarken:2000:BNR

[CCK00] Mary Croarken and Martin Campbell-Kelly. Beautiful numbers: the rise and decline of the British Association Mathematical Tables Committee, 1871–1965. IEEE Annals of the History of Computing, 22(4):44–61, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4044.pdf;   http://ieeexplore.
ieee.org/iel5/85/19196/00887989.
pdf;   http://www.computer.
org/annals/an2000/a4044abs.
htm.

Carpenter:2014:JWA

[CD14] Brian E. Carpenter and Robert W. Doran. John Womersley: Applied mathematician and pioneer of modern computing. IEEE Annals of the History of Computing, 36(2):60–70, April/June 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cukierman:2022:WBU

[CdLF^{+}22] Henrique Luiz Cukierman, Alberto Jorge Silva de Lima, Paulo Henrique Fidelis Feitosa, Fernando Gonçalves Severo, Ricardo Jullian da Silva Graça, and André Vinícius Leal Sobral. What1468371Whom is the Brazilian university for? The case of the Computing Projects Laboratory. IEEE Annals of the History of Computing, 44(3):17–29, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Croarken:2002:ES

[CE02] Mary Croarken and Nathan Ensmenger. Events & sightings. IEEE Annals of the History of Computing, 24(4):56– 59, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

computer.org/annals/an2002/
extras/a405602x.htm;   http:
//csdl.computer.org/dl/mags/
an/2002/04/a4056.htm;   http:
//csdl.computer.org/dl/mags/
an/2002/04/a4056.pdf.

Croarken:2003:ESa

[CE03a] Mary Croarken and Nathan Ensmenger. Events and sightings. IEEE Annals of the History of Computing, 25(1):65–69, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2003/extras/a1065.
htm;   http://csdl.computer.
org/dl/mags/an/2003/01/a1065.
htm;   http://csdl.computer.
org/dl/mags/an/2003/01/a1065.
pdf.

Croarken:2003:ESb

[CE03b] Mary Croarken and Nathan Ensmenger. Events and sightings. IEEE Annals of the History of Computing, 25(2):76–79, April/June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2076.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2076.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/02/a2076abs.
htm.

Croarken:2003:ESc

[CE03c] Mary Croarken and Nathan Ensmenger. Events and sightings. IEEE Annals of the History of Computing, 25(3):78–81, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2003/03/a3078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2003/03/a3078.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/an/2003/03/a3078abs.
                                        htm.

Croarken:2003:ESd

[CEG^{+}03] Mary Croarken, Nathan Ensmenger, David Alan Grier, Frederik Nebeker, and David Anderson. Events & sightings. IEEE Annals of the History of Computing, 25 (4):73–77, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2003/04/a4073.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2003/04/a4073.
                                        pdf.

Ceruzzi:1992:Ra

[Cer92a] Paul Ceruzzi. Reviews. IEEE Annals of the History of Computing, 14(1):69–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a1069.pdf.

Ceruzzi:1992:Rc

[Cer92b] Paul Ceruzzi. Reviews. IEEE Annals of the History of Computing, 14(3):77–??, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3077.pdf.

Ceruzzi:1992:Rd

[Cer92c] Paul Ceruzzi. Reviews. IEEE Annals of the History of Computing, 14(4):76–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4076.pdf.

Ceruzzi:1993:R

[Cer93] Paul Ceruzzi. Reviews. IEEE Annals of the History of Computing, 15(1):70–??, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a1070.pdf.

Ceruzzi:1997:CDA

[Cer97] P. Ceruzzi. Crossing the divide: architectural issues and the emergence of the stored program computer, 1935–1955. IEEE Annals of the History of Computing, 19(1): 5–12, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560724.pdf.

Ceruzzi:2001:VYH

[Cer01] Paul E. Ceruzzi. A view from 20 years as a historian of computing. IEEE Annals of the History of Computing, 23(4):49–55, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a4049abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a4049.
pdf.

Ceruzzi:2002:DAS

[Cer02] Paul Ceruzzi. Datamation, Annals, Slashdot, and tomorrow’s history. IEEE Annals of the History of Computing, 24 (4):72, 71, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4072.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4072.
pdf.

Ceruzzi:2003:BRD

[Cer03] Paul E. Ceruzzi. Book review: Diana H. Hook and Jeremy M. Norman, with contributions by Michael R. Williams, Origins of Cyberspace: A Library History of Computing, Networking, and Telecommunications, historyofscience.com, 2002, ISBN 0-930405-85-4, 670 pp. IEEE Annals of the History of Computing, 25(4):93, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ceruzzi:2006:MI

[Cer06] Paul E. Ceruzzi. The materiality of the Internet. IEEE Annals of the History of Computing, 28(3):96, c3, July/ September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ceruzzi:2014:HFT

[Cer14] P. E. Ceruzzi. Are historians failing to tell the real story about the history of computing? IEEE Annals of the History of Computing, 36(3):94– 95, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ceruzzi:2025:RKZ

[Cer25]  Paul Ceruzzi.  Review of

Konrad Zuse’s Early Computers: The Quest for the Computer in Germany, by Raúl Rojas. IEEE Annals of the History of Computing, 47(2): 60–61, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Crawford:1992:ASA

[CF92] David J. Crawford and Philip E.

Fox. The Autoscritcher and the Superscritcher: aids to cryptanalysis of the German Enigma cipher machine, 1944– 6. IEEE Annals of the History of Computing, 14(3):9–22, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3009.pdf;   http://www.
computer.org/annals/an1992/
a3009abs.htm.

Ceruzzi:1992:Rb

[CFSL92] Paul Ceruzzi, P. Forman, K. W. Smillie, and J. A. N. Lee. Reviews. IEEE Annals of the History of Computing, 14(2):60–63, April/ June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an1992/pdf/a2060.
pdf.

Crowe:1994:LBS

[CG94] Gregory D. Crowe and Seymour E. Goodman. S. A. Lebedev and the birth of Soviet computing. IEEE Annals of the History of Computing, 16 (1):4–24, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1004.pdf;   http://www.
computer.org/annals/an1994/
a1004abs.htm.

Ceruzzi:2006:GEI

[CG06] Paul E. Ceruzzi and Burton Grad. Guest Editors’ introduction: PC software– word processing for everyone. IEEE Annals of the History of Computing, 28(4):4–5, October/December 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ceruzzi:2007:GEI

[CG07] Paul Ceruzzi and Burton Grad. Guest Editors’ introduction: PC software — spreadsheets for everyone. IEEE Annals of the History of Computing, 29(3):4–5, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2007/03/
man2007030004.pdf.

Cignoni:2020:PAK

[CG20] G. A. Cignoni and F. Gadducci. Pisa, 1954–1961: Assessing key stages of a seminal Italian project. IEEE Annals of the History of Computing, 42 (2):6–19, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Crowther-Heyck:2005:MN

[CH05] Hunter Crowther-Heyck. Mind and network. IEEE Annals of the History of Computing, 27 (3):104, 103, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2015:AVH

[CH15] J. Copeland and A. A. Haeff.

Andrew V. Haeff: Enigma of the tube era and forgotten computing pioneer. IEEE Annals of the History of Computing, 37(1):67–74, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Chamberlin:2012:EHS

[Cha12] Donald D. Chamberlin. Early history of SQL. IEEE Annals of the History of Computing, 34 (4):78–82, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Chen:2025:MNT

[Che25] Mark Yu-Shan Chen. Mechanical numeracy: Thinking with the Soroban abacus in Modern Japan. IEEE Annals of the History of Computing, 47(1): 36–49, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2017:SHH

[CHGW17] B. Jack Copeland, Andre A. Haeff, Peter Gough, and Cameron Wright. Screen history: The Haeff memory and graphics tube. IEEE Annals of the History of Computing, 39(1):9–28, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/01/
                                        man2017010009-abs.html.

Chiappa:2014:ETR

[Chi14] J. Noel Chiappa. Early token ring work at MIT. IEEE Annals of the History of Computing, 36(2):80–85, April/June 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Choi:2016:HCE

[Cho16] Hyungsub Choi. History of computing in East Asia. IEEE Annals of the History of Computing, 38(2):8–10, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:1992:ATE

[CK92a] Martin Campbell-Kelly. The Airy tape: An early chapter in the history of debugging. IEEE Annals of the History of Computing, 14(4):16–26, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4016.pdf;   http://www.
computer.org/annals/an1992/
a4016abs.htm.

Campbell-Kelly:1992:ICU

[CK92b] Martin Campbell-Kelly. Introduction: Computing at the University of Cambridge. IEEE Annals of the History of Computing, 14(4):8–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4008.pdf.

Campbell-Kelly:1992:PEE

[CK92c] Martin Campbell-Kelly. Programming the EDSAC: early programming activity at the University of Cambridge. IEEE Annals of the History of Computing, 14(4):46–67, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a4008.pdf.

Campbell-Kelly:1994:CBA

[CK94] Martin Campbell-Kelly. Charles Babbage and the assurance of lives. IEEE Annals of the History of Computing, 16(3): 5–14, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3005.pdf;   http://www.
                                        computer.org/annals/an1994/
                                        a3005abs.htm.

Campbell-Kelly:1997:ICS

[CK97] M. Campbell-Kelly. The international computer software industry: a comparative study of industry evolution and structure [book reviews]. IEEE Annals of the History of Computing, 19(1):79, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/12228/00560750.pdf.

Campbell-Kelly:1998:PEE

[CK98] M. Campbell-Kelly. Programming the EDSAC: early programming activity at the University of Cambridge. IEEE Annals of the History of Computing, 20(4):46–67, October/December 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728229.pdf.

Campbell-Kelly:2000:CBE

[CK00] Martin Campbell-Kelly. Charles Babbage eulogy. IEEE Annals of the History of Computing, 22(4):20–21, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4020.pdf;   http://ieeexplore.
ieee.org/iel5/85/19196/00887987.
pdf.

Campbell-Kelly:2001:CBE

[CK01a] M. Campbell-Kelly. Charles Babbage and the engines of perfection [reviews]. IEEE Annals of the History of Computing, 23(3):78, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3078abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a3078.
pdf.

Campbell-Kelly:2001:CBI

[CK01b] Martin Campbell-Kelly. The Charles Babbage Institute reprint series for the history of computing. IEEE Annals of the History of Computing, 23(4):44–48, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2001/a4044abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2001/pdf/a4044.
                                        pdf.

Campbell-Kelly:2002:TP

[CK02a] Martin Campbell-Kelly. Think piece. IEEE Annals of the History of Computing, 24(1): 96, 95, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        computer.org/annals/an2002/
                                        a1096abs.htm;   http://dlib.
                                        computer.org/an/books/an2002/
                                        pdf/a1096.pdf.

Campbell-Kelly:2002:TPC

[CK02b] Martin Campbell-Kelly. Twin pillars of computing. IEEE Annals of the History of Computing, 24(3):88, 87, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a3088.pdf.

Campbell-Kelly:2007:HHS

[CK07a] Martin Campbell-Kelly. The history of the history of software. IEEE Annals of the History of Computing, 29(4):40– 51, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:2007:NCP

[CK07b] Martin Campbell-Kelly. Number crunching without programming: The evolution of spreadsheet usability. IEEE Annals of the History of Computing, 29(3):6–19, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Choi:2008:HWP

[CK08] Hyungsub Choi and Chigusa Kita. Hiroshi Wada: Pioneering electronics and computer technologies in postwar Japan. IEEE Annals of the History of Computing, 30(3): 84–89, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:2012:RIN

[CK12] Martin Campbell-Kelly. The RDBMS industry: A Northern California perspective. IEEE Annals of the History of Computing, 34(4):18–29, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:2013:RMM

[CK13] Martin Campbell-Kelly. Remembering Michael S. Mahoney. Perspectives on Science, 21(3):379–383, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:

                                        //muse.jhu.edu/journals/
                                        perspectives_on_science/
                                        v021/21.3.campbell-kelly.
                                        html;   http://muse.jhu.
                                        edu/journals/perspectives_
                                        on_science/v021/21.3.campbell-
                                        kelly.pdf.

Campbell-Kelly:2014:KSH

[CK14] M. Campbell-Kelly. Knuth and the spectrum of history. IEEE Annals of the History of Computing, 36(3):96, July/ September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Campbell-Kelly:2008:EPH

[CKGS08] Martin Campbell-Kelly and Daniel D. Garcia-Swartz. Economic perspectives on the history of the computer timesharing industry, 1965–1985. IEEE Annals of the History of Computing, 30(1):16–36, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Clark:2010:SSE

[Cla10] Martyn Clark. State support for the expansion of UK university computing in the 1950s. IEEE Annals of the History of Computing, 32(1):23–33, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Clymer:1993:MAC

[Cly93] A. Ben Clymer. The mechanical analog computers of Hannibal Ford and William Newell. IEEE Annals of the History of Computing, 15(2):19–34, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2019.pdf;   http://www.
computer.org/annals/an1993/
a2019abs.htm.

Carpenter:2025:ITP

[CM25] Brian E. Carpenter and Sathiamoorthy Manoharan. Information technology pioneers of Aotearoa New Zealand. IEEE Annals of the History of Computing, 47(2):44–58, April/ June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:1995:SDI

[CMM95] D. G. Copeland, R. O. Mason, and J. L. McKenney. Sabre: the development of information-based competence and execution of informationbased competition. IEEE Annals of the History of Computing, 17(3):30–57, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397059.pdf.

Choi:2012:FLT

[CO12] Hyungsub Choi and Takushi Otani. Failure to launch: Tarui Yasuo, the quadrupole transistor, and the meanings of the IC in postwar Japan. IEEE Annals of the History of Computing, 34(1):48–59, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cobb:2023:RKR

[Cob23] Matthew Cobb. The representation of knowledge and the relevance of biological models at the Symposium on the Mechanization of Thought Processes, 1958. IEEE Annals of the History of Computing, 45(3):32– 47, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cohen:1998:HAN

[Coh98] I. B. Cohen. Howard Aiken on the number of computers needed for the nation. IEEE Annals of the History of Computing, 20(3): 27–32, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/15262/00707572.pdf.

Cohn:2015:TAD

[Coh15] Julie Cohn. Transitions from analog to digital computing in electric power systems. IEEE Annals of the History of Computing, 37(3):32–43, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030032-abs.html.

Collopy:2014:VSA

[Col14] P. S. Collopy. Video synthesizers: From analog computing to digital art. IEEE Annals of the History of Computing, 36 (4):74–86, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

ConDiaz:2015:ESP

[Con15] Gerardo Con Diaz. Embodied software: Patents and the history of software development, 1946–1970. IEEE Annals of the History of Computing, 37(3):8–19, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030008-abs.html.

ConDiaz:2017:R

[Con17] Gerardo Con Dıaz. Reviews.

IEEE Annals of the History of Computing, 39(2):9–11, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

ConDiaz:2019:ED

[Con19] G. Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 41 (4):5, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

ConDiaz:2020:EDa

[Con20a] Gerardo Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 42(1):5, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

ConDiaz:2020:EDb

[Con20b] Gerardo Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 42 (2):5, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Coopersmith:2002:MP

[Coo02] Jonathan Coopersmith. Membership has its partisans. IEEE Annals of the History of Computing, 24(2):104, 103, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a2104.pdf;   http://www.
                                        computer.org/annals/an2002/
                                        a2104abs.htm.

Copeland:2004:COO

[Cop04a] B. Jack Copeland. Colossus: Its origins and originators. IEEE Annals of the History of Computing, 26 (4):38–45, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/04/a4038.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4038.
pdf.

Copeland:2004:UA

[Cop04b] B. Jack Copeland. Unfair to Aiken. IEEE Annals of the History of Computing, 26 (4):35–37, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/04/a4035.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4035.
pdf.

Copeland:2011:MCRa

[Cop11a] B. Jack Copeland. The Manchester computer: a revised history part 1: The memory. IEEE Annals of the History of Computing, 33(1):4–21, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2011:MCRb

[Cop11b] B. Jack Copeland. The Manchester computer: a revised history part 2: The baby computer. IEEE Annals of the History of Computing, 33(1):22– 37, January/March 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Copeland:2023:EAB

[Cop23]  B. Jack Copeland.  Early

AI in Britain: Turing et al. IEEE Annals of the History of Computing, 45(3):19–31, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:1996:CAD

[Cor96] J. W. Cortada. Commercial applications of the digital computer in American corporations, 1945–1995. IEEE Annals of the History of Computing, 18(2):18–29, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489721.pdf.

Cortada:1997:EPM

[Cor97] J. W. Cortada. Economic preconditions that made possible application of commercial computing in the United States. IEEE Annals of the History of Computing, 19(3): 27–40, July/September 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13172/00601734.pdf.

Cortada:1998:TCC

[Cor98] J. W. Cortada. The timetable of computers: a chronology of the most important people and events in the history of computers [review]. IEEE Annals of the History of Computing, 20(1):82–83, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646213.pdf.

Cortada:2001:TP

[Cor01a] James Cortada. Think piece.

IEEE Annals of the History of Computing, 23(2):88–??, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2088.pdf.

Cortada:2001:UTD

[Cor01b]  James W. Cortada.  Us-

ing textual demographics to understand computer use: 1950–1990. IEEE Annals of the History of Computing, 23(1):34–56, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1034.pdf;   http://www.
computer.org/annals/an2001/
a1034abs.htm.

Cortada:2002:RHS

[Cor02]  James W. Cortada.  Re-

searching the history of software from the 1960s. IEEE Annals of the History of Computing, 24(1):72–79, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2002/a1072abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2002/pdf/a1072.
                                        pdf.

Cortada:2004:HDC

[Cor04] James W. Cortada. How did computing go global? The need for an answer and a research agenda. IEEE Annals of the History of Computing, 26(1):53–58, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2004/01/a1053abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/01/a1053.
                                        pdf.

Cortada:2006:EIB

[Cor06a]  James W. Cortada.  The

ENIAC’s influence on business computing, 1940s–1950s. IEEE Annals of the History of Computing, 28(2):26–28, April/ June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2006:LTT

[Cor06b] James W. Cortada. Looking at technology through industry eyes. IEEE Annals of the History of Computing, 28 (2):88, April/June 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2007:SRI

[Cor07] James W. Cortada. Studying the role of IT in the evolution of American business practices: a way forward. IEEE Annals of the History of Computing, 29 (4):28–39, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Corry:2008:FMS

[Cor08a] Leo Corry. Fermat meets SWAC: Vandiver, the Lehmers, computers, and number theory. IEEE Annals of the History of Computing, 30(1):38– 49, January/March 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2008:PPH

[Cor08b] James W. Cortada. Patterns and practices in how information technology spread around the world. IEEE Annals of the History of Computing, 30(4): 4–25, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2009:TPP

[Cor09] James W. Cortada. Think piece: Preserving records of the past, today. IEEE Annals of the History of Computing, 31(2):87–89, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2012:ITG

[Cor12] James W. Cortada. Information technologies in the German Democratic Republic (GDR), 1949–1989. IEEE Annals of the History of Computing, 34(2):34–48, April/June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2013:XCB

[Cor13] James W. Cortada. “Carrying a Bag”: Memoirs of an IBM salesman, 1974–1981. IEEE Annals of the History of Computing, 35(4):32–47, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2014:WSI

[Cor14] James W. Cortada. War stories of an IBM salesman, 1974– 1981. IEEE Annals of the History of Computing, 36(1):74– 78, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2015:SNS

[Cor15a] J. W. Cortada. “There is no saturation point in education”: Inside IBM’s Sales School, 1970s–1980s. IEEE Annals of the History of Computing, 37(1):56–66, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2015:DHO

[Cor15b] James W. Cortada. “Doing History” in the olden days, 1970s–1990s. IEEE Annals of the History of Computing, 37(2):84–88, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//www.computer.org/csdl/
mags/an/2015/02/man2015020084-
abs.html.

Cortada:2015:SHI

[Cor15c] James W. Cortada. Studying history as it unfolds, Part 1: Creating the history of information technologies. IEEE Annals of the History of Computing, 37(3):20–31, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030020-abs.html.

Cortada:2016:SHI

[Cor16] James W. Cortada. Studying history as it unfolds, part 2: Tooling up the historians. IEEE Annals of the History of Computing, 38(1):48– 59, January/March 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2017:IBO

[Cor17]  James W. Cortada.  IBM

branch offices: What they were, how they worked, 1920s– 1980s. IEEE Annals of the History of Computing, 39(3):9–23, July/September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/03/
                                        man2017030009-abs.html.

Cortada:2018:SIH

[Cor18] James W. Cortada. Studying IBM’s history when you are part of it. IEEE Annals of the History of Computing, 40 (4):64–68, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8511045/.

Cortada:2019:RWHa

[Cor19a] J. W. Cortada. Rethinking why and how organizations acquire information technologies, part 1. IEEE Annals of the History of Computing, 41(4):48–60, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See correction [Cor20].

Cortada:2019:RWHb

[Cor19b] J. W. Cortada. Rethinking why and how organizations acquire information technologies, part 2. IEEE Annals of the History of Computing, 41(4):61–70, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2020:CRW

[Cor20] J. W. Cortada. Correction to Rethinking Why and How Organizations Acquire Information Technologies, Part 1: Conventional Explanations and Their Limits. IEEE Annals of the History of Computing, 42(4):127, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See [Cor19a].

Cortada:2021:LPC

[Cor21a] J. W. Cortada. From lapel pins to coffee cups: Links between corporate and material culture, lessons from IBM. IEEE Annals of the History of Computing, 43(1):74–88, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2021:PPI

[Cor21b] James W. Cortada. Preserving the past by industry participants. IEEE Annals of the History of Computing, 43(2): 42–44, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2022:WHS

[Cor22] James W. Cortada. Why historians should pay more attention to the social histories of objects and what they can learn from these. IEEE Annals of the History of Computing, 44(3):87– 90, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Corry:2024:CPC

[Cor24a]  Frances Corry.  Computer

pasts/computer futures: April 4 2024. IEEE Annals of the History of Computing, 46(2): 89–90, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2024:HIB

[Cor24b] James W. Cortada. A hidden IT “Best seller”: IBM customer engineering maintenance manuals, 1930s–1980s. IEEE Annals of the History of Computing, 46(1):48–58, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2025:PID

[Cor25a] James W. Cortada. Presence of IBM data processing equipment in Latin America, 1920s– 1980s, and insights for historians. IEEE Annals of the History of Computing, 47(1):50– 63, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cortada:2025:TCC

[Cor25b] James W. Cortada. Twentieth Century computer product proposals: a wealth of information about information technologies and their uses. IEEE Annals of the History of Computing, 47(2):33–43, April/June 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Costello:1996:TBE

[Cos96]  J. Costello.  As the twig

is bent: the early life of John Mauchly. IEEE Annals of the History of Computing, 18(1):45–50, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476560.pdf.

Couleur:1995:CBC

[Cou95]  J. Couleur.  The core of

the Black Canyon Computer Corporation. IEEE Annals of the History of Computing, 17(4):56–60, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477436.pdf.

Cowlishaw:1994:EHR

[Cow94] Mike Cowlishaw. The early history of REXX. IEEE Annals of the History of Computing, 16(4):15–24, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an1994/pdf/a4015.
pdf;   http://www.computer.
org/annals/an1994/a4015abs.
htm.

Cignoni:2023:TOF

[CP23] Giovanni A. Cignoni and Sergei P. Prokhorov. Tracing the origins of the first Soviet computers, beyond legends. IEEE Annals of the History of Computing, 45(4):85– 91, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Coghlan:2024:HPL

[CRB24] Brian Coghlan, Brian Randell, and Ralf Buelow. How Percy Ludgate’s 1909 paper (and IBM) helped thwart Konrad Zuse’s computer patent in 1960. IEEE Annals of the History of Computing, 46(3): 20–35, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https:

//treasures.scss.tcd.ie/
ludgate/.

Coghlan:2021:IWL

[CRH^{+}21] B. Coghlan, B. Randell, P. Hockie, T. Gonzalez, D. Mc-Quillan, and R. O’Regan. Investigating the work and life of Percy Ludgate. IEEE Annals of the History of Computing, 43(1):19–37, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Croarken:2004:ES

[CRM^{+}04] Mary Croarken, Dorian Rutter, Tony Mann, Alex Fedorec, and Brian Wichmann. Events & sightings. IEEE Annals of the History of Computing, 26(1): 68–70, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/01/a1068.pdf.

Croarken:1992:ECS

[Cro92] Mary G. Croarken. The emergence of computing science research and teaching at Cambridge, 1936–1949. IEEE Annals of the History of Computing, 14(4):10–15, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4010.pdf;   http://www.
computer.org/annals/an1992/
a4010abs.htm.

Croarken:1993:BMC

[Cro93]  Mary Croarken.  The be-

ginnings of the Manchester computer phenomenon: People and influences. IEEE Annals of the History of Computing, 15(3):9–16, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a3009.pdf;   http://www.
                                        computer.org/annals/an1993/
                                        a3009abs.htm.

Croarken:1999:CJC

[Cro99] M. Croarken. Case 5,656: L. J. Comrie and the origins of the Scientific Computing Service Ltd. IEEE Annals of the History of Computing, 21 (4):70–71, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17389/00801536.pdf.

Croarken:2000:DHL

[Cro00a]  M. Croarken.  Dorothy

Hodgkin: a life. IEEE Annals of the History of Computing, 22(4):65, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/19196/00887992.pdf.

Croarken:2000:HAP

[Cro00b] M. Croarken. Howard Aiken: portrait of a computer pioneer. IEEE Annals of the History of Computing, 22 (1):89, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17665/00815471.pdf.

Croarken:2003:MEC

[Cro03a] Mary Croarken. Mary Edwards: Computing for a living in 18th-Century England. IEEE Annals of the History of Computing, 25(4):9–15, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4009abs.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4009.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4009.pdf.

Croarken:2003:THC

[Cro03b] Mary Croarken. Tabulating the heavens: Computing the Nautical Almanac in 18th-Century England. IEEE Annals of the History of Computing, 25(3):48–61, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3048.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3048.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/03/a3048abs.
htm.

Crooks:2018:CFC

[Cro18]  Roderic Crooks.  Critical

failure: Computer-aided instruction and the fantasy of information. IEEE Annals of the History of Computing, 40(2):85–88, ????

2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8415738/.

Crocker:2019:LN

[Cro19] S. D. Crocker. Learning to network. IEEE Annals of the History of Computing, 41(2): 42–47, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cervantes-Villagomez:2015:SSM

[CVRM15] Ofelia Cervantes-Villagomez and Leandro Rodriguez-Medina. Small step for machines, giant leap for Mexico: A local history of computing. IEEE Annals of the History of Computing, 37(4):14–28, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

daCostaMarques:2015:BCM

[da 15a] Ivan da Costa Marques. Brazil’s computer market reserve: Democracy, authoritarianism, and ruptures. IEEE Annals of the History of Computing, 37(4):64–75, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

daCostaMarques:2015:HCL

[da 15b] Ivan da Costa Marques. History of computing in Latin America [Guest Editors’ introduction]. IEEE Annals of the History of Computing, 37 (4):10–12, ???? 2015. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durnova:2014:TPW

[DA14] H. Durnova and G. Alberts.

Think piece: Was Algol 60 the first algorithmic language? IEEE Annals of the History of Computing, 36(4):104, 102– 103, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

delAguila:2024:MIM

[dÁ24] Verónica Uribe del Águila.

Making innovation in the Mexican Silicon Valley: The early years of El Centro de Tecnología de Semiconductores (1981–2001). IEEE Annals of the History of Computing, 46(2):24–37, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

DominguezNava:2015:BHC

[DAGVR^{+}15] Jaquelinne Dominguez Nava, Juan C. Acosta-Guadarrama, Rosa M. Valdovinos-Rosas, Victor H. Solis Ramos, Nely Plata Cesar, Leticia Quintanar Rebollar, and Rogelio Davila Perez. A brief history of computing in Mexico. IEEE Annals of the History of Computing, 37(4):76–86, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Darwen:2012:RMB

[Dar12] Hugh Darwen. The relational model: Beginning of an era. IEEE Annals of the History of Computing, 34(4):30–37, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durno:2025:BBS

[DAT25] John Durno, John Aycock, and Alexandra Tenney. Big baggy shorts: Recovering data from Canada’s Telidon System. IEEE Annals of the History of Computing, 47(1):64– 71, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Davies:2017:ECP

[Dav17] Stephen M. Davies. The experimental computer programme: The first computing initiative for the National Health Service in England. IEEE Annals of the History of Computing, 39 (2):65–79, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

                                        URL http://ieeexplore.
                                        ieee.org/document/7974720/

.

Day:2016:COI

[Day16] John Day. The clamor outside as INWG debated: Economic war comes to networking. IEEE Annals of the History of Computing, 38(3):58–77, ???? 2016.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Daylight:2021:AQW

[Day21] Edgar G. Daylight. Addressing the question what is a program text? via Turing scholarship. IEEE Annals of the History of Computing, 43(4):87–91, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dayasindhu:2023:WIP

[Day23] N. Dayasindhu. Wipro Infotech’s pivot from computer hardware to R&D services for IT. IEEE Annals of the History of Computing, 45(1):6– 17, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dayasindhu:2024:OTE

[Day24] N Dayasindhu. The origins of training and education in the Indian information technology sector. IEEE Annals of the History of Computing, 46(3):7– 19, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

deSollaPrice:1959:AGC

[de 59] Derek J. de Solla Price. An ancient Greek computer. Scientific American, 200(6):60– 67, June 1959. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.
com/scientificamerican/journal/
v200/n6/pdf/scientificamerican0659-
60.pdf.

deSollaPrice:1974:GGA

[de 74] Derek John de Solla Price. Gears from the Greeks: the Antikythera mechanism, a calendar computer from ca. 80 B.C. Transactions of the American Philosophical Society, New Series, 64(7):1–70, November 1974. CODEN TAPSAY. ISBN 0-87169-647-9. ISSN 0065-9746. URL http://www.

jstor.org/stable/1006146.

DeWit:1995:CBH

[De 95]  D. De Wit.  Caught be-

tween historical experience and high hopes: automation at the Dutch Postal Cheque and Clearing Service, 1950–1965. IEEE Annals of the History of Computing, 17(2):9–21, Summer 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8635/00380268.pdf.

DeMol:2021:LPC

[De 21]  Liesbeth De Mol.  Logic,

programming, and computer science: Local perspectives. IEEE Annals of the History of Computing, 43(4):5–9, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Delve:2008:TML

[Del08] Janet Delve. A trip down memory lane? New challenges from other disciplines. IEEE Annals of the History of Computing, 30 (4):90–92, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Denning:2001:OVM

[Den01] Peter J. Denning. Origin of virtual machines and other virtualities. IEEE Annals of the History of Computing, 23(3):73, July/September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3073abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3073.
pdf.

DeNardis:2015:IDT

[DeN15a] Laura DeNardis. The Internet design tension between surveillance and security. IEEE Annals of the History of Computing, 37(2):72– 83, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020072-abs.html.

Dennis:2015:MLC

[Den15b] Jack B. Dennis. Machines, languages, and computation at MIT. IEEE Annals of the History of Computing, 37(3):76–77, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030076-abs.html.

Deringer:2020:MMS

[Der20] W. Deringer. Michael Milken’s spreadsheets: Computation and charisma in finance in the go-go 80s. IEEE Annals of the History of Computing, 42(3): 53–69, July/September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Deutsch:2013:SSH

[Deu13]  Donald R. Deutsch.  The

SQL standard: How it happened. IEEE Annals of the History of Computing, 35(2): 72–75, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

DuPont:2016:ECC

[DF16] Quinn DuPont and Bradley Fidler. Edge cryptography and the codevelopment of computer networks and cybersecurity. IEEE Annals of the History of Computing, 38(4): 55–73, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dujnic:1999:HCS

[DFM^{+}99] J. Dujnic, N. Fristacky, L. Molnar, I. Plander, and B. Rovan. On the history of computer science, computer engineering, and computer technology development in Slovakia. IEEE Annals of the History of Computing, 21(3): 38–48, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778981.pdf.

Dongarra:2008:NNB

[DGG^{+}08] Jack Dongarra, Gene H. Golub, Eric Grosse, Cleve Moler, and Keith Moore. Netlib and NA-Net: Building a scientific computing community. IEEE Annals of the History of Computing, 30(2):30–41, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=4544554.

Dongarra:2008:B

[DH08] Jack Dongarra and Thomas Haigh. Biographies. IEEE Annals of the History of Computing, 30(2):74–81, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Diaz:2012:OHA

[Dia12] Gerardo Con Diaz. Ownership and the history of American computing. IEEE Annals of the History of Computing, 34 (2):88, 86–87, April/June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Diaz:2016:COS

[Dia16] Gerardo Con Diaz. Contested ontologies of software: The story of Gottschalk v. Benson, 1963–1972. IEEE Annals of the History of Computing, 38(1): 23–33, January/March 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Diaz:2019:EDa

[Dia19a] Gerardo Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 41(1):5–6, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8667951/.

Diaz:2019:EDb

[Dia19b] Gerardo Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 41(2):5, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Diaz:2021:ED

[Dia21] Gerardo Con Diaz. From the Editor’s desk. IEEE Annals of the History of Computing, 43(3):5, July/September 2021.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dick:2013:MWW

[Dic13] Stephanie Dick. Machines who write [think piece]. IEEE Annals of the History of Computing, 35(2):88, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dzardanova:2023:VRJ

[DK23] Elena Dzardanova and Vlasios Kasapakis. Virtual reality: a journey from vision to commodity. IEEE Annals of the History of Computing, 45(1): 18–30, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

DeMol:2018:LMF

[DMD18] Liesbeth De Mol, Bullynck Maarten, and Edgar G. Daylight. Less is more in the fifties: Encounters between logical minimalism and computer design during the 1950s. IEEE Annals of the History of Computing, 40(1):19–45, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010019.pdf.

DeMarco:1999:ECI

[DMPS99] G. De Marco, G. Mainetto, S. Pisani, and P. Savino.

The early computers of Italy. IEEE Annals of the History of Computing, 21(4):28– 36, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801530.pdf.

Dritsa:2018:AHC

[DMS18] Konstantina Dritsa, Dimitris Mitropoulos, and Diomidis Spinellis. Aspects of the history of computing in modern Greece. IEEE Annals of the History of Computing, 40(1):47–60, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010047.pdf.

Dodge:2000:CBE

[Dod00] Nathaniel Shatswell Dodge. Charles Babbage eulogy (from the Smithsonian annual report of 1873). IEEE Annals of the History of Computing, 22(4):22–43, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4022.pdf;   http://ieeexplore.
ieee.org/iel5/85/19196/00887988.
pdf;   http://www.computer.
org/annals/an2000/a4022abs.
htm.

Donig:2010:AAT

[Don10] Simon Donig. Appropriating American technology in the 1960s: Cold War politics and the GDR computer industry. IEEE Annals of the History of Computing, 32(2): 32–45, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dornian:1994:RTC

[Dor94]  Alan Dornian.  ReserVec:

Trans-Canada Air Lines’ computerized reservation system. IEEE Annals of the History of Computing, 16(2):31– 42, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a2031.pdf;   http://www.
computer.org/annals/an1994/
a2031abs.htm.

Downey:2004:JCS

[Dow04] Greg Downey. Jumping contexts of space and time. IEEE Annals of the History of Computing, 26(2):96, 94–95, April/ June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/02/a2096.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2096.
pdf.

Downey:2009:RCH

[Dow09] Greg Downey. Reviews: a complex history of the commercial Internet. IEEE Annals of the History of Computing, 31(2): 80–81, April/June 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Doyle:2001:UNF

[Doy01] Robert Doyle. The US Navy’s first online crypto system. IEEE Annals of the History of Computing, 23(1):17–21, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1017.pdf;   http://www.
computer.org/annals/an2001/
a1017abs.htm.

dePosada:2021:LTQ

[dPRG21] Francisco González de Posada, Francisco A. González Redondo, and Alfonso Hernando González. Leonardo Torres Quevedo: Pioneer of computing, automatics, and artificial intelligence. IEEE Annals of the History of Computing, 43(3):22–43, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Drachmann:1976:BRB

[Dra76a] A. G. Drachmann. Book review: Gears from the Greeks: The Antikythera Mechanism: A Calendar Computer from ca. 80 B.C. by Derek de Solla Price. Technology and Culture, 17(1):112–116, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://www.
jstor.org/stable/3103259.

Drake:1976:GFM

[Dra76b]  Stillman Drake.  Galileo

and the first mechanical computing device. Scientific American, 234(4):104– 113, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.
com/scientificamerican/journal/
v234/n4/pdf/scientificamerican0476-
104.pdf.

Drane:2018:HAH

[Dra18] Douglas Drane. How Atex helped an industry change the world. IEEE Annals of the History of Computing, 40(3):25–29, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509527/.

Dryer:2018:AUR

[Dry18] Theodora Dryer. Algorithms under the reign of probability. IEEE Annals of the History of Computing, 40(1):93–96, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010093.pdf.

Di:1992:JPB

[DSW92] Li Di, Bai Shangshu, and Michael R. Williams. John E. Parker (biographies). IEEE Annals of the History of Computing, 14(4):67–75, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Denning:2021:MLE

[DT21] Peter J. Denning and Matti Tedre. Mathematics, logic, and engineering in computing. IEEE Annals of the History of Computing, 43(4):97–100, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dumey:1994:CQD

[Dum94]  Arnold I. Dumey.  Com-

ments, queries, and debate: Scritchers. IEEE Annals of the History of Computing, 16 (3):4, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3004.pdf.

Dunning:2021:WWP

[Dun21] David E. Dunning. The work of writing programs: Logic and inscriptive practice in the history of computing. IEEE Annals of the History of Computing, 43(4):27–42, October/ December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durnova:2010:SCC

[Dur10] Helena Durnova. Sovietization of Czechoslovakian computing: The rise and fall of the SAPO project. IEEE Annals of the History of Computing, 32(2): 21–31, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durnova:2014:EAE

[Dur14] H. Durnova. Embracing the Algol effort in Czechoslovakia. IEEE Annals of the History of Computing, 36(4):26–37, October/December 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dick:2018:DHS

[DV18] Stephanie Dick and Daniel Volmar. DLL Hell: software dependencies, failure, and the maintenance of Microsoft Windows. IEEE Annals of the History of Computing, 40 (4):28–51, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509170/.

Dionne:2020:II

[DW20] Mark Dionne and David Walden. Interleaf, Inc. 1981 to 2000. IEEE Annals of the History of Computing, 42(1):63– 74, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Elzway:2017:BR

[EA17] Salem Elzway and Kera Jones Allen. Book reviews. IEEE Annals of the History of Computing, 39(3):69–71, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/03/
                                        man2017030069.html.

Eckstein:1996:JPE

[Eck96] P. Eckstein. J. Presper Eckert. IEEE Annals of the History of Computing, 18(1): 25–44, Spring 1996. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476559.pdf.

Eckert:1998:SDC

[Eck98] J. P. Eckert, Jr. A survey of digital computer memory systems. IEEE Annals of the History of Computing, 20 (4):15–28, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/15706/00728227.pdf.

Eckstein:2007:JK

[Eck07a] Peter Eckstein. Jack Kilby (1923–2005). IEEE Annals of the History of Computing, 29(1):90–95, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Eckstein:2007:WCB

[Eck07b]  Peter Eckstein.  William

Charles (‘Bill’) Norris. IEEE Annals of the History of Computing, 29(2):80–86, April/ June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Edwards:2001:TP

[Edw01] Paul N. Edwards. Think piece.

IEEE Annals of the History of Computing, 23(1):88–87, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1088.pdf.

Edmunds:2011:UCD

[EF11] Michael Edmunds and Tony Freeth. Using computation to decode the first known computer. Computer, 44(7):32–39, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Emigh:2022:PPS

[EHP22] Rebecca Jean Emigh and Johanna Hernández-Pérez. The present of the past: A sociotechnological framework for understanding the availability of research materials. IEEE Annals of the History of Computing, 44(4):16–27, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Eustace:2024:SCH

[EK24] Kevin Eustace and Albert Klosterman. SDRC: Company history and impact on the CAD/MCAE industry in the 20th Century. IEEE Annals of the History of Computing, 46 (4):84–101, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Eklund:1992:CQD

[Ekl92]  Jon Eklund.  Comments,

queries, and debate: The final word on ‘The Bug’. IEEE Annals of the History of Computing, 14(3):6–7, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3006.pdf.

Eklund:1994:RAA

[Ekl94] Jon Eklund. The Reservisor Automated Airline Reservation System: Combining communications and computing. IEEE Annals of the History of Computing, 16(1):62– 69, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1062.pdf;   http://www.
computer.org/annals/an1994/
a1062abs.htm.

Elzen:1994:SLS

[EM94] Boelie Elzen and Donald MacKenzie. The social limits of speed: The development and use of supercomputers. IEEE Annals of the History of Computing, 16(1): 46–61, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1046.pdf;   http://www.
computer.org/annals/an1994/
a1046abs.htm.

Erskine:2002:LER

[EMW02] Ralph Erskine, Philip Marks, and Frode Weierud. Letter to the editor: Review of US bombes. IEEE Annals of the History of Computing, 24(3):85–87, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a3letter_web/
a3lettertoeditor_web.htm;
http://dlib.computer.org/
an/books/an2002/pdf/a3085.
pdf;   http://www.computer.
org/annals/an2002/a3085abs.
htm.

Endres:2004:IBE

[End04] Albert Endres. IBM Boeblingen’s early software contributions. IEEE Annals of the History of Computing, 26(3):31–41, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/03/a3031.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/03/a3031.
                                        pdf.

Endres:2013:ELC

[End13] Albert Endres. Early language and compiler developments at IBM Europe: A personal retrospection. IEEE Annals of the History of Computing, 35 (4):18–30, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2001:QPC

[Ens01] Nathan L. Ensmenger. The ‘question of professionalism’ in the computer fields. IEEE Annals of the History of Computing, 23(4):56–74, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2001/a4056abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2001/pdf/a4056.
                                        pdf.

Ensmenger:2004:OSL

[Ens04a]  Nathan Ensmenger.  Open

source’s lessons for historians. IEEE Annals of the History of Computing, 26(4):104, 102–103, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/

dl/mags/an/2004/04/a4104.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4104.
pdf.

Ensmenger:2004:PPT

[Ens04b] Nathan Ensmenger. Power to the people: Toward a social history of computing. IEEE Annals of the History of Computing, 26(1):96, 94– 95, January/March 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/01/a1096.pdf.

Ensmenger:2007:CEA

[Ens07] Nathan Ensmenger. Computers as ethical artifacts. IEEE Annals of the History of Computing, 29(3):88, 86–87, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://bell.computer.
org/dlcomments/.

Ensmenger:2009:SHE

[Ens09] Nathan Ensmenger. Software as history embodied. IEEE Annals of the History of Computing, 31(2):88–91, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2013:CMG

[Ens13] Nathan Ensmenger. Computation, materiality, and the global environment. IEEE Annals of the History of Computing, 35 (3):80, July/September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2015:EDa

[Ens15a] Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 37(1):2, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2015:EDb

[Ens15b]  Nathan Ensmenger.  From

the Editor’s desk. IEEE Annals of the History of Computing, 37(3):2–3, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/an/2015/03/
                                        man2015030002-abs.html.

Ensmenger:2016:EDa

[Ens16a] Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 38(1):2, January/March 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2016:EDb

[Ens16b] Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 38(3):4–5, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ensmenger:2017:EDa

[Ens17a]  Nathan Ensmenger.  From

the Editor’s desk. IEEE Annals of the History of Computing, 39(1):4–5, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010004.html.

Ensmenger:2017:EDb

[Ens17b] Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 39(2):4–5, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ensmenger:2017:EUD

[Ens17c]  Nathan Ensmenger.  From

the Editor’s desk. IEEE Annals of the History of Computing, 39(3):4–5, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030004.html.

Ensmenger:2017:EDc

[Ens17d] Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 39(4):3–4, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ensmenger:2018:HDC

[Ens18a] Nathan Ensmenger. The history of design in computing. IEEE Annals of the History of Computing, 40(1):3–4, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010003.pdf.

Ensmenger:2018:LGH

[Ens18b] Nathan Ensmenger. The local and the global in the history of computing. IEEE Annals of the History of Computing, 40(2):3–4, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8415730/.

Epstein:2013:HS

[Eps13]  Bob Epstein.  History of

Sybase. IEEE Annals of the History of Computing, 35(2): 31–41, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Eckdahl:2003:WCC

[ERS03] Donald E. Eckdahl, Irving S. Reed, and Hrant H. (Harold) Sarkissian. West Coast contributions to the development of the general-purpose computer: Building Maddida and the founding of Computer Research Corporation. IEEE Annals of the History of Computing, 25(1):4–33, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/01/a1004.
htm;   http://csdl.computer.
org/dl/mags/an/2003/01/a1004.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/01/a1004abs.
htm.

Estrada:2005:YAS

[ES05] Richard F. Estrada and Edward A. Starr. 50 years of acoustic signal processing for detection: Coping with the digital revolution. IEEE Annals of the History of Computing, 27(2):65–78, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2065.
htm;   http://csdl.computer.
org/dl/mags/an/2005/02/a2065.
pdf.

Estrin:1996:WSC

[Est96] T. Estrin. Women’s studies and computer science: their intersection. IEEE Annals of the History of Computing, 18(3):43–46, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511943.pdf.

Estrin:2000:CNB

[Est00] G. Estrin. Computer networkbased scientific collaboration in the energy research community, 1973–1977: a memoir. IEEE Annals of the History of Computing, 22 (2):42–52, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841136.pdf.

Estrin:2002:RCO

[Est02] Gerald Estrin. Reconfigurable computer origins: The UCLA fixed-plus-variable (F+V) structure computer. IEEE Annals of the History of Computing, 24(4):3–9, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2002/04/a4003.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2002/04/a4003.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/an/2002/04/a4003abs.
                                        htm.

Estruth:2022:GRC

[Est22] Jeannette Alden Estruth. From the Gold Rush to the colonization of Mars: How Silicon Valley imagines away the working class. IEEE Annals of the History of Computing, 44(2): 69–73, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Evans:2024:HCU

[Eva24] Michael A. Evans. A history of CAD in the United Kingdom: The Cambridge phenomenon. IEEE Annals of the History of Computing, 46(4):102–113, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fairbairn:2024:OHA

[Fai24] Douglas Fairbairn. Oral history of Alvar Green. IEEE Annals of the History of Computing, 46(4):142–147, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fallon:1998:ECG

[Fal98] K. K. Fallon. Early computer graphics developments in the architecture, engineering and construction industry. IEEE Annals of the History of Computing, 20 (2):20–29, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667293.pdf.

Fitzpatrick:2003:AHD

[FAP^{+}03] Anne Fitzpatrick, Janet Abbate, Frank S. Preston, Lucy Slater, Mary Lee Berners-Lee, Jean Sammet, Dame Stephanie Shirley, Ruzena Bajcsy, Ann Hardy, Adele Goldberg, and Anita Borg. Anecdotes: How did you first get into computing? IEEE Annals of the History of Computing, 25 (4):78–88, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2003/04/a4078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2003/04/a4078.
                                        pdf.

Fitzpatrick:2006:MBC

[FBK06] Anne Fitzpatrick, Simon Berkovich, and Tatiana Kazokova. MESM and the beginning of the computer era in the Soviet Union. IEEE Annals of the History of Computing, 28(3):4–16, July/ September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fitzpatrick:2000:ESa

[FC00a] A. Fitzpatrick and M. Croarken. Events and sightings. IEEE Annals of the History of Computing, 22(3):88–92, July/ September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://ieeexplore.
                                        ieee.org/iel5/85/18655/00859526.
                                        pdf.

Fitzpatrick:2000:ESb

[FC00b] Anne Fitzpatrick and Mary Croarken. Events and sightings. IEEE Annals of the History of Computing, 22 (4):77–??, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://dlib.computer.org/

an/books/an2000/pdf/a4077.
pdf.

Fitzpatrick:2001:ESa

[FC01a] Anne Fitzpatrick and Mary Croarken. Events and sightings. IEEE Annals of the History of Computing, 23(1):57–61, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an2001/pdf/a1057.
pdf.

Fitzpatrick:2001:ESb

[FC01b] Anne Fitzpatrick and Mary Croarken. Events and sightings. IEEE Annals of the History of Computing, 23(2):68– 73, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2068.pdf;   http://www.
computer.org/annals/an2001/
a2068x1.htm.

Fidler:2015:PIA

[FC15] Bradley Fidler and Morgan Currie. The production and interpretation of ARPANET maps. IEEE Annals of the History of Computing, 37(1):44– 55, January/March 2015. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fidler:2016:IRH

[FC16] Bradley Fidler and Morgan Currie. Infrastructure, representation, and historiography in BBN’s Arpanet maps. IEEE Annals of the History of Computing, 38(3):44–57, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fedorkow:2021:RSW

[Fed21] Guy C. Fedorkow. Recovering software for the Whirlwind computer. IEEE Annals of the History of Computing, 43(1):38–59, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//www.bitsavers.org/pdf/
mit/whirlwind/history/Wolf_
No_E_2005.pdf;   https://
www.historia-mollimercium.
com/whirlwind/Whirlwind-
Software-Recovery-Project-
v6d.pdf;   https://www.historia-
mollimercium.com/whirlwind/
Whirlwind-The-Physical-Artifact.
pdf; https://www.historia-
mollimercium.com/whirlwind/
ww-page.html.

Feinler:2010:NIC

[Fei10] Elizabeth Feinler. The Network Information Center and its archives. IEEE Annals of the History of Computing, 32(3): 83–89, July/September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Feinler:2011:HTT

[Fei11] Elizabeth (Jake) Feinler. Host tables, top-level domain names, and the origin of dot com. IEEE Annals of the History of Computing, 33(3):74–79, July/ September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Feichtinger:2022:CCH

[Fei22] Moritz J. Feichtinger. Computing counterinsurgency: The Hamlet Evaluation System (HES) and databasing during the Vietnam War. IEEE Annals of the History of Computing, 44(4):28–43, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Feldman:2011:CTB

[Fel11] Julian Feldman. Computers and thought — the back story. IEEE Annals of the History of Computing, 33(4):56–61, October/December 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fernandes:2022:RLD

[Fer22] Kim Fernandes. Review of Lives of Data: Essays on Computational Cultures from India by Sandeep Mertia. IEEE Annals of the History of Computing, 44(2):74–75, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Feurzeg:2006:ETB

[Feu06]  Wallace Feurzeg.  Educa-

tional technology at BBN.

IEEE Annals of the History of Computing, 28(1):18–31, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/portal/pages/annals/
articles/xtras/a1-2006/a1feurzeig.
html.

Fevolden:2013:BBW

[Fev13] Arne Martin Fevolden. The best of both worlds? A history of time-shared microcomputers, 1977–1983. IEEE Annals of the History of Computing, 35(1):23–34, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fuegi:2003:LBC

[FF03] John Fuegi and Jo Francis. Lovelace & Babbage and the creation of the 1843 ‘Notes’. IEEE Annals of the History of Computing, 25(4):16–26, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4016abs.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4016.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4016.pdf.

Fornes:2014:CTO

[FH14] Jordi Fornes and Nestor Herran. Computing in transition: The origins of Barcelona’s School of Informatics, 1976– 1984. IEEE Annals of the History of Computing, 36(1):18– 29, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fornes:2017:CDA

[FHD17] Jordi Fornés, Néstor Herran, and Luís Duque. Computing for democracy: The Asociación de Técnicos de Informática and the professionalization of computing in Spain. IEEE Annals of the History of Computing, 39 (2):30–48, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/7974689/

.

Fidler:2017:EOE

[Fid17] Bradley Fidler. Eternal October and the end of cyberspace. IEEE Annals of the History of Computing, 39(1):6–7, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010006-abs.html.

Fidler:2018:CCN

[Fid18] Bradley Fidler. Cryptography, capitalism, and national security. IEEE Annals of the History of Computing, 40 (4):80–84, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8620680/.

Field:2003:NWB

[Fie03] A. J. Field. Networking the world [book review]. IEEE Annals of the History of Computing, 25(3):83, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fitzpatrick:2002:A

[Fit02a]  Anne Fitzpatrick.  Anec-

dotes. IEEE Annals of the History of Computing, 24 (4):40–44, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2002/04/a4040.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2002/04/a4040.
                                        pdf.

Fitzpatrick:2002:AWI

[Fit02b]  Anne Fitzpatrick.  Anec-

dotes: a well-intentioned query and the Halloween Problem. IEEE Annals of the History of Computing, 24(2):86–89, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a2086.pdf;   http://www.
                                        computer.org/annals/an2002/
                                        a2086abs.htm.

Fitzpatrick:2003:A

[Fit03a] Anne Fitzpatrick. Anecdotes. IEEE Annals of the History of Computing, 25(1):72–79, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a1072abs.htm;   http:
//computer.org/annals/an2003/
extras/a1072.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1072.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1072.pdf.

Fitzpatrick:2003:ADT

[Fit03b] Anne Fitzpatrick. Anecdotes [developing Telnet’s negotiated options]. IEEE Annals of the History of Computing, 25(2):80–82, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2080.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2080.
pdf.

Fitzpatrick:2003:AO

[Fit03c]  Anne Fitzpatrick.  Anec-

dotes: On the origins of :-). IEEE Annals of the History of Computing, 25 (3):82–83, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3082.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3082.
pdf.

Fitzpatrick:2023:RSL

[Fit23] Anne Fitzpatrick. Review of Spies, Lies, and Algorithms, by Amy B. Zegart. IEEE Annals of the History of Computing, 45 (2):68, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Flamm:1998:C

[Fla98]  B. Flamm.  Calculators.

IEEE Annals of the History of Computing, 20(3): 72–73, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/15262/00707577.pdf.

Fisher:1993:DEB

[FM93] Amy Weaver Fisher and James L. McKenney. The development of the ERMA banking system: Lessons from history. IEEE Annals of the History of Computing, 15(1):44–57, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a1044.pdf;   http://www.
                                        computer.org/annals/an1993/
                                        a1044abs.htm.

Forest:1997:B

[For97]  R. B. Forest.  Biogra-

phies. IEEE Annals of the History of Computing, 19 (2):70–73, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586075.pdf.

Frana:2004:BWT

[Fra04]  Philip L. Frana.  Before

the Web there was Gopher. IEEE Annals of the History of Computing, 26(1): 20–41, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/01/a1020abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/01/a1020.
pdf.

Frailey:2013:CGD

[Fra13] Dennis J. Frailey. Computergenerated doctoral dissertations: An early experience. IEEE Annals of the History of Computing, 35(3):65–66, July/ September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Frana:2018:TEH

[Fra18] Philip L. Frana. Telematics and the early history of international digital information flows. IEEE Annals of the History of Computing, 40(2):32–47, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8415735/.

Frana:2023:DIC

[Fra23a] Philip L. Frana. Demographics, inc., computerized direct mail, and the rise of the digital attention economy. IEEE Annals of the History of Computing, 45(2):15–28, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Franz:2023:RCI

[Fra23b] Sam H. Franz. Review of Code: From Information Theory to French Theory, by Bernard Dionysius Geoghegan. IEEE Annals of the History of Computing, 45(3):105–107, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fitzpatrick:2004:A

[FRB04] Anne Fitzpatrick, Joe E. Rogers, and Bob Bemer. Anecdotes. IEEE Annals of the History of Computing, 26(1):59– 67, January/March 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/an/
                                        2004/01/a1059.pdf.

Freeman:2007:DEM

[Fre07]  Herbert Freeman.  Design

of an early minicomputer. IEEE Annals of the History of Computing, 29(3):68–71, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Fritz:1994:EPS

[Fri94] W. Barkley Fritz. ENIAC — a problem solver. IEEE Annals of the History of Computing, 16(1):25–45, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an1994/pdf/a1025.
pdf;   http://www.computer.
org/annals/an1994/a1025abs.
htm.

Fritz:1996:WE

[Fri96] W. Barkley Fritz. The women of ENIAC. IEEE Annals of the History of Computing, 18(3):13–28, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511940.pdf.

Fedorkow:2025:DMM

[FVW25] Guy Fedorkow, Jochen Viehoff, and David Woitkowski. In defense of model-makers: Adding a new point to the computer conservation continuum. IEEE Annals of the History of Computing, 47(1):72–75, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Genadek:2022:MLD

[GA22] Katie R. Genadek and J. Trent Alexander. The missing link: Data capture technology and the making of a longitudinal U.S. census infrastructure.

IEEE Annals of the History of Computing, 44(4):57–66, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Galler:2004:AHF

[Gal04]  Bernard A. Galler.  An-

nals: How the first issue came to be. IEEE Annals of the History of Computing, 26(1):4–7, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/01/a1004abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/01/a1004.
pdf.

Galloway:2011:PCM

[Gal11] Patricia Galloway. Personal computers, microhistory, and shared authority: Documenting the inventor–early adopter dialectic. IEEE Annals of the History of Computing, 33(2): 60–74, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ganzhorn:2004:BIB

[Gan04a]  Karl E. Ganzhorn.  The

buildup of the IBM Boeblingen Laboratory. IEEE Annals of the History of Computing, 26(3):4–19, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/03/a3004.
htm;   http://csdl.computer.
org/dl/mags/an/2004/03/a3004.
pdf.

Ganzhorn:2004:IBL

[Gan04b] Karl E. Ganzhorn. IBM Boeblingen Laboratory: Product development. IEEE Annals of the History of Computing, 26(3):20–30, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/03/a3020.
htm;   http://csdl.computer.
org/dl/mags/an/2004/03/a3020.
pdf.

Garvey:2021:GPS

[Gar21] S. C. Garvey. The general problem solver does not exist: Mortimer Taube and the art of AI criticism. IEEE Annals of the History of Computing, 43(1): 60–73, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gass:1999:PMM

[Gas99] S. I. Gass. Project Mercury’s man-in-space real-time computer system: “you have a go, at least seven orbits”. IEEE Annals of the History of Computing, 21(4):37– 48, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17389/00801531.pdf.

Gawrysiak:2019:PTE

[Gaw19] P. Gawrysiak. Polish text editors during the fall of the Iron Curtain. IEEE Annals of the History of Computing, 41(2):6–21, April 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2009:GEI

[GB09] Burton Grad and Thomas J. Bergin. Guest Editors’ introduction: History of database management systems. IEEE Annals of the History of Computing, 31(4):3–5, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gobbo:2013:AMC

[GB13] Federico Gobbo and Marco Benini. From ancient to modern computing: A history of information hiding. IEEE Annals of the History of Computing, 35(3):33–39, July/ September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2000:SHB

[GC00] D. A. Grier and M. Campbell. A social history of Bitnet and Listserv, 1985–1991. IEEE Annals of the History of Computing, 22(2):32–41, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841135.pdf.

Geoghegan:2008:HCI

[Geo08] Bernard Dionysius Geoghegan.

Historiographic conceptualization of information: a critical survey. IEEE Annals of the History of Computing, 30(1): 66–81, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gerovitch:2004:BHM

[Ger04] S. Gerovitch. Between Human and Machine: Feedback, Control, and Computing Before Cybernetics [book review]. IEEE Annals of the History of Computing, 26(1):71–73, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2006:BHN

[Ges06a] Michael N. Geselowitz. A busman’s holiday in New Jersey. IEEE Annals of the History of Computing, 28(1):104– 105, January/March 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2006:LAN

[Ges06b] Michael N. Geselowitz. Local area networking. IEEE Annals of the History of Computing, 28(3):84–85, July/September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2007:BEC

[Ges07a] Michael N. Geselowitz. A boat excursion into Canada’s early computer history — Toronto, Ottawa, and Montreal. IEEE Annals of the History of Computing, 29(4):88–89, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2007:OCA

[Ges07b] Michael N. Geselowitz. The Oxford–Cambridge arc. IEEE Annals of the History of Computing, 29(2):78–79, April/ June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2008:AAP

[Ges08] Michael N. Geselowitz. From Alcoa to Anacom–Pittsburgh. IEEE Annals of the History of Computing, 30(2):72–73, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2009:LAN

[Ges09a]  Michael Geselowitz.  Local

area networks: High (plains) tech. IEEE Annals of the History of Computing, 31(3): 77, July/September 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2009:FTO

[Ges09b] Michael N. Geselowitz. Fuhgedaboudit — technology in the Outer Boroughs of New York. IEEE Annals of the History of Computing, 31(1):74–75, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2010:I

[Ges10a] Michael Geselowitz. Israel.

IEEE Annals of the History of Computing, 32(2):69, April/ June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2010:LIN

[Ges10b] Michael Geselowitz. Long Island, New York. IEEE Annals of the History of Computing, 32 (4):88–89, October/December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2011:G

[Ges11a] Michael Geselowitz. Greece.

IEEE Annals of the History of Computing, 33(2):105, April/ June 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2011:V

[Ges11b] Michael Geselowitz. Vancouver. IEEE Annals of the History of Computing, 33(4):68– 69, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2012:I

[Ges12a] Michael Geselowitz. Indianapolis. IEEE Annals of the History of Computing, 34(2):73, April/ June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Geselowitz:2012:SS

[Ges12b] Michael Geselowitz. Stockholm, Sweden. IEEE Annals of the History of Computing, 34(4):71, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Goldstine:1946:ENI

[GG46] H. H. Goldstine and Adele Goldstine. The Electronic Numerical Integrator and Computer (ENIAC). Mathematical Tables and Other Aids to Computation, 2(15):97–110, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Grattan-Guinness:1992:CBA

[GG92] Ivor Grattan-Guinness. Charles Babbage as an algorithmic thinker. IEEE Annals of the History of Computing, 14(3):34–48, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3034.pdf;   http://www.
computer.org/annals/an1992/
a3034abs.htm.

Goldstine:1996:ENI

[GG96] H. H. Goldstine and A. Goldstine. The Electronic Numerical Integrator and Computer (ENIAC). IEEE Annals of the History of Computing, 18(1):10–16, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476557.pdf. Reprint

of 1946 technical report and [GG46].

Godfrey:1993:CNP

[GH93] M. D. Godfrey and D. F.

Hendry. The computer as von Neumann planned it. IEEE Annals of the History of Computing, 15(1):11–21, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a1011.pdf;   http://www.
computer.org/annals/an1993/
a1011abs.htm.

Grad:2018:DPL

[GH18a] Burton Grad and David Hemmendinger. Desktop publishing: Laying the foundation. IEEE Annals of the History of Computing, 40(3):3–6, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509552/.

Grier:2018:RRP

[GH18b] David Alan Grier and David Hemmendinger. Reviews of recent publications [two books reviewed: Thomas Haigh, Mark Priestley, and Crispin Rope, ENIAC in Action: Making and Remaking the Modern Computer. Cambridge: MIT Press, 2016. 366 pp. ISBN: 978-0-262-03398-5. Thomas J. Misa and Jeffrey R. Yost, FastLane: Managing Science in the Internet World, Baltimore: Johns Hopkins University Press, 2015. 224 pp. ISBN: 978-1-4214-1868-1]. IEEE Annals of the History of Computing, 40(2):80–84, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2019:DPP

[GH19] Burton Grad and David Hemmendinger. Desktop publishing, part 2. IEEE Annals of the History of Computing, 41(3):5–7, July 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2020:DPP

[GH20] Burton Grad and David Hemmendinger. Desktop publishing, part 3. IEEE Annals of the History of Computing, 42(1): 6, January/March 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Glusker:2005:TCM

[GHV05] Mark Glusker, David M.

Hogan, and Pamela Vass. The ternary calculating machine of Thomas Fowler. IEEE Annals of the History of Computing, 27(3):4–22, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Giloi:1997:KZP

[Gil97]  Wolfgang K. Giloi.  Kon-

rad Zuse’s Plankalkül: the first high-level, “non von Neumann” programming language. IEEE Annals of the History of Computing, 19 (2):17–24, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586068.pdf.

Gillmor:2007:SIF

[Gil07] C. Stewart Gillmor. Stanford, the IBM 650, and the first trials of computer date matching. IEEE Annals of the History of Computing, 29(1):74– 80, January/March 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Giordano:1993:ICR

[Gio93] Richard Giordano. Institutional change and regeneration: a biography of the Computer Science Department at the University of Manchester.

IEEE Annals of the History of Computing, 15(3):55–62, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3055.pdf;   http://www.
computer.org/annals/an1993/
a3055abs.htm.

Grad:2002:SSP

[GJ02] Burton Grad and Luanne Johnson. The start of the software products industry. IEEE Annals of the History of Computing, 24(1):3–4, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1003abs.htm;
 http://dlib.computer.org/
an/books/an2002/pdf/a1003.
pdf.

Grad:2012:CHS

[GJ12] Burton Grad and Luanne Johnson. Collecting the history of the software industry. IEEE Annals of the History of Computing, 34(4):88, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Griffiths:2002:UIR

[GK02] José-Marie Griffiths and Donald W. King. US information retrieval system evolution and evaluation. IEEE Annals of the History of Computing, 24(3):35–55, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3035.pdf;   http://www.
computer.org/annals/an2002/
a3035abs.htm.

Grad:2024:CFI

[GK24] Burton Grad and David Kasik.

CAD: Forming the industry. IEEE Annals of the History of Computing, 46(4):7–9, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Good:1996:CQD

[GM96] I. J. Good and Donald Michie.

Comments, queries, and debate: 50 years after breaking the codes: a disclaimer. IEEE Annals of the History of Computing, 18(4):65– 66, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539920.pdf.

Goetz:2002:MSPb

[Goe02a] Martin Goetz. Memoirs of a software pioneer: Part 2. IEEE Annals of the History of Computing, 24 (4):14–31, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/

                                        dl/mags/an/2002/04/a4014.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2002/04/a4014.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/an/2002/04/a4014abs.
                                        htm.

Goetz:2002:MSPa

[Goe02b] Martin Goetz. Memoirs of a software pioneer: Part I. IEEE Annals of the History of Computing, 24(1):43–56, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2002/a1043abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2002/pdf/a1043.
                                        pdf.

GonzalezRedondo:2020:ELT

[Gon20] F. A. González Redondo. Exhibition Leonardo Torres Quevedo: The universal engineer. IEEE Annals of the History of Computing, 42(4):119– 122, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Goncalves:2024:LLO

[Gon24a] Bernardo Gonçalves. Lady Lovelace’s objection: The Turing–Hartree disputes over the meaning of digital computers, 1946–1951. IEEE Annals of the History of Computing, 46 (1):6–18, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Goncalves:2024:TTB

[Gon24b] Bernardo Gonçalves. Turing’s test, a beautiful thought experiment. IEEE Annals of the History of Computing, 46(3):36– 49, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gotkin:2014:WCW

[Got14] Kevin Gotkin. When computers were amateur. IEEE Annals of the History of Computing, 36 (2):4–14, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gottwald:2023:SHX

[Got23] Emil J. Gottwald. A short history of Xerox’ Darwin Toolset for programming real-time machine control systems. IEEE Annals of the History of Computing, 45(2):63–67, April/ June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Goyal:1996:WCH

[Goy96] A. Goyal. Women in computing: historical roles, the perpetual glass ceiling, and current opportunities. IEEE Annals of the History of Computing, 18(3):36–42, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511942.pdf.

Graham:2002:GEI

[GR02] Rebecca Graham and W. Boyd Rayward. Guest Editors’ introduction: Computer applications in libraries, Part 1. IEEE Annals of the History of Computing, 24(2):3, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a2003.pdf;   http://www.
                                        computer.org/annals/an2002/
                                        a2003abs.htm.

Giacobazzi:2022:HAI

[GR22] Roberto Giacobazzi and Francesco Ranzato. History of abstract interpretation. IEEE Annals of the History of Computing, 44(2):33–43, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2002:PRI

[Gra02a] Burton Grad. A personal recollection: IBM’s unbundling of software and services. IEEE Annals of the History of Computing, 24(1):64–71, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2002/a1064abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2002/pdf/a1064.
                                        pdf.

Graham:2002:GEC

[Gra02b]  Rebecca Graham.  Guest

Editor’s conclusion: Reflections on the evolution of library computing. IEEE Annals of the History of Computing, 24(3):75–78, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3075.pdf.

Grad:2003:PSI

[Gra03]  Burton Grad.  Preserv-

ing the software industry’s past. IEEE Annals of the History of Computing, 25 (1):88, 3, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/01/a1088.
htm;   http://csdl.computer.
org/dl/mags/an/2003/01/a1088.
pdf.

Grad:2007:CDV

[Gra07] Burton Grad. The creation and the demise of VisiCalc. IEEE Annals of the History of Computing, 29(3):20–31, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2012:RDM

[Gra12]  Burton Grad.  Relational

database management systems: The formative years [Guest editor’s introduction]. IEEE Annals of the History of Computing, 34(4):7–8, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2013:RDM

[Gra13]  Burton Grad.  Relational

database management systems: The business explosion [Guest Editor’s introduction]. IEEE Annals of the History of Computing, 35(2):8–9, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2020:FSI

[Gra20a] B. Grad. Finding software industry history. IEEE Annals of the History of Computing, 42(3):83–91, July/September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2020:SSH

[Gra20b] B. Grad. In search of software history. IEEE Annals of the History of Computing, 42(3): 76–82, July/September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2021:CCH

[Gra21a] B. Grad. Collecting computing history is everyone’s job. IEEE Annals of the History of Computing, 43(1):97–98, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2021:SSB

[Gra21b] Burton Grad. Sterling Software: An $8 billion success story in the 1980s and 1990s. IEEE Annals of the History of Computing, 43(4):77–86, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2022:HCI

[Gra22] Burton Grad. History of computing industry infrastructure. IEEE Annals of the History of Computing, 44(1):131–132, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grad:2024:CAD

[Gra24] Burton Grad. Computer-aided design: CAD materials at the Computer History Museum and Charles Babbage Institute. IEEE Annals of the History of Computing, 46(4):148, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Greenstein:1995:LCS

[Gre95]  S. M. Greenstein.  Lock-

in and the costs of switching mainframe computer vendors in the US federal government in the 1970s. IEEE Annals of the History of Computing, 17(3):58–66, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397061.pdf.

Greenberg:2005:BEL

[Gre05] Joshua M. Greenberg. Between expert and lay. IEEE Annals of the History of Computing, 27(2):96, 95, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2096.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/02/a2096.
                                        pdf.

Green:2013:DVS

[Gre13] Mike Green. Dummer’s vision of solid circuits at the UK Royal Radar Establishment. IEEE Annals of the History of Computing, 35(1):56– 66, January/March 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Griffith:1994:MED

[Gri94] B. A. Griffith. My early days in Toronto. IEEE Annals of the History of Computing, 16 (2):55–64, Summer 1994. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:1996:EVP

[Gri96] D. A. Grier. The ENIAC, the verb “to program” and the emergence of digital computers. IEEE Annals of the History of Computing, 18(1):51–55, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476561.pdf.

Grier:1997:GBM

[Gri97] D. A. Grier. Gertrude Blanch of the Mathematical Tables Project. IEEE Annals of the History of Computing, 19 (4):18–27, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13659/00627896.pdf.

Grier:1998:MTP

[Gri98]  D. A. Grier.  The Math

Tables Project of the work projects administration: the reluctant start of the computing era. IEEE Annals of the History of Computing, 20(3):33–50, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707573.pdf.

Grier:1999:ES

[Gri99] D. A. Grier. Events and sightings. IEEE Annals of the History of Computing, 21(1): 71–75, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16436/00759372.pdf.

Grier:2000:ACC

[Gri00a] D. A. Grier. Agricultural computing and the context for John Atanasoff. IEEE Annals of the History of Computing, 22(1):48–61, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17665/00815466.pdf.

Grier:2000:TP

[Gri00b] David Alan Grier. Think piece. IEEE Annals of the History of Computing, 22(4):88–??, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2000/pdf/
                                        a4088.pdf.

Grier:2001:A

[Gri01a]  David Alan Grier.  Anec-

dotes. IEEE Annals of the History of Computing, 23(2):78– 79, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2001/pdf/
                                        a2078.pdf.

Grier:2001:RFC

[Gri01b] David Alan Grier. The rise and fall of the Committee on Mathematical Tables and Other Aids to Computation.

IEEE Annals of the History of Computing, 23(2):38–49, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2038.pdf;   http://www.
computer.org/annals/an2001/
a2038abs.htm.

Grier:2002:LE

[Gri02] David Alan Grier. Letters to the editor. IEEE Annals of the History of Computing, 24 (4):45–47, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4045.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4045.
pdf.

Grier:2003:CJN

[Gri03a]  David Alan Grier.  The

computer: From John von Neumann to Irma S. Rombauer. IEEE Annals of the History of Computing, 25(4): 104, 103, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/04/a4104.
htm;   http://csdl.computer.
org/dl/mags/an/2003/04/a4104.
pdf.

Grier:2003:GMT

[Gri03b]  David Alan Grier.  The

great machine theory of history. IEEE Annals of the History of Computing, 25 (3):96, 95, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2003/03/a3096.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2003/03/a3096.
                                        pdf.

Grier:2004:ES

[Gri04a] David Alan Grier. Events and sightings. IEEE Annals of the History of Computing, 26 (4):62–70, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/04/a4062.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/04/a4062.
                                        pdf.

Grier:2004:EDa

[Gri04b] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 26(1):2–3, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2004/01/a1002.pdf.

Grier:2004:EDb

[Gri04c] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 26(2):2, April/June 2004. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2002.pdf;
 http://csdl.computer.org/
dl/mags/an/2004/02/a2002.
htm.

Grier:2004:EDc

[Gri04d] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 26(4):2–3, October/ December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2004/04/
a4002.pdf;   http://csdl.
computer.org/dl/mags/an/
2004/04/a4002.htm.

Grier:2004:LE

[Gri04e] David Alan Grier. Letters to the editor. IEEE Annals of the History of Computing, 26(3):66–68, July/ September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2004/03/
a3066.pdf;   http://csdl.
computer.org/dl/mags/an/
2004/03/a3066.htm.

Grier:2005:EDa

[Gri05a] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 27(1):2–3, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/01/a1002.
                                        htm;   http://csdl2.computer.
                                        org/comp/mags/an/2005/01/
                                        a1002.pdf.

Grier:2005:EDb

[Gri05b] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 27(2):2–3, April/June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2002.
                                        htm;   http://csdl2.computer.
                                        org/comp/mags/an/2005/02/
                                        a2002.pdf.

Grier:2005:EDc

[Gri05c] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 27(3):2, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2005/03/a3002.pdf.

Grier:2005:EDd

[Gri05d] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 27(4):2–3, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2005/04/a4002.pdf.

Grier:2006:EDa

[Gri06a] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 28(1):2, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/01/a1002.pdf.

Grier:2006:EDb

[Gri06b] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 28(2):2–3, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/02/a2002.pdf.

Grier:2006:EDc

[Gri06c] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 28(3):2–3, July/September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/03/a3002.pdf.

Grier:2006:EDd

[Gri06d] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 28(4):2–3, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/04/a4002.pdf.

Grier:2007:CH

[Gri07a] David Alan Grier. The center of history. IEEE Annals of the History of Computing, 29(4):2–5, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2007/04/
                                        man2007040002.pdf.

Grier:2007:EDa

[Gri07b] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 29(1):2, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2007/01/a1002.pdf.

Grier:2007:EDb

[Gri07c] David Alan Grier. From the Editor’s desk. IEEE Annals of the History of Computing, 29(2):2, April/June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2007/02/a2002.pdf.

Grier:2007:EDc

[Gri07d]  David Alan Grier.  From

the Editor’s desk. IEEE Annals of the History of Computing, 29(3):2–3, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2007/03/
man2007030002.pdf.

Grier:2010:IYC

[Gri10] David A. Grier. The inconsistent youth of Charles Babbage. IEEE Annals of the History of Computing, 32(4):18–31, October/December 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2011:CCH

[Gri11a] David Alan Grier. From computer celebrities to historical biography. IEEE Annals of the History of Computing, 33(4): 88, 86–87, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2011:PP

[Gri11b] David Alan Grier. Programming and planning. IEEE Annals of the History of Computing, 33(1):86–88, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2012:RDC

[Gri12] David Alan Grier. The relational database and the concept of the information system. IEEE Annals of the History of Computing, 34(4):9–17, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2013:IEF

[Gri13]  David Alan Grier.  Inter-

view: Edward Feigenbaum. IEEE Annals of the History of Computing, 35(4):74–81, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2014:SFI

[Gri14] David Alan Grier. Steve Furber [interviews]. IEEE Annals of the History of Computing, 36(1):58–68, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2022:HAT

[Gri22a] David A Grier. HWMS audio theatre. IEEE Annals of the History of Computing, 44(3): 93, July/September 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grier:2022:IPF

[Gri22b] David A. Grier. Interview of Peter E. Friedland. IEEE Annals of the History of Computing, 44(1):109–122, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grindal:2022:ACO

[Gri22c] Karl Grindal. Artist collectives as the origins of DDoS: the Strano Network and Electronic Disturbance Theater.

IEEE Annals of the History of Computing, 44(3):30–42, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Grosch:1992:CQD

[Gro92]  Herb Grosch.  Comments,

queries, and debate: O, AFIPS! IEEE Annals of the History of Computing, 14(1):6–7, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1006.pdf.

Grosch:1993:CQD

[Gro93]  Herb Grosch.  Comments,

queries, and debate: a history of scientific computation. IEEE Annals of the History of Computing, 15(1):7–9, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a1007.pdf.

Grudin:2005:TFH

[Gru05] Jonathan Grudin. Three faces of human-computer interaction. IEEE Annals of the History of Computing, 27(4):46– 62, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gray:1998:SRT

[GS98] G. Gray and R. Smith. Sperry Rand’s transistor computers. IEEE Annals of the History of Computing, 20(3): 16–26, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707571.pdf.

Gray:2001:SRT

[GS01] George T. Gray and Ronald Q. Smith. Sperry Rand’s thirdgeneration computers 1964– 1980. IEEE Annals of the History of Computing, 23(1):3–16, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1003.pdf;   http://www.
computer.org/annals/an2001/
a1003abs.htm.

Gray:2003:BBB

[GS03] George T. Gray and Ronald Q. Smith. Before the B5000: Burroughs Computers, 1951– 1963. IEEE Annals of the History of Computing, 25(2): 50–61, April/June 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/02/a2050.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/02/a2050.pdf;   http:
//csdl.computer.org/dl/mags/
an/2003/02/a2050abs.htm.

Gray:2004:SRF

[GS04] George T. Gray and Ronald Q.

Smith. Sperry Rand’s first generation computers, 1955– 1960: Hardware and software. IEEE Annals of the History of Computing, 26 (4):20–34, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/04/a4020.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4020.
pdf.

Gray:2007:ACS

[GS07] George T. Gray and Ronald Q.

Smith. Against the current: The Sperry–Burroughs merger and the Unisys struggle to survive 1980–2001. IEEE Annals of the History of Computing, 29 (2):3–17, April/June 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gray:2009:ABB

[GS09] George T. Gray and Ronald Q.

Smith. After the B5000: Burroughs third-generation computers 1964–1980. IEEE Annals of the History of Computing, 31(2):44–55, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Guice:1999:JAE

[Gui99] J. Guice. James A. Anderson and Edward Rosenfeld, eds., Talking Nets: An Oral History of Neural Networks [book review]. IEEE Annals of the History of Computing, 21(3):75, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778990.pdf.

Gupta:2007:CSC

[Gup07] Gopal K. Gupta. Computer science curriculum developments in the 1960s. IEEE Annals of the History of Computing, 29(2):40–54, April/June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gurer:1996:WCE

[Gur96] D. W. Gurer. Women’s contributions to early computing at the National Bureau of Standards. IEEE Annals of the History of Computing, 18(3):29–35, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511941.pdf.

Hemmendinger:2024:EDb

[HA24] David Hemmendinger and Troy Kaighin Astarte. From the Editors’ desks. IEEE Annals of the History of Computing, 46(2):5, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See correction [Hem24a].

Haigh:2001:CPT

[Hai01] Thomas Haigh. The chromiumplated tabulator: Institutionalizing an electronic revolution, 1954–1958. IEEE Annals of the History of Computing, 23(4):75–104, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a4075abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a4075.
pdf.

Haigh:2002:SCS

[Hai02] Thomas Haigh. Software in the 1960s as concept, service, and product. IEEE Annals of the History of Computing, 24(1):5–13, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1005abs.htm;
http://dlib.computer.org/
an/books/an2002/pdf/a1005.
pdf.

Haigh:2003:BBM

[Hai03a] Thomas Haigh. Biographies [building the Manchester differential analyzers: a personal reflection]. IEEE Annals of the History of Computing, 25(2):86–92, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2086.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2086.
pdf.

Haigh:2003:BOR

[Hai03b] Thomas Haigh. Biographies: [obituaries: Rob Kling [1944– 2003]]. IEEE Annals of the History of Computing, 25(3):92–94, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3090.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3090.
pdf.

Haigh:2004:B

[Hai04a]  Thomas Haigh.  Biogra-

phies. IEEE Annals of the History of Computing, 26 (4):79–91, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/04/a4079.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4079.
pdf.

Haigh:2004:BAS

[Hai04b] Thomas Haigh. Biographies: ADAPSO and the service bureau industry, 1961–1968; Bernard “Bernie” Goldstein; Frank Lautenberg. IEEE Annals of the History of Computing, 26(1):78–93, January/March 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/01/a1078.pdf.

Haigh:2004:BRW

[Hai04c]  Thomas Haigh.  Biogra-

phies: Robert W Floyd, in memoriam. IEEE Annals of the History of Computing, 26(2):75–83, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/02/a2075.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2075.
pdf.

Haigh:2005:B

[Hai05a] Thomas Haigh. Biographies.

IEEE Annals of the History of Computing, 27(4):93–96, October/December 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2005:BAT

[Hai05b]  Thomas Haigh.  Biogra-

phies: ADAPSO, time-sharing firms, and software companies: 1968–1975. IEEE Annals of the History of Computing, 27(1):67–73, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/

                                        dl/mags/an/2005/01/a1067.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/01/a1067.
                                        pdf.

Haigh:2005:BLJ

[Hai05c] Thomas Haigh. Biographies: Lawrence J. Schoenberg. IEEE Annals of the History of Computing, 27(2):89–94, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2089.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/02/a2089.
                                        pdf.

Haigh:2006:ROF

[Hai06] Thomas Haigh. Remembering the office of the future: The origins of word processing and office automation. IEEE Annals of the History of Computing, 28(4):6–31, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2008:CMM

[Hai08] Thomas Haigh. Cleve Moler: Mathematical software pioneer and creator of Matlab. IEEE Annals of the History of Computing, 30(1):87–91, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2009:HDG

[Hai09] Thomas Haigh. How data got its base: Information storage software in the 1950s and 1960s. IEEE Annals of the History of Computing, 31(4):6–25, October/December 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2010:CAW

[Hai10a] Thomas Haigh. Computing the American way: Contextualizing the early US computer industry. IEEE Annals of the History of Computing, 32(2): 8–20, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2010:JRR

[Hai10b] Thomas Haigh. John R. Rice:

Mathematical software pioneer. IEEE Annals of the History of Computing, 32(4):72–81, October/December 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2011:CWB

[Hai11] Thomas Haigh. Charles W.

Bachman: Database software pioneer. IEEE Annals of the History of Computing, 33 (4):70–80, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2018:THB

[Hai18]  Thomas Haigh.  Thomas

Harold (“Tommy”) Flowers:

Designer of the Colossus codebreaking machines. IEEE Annals of the History of Computing, 40(1):72–82, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010072.pdf.

Hall:1992:CQD

[Hal92]  Charles Hall.  Comments,

queries, and debate: a mechanical delay line? IEEE Annals of the History of Computing, 14(1):9, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1006.pdf.

Hall:2000:FPD

[Hal00] E. C. Hall. From the farm to pioneering with digital control computers: an autobiography. IEEE Annals of the History of Computing, 22 (2):22–31, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841134.pdf.

Halsted:2018:OAM

[Hal18] David Halsted. The origins of the architectural metaphor in computing: Design and technology at IBM, 1957–1964. IEEE Annals of the History of Computing, 40(1):61–70, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010061.pdf.

Harris:1999:ESS

[Har99]  J. R. Harris.  The earli-

est solid-state digital computers. IEEE Annals of the History of Computing, 21(4):49– 54, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801532.pdf.

Hartley:2013:CFV

[Har13] David Hartley. CPL: Failed venture or noble ancestor? IEEE Annals of the History of Computing, 35(3):55–63, July/ September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Harmon:2022:ESB

[Har22a] Paul Harmon. The expert systems business: How it grew and died. IEEE Annals of the History of Computing, 44(1):31– 43, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hart:2022:AIO

[Har22b] Peter E. Hart. An artificial intelligence odyssey: From the research lab to the real world. IEEE Annals of the History of Computing, 44(1):57– 72, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hasegawa:2016:HCD

[Has16] Yutaka Hasegawa. History of COMTRAC: Development of the innovative traffic-control system for Shinkansen. IEEE Annals of the History of Computing, 38(2):11–21, April/ June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hodges:2006:R

[HC06] Andrew Hodges and James W. Cortada. Reviews. IEEE Annals of the History of Computing, 28(3):81–83, July/ September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Henin:2020:FTM

[HC20] S. Hénin and S. Casonato. Fake but true: Model maker Roberto Guatelli, science museums and replicated artifacts of computing history. IEEE Annals of the History of Computing, 42(2): 20–32, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Head:2001:ELB

[Hea01a] Robert V. Head. ERMA’s lost battalion. IEEE Annals of the History of Computing, 23(3):64–72, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3064abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3064.
pdf.

Head:2001:UPS

[Hea01b] Robert V. Head. Univac: a Philadelphia story. IEEE Annals of the History of Computing, 23(3):60–63, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3060abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3060.
pdf.

Head:2002:GSG

[Hea02] Robert V. Head. Getting Sabre off the ground. IEEE Annals of the History of Computing, 24 (4):32–39, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4032.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4032.
pdf;   http://csdl.computer.
org/dl/mags/an/2002/04/a4032abs.
htm.

Head:2003:WHF

[Hea03]  Robert V. Head.  Work-

ing in the house that FACT built. IEEE Annals of the History of Computing, 25(1):60–64, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/01/a1060.
htm;   http://csdl.computer.
org/dl/mags/an/2003/01/a1060.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/01/a1060abs.
htm.

Head:2004:DGD

[Hea04] Robert V. Head. Datamation’s glory days. IEEE Annals of the History of Computing, 26(2):16–21, April/ June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2016abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2016.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2016.
pdf.

Heart:2005:LTN

[Hea05] Frank Heart. Leading a topnotch R&D group in the BBN environment. IEEE Annals of the History of Computing, 27(2):39–51, April/June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2039.
htm;   http://csdl.computer.
org/dl/mags/an/2005/02/a2039.
pdf.

Heide:1997:STA

[Hei97] L. Heide. Shaping a technology:

American punched card systems 1880–1914. IEEE Annals of the History of Computing, 19 (4):28–41, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13659/00627897.pdf.

Heide:2011:SSA

[Hei11] Lars Heide. Scale and scope in American key-set office machine dynamics, 1880s–1930s. IEEE Annals of the History of Computing, 33(3):22–31, July/ September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2012:EDa

[Hei12a] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 34(3):2– 3, July/September 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2012:EDb

[Hei12b] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 34(4):2–3, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2013:EDa

[Hei13a] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 35(1):2– 3, January/March 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2013:EDb

[Hei13b] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 35 (2):2–3, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2013:EDc

[Hei13c] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 35(3):2– 3, July/September 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2013:EDd

[Hei13d] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 35(4):2–3, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2014:EDa

[Hei14a] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 36(1):2– 3, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heide:2014:EDb

[Hei14b] Lars Heide. From the Editor’s desk. IEEE Annals of the History of Computing, 36 (2):2–3, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2014:CAC

[Hem14a] D. Hemmendinger. COMIC:

An analog computer in the colorant industry. IEEE Annals of the History of Computing, 36(3):4–18, July/September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2014:MES

[Hem14b] David Hemmendinger. Messaging in the early SDC timesharing system. IEEE Annals of the History of Computing, 36(1):52–57, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2016:TEI

[Hem16] David Hemmendinger. Two early interactive computer network experiments. IEEE Annals of the History of Computing, 38(3):12–24, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2019:NPP

[Hem19]  David Hemmendinger.  A

novel program for philosophers to study computer science. IEEE Annals of the History of Computing, 41(1):58–60, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8667953/.

Hemmendinger:2020:ED

[Hem20] D. Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 42 (4):5, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2021:ED

[Hem21] D. Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 43(1):5, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2022:EDa

[Hem22a] David Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 44(2):5, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2022:EDb

[Hem22b] David Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 44(3):5, July/September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2022:EDc

[Hem22c] David Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 44 (4):5, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2023:EDa

[Hem23a] David Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 45(3):5, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2023:EDb

[Hem23b] David Hemmendinger. From the Editor’s desk. IEEE Annals of the History of Computing, 45 (4):5, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hemmendinger:2024:CED

[Hem24a] David Hemmendinger. Correction to From the Editor’s Desk. IEEE Annals of the History of Computing, 46(2):90, April/ June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See [HA24].

Hemmendinger:2024:EDa

[Hem24b] David Hemmendinger. From the editor’s desk. IEEE Annals of the History of Computing, 46(1):5, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hensch:2005:IHFa

[Hen05]  Kurt Hensch.  IBM his-

tory of Far Eastern languages in computing, part 1: Requirements and initial phonetic product solutions in the 1960s. IEEE Annals of the History of Computing, 27(1):17–26, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1017.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1017.
pdf.

Henin:2010:TEI

[Hen10]  Silvio Henin.  Two early

Italian key-driven calculators. IEEE Annals of the History of Computing, 32(1):34–43, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heyck:2008:DCHa

[Hey08a] Hunter Heyck. Defining the computer: Herbert Simon and the bureaucratic mind–part 1. IEEE Annals of the History of Computing, 30(2):42– 51, April/June 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heyck:2008:DCHb

[Hey08b] Hunter Heyck. Defining the computer: Herbert Simon and the bureaucratic mind–part 2. IEEE Annals of the History of Computing, 30(2):52– 63, April/June 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heyck:2009:R

[Hey09]  Hunter Heyck.  Reviews.

IEEE Annals of the History of Computing, 31(1):72–73, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

House:2024:AOH

[HG24] Chuck House and Burt Grad.

Autodesk oral history. IEEE Annals of the History of Computing, 46(4):137–141, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hicks:2008:RTH

[Hic08] Marie Hicks. Repurposing Turing’s ‘human brake’. IEEE Annals of the History of Computing, 30(4):108, October/ December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hicks:2010:OCC

[Hic10] Marie Hicks. Only the clothes changed: Women operators in British computing and advertising, 1950–1970. IEEE Annals of the History of Computing, 32(4):5–17, October/ December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hicks:2013:BHC

[Hic13] Marie Hicks. De-brogramming the history of computing. IEEE Annals of the History of Computing, 35(1):88, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hicks:2019:HCT

[Hic19] Marie Hicks. Hacking the cis-tem. IEEE Annals of the History of Computing, 41(1):20–33, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8634814/.

Hiede:2012:EDa

[Hie12a] Lars Hiede. From the Editor’s desk. IEEE Annals of the History of Computing, 34(1):2– 3, January/March 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hiede:2012:EDb

[Hie12b] Lars Hiede. From the Editor’s desk. IEEE Annals of the History of Computing, 34 (2):2–3, April/June 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hensch:2005:IHFc

[HII^{+}05] Kurt Hensch, Toshiaki Igi, Masumi Iwao, Akira Oda, and Toru Takeshita. IBM history of Far Eastern languages in computing, part 3: IBM Japan taking the lead, accomplishments through the 1990s. IEEE Annals of the History of Computing, 27(1):38–55, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1038.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1038.
pdf.

Hensch:2005:IHFb

[HIIT05] Kurt Hensch, Toshiaki Igi, Masumi Iwao, and Toru Takeshita. IBM history of Far Eastern languages in computing, part 2: Initial efforts for full Kanji solutions, early 1970s. IEEE Annals of the History of Computing, 27(1):27–37, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1027.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1027.
pdf.

Hintz:2018:SKD

[Hin18]  Eric Hintz.  Susan Kare:

Design icon. IEEE Annals of the History of Computing, 40(2):48–61, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8415733/.

Hudson:2023:CTR

[HmWK23] Petera Hudson, H mi Whaanga, and Te Taka Keegan. Computing technologies for resilience, sustainability, and resistance. IEEE Annals of the History of Computing, 45(4):27–38, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Harris:2013:SDI

[HN13] Hershel Harris and Bert Nicol. SQL/DS: IBM’s first RDBMS. IEEE Annals of the History of Computing, 35(2):69–71, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Head:2023:LMK

[HO23] Jennifer Head and Dianne P. O’Leary. The legacy of Mary Kenneth Keller, first U.S. Ph.D. in computer science. IEEE Annals of the History of Computing, 45(1):55–63, January/March 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL https://arxiv.
                                        org/abs/2208.01765.

Hoare:2003:APP

[Hoa03] C. A. R. Hoare. Assertions:

a personal perspective. IEEE Annals of the History of Computing, 25(2):14–25, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2014.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2014.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/02/a2014abs.
htm.

Hockenberry:2022:LHC

[Hoc22] Matthew Hockenberry. On logistical histories of computing. IEEE Annals of the History of Computing, 44(1):135– 136, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hodges:2019:TFL

[Hod19] J. A. Hodges. Technical fixes for legal uncertainty in the 1980s software cracking scene. IEEE Annals of the History of Computing, 41(4):20–33, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Holmevik:1994:CSH

[Hol94] Jan Rune Holmevik. Compiling SIMULA: a historical study of technological genesis. IEEE Annals of the History of Computing, 16(4):25– 37, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4025.pdf;   http://www.
computer.org/annals/an1994/
a4025abs.htm.

Holst:1996:SRO

[Hol96] P. A. Holst. Svein Rosseland and the Oslo analyzer. IEEE Annals of the History of Computing, 18(4):16– 26, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539912.pdf;
https://en.wikipedia.org/
wiki/Oslo_Analyzer.

Holmes:1997:TA

[Hol97] N. Holmes. A tale of assembly. IEEE Annals of the History of Computing, 19(4):47– 49, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13659/00627899.pdf.

Hossain:2024:TSR

[Hos24] Anushah Hossain. Text standards for the “Rest of world”: The making of the Unicode standard and the OpenType format. IEEE Annals of the History of Computing, 46(1): 20–33, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

House:2021:RIR

[Hou21] Charles H. House. Review of IBM: The Rise and Fall and Reinvention of a Global Icon. IEEE Annals of the History of Computing, 43(2): 45–47, April/June 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Howlett:1999:ACL

[How99] J. Howlett. The Atlas Computer Laboratory. IEEE Annals of the History of Computing, 21(1):17–23, January/March 1999. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759365.pdf.

Holzmann:1996:EHD

[HP96] Gerald J. Holzmann and Bjorn Pehrson. The early history of data networks. IEEE Annals of the History of Computing, ??(??):??, ???? 1996. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2018:CP

[HP18] Thomas Haigh and Mark Priestley. Colossus and programmability. IEEE Annals of the History of Computing, 40(4):5–27, October/December 2018. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8509146/.

Hockenberry:2024:LHC

[HP24] Matthew Hockenberry and Miriam Posner. Logistical histories of computing. IEEE Annals of the History of Computing, 46(2):6–12, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2014:ABE

[HPR14a] T. Haigh, M. Priestley, and C. Rope. Los Alamos bets on ENIAC: Nuclear Monte Carlo simulations, 1947–1948. IEEE Annals of the History of Computing, 36(3):42–63, July/ September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2014:EXM

[HPR14b] Thomas Haigh, Mark Priestley, and Crispin Rope. Engineering “The miracle of the ENIAC”: Implementing the modern code paradigm. IEEE Annals of the History of Computing, 36(2): 41–59, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2014:RSP

[HPR14c] Thomas Haigh, Mark Priestley, and Crispin Rope. Reconsidering the stored-program concept. IEEE Annals of the History of Computing, 36(1):4– 17, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haderle:2013:HGI

[HS13] Donald J. Haderle and Cynthia M. Saracco. The history and growth of IBM’s DB2. IEEE Annals of the History of Computing, 35(2): 54–66, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hepler-Smith:2014:SP

[HS14] Evan Hepler-Smith. Simple problems. IEEE Annals of the History of Computing, 36(1): 88, January/March 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hsu:2020:OHA

[Hsu20]  H. Hsu.  Oral history of

Alan Cooper. IEEE Annals of the History of Computing, 42(4):100–118, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hsu:2023:OHR

[Hsu23] Hansen Hsu. Oral history of Rodney Brooks. IEEE Annals of the History of Computing, 45 (4):92–103, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Henin:2012:OID

[HT12] Silvio Henin and Massimo Temporelli. An original Italian dial adder rediscovered. IEEE Annals of the History of Computing, 34(2):49–59, April/ June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Huskey:1997:SDF

[HTAY97] H. D. Huskey, R. Thorensen, B. F. Ambrosio, and E. C. Yowell. The SWAC design features and operating experience. IEEE Annals of the History of Computing, 19 (2):46–50, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/12737/00586072.pdf.

Hayakawa:2023:MHI

[HUA23] Rieko Hayakawa, Robert Underwood, and Jennifer Anson. The modern history of ICT in Oceania PEACESAT and USP-Net. IEEE Annals of the History of Computing, 45(4):11– 26, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hume:1994:DSS

[Hum94] J. N. Patterson Hume. Development of systems software for the Ferut computer at the University of Toronto, 1952 to 1955. IEEE Annals of the History of Computing, 16(2):13– 19, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a2013.pdf;   http://www.
computer.org/annals/an1994/
a2013abs.htm.

Humphrey:2002:SUP

[Hum02] Watts S. Humphrey. Software unbundling: a personal perspective. IEEE Annals of the History of Computing, 24(1):59–63, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1059abs.htm;
http://dlib.computer.org/
an/books/an2002/pdf/a1059.
pdf.

Hunt:2012:UGC

[Hun12] Edward Hunt. US Government computer penetration programs and the implications for cyberwar. IEEE Annals of the History of Computing, 34(3):4–21, July/September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Huskey:1997:SSW

[Hus97]  H. D. Huskey.  SWAC

— Standards Western Automatic Computer: the Pioneer Day session at NCC July 1978. IEEE Annals of the History of Computing, 19 (2):51–61, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586073.pdf.

Henry:2004:SVD

[HVB04] Albert J. Henry, Henry Venture, and Glenn E. Bugos. Sydis and the voice/data terminal craze of 1984. IEEE Annals of the History of Computing, 26(2):22–33, April/ June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2022abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2022.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2022.
pdf.

Haigh:2003:B

[HW03a] Thomas Haigh and Maurice Wilkes. Biographies. IEEE Annals of the History of Computing, 25(4):89–92, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/04/a4089.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4089.pdf.

Haigh:2003:BOB

[HW03b] Thomas Haigh and Maurice Wilkes. Biographies: Obituary: I. Bernard Cohen. IEEE Annals of the History of Computing, 25 (4):89–92, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/04/a4089.
htm;   http://csdl.computer.
org/dl/mags/an/2003/04/a4089.
pdf.

Hocquet:2017:OIW

[HW17] Alexandre Hocquet and Frédéric Wieber. “Only the initiates will have the secrets revealed”: Computational chemists and the openness of scientific software. IEEE Annals of the History of Computing, 39(4): 40–58, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/8268025/

.

Hyman:1992:CQD

[Hym92]  Anthony Hyman.  Com-

ments, queries, and debate: The private Babbage. IEEE Annals of the History of Computing, 14(1):7, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1006.pdf.

Ibbett:1999:UMM

[Ibb99]  R. N. Ibbett.  The Uni-

versity of Manchester MU5 project. IEEE Annals of the History of Computing, 21(1): 24–33, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16436/00759366.pdf.

Ichikawa:2006:SFS

[Ich06] Hiroshi Ichikawa. Strela-1, the first Soviet computer: Political success and technological failure. IEEE Annals of the History of Computing, 28(3):18– 31, July/September 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Impagliazzo:1999:HCC

[ICKDL99] J. Impagliazzo, M. Campbell-Kelly, G. Davies, and J. A. N. Lee. History in the computing curriculum. IEEE Annals of the History of Computing, 21 (1):4–16, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16436/00759364.pdf.

Irvine:2001:EDC

[Irv01] M. M. Irvine. Early digital computers at Bell Telephone Laboratories. IEEE Annals of the History of Computing, 23(3):22–42, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/

annals/an2001/a3022abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3022.
pdf.

Jacobi:1993:CQD

[Jac93] George T. Jacobi. Comments, queries, and debate: Joe Desch. IEEE Annals of the History of Computing, 15(2):6, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2006.pdf.

Jack:2016:MPP

[Jac16] Caroline Jack. Meaning and persuasion: The personal computer and economic education. IEEE Annals of the History of Computing, 38(3):6–9, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jiuchun:2007:FCA

[JB07] Zhang Jiuchun and Zhang Baichun. Founding of the Chinese Academy of Sciences’ Institute of Computing Technology. IEEE Annals of the History of Computing, 29(1):16– 33, January/March 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jesiek:2007:PBH

[Jes07] Brent Jesiek. Pushing boundaries in the history of computing. IEEE Annals of the History of Computing, 29 (4):112, 110–111, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://bell.computer.
org/dlcomments/.

Jesiek:2013:OEH

[Jes13] Brent K. Jesiek. The origins and early history of computer engineering in the United States. IEEE Annals of the History of Computing, 35(3):6– 18, July/September 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jessen:2010:ATT

[JMSV10] Eike Jessen, Dieter Michel, Hans-Juergen Siegert, and Heinz Voigt. The AEG-Telefunken TR 440 computer: Company and large-scale computer strategy. IEEE Annals of the History of Computing, 32(3):20–29, July/September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jessen:2010:STD

[JMV10] Eike Jessen, Dieter Michel, and Heinz Voigt. Structure, technology, and development of the AEG-Telefunken TR 440 computer. IEEE Annals of the History of Computing, 32(3):30– 39, July/September 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Johansson:1996:EAC

[Joh96] M. Johansson. Early analog computers in Sweden — with examples from Chalmers University of Technology and the Swedish aerospace industry. IEEE Annals of the History of Computing, 18(4):27– 33, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539913.pdf.

Johnson:1998:VHS

[Joh98] L. Johnson. A view from the 1960s: how the software industry began. IEEE Annals of the History of Computing, 20(1):36–42, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646207.pdf.

Johansson:1999:BBG

[Joh99] M. Johansson. Big Blue gets beaten: the technological and political controversy of the first large Swedish computerization project in a rhetoric of technology perspective. IEEE Annals of the History of Computing, 21(2):14–30, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16485/00761791.pdf.

Johnson:2002:CSI

[Joh02] Luanne Johnson. Creating the software industry: Recollections of software company founders of the 1960s. IEEE Annals of the History of Computing, 24(1):14–42, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://computer.org/
                                        annals/an2002/a1014abs.htm;
                                         http://dlib.computer.org/
                                        an/books/an2002/pdf/a1014.
                                        pdf.

Johnson:2006:CGU

[Joh06] Lyle R. Johnson. Coming to grips with Univac. IEEE Annals of the History of Computing, 28(2):32–42, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Johnson:2020:PHS

[Joh20] L. Johnson. Preserving the history of the software industry. IEEE Annals of the History of Computing, 42(3):70– 75, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jones:2003:EST

[Jon03] Cliff B. Jones. The early search for tractable ways of reasoning about programs. IEEE Annals of the History of Computing, 25(2):26–49, April/ June 2003. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2026.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2026.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/02/a2026abs.
htm.

Jones:2004:FBB

[Jon04] Allan Jones. Five 1951 BBC broadcasts on automatic calculating machines. IEEE Annals of the History of Computing, 26(2):3–15, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2003abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2003.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2003.
pdf.

Jones:2019:DCC

[Jon19] M. L. Jones. The development of consent to computing. IEEE Annals of the History of Computing, 41(4):34–47, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Jorgensen:2009:CDF

[Jor09] Anker Helms Jorgensen. Context and driving forces in the development of the early computer game Nimbi. IEEE Annals of the History of Computing, 31(3):44–53, July/ September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kane:2021:RAW

[Kan21] Carolyn L. Kane. Review of Artificial Whiteness: Politics and Ideology in Artificial Intelligence. IEEE Annals of the History of Computing, 43(4):103– 104, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kaplan:1997:A

[Kap97]  E. Kaplan.  Anecdotes.

IEEE Annals of the History of Computing, 19(2):62– 69, April/June 1997. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/12737/00586074.pdf.

Kapor:2007:RLB

[Kap07] Mitch Kapor. Recollections on Lotus 1-2-3: Benchmark for spreadsheet software. IEEE Annals of the History of Computing, 29(3):32–40, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kasik:2024:DUC

[Kas24a] David J. Kasik. Developing and using CAD/CAM/CAE systems in Boeing. IEEE Annals of the History of Computing, 46 (4):38–49, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kasik:2024:IDC

[Kas24b] David J. Kasik. Industrydeveloped CAD/CAM software. IEEE Annals of the History of Computing, 46(4):27– 37, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Katz:1997:HCC

[Kat97] Kaila Katz. Historical content in computer science texts: a concern. IEEE Annals of the History of Computing, 19(1): 16–19, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560727.pdf.

Katz:2010:RPR

[Kat10] Randy H. Katz. RAID: a personal recollection of how storage became a system. IEEE Annals of the History of Computing, 32(4):82–87, October/ December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Katz:2018:OHD

[Kat18] Jeff Katz. Oral history of Dov Frohman. IEEE Annals of the History of Computing, 40 (4):52–63, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8540014/.

Kaufmann-Buhler:2021:CCD

[KB21] Jennifer Kaufmann-Buhler. Computing the cubicle: Design for the high-tech office, 1970–1990. IEEE Annals of the History of Computing, 43(3):6– 21, July/September 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Koomey:2011:IHT

[KBSW11] Jonathan G. Koomey, Stephen Berard, Marla Sanchez, and Henry Wong. Implications of historical trends in the electrical efficiency of computing. IEEE Annals of the History of Computing, 33(3):46–54, July/ September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kennedy:2020:ICC

[KC20] D. Kennedy and G. Con Diaz. Introduction to computing capitalisms. IEEE Annals of the History of Computing, 42(3):5– 10, July/September 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kennedy:2021:CCP

[KD21] Devin Kennedy and Gerardo Con Diaz. Computing capitalisms Part 2. IEEE Annals of the History of Computing, 43(2):5, April/June 2021.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Keet:2004:PRS

[Kee04]  Ernest E. Keet.  A per-

sonal recollection of software’s early days (1960–79), part 1. IEEE Annals of the History of Computing, 26 (4):46–61, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2004/04/a4046.
htm;   http://csdl.computer.
org/dl/mags/an/2004/04/a4046.
pdf.

Keet:2005:PRS

[Kee05] Ernest E. (“Lee”) Keet. A personal recollection of software’s early days (1960–1979): Part 2. IEEE Annals of the History of Computing, 27(4):31–45, October/December 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kelkar:2022:BAL

[Kel22] Shreeharsh Kelkar. Between AI and learning science: The evolution and commercialization of intelligent tutoring systems. IEEE Annals of the History of Computing, 44(1):20– 30, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kennedy:2022:RPM

[Ken22] Devin Kennedy. Review of Productivity Machines: German Appropriations of American Technology From Mass Production to Computer Automation by Corinna Schlombs. IEEE Annals of the History of Computing, 44(4):84–85, October/December 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2007:ESd

[KFS^{+}07] Chigusa Kita, David L. Ferro, Dag Spicer, Kelly Rodoski, Hiroyuki Seki, Fred R. Shapiro, and Thomas J. (Tim) Bergin. Events and sightings. IEEE Annals of the History of Computing, 29(4):76–81, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kahrs:2021:CTC

[KH21a] Mark Kahrs and Ann Hardy. Correction to Tymshare’s Changes to the Project Genie SDS-940 Operating System: a Historical Review. IEEE Annals of the History of Computing, 43(4):111, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See [KH21b].

Kahrs:2021:TCP

[KH21b] Mark Kahrs and Ann Hardy. Tymshare’s changes to the Project Genie SDS-940 operating system: a historical review. IEEE Annals of the History of Computing, 43(3):57–67, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). See correction [KH21a].

Kita:2008:ESb

[KHS^{+}08] Chigusa Kita, Thomas Haigh, Dag Spicer, Akihiko Yamada, and Thomas J. Misa. Events and sightings. IEEE Annals of the History of Computing, 30(2):64–67, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1993:Ra

[Kid93a] Peggy A. Kidwell. Reviews.

IEEE Annals of the History of Computing, 15(2):62–??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2062.pdf.

Kidwell:1993:Rb

[Kid93b] Peggy A. Kidwell. Reviews.

IEEE Annals of the History of Computing, 15(3):72–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3072.pdf.

Kidwell:1993:Rc

[Kid93c] Peggy A. Kidwell. Reviews. IEEE Annals of the History of Computing, 15(4):87–??, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a4087.pdf.

Kidwell:1994:Ra

[Kid94a] Peggy A. Kidwell. Reviews. IEEE Annals of the History of Computing, 16(1):76– ??, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a1076.pdf.

Kidwell:1994:Rb

[Kid94b] Peggy A. Kidwell. Reviews. IEEE Annals of the History of Computing, 16(2):73– ??, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a2073.pdf.

Kidwell:1994:Rc

[Kid94c]  Peggy A. Kidwell.  Re-

views. IEEE Annals of the History of Computing, 16(3): 70–??, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3070.pdf.

Kidwell:1994:Rd

[Kid94d] Peggy A. Kidwell. Reviews.

IEEE Annals of the History of Computing, 16(4):88– ??, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4088.pdf.

Kidwell:1995:R

[Kid95] Peggy Kidwell. Reviews. IEEE Annals of the History of Computing, 17(1):73–78, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1996:BAR

[Kid96a] Peggy Kidwell. BYTE, 20th anniversary report [reviews]. IEEE Annals of the History of Computing, 18(2):77, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489727.pdf.

Kidwell:1996:CWM

[Kid96b] Peggy Kidwell. Collected works of A. M. Turing — morphogenesis. IEEE Annals of the History of Computing, 18(4): 69, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539923.pdf.

Kidwell:1996:CPR

[Kid96c] Peggy Kidwell. Computing perspectives [reviews]. IEEE Annals of the History of Computing, 18(2):77, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489726.pdf.

Kidwell:1996:EHD

[Kid96d] Peggy Kidwell. The early history of data networks [reviews]. IEEE Annals of the History of Computing, 18(3):77, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511945.pdf.

Kidwell:1996:ER

[Kid96e] Peggy Kidwell. Enigma [reviews]. IEEE Annals of the History of Computing, 18(3):77, Fall 1996. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511946.pdf.

Kidwell:1996:ISV

[Kid96f] Peggy Kidwell. Information and secrecy: Vannevar Bush, Ultra and the other Memex. IEEE Annals of the History of Computing, 18(4):70, October/December 1996. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1996:TCU

[Kid96g] Peggy Kidwell. The trouble with computers: Usefulness, usability and productivity. IEEE Annals of the History of Computing, 18(4): 69, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539922.pdf.

Kidwell:1997:RDG

[Kid97a] Peggy Kidwell. Review: David G. Stork, Ed., HAL’s Legacy: 2001’s Computer as Dream and Reality. IEEE Annals of the History of Computing, 19(3): 78, July/September 1997. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1997:RNB

[Kid97b] Peggy Kidwell. Review: Nicolas Bion: The Construction and Principal Uses of Mathematical Instruments. IEEE Annals of the History of Computing, 19(2):80–83, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1998:SEC

[Kid98]  Peggy A. Kidwell.  Stalk-

ing the elusive computer bug. IEEE Annals of the History of Computing, 20(4):5– 9, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728224.pdf.

Kidwell:1999:BRM

[Kid99a] Peggy Kidwell. Book review: Michael Riordan and Lillian Doddeson, Crystal fire: the birth of the information age. IEEE Annals of the History of Computing, 21(1):78, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:1999:DAM

[Kid99b] Peggy Kidwell. Darwin among the machines: The evolution of global intelligence. IEEE Annals of the History of Computing, 21(1): 77–78, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759374.pdf.

Kidwell:1999:EME

[Kid99c] Peggy Kidwell. Engines of the mind: the evolution of computer from mainframes to microprocessors. IEEE Annals of the History of Computing, 21(1):76–77, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759373.pdf.

Kidwell:1999:JMH

[Kid99d] Peggy Kidwell. Journey to the moon: the history of the Apollo Guidance Computer. IEEE Annals of the History of Computing, 21(1): 78–79, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759376.pdf.

Kidwell:1999:DRE

[Kid99e] Peggy Aldrich Kidwell. Departments — reviews — electronic genie rescuing Prometheus; Forbes’ greatest technology stories; information ages; the mathematical theory of communication; LEO: The incredible story of the world’s first business; computer architects of the information society; Cryptologia; the supply of information technology workers in the US. IEEE Annals of the History of Computing, 21 (4):81–84, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801538.pdf.

Kidwell:2000:AMF

[Kid00a] Peggy A. Kidwell. The adding machine fraternity at St. Louis: creating a center of invention, 1880–1920. IEEE Annals of the History of Computing, 22(2):4–21, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841133.pdf.

Kidwell:2000:CR

[Kid00b] Peggy A. Kidwell. Cryptonomicon [reviews]. IEEE Annals of the History of Computing, 22(2):79, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841144.pdf.

Kidwell:2000:ICR

[Kid00c] Peggy A. Kidwell. Illustrators on the computer [reviews]. IEEE Annals of the History of Computing, 22(2):80, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841149.pdf.

Kidwell:2000:BRS

[Kid00d] Peggy A. Kidwell. In the beginning: recollections of software pioneers [reviews]. IEEE Annals of the History of Computing, 22(2):76–77, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841141.pdf.

Kidwell:2000:LBS

[Kid00e] Peggy A. Kidwell. Legal battles that shaped the computer industry [reviews]. IEEE Annals of the History of Computing, 22(2):77, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841142.pdf.

Kidwell:2000:MMP

[Kid00f] Peggy A. Kidwell. Making microchips: policy, globalization, and economic restructuring in the semiconductor industry [reviews]. IEEE Annals of the History of Computing, 22(2):75–76, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841140.pdf.

Kidwell:2000:NIM

[Kid00g] Peggy A. Kidwell. The near impossibility of making a microchip [reviews]. IEEE Annals of the History of Computing, 22(2):80, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:2000:NBH

[Kid00h]  Peggy A. Kidwell.  Nerds

2.0.1: a brief history of the Internet [reviews]. IEEE Annals of the History of Computing, 22(2):77–79, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841143.pdf.

Kidwell:2000:PLL

[Kid00i] Peggy A. Kidwell. Poor little Lisa, invention and technology [reviews]. IEEE Annals of the History of Computing, 22 (2):80, April/June 2000. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:2000:STC

[Kid00j]  Peggy A. Kidwell.  Sort-

ing things out: classification and its consequences [reviews]. IEEE Annals of the History of Computing, 22(2):80, April/ June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:2000:SNC

[Kid00k] Peggy A. Kidwell. The Swiss Nema cipher machine [reviews]. IEEE Annals of the History of Computing, 22(2):80, April/ June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:2000:WYN

[Kid00l]  Peggy A. Kidwell.  Why

you need to understand Y2k [reviews]. IEEE Annals of the History of Computing, 22(2):80, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/18190/00841150.pdf.

Kidwell:2001:Rc

[Kid01a]  Peggy Kidwell.  Reviews.

IEEE Annals of the History of Computing, 23(3):78–81, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3078abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3078.
pdf.

Kidwell:2001:Ra

[Kid01b] Peggy Aldrich Kidwell. Reviews. IEEE Annals of the History of Computing, 23(1):62–63, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an2001/pdf/a1062.
pdf.

Kidwell:2001:Rb

[Kid01c] Peggy Aldrich Kidwell. Reviews. IEEE Annals of the History of Computing, 23(2):74– 75, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2074.pdf.

Kidwell:2001:YIA

[Kid01d] Peggy Aldrich Kidwell. ‘Yours for improvement’ — the adding machines of Chicago, 1884– 1930. IEEE Annals of the History of Computing, 23(3):3–21, July/September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3003abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a3003.
pdf.

Kidwell:2002:R

[Kid02]  Peggy Kidwell.  Reviews.

IEEE Annals of the History of Computing, 24(4):48–52, October/December 2002. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2002/04/a4048.htm;   http:
//csdl.computer.org/dl/mags/
an/2002/04/a4048.pdf.

Kidwell:2015:UIP

[Kid15] Peggy Aldrich Kidwell. Useful instruction for practical people: Early printed discussions of the slide rule in the US. IEEE Annals of the History of Computing, 37(1):36–43, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kidwell:2019:FT

[Kid19a] Peggy Aldrich Kidwell. Flowcharting templates. IEEE Annals of the History of Computing, 41(1):55–57, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8667955/.

Kidwell:2019:RGS

[Kid19b] Peggy Aldrich Kidwell. The role of governments in the spread of novel computing devices in the Nineteenth and early Twentieth Century United States. IEEE Annals of the History of Computing, 41(1):7–19, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8632739/.

Kidwell:2023:PCP

[Kid23] Peggy Aldrich Kidwell. Promoting computing in the postwar United States: The case of UCLA. IEEE Annals of the History of Computing, 45(2): 43–52, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kirsch:1998:SSI

[Kir98] R. A. Kirsch. SEAC and the start of image processing at the National Bureau of Standards. IEEE Annals of the History of Computing, 20(2):7–13, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667290.pdf.

Kirstein:1999:EEA

[Kir99] P. T. Kirstein. Early experiences with the Arpanet and Internet in the United Kingdom. IEEE Annals of the History of Computing, 21(1): 38–44, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759368.pdf.

Kirsch:2020:TPT

[Kir20] C. J. Kirsch. A three-person teaching machine designed for crisis: The Geromat III in Berlin and Ulm. IEEE Annals of the History of Computing, 42 (4):39–52, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kistermann:1992:CQDb

[Kis92a] F. W. Kistermann. Comments, queries, and debate: More on mechanical delay lines. IEEE Annals of the History of Computing, 14(4):7, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4007.pdf.

Kistermann:1992:CQDa

[Kis92b] F. W. Kistermann. Comments, queries, and debate: The size of the Hollerith punched card. IEEE Annals of the History of Computing, 14(3):7, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3006.pdf.

Kistermann:1995:WFA

[Kis95]  F. W. Kistermann.  The

way to the first automatic sequence-controlled calculator: the 1935 DEHOMAG D 11 tabulator. IEEE Annals of the History of Computing, 17(2):33–49, Summer 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8635/00380270.pdf.

Kistermann:1997:LVN

[Kis97] F. W. Kistermann. Locating the victims: the nonrole of punched card technology and census work. IEEE Annals of the History of Computing, 19(2):31–45, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586070.pdf.

Kistermann:1998:BPA

[Kis98] F. W. Kistermann. Blaise Pascal’s adding machine: new findings and conclusions. IEEE Annals of the History of Computing, 20(1):69–76, January/March 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646211.pdf.

Kistermann:1999:LWP

[Kis99a] F. W. Kistermann. Leo Wenzel Pollak (1888–1964): Czechoslovakian pioneer in scientific data processing. IEEE Annals of the History of Computing, 21 (4):62–68, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17389/00801534.pdf.

Kistermann:1999:WCA

[Kis99b] F. W. Kistermann. When could anyone have seen Leibniz’s stepped wheel? IEEE Annals of the History of Computing, 21(2):68–72, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16485/00761796.pdf.

Kistermann:2001:CHU

[Kis01] Friedrich W. Kistermann. Calculators: How to use the Schickard calculator. IEEE Annals of the History of Computing, 23(1):80–85, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2001/pdf/
                                        a1080.pdf.

Kistermann:2005:HPC

[Kis05] Friedrich W. Kistermann. Hollerith punched card system development (1905–1913).

IEEE Annals of the History of Computing, 27(1):56–66, January/March 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2005/01/a1056.htm;   http:
//csdl.computer.org/dl/mags/
an/2005/01/a1056.pdf.

Kita:2003:JCR

[Kit03] Chigusa Ishikawa Kita. J.

C. R. Licklider’s vision for the IPTO. IEEE Annals of the History of Computing, 25(3):62–77, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3062.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3062.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/03/a3062abs.
htm.

Kita:2005:ESa

[Kit05a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 27(1):74–77, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1074.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1074.
pdf.

Kita:2005:ESb

[Kit05b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 27(2):79–81, April/June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2079.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/02/a2079.
                                        pdf.

Kita:2005:ESc

[Kit05c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 27(3): 89–92, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2005:ESd

[Kit05d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 27 (4):81–85, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2006:ESa

[Kit06a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 28(1): 89–92, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2006:ESb

[Kit06b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 28 (2):63–69, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2006/02/a2063.pdf.

Kita:2006:ESc

[Kit06c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 28 (3):86–88, 95, July/September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2006:ESd

[Kit06d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 28 (4):92–95, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2007:ESa

[Kit07] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 29(1): 81–84, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2008:ESc

[Kit08a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 30(3): 74–77, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2008:LC

[Kit08b] Chigusa Kita. The LINC computer at 45. IEEE Annals of the History of Computing, 30 (4):92–97, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2009:ESa

[Kit09a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 31(1): 62–67, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2009:ESb

[Kit09b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 31(2): 76–79, April/June 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2009:ESc

[Kit09c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 31(3): 78–82, July/September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2009:ESd

[Kit09d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 31(4): 107–109, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2010:ESa

[Kit10a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 32(1): 80–81, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2010:ESb

[Kit10b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 32(2): 84–86, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2010:ESc

[Kit10c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 32(3): 90–93, July/September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2010:ESd

[Kit10d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 32 (4):90–93, October/December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2011:ESa

[Kit11a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 33(1):

78–82, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2011:ESb

[Kit11b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 33(2): 106–110, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2011:ESc

[Kit11c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 33(3): 82–85, July/September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2011:ESd

[Kit11d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 33 (4):84–85, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2012:ESa

[Kit12a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 34(1): 88–91, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2012:ESb

[Kit12b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 34(2): 60–69, April/June 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2012:ESO

[Kit12c] Chigusa Kita. Events and sightings [obituary of Lawrence A. Welke]. IEEE Annals of the History of Computing, 34 (4):83–85, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2013:ESa

[Kit13a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 35(1): 67–71, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2013:ESb

[Kit13b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 35(2): 76–81, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2013:ESc

[Kit13c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 35(3): 72–76, July/September 2013.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2013:ESd

[Kit13d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 35 (4):82–85, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2014:ESc

[Kit14a] C. Kita. Events and sightings. IEEE Annals of the History of Computing, 36(3):85– 89, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2014:ESa

[Kit14b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 36(1): 82–84, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2014:ESb

[Kit14c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 36(2): 77–78, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2014:ESd

[Kit14d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 36 (4):92–96, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2015:ESa

[Kit15a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 37(1): 83–86, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2015:ESb

[Kit15b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 37 (2):94–96, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//www.computer.org/csdl/
mags/an/2015/02/man2015020094.
html.

Kita:2015:ESc

[Kit15c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 37 (3):4–7, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//www.computer.org/csdl/
mags/an/2015/03/man2015030004-
abs.html.

Kita:2015:ESd

[Kit15d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 37 (4):87–89, ???? 2015. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2016:ESa

[Kit16a] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 38(1): 89–93, January/March 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kita:2016:ESb

[Kit16b] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 38(2): 6–7, April/June 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2016:ESc

[Kit16c] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 38 (3):88, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2016:ESd

[Kit16d] Chigusa Kita. Events and sightings. IEEE Annals of the History of Computing, 38 (4):3–5, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kunze:2022:HFL

[KK22] Fritz Kunze and Lauren Kunze. The history of Franz and Lisp.

IEEE Annals of the History of Computing, 44(1):103–108, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Koanantakool:2009:CTL

[KKW09] Hugh Thaweesak Koanantakool, Theppitak Karoonboonyanan, and Chai Wutiwiwatchai. Computers and the Thai language. IEEE Annals of the History of Computing, 31(1):46–61, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Klimenko:1999:CSR

[Kli99] S. V. Klimenko. Computer science in Russia: a personal view. IEEE Annals of the History of Computing, 21(3): 16–30, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778979.pdf.

Kline:2011:CAS

[Kli11] Ronald R. Kline. Cybernetics, automata studies, and the Dartmouth Conference on Artificial Intelligence. IEEE Annals of the History of Computing, 33(4):5–16, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kluver:1999:RIC

[Klu99] P. V. Kluver. From research institute to computer company: Regnecentralen 1946– 1964. IEEE Annals of the History of Computing, 21 (2):31–43, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16485/00761793.pdf.

Kita:2008:ESa

[KMA^{+}08] Chigusa Kita, Thomas J. Misa, William Aspray, David Walden, Robert M. Price, and Eric A. Weiss. Events and sightings. IEEE Annals of the History of Computing, 30(1): 82–86, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kneuper:2017:SYS

[Kne17] Ralf Kneuper. Sixty years of software development life cycle models. IEEE Annals of the History of Computing, 39(3):41–54, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/03/
                                        man2017030041-abs.html.

Kneese:2021:RYC

[Kne21] Tamara Kneese. Review of Your Computer is on Fire. IEEE Annals of the History of Computing, 43(3):81–82, July/ September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Koss:2003:PBP

[Kos03a] Adele Mildred Koss. Programming at Burroughs and Philco in the 1950s. IEEE Annals of the History of Computing, 25(4):40–50, October/ December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4040abs.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4040.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4040.pdf.

Koss:2003:PUW

[Kos03b] Adele Mildred Koss. Programming on the Univac 1: a woman’s account. IEEE Annals of the History of Computing, 25(1):48–59, January/March 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/01/a1048.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1048.pdf;   http:
//csdl.computer.org/dl/mags/
an/2003/01/a1048abs.htm.

Koss:2007:CRH

[Kos07] Adele Mildred Koss. CDC, Raytheon, and Harvard University: Three early data management systems. IEEE Annals of the History of Computing, 29(2):55–65, April/June 2007.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Krapp:2022:RMA

[Kra22] Peter Krapp. Review of Milestones in Analog and Digital Computing, 3rd ed by Herbert Bruderer. IEEE Annals of the History of Computing, 44(2): 76–77, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Krull:1994:OCG

[Kru94] Fred N. Krull. The origin of computer graphics within General Motors. IEEE Annals of the History of Computing, 16 (3):40, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3040.pdf;   http://www.
                                        computer.org/annals/an1994/
                                        a3040abs.htm.

Krueger:2007:MCI

[Kru07] Tara Krueger. Mapping cyberspace: The image of the Internet. IEEE Annals of the History of Computing, 29(1): 100, c3, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Kruntorad:2009:HCI

[Kru09] Judith M. Kruntorad. History of the CA IDMS Database Management System. IEEE Annals of the History of Computing, 31(4):100–106, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

King:1997:FHH

[KS97] J. King and W. A. Shelly.

A family history of Honeywell’s large-scale computer systems. IEEE Annals of the History of Computing, 19(4):42– 46, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13659/00627898.pdf.

Kita:2007:ESb

[KSFY07] Chigusa Kita, Dag Spicer, Anne Fitzpatrick, and Akihiko Yamada. Events and sightings. IEEE Annals of the History of Computing, 29(2):66– 69, April/June 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kirkpatrick:2019:RRP

[KW19] Bill Kirkpatrick and Jacob Ward. Reviews of recent publications [two books reviewed]. IEEE Annals of the History of Computing, 41(1):51–54, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kita:2007:ESc

[KYG^{+}07] Chigusa Kita, Jeffrey R. Yost, Burt Grad, Thomas J. (Tim) Bergin, and Mike Marcus. Events and sightings. IEEE Annals of the History of Computing, 29(3):62–66, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lawson:2012:R

[LA12] Sean Lawson and Jon Agar. Reviews. IEEE Annals of the History of Computing, 34(3): 67–69, July/September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Land:2000:FBC

[Lan00] F. Land. The first business computer: a case study in user-driven innovation. IEEE Annals of the History of Computing, 22(3): 16–26, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18655/00859523.pdf.

Laprise:2009:TPP

[Lap09] John Laprise. Think piece: The purloined mainframe: Hiding the history of computing in plain sight. IEEE Annals of the History of Computing, 31(3): 83–84, July/September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Laprise:2011:KCN

[Lap11] John Laprise. Kissinger’s computer: National Security Council computerization, 1969– 1972. IEEE Annals of the History of Computing, 33(1):38– 51, January/March 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lathrop:2013:DOF

[Lat13] Jay W. Lathrop. The Diamond Ordnance Fuze Laboratory’s photolithographic approach to microcircuits. IEEE Annals of the History of Computing, 35(1):48–55, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lavington:1993:MCA

[Lav93] Simon H. Lavington. Manchester computer architectures, 1948–75. IEEE Annals of the History of Computing, 15(3):44–54, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3044.pdf;   http://www.
computer.org/annals/an1993/
a3044abs.htm.

Lavington:2006:FCD

[Lav06] Simon Lavington. In the footsteps of Colossus: a description of Oedipus. IEEE Annals of the History of Computing, 28(2):44–55, April/June 2006.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lavington:2022:TMC

[Lav22a] Simon Lavington. Two Manchester computer milestones. IEEE Annals of the History of Computing, 44(3):91–92, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lavington:2022:EDC

[Lav22b] Simon H. Lavington. Early days of computing at Manchester: Max Newman’s Royal Society Project, 1946–1951. IEEE Annals of the History of Computing, 44(2):20–32, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

LaViolette:2024:WTF

[LaV24] Marc LaViolette. Was there a French engine before Babbage’s Difference Engine? IEEE Annals of the History of Computing, 46(1):62–65, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Laws:2010:CLL

[Law10] David A. Laws. A company of legend: The legacy of Fairchild Semiconductor. IEEE Annals of the History of Computing, 32(1):60–74, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lazear:2024:CFT

[Laz24] Tom Lazear. CAD at Fluor, T&W/Versacad, Computervision, and Archway. IEEE Annals of the History of Computing, 46(4):114–123, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lecuyer:2006:B

[LB06] Christophe Lecuyer and David C.

Brock. Biographies. IEEE Annals of the History of Computing, 28(3):89–95, July/ September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:2000:UBN

[LBA00] J. A. N. Lee, C. Burke, and D. Anderson. The US Bombes, NCR, Joseph Desch, and 600 WAVES: the first reunion of the US Naval Computing Machine Laboratory. IEEE Annals of the History of Computing, 22(3):27–41, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18655/00859524.pdf.

Lea:2023:DDH

[Lea23] Andrew S. Lea. Dissecting data: History of data as history of the body. IEEE Annals of the History of Computing, 45(4):104–106, October/December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lecuyer:2021:RLO

[Léc21] C. Lécuyer. Review of Luc Olivier Bauer and E. Marshall Wilder, The Microchip Revolution: A Brief History (Amazon, 2020). IEEE Annals of the History of Computing, 43(1):95– 96, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:1992:AIa

[Lee92a] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 14(1):3–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a1003.pdf.

Lee:1992:AIb

[Lee92b] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 14(2):3– ??, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a2003.pdf.

Lee:1992:AIc

[Lee92c] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 14(3):3–??, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.

org/an/books/an1992/pdf/
a3003.pdf.

Lee:1992:AId

[Lee92d] J. A. N. Lee. About this issue.

IEEE Annals of the History of Computing, 14(4):3–??, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4003.pdf.

Lee:1992:CTT

[Lee92e] J. A. N. Lee. Claims to the term “Time-Sharing”. IEEE Annals of the History of Computing, 14(1):16–54, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4016.pdf;   http://www.
computer.org/annals/an1992/
a1016abs.htm.

Lee:1992:CCT

[Lee92f]  J. A. N. Lee.  CTSS

— the Compatible Time-Sharing System. IEEE Annals of the History of Computing, 14(1):31–54, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1031.pdf.

Lee:1992:ITL

[Lee92g]  J. A. N. Lee.  Introduc-

tion: Time line. IEEE Annals of the History of Computing, 14(1):13–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a1013.pdf.

Lee:1992:PF

[Lee92h] J. A. N. Lee. Prolog to the future. IEEE Annals of the History of Computing, 14(2):42– 50, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a2042.pdf.

Lee:1992:SI

[Lee92i] J. A. N. Lee. The social impact. IEEE Annals of the History of Computing, 14(2):36– 41, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a2036.pdf.

Lee:1993:AIa

[Lee93a] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 15(1):3–??, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a1003.pdf.

Lee:1993:AIb

[Lee93b] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 15(2):3– ??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2003.pdf.

Lee:1993:AIc

[Lee93c] J. A. N. Lee. About this issue.

IEEE Annals of the History of Computing, 15(4):3–??, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4003.pdf.

Lee:1993:BWM

[Lee93d]  J. A. N. Lee.  Birthday

wishes to Maurice V. Wilkes. IEEE Annals of the History of Computing, 15(3):5–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3005.pdf.

Lee:1994:AIa

[Lee94a] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 16(1): 3–??, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.

                                        org/an/books/an1994/pdf/
                                        a1003.pdf.

Lee:1994:AIb

[Lee94b] John A. N. Lee. About this issue. IEEE Annals of the History of Computing, 16(3): 3–??, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3003.pdf.

Lee:1995:AIa

[Lee95a] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 17(1):3, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:1995:AIb

[Lee95b] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 17(4):3, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10129/00477429.pdf.

Lee:1995:JPE

[Lee95c] J. A. N. Lee. J. Presper Eckert, 1919–1995. IEEE Annals of the History of Computing, 17(3):3, 5, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397053.pdf.

Lee:1995:JVA

[Lee95d] J. A. N. Lee. John V. Atanasoff, 1903–1995. IEEE Annals of the History of Computing, 17(3):3, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397052.pdf.

Lee:1995:BKH

[Lee95e] J. A. N. Lee. On “Babbage and Kings” and “How sausage was made”: and now for the rest of the story. IEEE Annals of the History of Computing, 17(4):7–23, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477433.pdf.

Lee:1995:RFG

[Lee95f]  J. A. N. Lee.  The rise

and fall of the General Electric Corporation computer department. IEEE Annals of the History of Computing, 17(4):24–45, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477434.pdf.

Lee:1996:WFL

[Lee96] J. A. N. Lee. “Those who forget the lessons of history are doomed to repeat it”: or, why I study the history of computing. IEEE Annals of the History of Computing, 18(2):54–62, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489724.pdf.

Lee:1997:KSD

[Lee97] J. A. N. Lee. King of the Seven Dwarfs, General Electric’s ambiguous challenge to the computer industry [book reviews]. IEEE Annals of the History of Computing, 19(1): 78–79, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560749.pdf.

Lee:1998:RFC

[Lee98a] J. Lee. Richard F. Clippinger: 1913–1997. IEEE Annals of the History of Computing, 20(2):59, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667308.pdf.

Lee:1998:RWH

[Lee98b] Jan Lee. Richard Wesley Hamming: 1915–1998. IEEE Annals of the History of Computing, 20(2):60–62, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667309.pdf.

Lee:1999:B

[Lee99a]  J. A. N. Lee.  Biogra-

phies. IEEE Annals of the History of Computing, 21(1): 48–65, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759370.pdf.

Lee:1999:JYF

[Lee99b]  J. A. N. Lee.  Jeffrey

Young, Forbes’ Greatest Technology Stories [book review]. IEEE Annals of the History of Computing, 21(4):82– 83, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801541.pdf.

Lee:2000:AAC

[Lee00a]  J. A. N. Lee.  Ada Au-

gusta, Countess of Lovelace, 1815–1852 what was her family name? IEEE Annals of the History of Computing, 22 (4):72–73, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887999.pdf.

Lee:2000:B

[Lee00b]  J. A. N. Lee.  Biogra-

phies. IEEE Annals of the History of Computing, 22(1): 86–87, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17665/00815469.pdf.

Lee:2000:HAT

[Lee00c] J. A. N. Lee. Howard Aiken’s third machine: the Howard Mark III calculator or Aiken-Dahlgren electronic calculator. IEEE Annals of the History of Computing, 22(1): 62–81, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17665/00815467.pdf.

Lee:2000:O

[Lee00d]  J. A. N. Lee.  Obitu-

ary. IEEE Annals of the History of Computing, 22 (2):67–68, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841139.pdf.

Lee:2001:Ba

[Lee01a] J. A. N. Lee. Biographies. IEEE Annals of the History of Computing, 23(2):80–??, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2080.pdf.

Lee:2001:Bb

[Lee01b] J. A. N. Lee. Biographies.

IEEE Annals of the History of Computing, 23(3):82–86, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3082abs.htm;
http://dlib.computer.org/
an/books/an2001/pdf/a3082.
pdf.

Lee:2001:BOJb

[Lee01c]  J. A. N. Lee.  Biogra-

phies: Obituaries: [John W. Haanstra]. IEEE Annals of the History of Computing, 23(1):82–87, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1067.pdf.

Lee:2001:BOJa

[Lee01d]  J. A. N. Lee.  Biogra-

phies: [obituaries: John Weber Carr III & Joan Elisabeth Lowther Clarke Murray]. IEEE Annals of the History of Computing, 23(1):67–75, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a1067.pdf.

Lee:2002:B

[Lee02]  J. A. N. Lee.  Biogra-

phies. IEEE Annals of the History of Computing, 24 (4):53–55, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4053.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4053.
pdf.

Leimbach:2008:SSE

[Lei08] Timo Leimbach. The SAP story: Evolution of SAP within the German software industry. IEEE Annals of the History of Computing, 30(4):60–76, October/December 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Leimbach:2017:PCS

[Lei17] Timo Leimbach. Path creation in the software industry — the case of Software AG. IEEE Annals of the History of Computing, 39(4):59– 80, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/8267986/

.

Lenat:2022:CMR

[Len22] Douglas Lenat. Creating a 30million-rule system: MCC and Cycorp. IEEE Annals of the History of Computing, 44(1): 44–56, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Levy:2005:HTT

[Lev05] Stephen Levy. History of technology transfer at BBN. IEEE Annals of the History of Computing, 27(2):30–38, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2030.
htm;   http://csdl.computer.
org/dl/mags/an/2005/02/a2030.
pdf.

Lewendal:2014:MCT

[Lew14] B. Lewendal. My corner of the time-sharing innovation world. IEEE Annals of the History of Computing, 36(4):97–101, October/December 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lewis:2016:PTC

[Lew16]  Nicholas Lewis.  Peering

through the curtain: Soviet computing through the eyes of Western experts. IEEE Annals of the History of Computing, 38(1):34–47, January/March 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lewis:2017:PPR

[Lew17]  Nicholas Lewis.  Purchas-

ing power: Rivalry, dissent, and computing strategy in supercomputer selection at Los Alamos. IEEE Annals of the History of Computing, 39(3):25–40, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/03/
                                        man2017030025-abs.html.

Lee:1992:PM

[LFS^{+}92] J. A. N. Lee, R. M. Fano, A. L. Scherr, F. J. Corbató, and V. A. Vyssotsky. Project MAC. IEEE Annals of the History of Computing, 14(2):9– 13, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a2009.pdf.

Lee:1995:YAB

[LH95] John A. N. Lee and Golde Holtzman. 50 years after breaking the codes: Interviews with two of the Bletchley Park scientists. IEEE Annals of the History of Computing, 17(1): 32–43, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:2003:BPB

[LH03] J. A. N. Lee and Thomas Haigh. Biographies: Per Brinch Hansen. IEEE Annals of the History of Computing, 25(1): 80–83, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lane:2025:ERR

[LHS^{+}25] Rupert Lane, Anthony Hay, Arthur Schwarz, David M. Berry, and Jeff Shrager. ELIZA reanimated: Restoring the mother of all chatbots to one of the world’s first time-sharing systems. IEEE Annals of the History of Computing, 47(2): 68–76, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Light:2003:BIT

[Lig03] Jennifer S. Light. Before the Internet, there was cable. IEEE Annals of the History of Computing, 25(2):96, 94–95, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2096.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2096.
pdf.

Lin:2016:DEF

[Lin16] Ling-Fei Lin. Design engineering or factory capability? Building laptop contract manufacturing in Taiwan. IEEE Annals of the History of Computing, 38(2):22–39, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lipner:2015:BDO

[Lip15] Steven B. Lipner. The birth and death of the Orange Book. IEEE Annals of the History of Computing, 37(2):19–31, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020019-abs.html.

Lison:2024:HSS

[Lis24] Andrew Lison. Hardware standardization and state-socialist piracy: The global reach of the Zilog Z80. IEEE Annals of the History of Computing, 46(2): 38–51, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Llach:2024:CAD

[Lla24] Daniel Cardoso Llach. Computeraided design in the United States, 1949–1984: Designing in a closed world. IEEE Annals of the History of Computing, 46 (4):10–26, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Luebke:1994:LVO

[LM94] David Martin Luebke and Sybil Milton. Locating the victim: An overview of censustaking, tabulation technology, and persecution in Nazi Germany. IEEE Annals of the History of Computing, 16 (3):25, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a3025.pdf;   http://www.
computer.org/annals/an1994/
a3025abs.htm.

Laplante:2006:A

[LM06] Phillip A. Laplante, Jr. and Stanley Mazor. Anecdotes. IEEE Annals of the History of Computing, 28(2):70–76, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:1992:B

[LML92] J. A. N. Lee, John Mc-Carthy, and J. C. R. Licklider. The beginnings at MIT. IEEE Annals of the History of Computing, 14(1):18–54, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/

a1018.pdf. See correction [Ros92].

Logue:1998:VTV

[Log98] J. C. Logue. From vacuum tubes to very large scale integration: a personal memoir. IEEE Annals of the History of Computing, 20(3): 55–68, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707575.pdf.

Longo:2004:EBC

[Lon04]  Bernadette Longo.  Ed-

mund Berkeley, computers, and modern methods of thinking. IEEE Annals of the History of Computing, 26 (4):4–18, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/04/a4004.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/04/a4004.
                                        pdf.

Longo:2008:MCD

[Lon08] Bernadette Longo. Mathematics, computer development, and science policy debates after World War II. IEEE Annals of the History of Computing, 30(3):64–72, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lorenzen:2019:VS

[Lor19] Lee Lorenzen. The Ventura story. IEEE Annals of the History of Computing, 41(3): 75–88, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lowood:2009:VCS

[Low09] Henry Lowood. Videogames in computer space: The complex history of Pong. IEEE Annals of the History of Computing, 31(3):5–19, July/September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:1992:PMI

[LR92] J. A. N. Lee and Robert Rosin. The Project MAC interviews. IEEE Annals of the History of Computing, 14(2):14– 35, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a2014.pdf;   http://www.
computer.org/annals/an1992/
a2014abs.htm.

Laws:2012:MMD

[LR12] David A. Laws and Michael Riordan. Making micrologic: The development of the planar IC at Fairchild Semiconductor, 1957– 1963. IEEE Annals of the History of Computing, 34(1):20– 36, January/March 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lee:2000:RSG

[LS00] J. A. N. Lee and G. E. Snively.

The rise and sale of the General Electric Computer Department: a further look. IEEE Annals of the History of Computing, 22(2):53–60, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841137.pdf.

Larsson:2008:VSC

[LS08] Thomas Larsson and Ulf Stenlund. Vignettes of Sweden’s computer graphics history. IEEE Annals of the History of Computing, 30(1):50– 58, January/March 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lukasik:2011:WAW

[Luk11] Stephen J. Lukasik. Why the Arpanet was built. IEEE Annals of the History of Computing, 33(3):4–21, July/ September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lundh:2018:SNC

[Lun18] Yngvar Lundh. A slice of Norway’s computing history. IEEE Annals of the History of Computing, 40(2):62–72, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8415734/.

Lussier:2021:PPC

[Lus21] Kira Lussier. From papers to programs: Courts, corporations, clinics, and the battle over computerized psychological testing. IEEE Annals of the History of Computing, 43(2):30–41, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Marcus:1996:EAE

[MA96] Mitchell Marcus and Atsushi Akera. Exploring the architecture of an early machine: The historical relevance of the ENIAC machine architecture. IEEE Annals of the History of Computing, 18(1):17–24, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476558.pdf.

Mackenzie:1995:APH

[Mac95] D. Mackenzie. The automation of proof: a historical and sociological exploration. IEEE Annals of the History of Computing, 17(3):7–29, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397057.pdf.

Machover:1998:IAV

[Mac98]  C. Machover.  An inter-

view with Andries Van Dam. IEEE Annals of the History of Computing, 20(2):81– 84, April/June 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667314.pdf.

Macduff:2003:HLC

[Mac03] Valerie A. G. Macduff. Half a lifetime in computing: Experiences in Zambia, Britain, and Australia. IEEE Annals of the History of Computing, 25(4):64–66, October/ December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4064abs.htm;   http:

                                        //csdl.computer.org/dl/mags/
                                        an/2003/04/a4064.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/04/a4064.pdf.

Maguire:2009:ASA

[Mag09] John Norris Maguire. Adabas: Software AG of North America. IEEE Annals of the History of Computing, 31(4):92–99, October/December 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Magoun:2020:ROI

[Mag20] A. B. Magoun. Resources and opportunities of the IEEE History Center and ETHW. IEEE Annals of the History of Computing, 42(4):122–123, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mahoney:2004:FHS

[Mah04] Michael S. Mahoney. Finding a history for software engineering. IEEE Annals of the History of Computing, 26 (1):8–19, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2004/01/a1008abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/01/a1008.
                                        pdf.

Mahoney:2008:WMH

[Mah08] Michael S. Mahoney. What makes the history of software hard. IEEE Annals of the History of Computing, 30(3):8– 18, July/September 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mailland:2016:OAM

[Mai16]  Julien Mailland.  101 on-

line: American Minitel network and lessons from its failure. IEEE Annals of the History of Computing, 38(1):6– 22, January/March 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Makhoul:2006:SPB

[Mak06] John Makhoul. Speech processing at BBN. IEEE Annals of the History of Computing, 28(1):32–45, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Malone:2002:IIR

[Mal02] Cheryl Knott Malone. Imagining information retrieval in the library: Desk set in historical context. IEEE Annals of the History of Computing, 24(3):14–22, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3014.pdf;   http://www.
computer.org/annals/an2002/
a3014abs.htm.

Marcum:2002:ALC

[Mar02] Deanna B. Marcum. Automating the library: The Council on Library Resources. IEEE Annals of the History of Computing, 24(3):2–13, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2002/pdf/
                                        a3002.pdf;   http://www.
                                        computer.org/annals/an2002/
                                        a3002abs.htm.

Marchant:2010:DHY

[Mar10a] Josephine Marchant. Decoding the heavens: a 2,000year-old-computer — and the century-long search to discover its secrets. Da Capo Press, Cambridge, MA, USA, 2010. ISBN 0-306-81861-2 (paperback). 328 + 8 pp. LCCN QB107 .M38 2010. URL http:

                                        //www.decodingtheheavens.
                                        com/reviews.aspx.

Marchant:2010:DHS

[Mar10b]  Josephine Marchant.  De-

coding the heavens: solving the mystery of the world’s first computer. William Heinemann, London, UK, 2010. ISBN 0-434-01835-X (hardcover). 328 + 8 pp. LCCN QB107 .M37 2010.

                                        URL http://site.ebrary.
                                        com/id/10372916;   http:/
                                        /www.decodingtheheavens.
                                        com/reviews.aspx.

Martin:2012:TFA

[Mar12] Ian Martin. Too far ahead of its time: Barclays, Burroughs, and real-time banking. IEEE Annals of the History of Computing, 34(2):5–19, April/ June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mazor:2007:ICC

[Maz07] Stanley Mazor. Intel 8080 CPU chip development. IEEE Annals of the History of Computing, 29(2):70–73, April/June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mazor:2008:FSC

[Maz08] Stanley Mazor. Fairchild symbol computer. IEEE Annals of the History of Computing, 30(1):92–95, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mazor:2010:I

[Maz10] Stanley Mazor. Intel’s 8086.

IEEE Annals of the History of Computing, 32(1):75–79, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mazor:2024:ER

[Maz24] Stanley Mazor. Easy reciprocals. IEEE Annals of the History of Computing, 46(1):60– 61, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Montfort:2009:RRD

[MB09] Nick Montfort and Ian Bogost. Random and raster: Display technologies and the development of videogames. IEEE Annals of the History of Computing, 31(3):34–43, July/ September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McCallum:2002:MKL

[McC02] Sally H. McCallum. MARC: Keystone for library automation. IEEE Annals of the History of Computing, 24(2):34– 49, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2034.pdf;   http://www.
computer.org/annals/an2002/
a2034abs.htm.

McCorduck:2022:SLE

[McC22] Pamela McCorduck. The scientific life of Edward A. Feigenbaum. IEEE Annals of the History of Computing, 44(1):123– 128, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McDonald:2010:AFE

[McD10a] Christopher McDonald. From art form to engineering discipline? A history of US military software development standards, 1974–1998. IEEE Annals of the History of Computing, 32(4):32–47, October/ December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McDonald:2010:TPL

[McD10b] Christopher McDonald. Technology in the political landscape. IEEE Annals of the History of Computing, 32(2): 87–88, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McGee:2009:IMS

[McG09]  William C. McGee.  The

Information Management System (IMS) program product. IEEE Annals of the History of Computing, 31(4):66–75, October/December 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McGee:2012:SFI

[McG12] Andrew Meade McGee. Stating the field: Institutions and outcomes in computer history. IEEE Annals of the History of Computing, 34(1):104, 102– 103, January/March 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McGovern:2021:RBS

[McG21] Michael F. McGovern. Review of Black Software: The Internet & Racial Justice, from the AfroNet to Black Lives Matter. IEEE Annals of the History of Computing, 43(4):101– 102, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McJones:2017:SOF

[McJ17a] Paul McJones. In search of the original Fortran compiler. IEEE Annals of the History of Computing, 39(2):81–88, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

                                        www.softwarepreservation.
                                        org/projects/FORTRAN/.

McJones:2017:LP

[McJ17b] Paul McJones. The LISP 2 project. IEEE Annals of the History of Computing, 39(4): 85–92, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

McJones:2020:ADT

[McJ20] Paul McJones. The advent of digital typography: Interview with Liz Bond Crews. IEEE Annals of the History of Computing, 42(1):41–50, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McKenney:1995:DCM

[McK95] J. L. McKenney. Developing a common machine language for banking: the ABA technical subcommittee story. IEEE Annals of the History of Computing, 17(4):61–75, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477621.pdf.

McKenzie:2011:ICI

[McK11] Alexander McKenzie. INWG and the conception of the Internet: An eyewitness account. IEEE Annals of the History of Computing, 33(1):66–71, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McKelvey:2022:WNM

[McK22a] Fenwick McKelvey. When the new magic was new: The Claritas Corporation and the clustering of America. IEEE Annals of the History of Computing, 44(4):44–56, October/ December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McKenzie:2022:DCW

[McK22b] Alexander A. McKenzie. David Corydon Walden’s five careers. IEEE Annals of the History of Computing, 44(3):70–79, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McPherson:1995:NWM

[McP95] John McPherson. New ways of multiplying. IEEE Annals of the History of Computing, 17 (1):44–46, Spring 1995. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McQuillan:2009:BLS

[McQ09] John M. McQuillan. The birth of link-state routing. IEEE Annals of the History of Computing, 31(1):68–71, January/March 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McTiernan:1998:EP

[McT98] C. E. McTiernan. The ENIAC patent. IEEE Annals of the History of Computing, 20(2): 54–58, 80, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667303.pdf.

Medina:2005:CMC

[Med05] Eden Medina. Computer memory, collective memory: Recovering history through Chilean computing. IEEE Annals of the History of Computing, 27(4):104, 102–103, October/December 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Medina:2008:BBB

[Med08] Eden Medina. Big Blue in the bottomless pit: The early years of IBM Chile. IEEE Annals of the History of Computing, 30 (4):26–41, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mendelsohn:2013:OS

[Men13] Andrew Mendelsohn. The Oracle story: 1984–2001. IEEE Annals of the History of Computing, 35(2):10–23, April/ June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Merrin:2006:SPS

[Mer06] Seymour Merrin. Selling PC software: a once-irresistible opportunity. IEEE Annals of the History of Computing, 28 (4):80–86, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Meyers:1999:IFE

[Mey99] G. E. Meyers. IBM field engineering experiences: a personal memoir. IEEE Annals of the History of Computing, 21 (4):72–76, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17389/00801537.pdf.

McKenney:1993:MEB

[MF93] James L. McKenney and Amy Weaver Fisher. Manufacturing the ERMA banking system: Lessons from history. IEEE Annals of the History of Computing, 15(4):7–26, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4007.pdf;   http://www.
computer.org/annals/an1993/
a4007abs.htm.

Merchant:2022:ID

[MG22] Emily Klancher Merchant and Myron P. Gutmann. The IT of demography. IEEE Annals of the History of Computing, 44(4):6–15, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mahoney:2011:HC

[MH11] Michael S. (Michael Sean) Mahoney and Thomas Haigh, editors. Histories of computing. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05568-3. 250 pp. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [CK13] for comments on the author and this book.

Miller:2006:GSE

[Mil06] Raymond E. Miller. Graduate student experiences at Illinois. IEEE Annals of the History of Computing, 28(2):56– 62, April/June 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mills:2011:HAH

[Mil11] Mara Mills. Hearing aids and the history of electronics miniaturization. IEEE Annals of the History of Computing, 33(2): 24–45, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Minker:2007:DCS

[Min07a] Jack Minker. Developing a Computer Science Department at the University of Maryland. IEEE Annals of the History of Computing, 29(4):64–75, October/December 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Minker:2007:FCS

[Min07b]  Jack Minker.  Forming a

Computer Science Center at the University of Maryland. IEEE Annals of the History of Computing, 29(1):49–64, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mirowski:2017:ECO

[Mir17]  Alexander Mirowski.  At

the electronic crossroads once again: The myth of the modern computer utility in the United States. IEEE Annals of the History of Computing, 39(2): 13–29, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.

ieee.org/document/7974719/

.

Misa:1999:SKS

[Mis99] T. J. Misa. Susanne K. Schmidt and Raymund Werle, Coordinating Technology: Studies in the International Standardization of Telecommunications [book review]. IEEE Annals of the History of Computing, 21(3):76–77, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778992.pdf.

Misa:2007:ANC

[Mis07a] Thomas J. Misa. Arthur Norberg, the Charles Babbage Institute, and the history of computing. IEEE Annals of the History of Computing, 29(4): 8–15, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Misa:2007:GEI

[Mis07b]  Thomas J. Misa.  Guest

Editors’ introduction: New directions in the history of computing. IEEE Annals of the History of Computing, 29(4):6–7, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2007/04/
man2007040006.pdf.

Misa:2007:UHC

[Mis07c]  Thomas J. Misa.  Under-

standing ‘How Computing Has Changed the World’. IEEE Annals of the History of Computing, 29(4):52–63, October/ December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Misa:2016:CSD

[Mis16] Thomas J. Misa. Computer security discourse at RAND, SDC, and NSA (1958–1970). IEEE Annals of the History of Computing, 38(4):12–25, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mittermeir:2008:SLT

[Mit08] Roland T. Mittermeir. A seminar looking for traces of history. IEEE Annals of the History of Computing, 30(3):44– 54, July/September 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mounier-Kuhn:2010:JC

[MK10] Pierre-E. Mounier-Kuhn. Jean Carteron. IEEE Annals of the History of Computing, 32(1): 82–89, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mounier-Kuhn:2014:AFU

[MK14] P. Mounier-Kuhn. Algol in France: From universal project to embedded culture. IEEE Annals of the History of Computing, 36(4):6–25, October/ December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mori:2020:ISI

[MK20] E. Mori and S. Kelkar. Introduction to the special issue on interface architects: The evolution of human computer interaction. IEEE Annals of the History of Computing, 42(4): 6–7, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mollick:2006:EML

[Mol06] Ethan Mollick. Establishing Moore’s Law. IEEE Annals of the History of Computing, 28(3):62–75, July/September 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Molina:2014:TRC

[Mol14] E. Perez Molina. The technological roots of computer graphics. IEEE Annals of the History of Computing, 36(3): 30–41, July/September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mooers:2001:CPN

[Moo01]  Calvin N. Mooers.  The

computer project at the Naval Ordnance Laboratory. IEEE Annals of the History of Computing, 23(2):51–??, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2051.pdf;   http://www.
computer.org/annals/an2001/
a2051abs.htm.

Moore:2003:CPC

[Moo03]  Joe F. Moore.  Creating

profit with computers: My life as CEO of Bonner & Moore Associates. IEEE Annals of the History of Computing, 25(3):30–47, July/ September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a3030abs.htm;   http:
//computer.org/annals/an2003/
extras/a3030.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/03/a3030.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/03/a3030.pdf.

Morris:1984:PDD

[Mor84] L. R. Morris. Price, Derek, Desolla and the Antikythera mechanism — an appreciation 1922–1983 — obituary. IEEE Micro, 4(1):15–21, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morris:2003:BRR

[Mor03] Robin Morris. Book review:

Ross Knox Bassett, To the Digital Age: Research Labs, Startup Companies and the Rise of MOS Technology, Johns Hopkins University Press, 2002, ISBN 0-8018-6809-2, 421 pp., $44.95. IEEE Annals of the History of Computing, 25 (4):94–96, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mori:2019:CAG

[Mor19] E. Mori. Coping with the “American giants”. IEEE Annals of the History of Computing, 41(4):83–96, October/ December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mori:2020:OES

[Mor20] E. Mori. Olivetti ELEA Sign System: Interfaces before the advent of HCI. IEEE Annals of the History of Computing, 42 (4):24–38, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

MacKenzie:1997:MTT

[MP97] D. MacKenzie and G. Pottinger. Mathematics, technology, and trust: formal verification, computer security, and the U.S. military. IEEE Annals of the History of Computing, 19(3):41–59, July/September 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/13172/00601735.pdf.

Milivojevic:2012:HCC

[MPDT12] Dragan R. Milivojevic, Marijana Pavlov, Vladimir Despotovic, and Visa Tasic. Half a century of computing in the Serbian copper mining and metallurgy industry. IEEE Annals of the History of Computing, 34(3):34–43, July/ September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Magnuson:2022:CLS

[MR22] Diana L. Magnuson and Steven Ruggles. Challenges of largescale data processing in the 1990s: The IPUMS Experience. IEEE Annals of the History of Computing, 44(4):71– 83, October/December 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McJones:2023:HCT

[MR23] Paul McJones and David Redell. History of the CAL timesharing system. IEEE Annals of the History of Computing, 45(3):80–91, July/September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mazor:2009:DAM

[MS09] Stanley Mazor and Peter Salmon. Departments: Anecdotes: Magnavox and Intel: An odyssey. IEEE Annals of the History of Computing, 31(3): 64–67, July/September 2009.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Marcus:2006:TNH

[MT06] Mike Marcus and George Trimble. Taking newspapers from hot lead into the electronic age. IEEE Annals of the History of Computing, 28(4):96–100, October/December 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mullin:2003:DCT

[Mul03] Sherman N. Mullin. Into digital computing through the back door. IEEE Annals of the History of Computing, 25(3):20–28, July/September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/03/a3020.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3020.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/03/a3020abs.
htm.

Mulvin:2020:DLL

[Mul20] D. Mulvin. Distributing liability: The legal and political battles of Y2K. IEEE Annals of the History of Computing, 42(3):26–37, July/September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Murphy:2009:BT

[Mur09] Dan Murphy. The beginnings of TECO. IEEE Annals of the History of Computing, 31(4):110–115, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Murphy:2015:TT

[Mur15] Dan Murphy. TENEX and TOPS-20. IEEE Annals of the History of Computing, 37(1): 75–82, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Murray:2019:FTF

[Mur19] David J. Murray. Frame Technology and FrameMaker. IEEE Annals of the History of Computing, 41(3):62–74, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nandasara:2009:PPE

[Nan09] S. T. Nandasara. From the past to the present: Evolution of computing in the Sinhala language. IEEE Annals of the History of Computing, 31(1):32– 45, January/March 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nanez:2022:ACC

[Ñañ22] Fabian Prieto Ñañez. Assembling a Colombian-cloned computer: National development and the transnational trade of electronics parts in the 1980s. IEEE Annals of the History of Computing, 44(2): 55–64, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nash:1975:BRB

[Nas75] Robert A. Nash. Book review: Gears from the Greeks: The Antikythera Mechanism — A Calendar Computer from ca. 80 B.C. by Derek de Solla Price. American Scientist, 63 (5):594, September 1975. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.
jstor.org/stable/27845769.

Nebeker:1998:EID

[Neb98]  Frederick Nebeker.  Elec-

tronic inventions and discoveries: Electronics from its earliest beginnings to the present day [reviews]. IEEE Annals of the History of Computing, 20 (3):79, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Needham:1992:LDC

[Nee92] Roger M. Needham. Later developments at Cambridge: Titan, CAP, and the Cambridge Ring. IEEE Annals of the History of Computing, 14(4):57–58, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4057.pdf;   http://www.
computer.org/annals/an1992/
a4057abs.htm.

Nelsen:2017:RCP

[Nel17]  R. Arvid Nelsen.  Race

and computing: The problem of sources, the potential of prosopography, and the lesson of Ebony Magazine. IEEE Annals of the History of Computing, 39(1):29–51, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010029-abs.html.

Nel:2022:WPW

[Nel22] Philip Nel. Well paid for a woman: Gloria Hardman’s 50-year career in computing. IEEE Annals of the History of Computing, 44(3):80–86, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Neukom:2004:EUC

[Neu04]  Hans Neukom.  Early use

of computers in Swiss banks. IEEE Annals of the History of Computing, 26(3):50–59, July/September 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/03/a3050.htm;   http:

                                        //csdl.computer.org/dl/mags/
                                        an/2004/03/a3050.pdf.

Neukom:2005:EFS

[Neu05] Hans Neukom. ERMETH: The first Swiss computer. IEEE Annals of the History of Computing, 27(4):5–22, October/ December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Neukom:2006:SLE

[Neu06] Hans Neukom. The second life of ENIAC. IEEE Annals of the History of Computing, 28(2):4– 16, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/portal/pages/annals/
                                        articles/xtras/a2-2006/a2neukom.
                                        html.

Neukom:2009:UCA

[Neu09] Hans Neukom. Ubisco and CDC: Analysis of a failure. IEEE Annals of the History of Computing, 31(2):31–43, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nies:2009:CST

[Nie09] Thomas M. (Tom) Nies. Cincom Systems’ Total. IEEE Annals of the History of Computing, 31(4):55–61, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nikivincze:2017:ALG

[Nik17a]  Irina Nikivincze.  Anne-

Louise Guichard Radimsky: An educator and a champion for diversity in computing. IEEE Annals of the History of Computing, 39(4):81– 84, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/8267997/

.

Nikivincze:2017:DUP

[Nik17b] Irina Nikivincze. Dana Ulery:

Pioneer of statistical computing and architect of large, complex systems. IEEE Annals of the History of Computing, 39 (2):91–95, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/7974696/

.

Nisbet:2023:ECN

[Nis23] Alastair Nisbet. Educational computers in New Zealand schools: 1977 to 1983. IEEE Annals of the History of Computing, 45(4):39–52, October/ December 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nunemacher:1998:SPC

[NK98] J. Nunemacher and P. Kidwell. Selected papers on computer science. IEEE Annals of the History of Computing, 20(1):82–83, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/14098/00646215.pdf.

Nyman:2016:NHF

[NL16] Linus Nyman and Mikael Laakso. Notes on the history of fork and join. IEEE Annals of the History of Computing, 38(3):84–87, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nandasara:2009:ALP

[NM09] S. T. Nandasara and Yoshiki Mikami. Asian language processing: History and perspectives. IEEE Annals of the History of Computing, 31(1):4– 7, January/March 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nofre:2010:UAU

[Nof10] David Nofre. Unraveling Algol: US, Europe, and the creation of a programming language. IEEE Annals of the History of Computing, 32(2): 58–68, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nofre:2023:CMS

[Nof23] David Nofre. Content is meaningless, and structure is allimportant: Defining the nature of computer science in the age of high modernism, c. 1950 c. 1965. IEEE Annals of the History of Computing, 45(2): 29–42, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nofre:2025:CIT

[Nof25] David Nofre. A compelling image: The Tower of Babel and the proliferation of programming languages during the 1960s. IEEE Annals of the History of Computing, 47(1):22– 35, January/March 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

North:1977:BRD

[Nor77] J. D. North. Book review:

Derek de Solla Price: Gears from the Greeks: The Antikythera Mechanism. A Calendar Computer from ca. 80 B.C. Isis, 68(1):142– 143, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/
stable/230405.

Norberg:1996:CCC

[Nor96] A. L. Norberg. Changing computing: the computing community and DARPA. IEEE Annals of the History of Computing, 18(2):40–53, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489723.pdf.

Norberg:1998:TNU

[Nor98] A. L. Norberg. Trapped in the net: The unanticipated consequences of computerization [reviews]. IEEE Annals of the History of Computing, 20(3):79–80, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707583.pdf.

Norberg:2001:PHC

[Nor01] Arthur L. Norberg. A perspective on the history of the Charles Babbage Institute and the Charles Babbage Foundation. IEEE Annals of the History of Computing, 23(4):12–23, October/ December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a4012.pdf;   http://www.
computer.org/annals/an2001/
a4012abs.htm.

Norton:2023:RTD

[Nor23] Emma Rae Bruml Norton. Review of Transparent Designs: Personal Computing and the Politics of User-Friendliness, by Michael L. Black. IEEE Annals of the History of Computing, 45(1):86–87, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Notarnicola:2020:VCF

[Not20] D. Notarnicola. Vintage Computer Federation. IEEE Annals of the History of Computing, 42(4):123–124, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

November:2011:EBC

[Nov11] Joseph A. November. Early biomedical computing and the roots of evidence-based medicine. IEEE Annals of the History of Computing, 33(2): 9–23, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

November:2016:AYD

[Nov16] Joseph November. Ask your doctor … about computers. IEEE Annals of the History of Computing, 38(1):3–5, January/March 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nyce:1996:GEI

[Nyc96] J. M. Nyce. Guest Editor’s introduction. IEEE Annals of the History of Computing, 18 (4):3, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539910.pdf.

Nyitray:2011:WAH

[Nyi11] Kristen J. Nyitray. William Alfred Higinbotham: Scientist, activist, and computer game pioneer. IEEE Annals of the History of Computing, 33(2): 96–101, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ortiz-Arroyo:2010:TPD

[OARHC10] Daniel Ortiz-Arroyo, Francisco Rodriguez-Henriquez, and Carlos A. Coello Coello. The Turing-850 project: Developing a personal computer in the early 1980s in Mexico. IEEE Annals of the History of Computing, 32(4):60–71, October/December 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Oberquelle:2008:BSC

[OB08] Horst Oberquelle and Oskar Beckmann. Beckmann’s Studio Computers specified for early computer art. IEEE Annals of the History of Computing, 30(3):20–31, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Obayan:2022:EHR

[Oba22]  Kanyinsola Obayan.  En-

trepreneurial hustle and the rise (and fall) of personal computer companies in Lagos, Nigeria: 1960–1999. IEEE Annals of the History of Computing, 44(3):6–16, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

ODonnell:2009:PPC

[O’D09] Casey O’Donnell. Production protection to copy(right) protection: From the 10NES to DVDs. IEEE Annals of the History of Computing, 31(3): 54–63, July/September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Oinas-Kukkonen:2007:BES

[OK07] Henry Oinas-Kukkonen. From Bush to Engelbart: ‘Slowly, Some Little Bells Were Ringing’. IEEE Annals of the History of Computing, 29(2):31– 39, April/June 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Oldfield:1995:GEE

[Old95] H. R. Oldfield. General Electric enters the computer business — revisited. IEEE Annals of the History of Computing, 17(4):46–55, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477435.pdf.

Oleson:1977:BRB

[Ole77] John Peter Oleson. Book review: Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price. Phoenix, 31(3):271–273, Autumn 1977. ISSN 0031-8299 (print), 1929-4883 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1087110.

ONeill:1992:Ha

[O’N92a] Judy O’Neill. Happenings. IEEE Annals of the History of Computing, 14(1):55–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a1055.pdf.

ONeill:1992:Hc

[O’N92b] Judy O’Neill. Happenings. IEEE Annals of the History of Computing, 14(3):49–56, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a3049.pdf.

ONeill:1993:Ha

[O’N93a] Judy E. O’Neill. Happenings. IEEE Annals of the History of Computing, 15(1):58–??, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a1058.pdf.

ONeill:1993:Hb

[O’N93b] Judy E. O’Neill. Happenings. IEEE Annals of the History of Computing, 15(2):53– ??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2053.pdf.

ONeill:1993:Hc

[O’N93c] Judy E. O’Neill. Happenings.

IEEE Annals of the History of Computing, 15(4):76–??, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4076.pdf.

ONeill:1994:H

[O’N94] Judy E. O’Neill. Happenings. IEEE Annals of the History of Computing, 16(1): 70–??, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1070.pdf.

ONeill:1995:RAD

[O’N95] Judy E. O’Neill. The role of ARPA in the development of the ARPANET, 1961–1972. IEEE Annals of the History of Computing, 17(4):76–81, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477437.pdf.

ONeill:1992:HRB

[ORB^{+}92] Judy O’Neill, Brian Randell, Geoffrey C. Bowker, J.- P. Poitou, Pierre-E. Mounier-Kuhn, and Brian Crozier. Happenings: Reunion at Bletchley Park; UK National Archive for the History of Computing; Center for Research, Information, and Archives on Computer-Aided Design; DPMA donates records; European Conference on the History of Computing; Television Series on the History of Computing; Hartley computers. IEEE Annals of the History of Computing, 14(2):51– 55, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a2051.pdf.

Otnes:1998:C

[Otn98]  R. Otnes.  The calcume-

ter. IEEE Annals of the History of Computing, 20(1): 67–69, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646210.pdf.

Outterson:2003:OEL

[Out03] Sara Outterson. Obituary: Edward Louis Glaser [1929–1991]. IEEE Annals of the History of Computing, 25(1):83–87, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Owens:1996:WWG

[Owe96] L. Owens. Where are we going, Phil Morse? Changing agendas and the rhetoric of obviousness in the transformation of computing at MIT, 1939–1957. IEEE Annals of the History of Computing, 18 (4):34–41, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539914.pdf.

Pugh:1996:CCI

[PA96] E. W. Pugh and W. Aspray. Creating the computer industry. IEEE Annals of the History of Computing, 18(2):7–17, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489720.pdf.

Page:2008:ETF

[Pag08] John Page. The engineer’s tale:

The founding of Software Publishing Corporation. IEEE Annals of the History of Computing, 30(2):68–70, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Paju:2008:NPI

[Paj08] Petri Paju. National projects and international users: Finland and early European computerization. IEEE Annals of the History of Computing, 30 (4):77–91, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Parker:2007:DSC

[Par07] Donn B. Parker. The dark side of computing: SRI International and the study of computer crime. IEEE Annals of the History of Computing, 29 (1):3–15, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Partridge:2008:TDI

[Par08] Craig Partridge. The technical development of Internet email. IEEE Annals of the History of Computing, 30(2):3–29, April/ June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Park:2015:SLP

[Par15] Dongoh Park. Social life of PKI: Sociotechnical development of Korean public-key infrastructure. IEEE Annals of the History of Computing, 37(2):59–71, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//www.computer.org/csdl/
mags/an/2015/02/man2015020059-
abs.html.

Park:2016:KCC

[Par16a] Dongoh Park. The Korean character code: A national controversy, 1987–1995. IEEE Annals of the History of Computing, 38(2):40–53, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Partridge:2016:RIS

[Par16b] Craig Partridge. The restructuring of Internet standards governance: 1987–1992. IEEE Annals of the History of Computing, 38(3):25–43, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Patrick:2006:A

[Pat06] Robert L. Patrick. Anecdotes.

IEEE Annals of the History of Computing, 28(1):93–95, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Patrick:2009:IC

[Pat09] Robert L. Patrick. IMS @ Conception. IEEE Annals of the History of Computing, 31 (4):62–65, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Patrick:2014:TC

[Pat14] Robert L. Patrick. Times have changed. IEEE Annals of the History of Computing, 36(1): 79–81, January/March 2014.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Patterson:2024:IMC

[Pat24] Alan Patterson. Interview of Morris Chang. IEEE Annals of the History of Computing, 46(1):66–73, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Paulsen:2014:WSB

[Pau14] G. Paulsen. When switches became programs: Programming languages and telecommunications, 1965–1980. IEEE Annals of the History of Computing, 36 (4):38–50, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pawson:2022:MHA

[Paw22] Richard Pawson. The myth of the Harvard architecture. IEEE Annals of the History of Computing, 44(3):59–69, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Payette:2014:HDC

[Pay14] S. Payette. Hopper and Dijkstra: Crisis, revolution, and the future of programming. IEEE Annals of the History of Computing, 36(4):64–73, October/December 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pickstone:1993:MH

[PB93] J. V. Pickstone and Geof Bowker. The Manchester heritage. IEEE Annals of the History of Computing, 15(3):7–8, July/September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3007.pdf;   http://www.
computer.org/annals/an1993/
a3007abs.htm.

Partridge:2006:DNB

[PB06] Craig Partridge and Steven Blumenthal. Data networking at BBN. IEEE Annals of the History of Computing, 28(1): 56–71, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Paloque-Berges:2022:HEH

[PB22] Camille Paloque-Bergès. How EUNET hacked European digital networks and disappeared. IEEE Annals of the History of Computing, 44(3):43–58, July/ September 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Peeters:2009:EMS

[Pee09] Jos Peeters. Early MRP systems at Royal Philips Electronics in the 1960s and 1970s. IEEE Annals of the History of Computing, 31(2):56–69, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Persson:1999:TAC

[Per99] P.-A. Persson. Transformation of the analog: the case of the Saab BT 33 artillery fire control simulator and the introduction of the digital computer as control technology. IEEE Annals of the History of Computing, 21(2):52–64, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16485/00761795.pdf.

Perold:2020:IWC

[Per20] C. Perold. IBM’s world citizens: Valentim Bouças and the politics of IT expansion in authoritarian Brazil. IEEE Annals of the History of Computing, 42(3):38–52, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Petiot:1994:DGD

[Pet94]  Linda Petiot.  Dirty Ger-

tie: The DRTE computer. IEEE Annals of the History of Computing, 16(2):43– 52, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a2043.pdf;   http://www.
                                        computer.org/annals/an1994/
                                        a2043abs.htm.

Petersson:2005:FBB

[Pet05] Tom Petersson. Facit and the BESK boys: Sweden’s computer industry (1956–1962). IEEE Annals of the History of Computing, 27(4):23–30, October/December 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Petrick:2017:IPC

[Pet17] Elizabeth Petrick. Imagining the personal computer: Conceptualizations of the Homebrew Computer Club 1975– 1977. IEEE Annals of the History of Computing, 39(4): 27–39, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/8267600/

.

Petrick:2020:HHC

[Pet20] E. R. Petrick. A historiography of human computer interaction. IEEE Annals of the History of Computing, 42(4): 8–23, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Phipps:2012:EHI

[Phi12] Charles Phipps. The early history of ICs at Texas Instruments: a personal view. IEEE Annals of the History of Computing, 34(1):37–47, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pinkerton:1992:ICM

[PHL92] John M. M. Pinkerton, Derek Hemy, and Ernest H. Lenaerts. The influence of the Cambridge Mathematical Laboratory on the LEO project. IEEE Annals of the History of Computing, 14(4):41–48, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4041.pdf;   http://www.
computer.org/annals/an1992/
a4041abs.htm.

Plasek:2016:CRW

[Pla16] Aaron Plasek. On the cruelty of really writing a history of machine learning. IEEE Annals of the History of Computing, 38(4):6–8, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pennino:2012:RRP

[PM12] Anthony Pennino and Francesca Musiani. Reviews [review of “Prometheus (motion picture)”, Scott, R., Director; 2012 and “The Networked France, Vol. 1: The Meeting of Telecommunications and Computing” (Schafer, V.; 2012)]. IEEE Annals of the History of Computing, 34(4):4–6, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Prieto-Nanez:2016:PHC

[PN16] Fabian Prieto-Nanez. Postcolonial histories of computing. IEEE Annals of the History of Computing, 38(2):2–4, April/ June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Policriti:2023:OMD

[POD23] Alberto Policriti, Eugenio Omodeo, and Liesbeth De Mol. Obituary of Martin Davis (1928–2023). IEEE Annals of the History of Computing, 45(1):93–94, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Poirier:2025:NVS

[Poi25] Lindsay Poirier. Neat versus scruffy: How early AI researchers classified epistemic cultures of knowledge representation. IEEE Annals of the History of Computing, 47(2): 7–19, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Polachek:1995:HJM

[Pol95] H. Polachek. History of the journal Mathematical Tables and other Aids to Computation, 1959–1965. IEEE Annals of the History of Computing, 17(3):67–74, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397062.pdf.

Polachek:1997:BEW

[Pol97]  H. Polachek.  Before the

ENIAC [weapons firing table calculations]. IEEE Annals of the History of Computing, 19(2):25–30, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586069.pdf.

Polachek:2003:BOH

[Pol03] Solomon Polachek. Biographies: Obituary: Harry Polachek [1913–2002]. IEEE Annals of the History of Computing, 25(3):90–92, July/ September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/dl/mags/an/2003/03/a3090.
htm;   http://csdl.computer.
org/dl/mags/an/2003/03/a3090.
pdf.

Postley:1998:MIE

[Pos98] J. A. Postley. Mark IV: evolution of the software product, a memoir. IEEE Annals of the History of Computing, 20(1):43–50, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646208.pdf.

Post:2000:MTB

[Pos00] R. C. Post. Meaning in technology [book review]. IEEE Annals of the History of Computing, 22(4):67–68, October/December 2000. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887995.pdf.

Posner:2022:SOD

[Pos22] Miriam Posner. The software at our doorsteps. IEEE Annals of the History of Computing, 44(1):133–134, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pow:2019:SM

[Pow19]  Whitney Pow.  Stored in

memory. IEEE Annals of the History of Computing, 41(1):64–65, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8667942/.

Priestley:2021:RHI

[PP21] Mark Priestley and Tomas Petricek. Report on HOPL IV ACM SIGPLAN History of Programming Languages Conference. IEEE Annals of the History of Computing, 43(3): 83–85, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pfau:2023:KRE

[PPM^{+}23] Dinah Pfau, Helen Piel, Florian Müller, Jakob Tschandl, and Rudolf Seising. The KI-Rundbrief, its editors, and its community: a perspective on West German AI, 1975– 1987. IEEE Annals of the History of Computing, 45(3):48– 65, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Preston:1996:C

[Pre96]  F. Preston.  Calculators.

IEEE Annals of the History of Computing, 18(4):62– 64, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11673/00539919.pdf.

Preger:2012:OSP

[Pre12] Robert Preger. The Oracle story, part 1: 1977–1986. IEEE Annals of the History of Computing, 34(4):51–57, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Priestley:2021:LCH

[Pri21] Mark Priestley. Logic, code, and the history of programming. IEEE Annals of the History of Computing, 43(4):92– 96, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Priestley:2024:WSS

[Pri24] Mark Priestley. Working in the scientific state: John Womersley’s early career. IEEE Annals of the History of Computing, 46(3):86–93, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Prokhorov:1999:CRS

[Pro99] S. P. Prokhorov. Computers in Russia: science, education, and industry. IEEE Annals of the History of Computing, 21(3): 4–15, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778978.pdf.

Piel:2023:PAI

[PS23] Helen Piel and Rudolf Seising. Perspectives on artificial intelligence in Europe. IEEE Annals of the History of Computing, 45(3):6–10, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Puchta:1996:RMM

[Puc96] S. Puchta. On the role of mathematics and mathematical knowledge in the invention of Vannevar Bush’s early analog computers. IEEE Annals of the History of Computing, 18 (4):49–59, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539916.pdf.

Pugh:2002:OSB

[Pug02]  Emerson W. Pugh.  Ori-

gins of software bundling. IEEE Annals of the History of Computing, 24(1):57–58, January/March 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2002/a1057abs.htm;
 http://dlib.computer.org/
an/books/an2002/pdf/a1057.
pdf.

Rojas:2004:RPJ

[RAG^{+}04] Raul Rojas, Jon Agar, Slava Gerovitch, Ross Bassett, Tim Bergin, Nathan L. Ensmenger, Kenneth Lipartito, and J. A. N. Lee. Reviews: Pamela J. Hinds and Sara Kiesler (eds.), Distributed Work, MIT Press, 2002, ISBN 0-262-08305-1; David Mindell, Between Human and Machine: Feedback, Control, and Computing Before Cybernetics, Johns Hopkins University Press, 2002, ISBN 0-801-86895-5; Jeffrey Zygmont, Microchip: An Idea, Its Genesis, and the Revolution It Created, Peresus Publishing, 2003, ISBN 0-7382-0561-3; Atsushi Akera and Frederick Nebeker, eds., From 0 to 1: An Authoritative History of Modern Computing, Oxford Univ. Press, 2002, ISBN 0-19-514025-7; Martin Campbell-Kelly, From Airline Reservations to Sonic the Hedgehog, MIT Press, 2003, ISBN 0-262-03303-8; Narain Gehani, Bell Labs: Life in the Crown Jewel, Silicon Press, 2003, ISBN 0-929306-27-9; Smart Computing, People in Computing: Biographies of Technology’s Pioneers (Computing Encyclopedia), vol. 5, Sandhill Publishing Company, 2002. IEEE Annals of the History of Computing, 26(1):71–77, January/March 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/01/a1071.pdf.

Rajaraman:2015:HCI

[Raj15] V. Rajaraman. History of computing in India: 1955–2010. IEEE Annals of the History of Computing, 37(1):24–35, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ralston:2004:FEE

[Ral04] Anthony Ralston. Four editions and eight publishers: a history of the Encyclopedia of Computer Science. IEEE Annals of the History of Computing, 26(1):42–52, January/March 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/

                                        mags/an/2004/01/a1042abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/01/a1042.
                                        pdf.

Raman:1997:KFM

[Ram97] Anand V. Raman. The Katapayadi formula and the modern hashing technique. IEEE Annals of the History of Computing, 19(4):49–52, October/ December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Randell:1994:OCP

[Ran94]  Brian Randell.  The ori-

gins of computer programming. IEEE Annals of the History of Computing, 16(4): 6–14, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a4006.pdf;   http://www.
                                        computer.org/annals/an1994/
                                        a4006abs.htm.

Rankin:2014:THS

[Ran14] Joy Rankin. Toward a history of social computing: Children, classrooms, campuses, and communities. IEEE Annals of the History of Computing, 36(2):88, April/June 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Randell:2015:RPA

[Ran15]  Brian Randell.  Reminis-

cences of Project Y and the ACS Project. IEEE Annals of the History of Computing, 37(3):56–66, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030056-abs.html.

Rawlings:2014:HNF

[Raw14] Nicholas Rawlings. The history of NOMAD: A fourth generation language. IEEE Annals of the History of Computing, 36(1):30–38, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rawn:2024:WVU

[Raw24] Eric Rawn. The work and vision of ubiquitous computing at Xerox PARC. IEEE Annals of the History of Computing, 46(3):50–60, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rayward:2002:HCA

[Ray02] W. Boyd Rayward. A history of computer applications in libraries: Prolegomena. IEEE Annals of the History of Computing, 24(2):4– 15, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2004.pdf;   http://www.
computer.org/annals/an2002/
a2004abs.htm.

RevoredoMenezesdeCarvalho:2015:DIB

[RC15] Marcelo Savio Revoredo Menezes de Carvalho and Henrique Luiz Cukierman. The dawn of the Internet in Brazil. IEEE Annals of the History of Computing, 37(4):54–63, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Randell:2023:CAF

[RC23] Brian Randell and Brian Coghlan. ChatGPT’s astonishing fabrications about Percy Ludgate. IEEE Annals of the History of Computing, 45(2): 71–72, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2013:RMS

[RCP^{+}13] Andrew Russell, Jonathan Coopersmith, Anthony P. Pennino, Giuditta Parolini, and Benjamin Gross. Reviews: Michael Sean Mahoney, Histories of Computing, ed. and introduction by Thomas Haigh, Harvard University Press, 2011, 250 pp.; Elysium, directed by Neill Blomkamp and produced by TriStar Pictures, 2013; Paul E. Ceruzzi, Computing: A Concise History, MIT Press, 2012, 199 pp.; James W. Cortada, The Digital Flood: The Diffusion of Information Technology Across the U.S., Europe, and Asia, Oxford University Press, 2012, 789 pp. IEEE Annals of the History of Computing, 35(4):86– 89, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rojas:2005:RKZ

[RDGH05] Raul Rojas, Frank Darius, Cuneyt Goktekin, and George Heyne. The reconstruction of Konrad Zuse’s Z3. IEEE Annals of the History of Computing, 27(3):23–32, July/ September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Reitwiesner:1997:FOS

[Rei97] G. W. Reitwiesner. The first operating system for the ED-VAC. IEEE Annals of the History of Computing, 19(1): 55–59, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560738.pdf.

Reitman:2008:OWS

[Rei08] Julian Reitman. The origin of the Winter Simulation Conferences. IEEE Annals of the History of Computing, 30(4):98– 102, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Reid:2023:CNI

[Rei23a]  T. Alex Reid.  Computer

networking initiatives in one of the world’s remote cities. IEEE Annals of the History of Computing, 45(4):72–80, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Reid:2023:MSC

[Rei23b] T. Alex Reid. Monte Sala’s cryptographic achievements. IEEE Annals of the History of Computing, 45(4):65–71, October/December 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Reid-Green:2002:TEA

[RG02] Keith S. Reid-Green. Three early algorithms: [Bresenham’s line-drawing algorithm; a square-root algorithm; Machin’s algorithm: computation of π]. IEEE Annals of the History of Computing, 24 (4):10–13, October/December 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2002/04/a4010.
htm;   http://csdl.computer.
org/dl/mags/an/2002/04/a4010.
pdf;   http://csdl.computer.
org/dl/mags/an/2002/04/a4010abs.
htm.

Richardson:2024:AED

[Ric24] Fontaine K. Richardson. Applicon: The early days of computer-aided electronics design. IEEE Annals of the History of Computing, 46(4):58– 62, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Riordan:2000:EFV

[Rio00] M. Riordan. Endless frontier: Vannevar Bush, engineer of the American century [book review]. IEEE Annals of the History of Computing, 22(4):64, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887991.pdf.

Rosin:1992:CI

[RL92] Robert Rosin and J. A. N.

Lee. The CTSS interviews. IEEE Annals of the History of Computing, 14(1):33–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1033.pdf;   http://www.
computer.org/annals/an1992/
a1033abs.htm.

Roberts:1993:CQD

[Rob93]  C. J. D. Roberts.  Com-

ments, queries, and debate: Babbage and Clement. IEEE Annals of the History of Computing, 15(2):6– ??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.

                                        org/an/books/an1993/pdf/
                                        a2006.pdf.

Robertson:2004:Aa

[Rob04a]  Laurie Robertson.  Anec-

dotes. IEEE Annals of the History of Computing, 26(2): 86–89, April/June 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/an/
                                        2004/02/a2086.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2004/02/a2086.pdf.

Robertson:2004:Ab

[Rob04b]  Laurie Robertson.  Anec-

dotes. IEEE Annals of the History of Computing, 26(3):70–77, July/September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/03/a3070.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/03/a3070.
                                        pdf.

Robertson:2004:Ac

[Rob04c]  Laurie Robertson.  Anec-

dotes. IEEE Annals of the History of Computing, 26 (4):71–73, October/December 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/04/a4071.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/04/a4071.
                                        pdf.

Roberston:2005:A

[Rob05a]  Laurie Roberston.  Anec-

dotes. IEEE Annals of the History of Computing, 27(1):78–81, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1078.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1078.
pdf.

Robertson:2005:ADE

[Rob05b] L. Robertson. Anecdotes [Difference Engine]. IEEE Annals of the History of Computing, 27(2):89–92, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2089.
htm;   http://csdl.computer.
org/dl/mags/an/2005/02/a2089.
pdf.

Robertson:2005:Aa

[Rob05c]  Laurie Robertson.  Anec-

dotes. IEEE Annals of the History of Computing, 27(2): 82–84, April/June 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2005/02/a2082.htm;   http:
//csdl.computer.org/dl/mags/
an/2005/02/a2082.pdf.

Robertson:2005:Ab

[Rob05d] Laurie Robertson. Anecdotes. IEEE Annals of the History of Computing, 27(3):96–97, July/ September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Robertson:2005:Ac

[Rob05e] Laurie Robertson. Anecdotes. IEEE Annals of the History of Computing, 27(4):89–92, October/December 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Robelen:2016:IAP

[Rob16] Russell J. Robelen. The IBM ACS project. IEEE Annals of the History of Computing, 38(1):60–74, January/March 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Robertson:2019:BHS

[Rob19] Stephen Robertson. A brief history of search results ranking. IEEE Annals of the History of Computing, 41(2):22–28, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Robbins:2024:SFI

[Rob24] Thomas K. Robbins. Stepping into the future — the invention and evolution of the stepping switch. IEEE Annals of the History of Computing, 46(3): 74–80, July/September 2024.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2008:EKD

[Roe08] Denis Roegel. An early (1844)

key-driven adding machine. IEEE Annals of the History of Computing, 30(1):59–65, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2009:APF

[Roe09] Denis Roegel. Anecdotes: Prototype fragments from Babbage’s first Difference Engine. IEEE Annals of the History of Computing, 31(2):70–75, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2012:LPR

[Roe12] Denis Roegel. The LOCOMAT Project: Recomputing mathematical and astronomical tables. IEEE Annals of the History of Computing, 34(2):74– 79, April/June 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2015:MCA

[Roe15] Denis Roegel. A mechanical calculator for arithmetic sequences (1844–1852): Part 1, historical context and structure. IEEE Annals of the History of Computing, 37(4): 90–96, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2016:BTS

[Roe16a] Denis Roegel. Before Torchi and Schwilgué, there was white. IEEE Annals of the History of Computing, 38(4): 92–93, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Roegel:2016:MCA

[Roe16b] Denis Roegel. A mechanical calculator for arithmetic sequences (1844–1852): Part 2, working details. IEEE Annals of the History of Computing, 38(1):80–88, January/March 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Roegel:2017:CSB

[Roe17] Denis Roegel. Carries stripped to the bone: Episodes in the history of coaxial modular digital counters. IEEE Annals of the History of Computing, 39(3):55–64, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030055-abs.html.

Rohl:1999:IPL

[Roh99] J. S. Rohl. The influence of programming languages on the design of MU5. IEEE Annals of the History of Computing, 21(1):34–37, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16436/00759367.pdf.

Rojas:1997:KZL

[Roj97]  R. Rojas.  Konrad Zuse’s

legacy: the architecture of the Z1 and Z3. IEEE Annals of the History of Computing, 19(2):5–16, April/June 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12737/00586067.pdf.

Rojas:1998:HMZ

[Roj98]  R. Rojas.  How to make

Zuse’s Z3 a universal computer. IEEE Annals of the History of Computing, 20(3): 51–54, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707574.pdf.

Rojas:2003:Ra

[Roj03a] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 25(2):83–85, April/ June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2083.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2083.
pdf.

Rojas:2003:Rb

[Roj03b] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 25(3):84–89, July/ September 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/dl/mags/an/2003/03/a3084.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2003/03/a3084.
                                        pdf.

Rojas:2004:KGI

[Roj04a] R. Rojas. Karl E. Ganzhorn, The IBM Laboratories Boeblingen Foundation and Build-Up: A Personal Review [book review]. IEEE Annals of the History of Computing, 26(2):90–91, April/ June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2004/02/a2090.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2004/02/a2090.
                                        pdf.

Rojas:2004:RRE

[Roj04b] Raul Rojas. Raul Rojas, Encyclopedia of Computers and Computer History [book review]. IEEE Annals of the History of Computing, 26(2): 90, April/June 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rojas:2004:Ra

[Roj04c] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 26(3):78–84, July/ September 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2004/03/
a3078.pdf;   http://csdl.
computer.org/dl/mags/an/
2004/03/a3078.htm.

Rojas:2004:Rb

[Roj04d] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 26(4):74–78, October/December 2004. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2004/04/a4074.htm;   http:
//csdl.computer.org/dl/mags/
an/2004/04/a4074.pdf.

Rojas:2005:Ra

[Roj05a] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 27(1):82–87, January/March 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2005/01/a1082.htm;   http:
//csdl2.computer.org/comp/
mags/an/2005/01/a1082.pdf.

Rojas:2005:Rb

[Roj05b] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 27(2):85–88, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2085.
                                        htm;   http://csdl2.computer.
                                        org/comp/mags/an/2005/02/
                                        a2085.pdf.

Rojas:2005:Rc

[Roj05c] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 27(3):93–95, July/ September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2005/03/
                                        a3093.pdf.

Rojas:2005:Rd

[Roj05d] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 27(4):86–88, October/ December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2005/04/
                                        a4086.pdf.

Rojas:2006:Ra

[Roj06a] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 28(1):96–98, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2006/01/
                                        a1096.pdf.

Rojas:2006:Rb

[Roj06b] Raul Rojas. Reviews. IEEE Annals of the History of Computing, 28(4):101–103, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2006/04/
a4101.pdf.

Rojas:2007:Ra

[Roj07] Raúl Rojas. Reviews. IEEE Annals of the History of Computing, 29(1):96–99, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2007/01/
a1096.pdf.

Rojas:2021:CPC

[Roj21]  R. Rojas.  The computer

programs of Charles Babbage. IEEE Annals of the History of Computing, 43(1):6–18, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rope:2007:ESP

[Rop07] Crispin Rope. ENIAC as a stored-program computer: a new look at the old records. IEEE Annals of the History of Computing, 29(4):82–87, October/December 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rosin:1992:CQD

[Ros92] Robert F. Rosin. Comments, queries, and debate: Timesharing at MIT (correction). IEEE Annals of the History of Computing, 14(3):8, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/

a3006.pdf. See [LML92].

Rosen:2004:RPT

[Ros04]  Saul Rosen.  Recollections

of the Philco Transac S-2000. IEEE Annals of the History of Computing, 26 (2):34–47, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2034abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2034.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2034.
pdf.

Rowe:2012:HIC

[Row12]  Lawrence A. Rowe.  His-

tory of the Ingres Corporation. IEEE Annals of the History of Computing, 34(4):58–70, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Royer:2011:FCU

[Roy11] George Royer. Familiar concepts, unfamiliar territory.

IEEE Annals of the History of Computing, 33(2):112, 111, April/June 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Ristanovic:2012:OUP

[RP12] Dejan Ristanovic and Jelica Protic. Once upon a pocket: Programmable calculators from the late 1970s and early 1980s and the social networks around them. IEEE Annals of the History of Computing, 34(3):55–66, July/ September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rice:2004:CSP

[RR04] John R. Rice and Saul Rosen. Computer sciences at Purdue University—1962 to 2000. IEEE Annals of the History of Computing, 26 (2):48–61, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2048abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2048.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2048.
pdf.

Robbins:2023:OJP

[RR23] Loring G. Robbins and Andrew L. Russell. Obituary for James Pelkey. IEEE Annals of the History of Computing, 45(2):76–77, April/June 2023.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rojas:2007:Rb

[RSCF07] Raúl Rojas, Dag Spicer, Rocky R. Clark, and Gerald Friedland. Reviews. IEEE Annals of the History of Computing, 29(2):74–77, April/ June 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/an/2007/02/a2074.pdf.

Rubinstein:2006:RRF

[Rub06] Seymour Rubinstein. Recollections: The rise and fall of WordStar. IEEE Annals of the History of Computing, 28 (4):64–72, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Runge:2008:GTE

[Run08] Chuck Runge. The genesis of the Tenet 210: An early timesharing system. IEEE Annals of the History of Computing, 30(3):78–81, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2006:RCR

[Rus06] Andrew L. Russell. ‘Rough consensus and running code’ and the Internet–OSI standards war. IEEE Annals of the History of Computing, 28(3): 48–61, July/September 2006.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2012:R

[Rus12a]  Andrew Russell.  Reviews.

IEEE Annals of the History of Computing, 34(2):70–72, April/June 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2012:SNC

[Rus12b] Andrew L. Russell. Standards, networks, and critique. IEEE Annals of the History of Computing, 34(3):80, July/ September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2013:BCD

[Rus13a] Andrew Russell. Biomedical computing: Digitizing life in the United States (November, J.; 2012) [reviews]. IEEE Annals of the History of Computing, 35(2):82–84, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2013:RFB

[Rus13b] Andrew Russell. Reviews [four books reviewed]. IEEE Annals of the History of Computing, 35 (1):4–10, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2013:RFW

[Rus13c] Andrew Russell. Reviews [“The future Was here: The Commodore Amiga” (Maher, J.; 2012) and “Recoding Gender: Women’s Changing Participation in Computing” (Abbate, J.; 2012)]. IEEE Annals of the History of Computing, 35(3): 68–70, July/September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2014:RBRb

[Rus14a] Andrew Russell. Reviews [3 books reviewed]. IEEE Annals of the History of Computing, 36(3):90–93, July/September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2014:RBRa

[Rus14b] Andrew Russell. Reviews [5 books reviewed]. IEEE Annals of the History of Computing, 36(1):69–73, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2015:Ra

[Rus15a]  Andrew Russell.  Reviews.

IEEE Annals of the History of Computing, 37(3):68–69, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030068-abs.html.

Russell:2015:Rb

[Rus15b]  Andrew Russell.  Reviews.

IEEE Annals of the History of Computing, 37(4):98–104, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2015:RBRb

[Rus15c] Andrew Russell. Reviews [2 book reviews; 2 film reviews]. IEEE Annals of the History of Computing, 37(2):89–93, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020089-abs.html.

Russell:2015:RBRa

[Rus15d] Andrew Russell. Reviews [4 books reviewed]. IEEE Annals of the History of Computing, 37(1):87–92, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2016:IHG

[Rus16a] Andrew Russell. The innovators: how a group of hackers, geniuses, and geeks created the digital revolution (Isaacson, W.; 2014) [book review]. IEEE Annals of the History of Computing, 38(1):94–c3, January/March 2016. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Russell:2016:Ra

[Rus16b]  Andrew Russell.  Reviews.

IEEE Annals of the History of Computing, 38(2):80–84, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2016:Rb

[Rus16c]  Andrew Russell.  Reviews.

IEEE Annals of the History of Computing, 38(3):10–11, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2016:RBR

[Rus16d] Andrew Russell. Reviews [book review and play review]. IEEE Annals of the History of Computing, 38(4):94–96, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Russell:2017:RBR

[Rus17] Andrew Russell. Reviews [book review]. IEEE Annals of the History of Computing, 39(1):8, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2017/01/
                                        man2017010008-abs.html.

Sachs:2007:RDL

[Sac07]  Jonathan Sachs.  Recollec-

tions: Developing Lotus 1-2-3. IEEE Annals of the History of Computing, 29(3):41– 48, July/September 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sale:2005:RCB

[Sal05] Anthony E. Sale. The rebuilding of Colossus at Bletchley Park. IEEE Annals of the History of Computing, 27(3):61– 69, July/September 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Saltzer:2020:OL

[Sal20] J. H. Saltzer. The origin of the MIT License. IEEE Annals of the History of Computing, 42(4):94–98, October/ December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.mit.edu/~
Saltzer/publications/MITLicense.
pdf.

Saltzer:2021:OK

[Sal21] J. H. Saltzer. On the origin of Kerberos. IEEE Annals of the History of Computing, 43(1): 89–91, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sandstrom:2011:HSD

[San11] Christian Sandstrom. Hasselblad and the shift to digital imaging. IEEE Annals of the History of Computing, 33(3): 55–66, July/September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sandstrom:2013:FDM

[San13] Christian Sandstrom. Facit and the displacement of mechanical calculators. IEEE Annals of the History of Computing, 35(3): 20–31, July/September 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sato:2011:IGN

[Sat11] Yasushi Sato. An inconspicuous giant: NTT’s role in the development of software engineering in Japan. IEEE Annals of the History of Computing, 33 (4):28–37, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Saunders:1993:ATB

[Sau93] P. T. Saunders. Alan Turing and biology. IEEE Annals of the History of Computing, 15(3):33–36, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a3033.pdf;   http://www.
computer.org/annals/an1993/
a3033abs.htm.

Saviers:2023:IFC

[Sav23] Grant Saviers. Interview of Finis Conner. IEEE Annals of the History of Computing, 45(3): 92–102, July/September 2023.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sicking:2024:HU

[SAV24] Paul O. Sicking, George A.

Allen, and Wilfredo R. Valenzuela. The history of Unigraphics, 1974–2001. IEEE Annals of the History of Computing, 46 (4):70–83, October/December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shelburne:1998:EPM

[SB98] B. J. Shelburne and C. P.

Burton. Early programs on the Manchester Mark I prototype. IEEE Annals of the History of Computing, 20(3): 4–15, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707570.pdf.

Scroggins:2020:OFA

[SB20] M. Scroggins and B. M. Boscoe.

Once FITS, always FITS? Astronomical infrastructure in transition. IEEE Annals of the History of Computing, 42(2): 42–54, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sproull:2023:IIS

[SB23] Robert Sproull and David C.

Brock. Interview of Ivan Sutherland. IEEE Annals of the History of Computing, 45(1):64–76, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Schantz:2006:BNC

[Sch06a] Richard E. Schantz. BBN’s network computing software infrastructure and distributed applications (1970–1990). IEEE Annals of the History of Computing, 28(1):72–88, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Schlombs:2006:TIC

[Sch06b] Corinna Schlombs. Toward international computing history. IEEE Annals of the History of Computing, 28(1):108, 107, January/March 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Schlombs:2008:EIE

[Sch08] Corinna Schlombs. Engineering international expansion: IBM and Remington Rand in European computer markets. IEEE Annals of the History of Computing, 30(4):42–58, October/ December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Schlombs:2017:IFA

[Sch17]  Corinna Schlombs.  The

“IBM Family”: American welfare capitalism, labor, and gender in postwar Germany.

IEEE Annals of the History of Computing, 39(4):12– 26, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://ieeexplore.
ieee.org/document/8267999/

.

Schmitt:2023:SAS

[Sch23]  Martin Schmitt.  Socialist

AI? societal use, economic implementation, and the tensions of applied computer science in late socialist GDR. IEEE Annals of the History of Computing, 45(3):66–79, July/ September 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stierhoff:1998:HIS

[SD98] G. C. Stierhoff and A. G. Davis.

A history of the IBM Systems Journal. IEEE Annals of the History of Computing, 20(1):29–35, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14098/00646206.pdf.

Seidel:1995:CBI

[Sei95] R. W. Seidel. Charles Babbage Institute Director’s column. IEEE Annals of the History of Computing, 17(4):5, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477431.pdf.

Seidel:1996:GEI

[Sei96] R. W. Seidel. Guest Editor’s introduction. IEEE Annals of the History of Computing, 18(2):6, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10525/00490111.pdf.

Seidel:1997:KME

[Sei97] R. W. Seidel. Knowing machines: Essays on technical change. IEEE Annals of the History of Computing, 19(1): 79–80, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/12228/00560751.pdf.

Seidel:2000:FCE

[Sei00] R. W. Seidel. Feynman and computation: exploring the limits of computers. IEEE Annals of the History of Computing, 22(1):90–91, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Seitz:2003:TGL

[Sei03] F. Seitz. True genius: The life and science of John Bardeen [book review]. IEEE Annals of the History of Computing, 25(2):83–84, April/June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Seybold:2018:MAA

[Sey18a]  Jonathan Seybold.  More

about Atex. IEEE Annals of the History of Computing, 40(3):30–36, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509547/.

Seybold:2018:RCP

[Sey18b] Jonathan W. Seybold. Rocappi: Computerizing the publishing industry. IEEE Annals of the History of Computing, 40(3):8–24, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509537/.

Seybold:2019:SPS

[Sey19] Jonathan Seybold. Seybold publications and seminars. IEEE Annals of the History of Computing, 41(3):8–23, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Scull:2019:KAT

[SH19] John Scull and Hansen Hsu.

The killer app that saved the Macintosh. IEEE Annals of the History of Computing, 41 (3):42–52, July 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shapiro:1997:SDH

[Sha97] S. Shapiro. Splitting the difference: the historical necessity of synthesis in software engineering. IEEE Annals of the History of Computing, 19(1):20–54, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/12228/00560729.pdf.

Shapiro:2000:CQD

[Sha00] Fred R. Shapiro. Comments, queries, and debates: Origin of the term “personal computer”: evidence from the JSTOR electronic journal archive. IEEE Annals of the History of Computing, 22 (4):70–71, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887997.pdf.

Sharples:2023:JCL

[Sha23] Mike Sharples. John Clark’s Latin verse machine: 19th century computational creativity. IEEE Annals of the History of Computing, 45(1):31–42, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shelburne:2012:ED

[She12]  Brian J. Shelburne.  The

ENIAC’s 1949 determination of π. IEEE Annals of the History of Computing, 34(3):44– 54, July/September 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shin:2019:SAI

[Shi19] Y. Shin. The spring of artificial intelligence in its global winter. IEEE Annals of the History of Computing, 41(4):71–82, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shoup:2001:SEF

[Sho01] Richard Shoup. SuperPaint:

An early frame buffer graphics system. IEEE Annals of the History of Computing, 23(2):32–37, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an2001/pdf/a2032.
pdf;   http://www.computer.
org/annals/an2001/a2032abs.
htm.

Shreiner:2024:RBC

[Shr24] Dave Shreiner. Review of The Big 50: Celebrating 50 ACM SIGGRAPH Conferences, by David J. Kasik, Mary C. Whitton, and Chris R. Johnson. IEEE Annals of the History of Computing, 46(4):150–151, October/December 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shustek:2006:WSW

[Shu06] Len Shustek. What should we collect to preserve the history of software? IEEE Annals of the History of Computing, 28(4):112, 110–111, October/December 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Siddiqui:2020:BRM

[Sid20] N. Siddiqui. Book review: Meryl Alper, Giving Voice: Mobile Communication, Disability, and Inequality. Cambridge, MA, USA: MIT Press, 2017. IEEE Annals of the History of Computing, 42(2): 87–88, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Siegert:2010:GTC

[Sie10] Hans-Juergen Siegert. The German TR 440 computer: Software and its development. IEEE Annals of the History of Computing, 32(3):40–73, July/ September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Simon:1998:AN

[Sim98] H. A. Simon. Allen Newell: 1927–1992. IEEE Annals of the History of Computing, 20 (2):63–76, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667311.pdf.

Simon:2025:OVA

[Sim25] Sara M. B. Simon. Operation voder: ATT, bell labs, and the labor of techno-utopia at the 1939 new york world’s fair. IEEE Annals of the History of Computing, 47(2):20– 31, April/June 2025. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sinha:2009:JIS

[Sin09] R. Mahesh K. Sinha. A journey from Indian scripts processing to Indian language processing. IEEE Annals of the History of Computing, 31(1):8– 31, January/March 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sippl:2013:IIM

[Sip13] Roger Sippl. Informix: Information management on Unix. IEEE Annals of the History of Computing, 35(2):42–53, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sjoblom:2011:CHC

[Sjö11] Gustav Sjöblom. Control in the history of computing: Making an ambiguous concept useful. IEEE Annals of the History of Computing, 33(3):88, 86– 87, July/September 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Solcova:2011:VVP

[SK11] Alena Solcova and Michal Krizek. Vladimir Vand (1911– 1968): Pioneer of computational methods in crystallography. IEEE Annals of the History of Computing, 33(4):38– 44, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sack:2012:WMP

[SL12] Edgar A. Sack and David A. Laws. Westinghouse: Microcircuit pioneer from molecular electronics to ICs. IEEE Annals of the History of Computing, 34(1):74–82, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Slayton:2008:RRR

[Sla08]  Rebecca Slayton.  Revolu-

tion and resistance: Rethinking power in computing history. IEEE Annals of the History of Computing, 30(1):96, c3, January/March 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Slayton:2015:MRC

[Sla15] Rebecca Slayton. Measuring risk: Computer security metrics, automation, and learning. IEEE Annals of the History of Computing, 37(2):32– 45, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020032-abs.html.

Small:1993:GPE

[Sma93] James S. Small. Generalpurpose electronic analog computing: 1945–1965. IEEE Annals of the History of Computing, 15(2):8–18, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://dlib.computer.org/
an/books/an1993/pdf/a2008.
pdf;   http://www.computer.
org/annals/an1993/a2008abs.
htm.

Smith:1978:BRA

[Smi78] Norman A. F. Smith. Book review: Antiquity Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price. New York: Science History Publications, 1975. Pp. 70. $8.50. British Journal for the History of Science, 11(1):77– 78, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025619.

Smilie:1999:DED

[Smi99a] K. Smilie. David Ewing Duncan, Calendar: Humanity’s Epic Struggle to Determine a True and Accurate Year [book review]. IEEE Annals of the History of Computing, 21(3):75, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778989.pdf.

Smith:1999:TPH

[Smi99b] T. M. Smith. Thomas P. Hughes, Rescuing Prometheus [book review]. IEEE Annals of the History of Computing, 21(4):82, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17389/00801539.pdf.

Smillie:2000:MTC

[Smi00] K. Smillie. Mapping time. the calendar and its history. IEEE Annals of the History of Computing, 22(1): 88–89, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/17665/00815470.pdf.

Smillie:2001:CSC

[Smi01a] K. Smillie. The complete story of codebreaking in World War II [reviews]. IEEE Annals of the History of Computing, 23(3):78–79, July/September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smith:2001:DPS

[Smi01b]  Alvy Ray Smith.  Digital

paint systems: An anecdotal and historical overview. IEEE Annals of the History of Computing, 23(2):4– 30, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2004.pdf;   http://www.
computer.org/annals/an2001/
a2004abs.htm.

Smillie:2002:BRT

[Smi02] K. Smillie. Book review: Turing and the universal machine: the making of the modern computer. IEEE Annals of the History of Computing, 24(2): 95, April/June 2002. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smillie:2003:BRS

[Smi03] Keith Smillie. Book review:

Severo M. Ornstein, Computing in the Middle Ages: A View from the Trenches 1955–1983, 1st Books Library, 2002, ISBN 1-4033-1517-5, 274 pp., $14.95. IEEE Annals of the History of Computing, 25(4):96, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smillie:2004:PLC

[Smi04] Keith Smillie. People, languages, and computers: a short memoir. IEEE Annals of the History of Computing, 26(2):62–74, April/ June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2004/02/a2062abs.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2062.
htm;   http://csdl.computer.
org/dl/mags/an/2004/02/a2062.
pdf.

Smillie:2007:ECU

[Smi07] Keith Smillie. Early computing at the University of Alberta and the introduction of the LGP-30. IEEE Annals of the History of Computing, 29(1): 65–73, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smillie:2011:DMG

[Smi11] Keith Smillie. Donald McIntyre: Geologist, historian, and array language advocate, 1923– 2009. IEEE Annals of the History of Computing, 33(1):72– 77, January/March 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smillie:2012:ESC

[Smi12] Keith Smillie. An early statistical calculation. IEEE Annals of the History of Computing, 34(3):74–76, July/September 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smith:2016:DDL

[Smi16a] Alvy Ray Smith. The dawn of digital light. IEEE Annals of the History of Computing, 38(4):74–91, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smith:2016:IOH

[Smi16b] Andrew Smith. The industrial organization of Hong Kong’s progression toward a cashless economy (1960s–2000s). IEEE Annals of the History of Computing, 38(2):54–65, April/ June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Southwick:2019:HNC

[Sou19] D. Southwick. High Noon on the creative frontier: Configuring human and machine expertise. IEEE Annals of the History of Computing, 41(4):97– 109, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Speiser:1998:IRL

[Spe98] A. P. Speiser. IBM Research Laboratory Zurich: The early years. IEEE Annals of the History of Computing, 20(1): 15–28, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/14098/00646205.pdf.

Spicer:2003:TDT

[Spi03] D. Spicer. They dream tomorrow: a portrait of Ada Lovelace [book review]. IEEE Annals of the History of Computing, 25 (2):83, April/June 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2005:IRP

[Spi05] Dag Spicer. The IBM 1620 restoration project. IEEE Annals of the History of Computing, 27(3):33–43, July/ September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spinellis:2008:AMC

[Spi08]  Diomidis Spinellis.  The

Antikythera mechanism: a computer science perspective. Computer, 41(5):22–27, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spicer:2014:IDE

[Spi14] Dag Spicer. [Interview]: Deborah Estrin. IEEE Annals of the History of Computing, 36(3): 64–69, July/September 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2015:FCT

[Spi15a] Dag Spicer. Fernando Corbató:

Time-sharing pioneer, Part 1. IEEE Annals of the History of Computing, 37(4):5–9, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2015:GB

[Spi15b] Dag Spicer. Gordon Bell.

IEEE Annals of the History of Computing, 37(1):4–11, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2016:FCT

[Spi16a] Dag Spicer. Fernando Corbató: Time-sharing pioneer, part 2. IEEE Annals of the History of Computing, 38(1):75– 79, January/March 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Spicer:2016:RTEa

[Spi16b] Dag Spicer. Raymond Tomlinson: Email pioneer, part 1. IEEE Annals of the History of Computing, 38(2):72– 79, April/June 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2016:RTEb

[Spi16c] Dag Spicer. Raymond Tomlinson: Email pioneer, part 2. IEEE Annals of the History of Computing, 38(3):78–83, ????

2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2017:CH

[Spi17] Dag Spicer. CHM happenings. IEEE Annals of the History of Computing, 39(2):89–90, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Spicer:2018:OHD

[Spi18]  Dag Spicer.  Oral history

of Dame Stephanie Shirley. IEEE Annals of the History of Computing, 40(1):6–17, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010006.pdf.

Spicer:2021:CHM

[Spi21] Dag Spicer. Computer History Museum, fall 2021 update. IEEE Annals of the History of Computing, 43(4):108– 110, October/December 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2023:CHM

[Spi23a] Dag Spicer. Computer History Museum update. IEEE Annals of the History of Computing, 45(1):88–93, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2023:ILC

[Spi23b] Dag Spicer. Interview of Lynn Conway. IEEE Annals of the History of Computing, 45(2): 53–61, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2024:CHM

[Spi24a] Dag Spicer. The Computer History Museum’s 2023 in review. IEEE Annals of the History of Computing, 46(1):75– 79, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spicer:2024:NWO

[Spi24b] Dag Spicer. Niklaus Wirth obituary. IEEE Annals of the History of Computing, 46(1): 74, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Spring:2016:RCC

[Spr16] Adam P. Spring. Retrospective computing and consumerled development. IEEE Annals of the History of Computing, 38(2):66–71, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sproull:2018:PCG

[Spr18a] Robert F. Sproull. Publishing a computer graphics book with prototype desktop publishing tools. IEEE Annals of the History of Computing, 40 (4):69–76, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8506443/.

Sproull:2018:XAP

[Spr18b]  Robert F. Sproull.  The

Xerox Alto publishing platform. IEEE Annals of the History of Computing, 40(3):38–54, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509528/.

Saarikoski:2014:GNM

[SR14] Petri Saarikoski and Markku Reunanen. Great northern machine wars: Rivalry between user groups in Finland. IEEE Annals of the History of Computing, 36(2): 16–26, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Segesta:2002:HTC

[SRG02] Jim Segesta and Keith Reid-Green. Harley Tillitt and computerized library searching. IEEE Annals of the History of Computing, 24(3):23–34, July/ September 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a3023.pdf;   http://www.
computer.org/annals/an2002/
a3023abs.htm.

Saarikoski:2009:CHG

[SS09] Petri Saarikoski and Jaakko Suominen. Computer hobbyists and the gaming industry in Finland. IEEE Annals of the History of Computing, 31(3): 20–33, July/September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smith:2022:EAA

[SST22] Reid G. Smith, Eric J. Schoen, and Jay M. Tenenbaum. Early AI applications at Schlumberger. IEEE Annals of the History of Computing, 44(1): 88–102, January/March 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2003:MMM

[Sta03]  Zbigniew Stachniak.  The

making of the MCM/70 microcomputer. IEEE Annals of the History of Computing, 25(2):62–75, April/June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2062.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2062.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/02/a2062abs.
htm.

Stachniak:2007:ISP

[Sta07]  Zbigniew Stachniak.  In-

tel SIM8-01: a proto-PC. IEEE Annals of the History of Computing, 29(1):34–48, January/March 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2010:MMM

[Sta10] Zbigniew Stachniak. The MIL MF7114 microprocessor. IEEE Annals of the History of Computing, 32(4):48–59, October/ December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2013:CNM

[Sta13] Zbigniew Stachniak. This is not a computer: Negotiating the microprocessor. IEEE Annals of the History of Computing, 35 (4):48–54, October/December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2014:ECE

[Sta14] Zbigniew Stachniak. Early commercial electronic distribution of software. IEEE Annals of the History of Computing, 36(1):39–51, January/March 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2015:RCS

[Sta15] Z. Stachniak. Red clones: The Soviet computer hobby movement of the 1980s. IEEE Annals of the History of Computing, 37(1):12–23, January/March 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2017:MPS

[Sta17] Zbigniew Stachniak. MCM on personal software. IEEE Annals of the History of Computing, 39(1):52–64, January/March 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/01/
man2017010052-abs.html.

Stachniak:2019:SRB

[Sta19] Z. Stachniak. Software recovery and beyond. IEEE Annals of the History of Computing, 41(4):110–118, October/December 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stachniak:2020:LP

[Sta20a] Z. Stachniak. Learning from prototypes. IEEE Annals of the History of Computing, 42(2): 63–71, April/June 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stark:2020:HCC

[Sta20b] L. Stark. Here come the computer people: Anthropomorphosis, command, and control in early personal computing. IEEE Annals of the History of Computing, 42(4):53–70, October/December 2020. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stepulevage:2003:CBO

[Ste03] Linda Stepulevage. Computerbased office work: Stories of gender, design, and use. IEEE Annals of the History of Computing, 25(4):67–72, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL a4067abs.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4067.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4067.pdf.

Sterling:2006:PRL

[Ste06] Christopher H. Sterling. Pioneering risk: Lessons from the US Teletext/Videotex failure. IEEE Annals of the History of Computing, 28(3):41– 47, July/September 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stevens:2009:HDD

[Ste09] Orrin Stevens. The history of Datacom/DB. IEEE Annals of the History of Computing, 31 (4):87–91, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stevens:2022:BMB

[Ste22] Hallam Stevens. The business machine in biology the commercialization of AI in the life sciences. IEEE Annals of the History of Computing, 44(1):8– 19, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stine:2024:LLL

[Sti24]  Kyle Stine.  The logistics

of labor and life at Signetics. IEEE Annals of the History of Computing, 46(2):52– 64, April/June 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stolyarov:1999:CBC

[Sto99] G. K. Stolyarov. Computers in Belarus: chronology of the main events. IEEE Annals of the History of Computing, 21(3):61–70, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/16907/00778984.pdf.

Subramanian:2012:MTF

[Sub12] Ramesh Subramanian. Murray Turoff: Father of computer conferencing. IEEE Annals of the History of Computing, 34(1):92–98, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Subramanian:2013:SRH

[Sub13] Ramesh Subramanian. Starr Roxanne Hiltz: Pioneer digital sociologist. IEEE Annals of the History of Computing, 35(1): 78–85, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Subramanian:2014:TPN

[Sub14] R. Subramanian. Technology policy and national identity: The microcomputer comes to India. IEEE Annals of the History of Computing, 36(3):19– 29, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Subramanian:2020:HCC

[Sub20] R. Subramanian. Historical consciousness of cyber security in India. IEEE Annals of the History of Computing, 42 (4):71–93, October/December 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Subramanian:2023:EDa

[Sub23a] Ramesh Subramanian. From the Editor’s desk. IEEE Annals of the History of Computing, 45(1):5, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Subramanian:2023:EDb

[Sub23b] Ramesh Subramanian. From the Editor’s desk. IEEE Annals of the History of Computing, 45(2):5, April/June 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Sumner:2007:WMP

[Sum07] James Sumner. What makes a PC? thoughts on computing platforms, standards, and compatibility. IEEE Annals of the History of Computing, 29(2): 88, 87, April/June 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Supnik:2015:SS

[Sup15]  Bob Supnik.  The story

of SimH. IEEE Annals of the History of Computing, 37(3):78–80, July/ September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/03/
man2015030078-abs.html.

Stout:1995:PWF

[SW95] Thomas M. Stout and Theodore J.

Williams. Pioneering work in the field of computer process control. IEEE Annals of the History of Computing, 17(1): 6–18, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Smillie:1996:HCS

[SW96] K. Smillie and E. Weiss. The history of computing science at the University of Alberta. IEEE Annals of the History of Computing, 18(1):70, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476568.pdf.

Swartzlander:1995:Ca

[Swa95a] E. Swartzlander. Calculators. IEEE Annals of the History of Computing, 17(3):75–77, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397063.pdf.

Swartzlander:1995:Cb

[Swa95b] E. Swartzlander. Calculators. IEEE Annals of the History of Computing, 17(4):4, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10129/00477430.pdf.

Swartzlander:1995:GC

[Swa95c] Earl Swartzlander. Generations of calculators. IEEE Annals of the History of Computing, 17(3):75–77, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397076.pdf.

Swade:2005:CCB

[Swa05a] Doron D. Swade. The construction of Charles Babbage’s Difference Engine No.

2. IEEE Annals of the History of Computing, 27(3):70– 88, July/September 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Swade:2005:HR

[Swa05b] Doron D. Swade. Historical reconstructions. IEEE Annals of the History of Computing, 27(3):3, July/September 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//csdl.computer.org/comp/
mags/an/2005/03/a3003.pdf.

Swerdlow:1998:CCR

[Swe98]  N. M. Swerdlow.  Calen-

drical calculations [reviews]. IEEE Annals of the History of Computing, 20(3): 78, July/September 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15262/00707580.pdf.

Swets:2005:ABA

[Swe05] John A. Swets. The ABC’s of BBN: From acoustics to behavioral sciences to computers. IEEE Annals of the History of Computing, 27(2):15–29, April/June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/02/a2015.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/an/2005/02/a2015.
                                        pdf.

Szabo:2018:HAB

[SWH18] Mate Szabo, Amanda Wick, and Hiromichi Hashizume. The House of Arts, Brno; Jean Sammet Collection; and IPSJ convention. IEEE Annals of the History of Computing, 40 (2):73–79, ???? 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Szabo:2018:WAE

[Sza18] Máté Szabó. Warhol and the Amiga: Exhibition review. IEEE Annals of the History of Computing, 40 (4):77–79, October/December 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8579523/.

Szentgyorgyi:1999:SHC

[Sze99] Z. Szentgyorgyi. A short history of computing in Hungary. IEEE Annals of the History of Computing, 21(3): 49–57, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778982.pdf.

Takahashi:1996:BHJ

[Tak96] S. Takahashi. A brief history of the Japanese computer industry before 1985. IEEE Annals of the History of Computing, 18(1):76, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476569.pdf.

Takahashi:2005:RFP

[Tak05] Shigeru Takahashi. The rise and fall of plug-compatible mainframes. IEEE Annals of the History of Computing, 27(1):4–16, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1004.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1004.
pdf.

Tatarchenko:2010:CWO

[Tat10]  Ksenia Tatarchenko.  Cold

War origins of the International Federation for Information Processing. IEEE Annals of the History of Computing, 32(2):46–57, April/June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tatnall:2013:AEC

[Tat13] Arthur Tatnall. The Australian educational computer that never was. IEEE Annals of the History of Computing, 35(1):35–47, January/March 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tesler:2018:HME

[Tes18] Lawrence G. Tesler. How modeless editing came to be. IEEE Annals of the History of Computing, 40 (3):55–67, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8509519/.

Thakor:2017:HLA

[Tha17] Mitali Thakor. How to look: Apprehension, forensic craft, and the classification of child exploitation images. IEEE Annals of the History of Computing, 39(2):6–8, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Tinn:2010:CWP

[Tin10] Honghong Tinn. Cold War politics: Taiwanese computing in the 1950s and 1960s. IEEE Annals of the History of Computing, 32(1):92–95, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tinn:2011:DCI

[Tin11] Honghong Tinn. From DIY computers to illegal copies: The controversy over tinkering with microcomputers in Taiwan, 1980–1984. IEEE Annals of the History of Computing, 33(2):75–88, April/June 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Turkel:2014:GDH

[TMS14] William J. Turkel, Shezan Muhammedi, and Mary Beth Start. Grounding digital history in the history of computing. IEEE Annals of the History of Computing, 36(2): 72–75, April/June 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Todd:1993:CQD

[Tod93]  John Todd.  Comments,

queries, and debate: Alwin Walther (1898–1967). IEEE Annals of the History of Computing, 15(4):5–6, October/ December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4005.pdf.

Toland:2021:CCT

[Tol21] Janet Toland. Consortium computing and time slicing in the banking sector: Databank Systems Ltd New Zealand. IEEE Annals of the History of Computing, 43(2):18–29, April/June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tomayko:1992:Aa

[Tom92a] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 14(1):56–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1056.pdf.

Tomayko:1992:Ab

[Tom92b] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 14(2):55– 56, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a2055.pdf.

Tomayko:1992:Ac

[Tom92c] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 14(3):52–??, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3052.pdf.

Tomayko:1993:Aa

[Tom93a] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 15(1):59–??, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a1059.pdf.

Tomayko:1993:Ab

[Tom93b] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 15(2):54– ??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a2054.pdf.

Tomayko:1993:Ac

[Tom93c] James E. Tomayko. Anecdotes.

IEEE Annals of the History of Computing, 15(4):79–83, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4079.pdf.

Tomayko:1994:A

[Tom94] James E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 16(3): 59–??, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a3059.pdf.

Tomayko:1995:Ab

[Tom95a] J. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 17(3):77, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397064.pdf.

Tomayko:1995:Aa

[Tom95b] James Tomayko. Anecdotes. IEEE Annals of the History of Computing, 17(1):61– 64, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tomayko:1995:Ac

[Tom95c] James Tomayko. Anecdotes. IEEE Annals of the History of Computing, 17(3):77, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tomayko:1996:CVH

[Tom96] J. Tomayko. Computers vs. the human race. IEEE Annals of the History of Computing, 18(4):60, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11673/00539917.pdf.

Tomayko:1999:A

[Tom99] J. E. Tomayko. Anecdotes. IEEE Annals of the History of Computing, 21(1): 45–47, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16436/00759369.pdf.

Tomayko:2000:Aa

[Tom00a]  J. E. Tomayko.  Anec-

dotes. IEEE Annals of the History of Computing, 22 (2):61–66, April/June 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/18190/00841138.pdf.

Tomayko:2000:IRD

[Tom00b] J. E. Tomayko. Ida Rhodes and the dreams of a human computer. IEEE Annals of the History of Computing, 22(1): 82–85, January/March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17665/00815468.pdf.

Tomayko:2000:Ab

[Tom00c] James Tomayko. Anecdotes.

IEEE Annals of the History of Computing, 22(4):72–73, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4072.pdf.

Toole:1996:ABL

[Too96] B. A. Toole. Ada Byron, Lady Lovelace, an analyst and metaphysician. IEEE Annals of the History of Computing, 18(3):4–12, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511939.pdf.

Toomey:2010:FEU

[Too10] Warren Toomey. First Edition Unix: Its creation and restoration. IEEE Annals of the History of Computing, 32(3):74– 82, July/September 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tran:2019:MHP

[Tra19] J. L. Tran. The myth of Hush-A-Phone v. United States. IEEE Annals of the History of Computing, 41(4):6–19, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tracy:2024:BLP

[Tra24] Kim W. Tracy. Bell Labs’ portrayal of switching as computing (or not). IEEE Annals of the History of Computing, 46(3):81–85, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tremblay:2001:A

[Tre01] Guy Tremblay. Anecdotes. IEEE Annals of the History of Computing, 23(1):76–79, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an2001/pdf/
                                        a1076.pdf.

Trimble:2001:BHC

[Tri01] George R. Trimble, Jr. A brief history of computing: Memoirs of living on the edge. IEEE Annals of the History of Computing, 23(3):44–59, July/ September 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3044abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a3044.
pdf.

Tropp:1993:CQD

[Tro93] Henry S. Tropp. Comments, queries, and debate: Turing’s visit to the United States. IEEE Annals of the History of Computing, 15(1):10, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/

a1007.pdf. See correction [Tro95].

Tropp:1995:CQD

[Tro95] Henry S. Tropp. Comments, queries, and debate: Correction: Turing’s visit to the United States. IEEE Annals of the History of Computing, 17(3):6, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8988/00397054.pdf. See

[Tro93].

Turdean:2011:CDS

[Tur11] Cristina Turdean. Casinos and the digitization of the slot machine, 1950–1989. IEEE Annals of the History of Computing, 33(2):46–59, April/June 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tweedale:1993:MCP

[Twe93] Geoffrey Tweedale. A Manchester computer pioneer: Ferranti in retrospect. IEEE Annals of the History of Computing, 15(3):37–43, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a3037.pdf;   http://www.
                                        computer.org/annals/an1993/
                                        a3037abs.htm.

Tatarchenko:2021:RLCa

[TYD21a] Ksenia Tatarchenko, Anya Yermakova, and Liesbeth De Mol. Russian logics and the culture of impossible: Part I. Recovering intelligentsia logics. IEEE Annals of the History of Computing, 43(4):43–56, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tatarchenko:2021:RLCb

[TYD21b] Ksenia Tatarchenko, Anya Yermakova, and Liesbeth De Mol. Russian logics and the culture of impossible: Part II. Reinterpreting algorithmic rationality. IEEE Annals of the History of Computing, 43(4):57–69, October/December 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Tympas:1996:DAB

[Tym96] A. Tympas. From digital to analog and back: the ideology of intelligent machines in the history of the electrical analyzer, 1870s–1960s. IEEE Annals of the History of Computing, 18(4):42– 48, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11673/00539915.pdf.

Tympas:1999:PLS

[Tym99] A. Tympas. Paul Levinson, The Soft Edge: A Natural History and Future of the Information Revolution [book review]. IEEE Annals of the History of Computing, 21(3): 76, July/September 1999. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Telksnys:1999:CL

[TZ99] L. Telksnys and A. Zilinskas. Computers in Lithuania. IEEE Annals of the History of Computing, 21(3): 31–37, July/September 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        85/16907/00778980.pdf.

Urquhart:2024:BRT

[Urq24] Neil Urquhart. The British Rail Total Operations Processing System and the birth of telematics. IEEE Annals of the History of Computing, 46(3): 61–73, July/September 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Usselman:1996:FCC

[Uss96] S. W. Usselman. Fostering a capacity for compromise: business, government, and the stages of innovation in American computing. IEEE Annals of the History of Computing, 18(2):30–39, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10525/00489748.pdf.

Vagts:2024:THI

[Vag24] Brian K. Vagts. Telenet, the 1983 hacking incidents, and the construction of network security in the United States. IEEE Annals of the History of Computing, 46(1):34–47, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

VanDenEnde:1995:CIO

[Van95] J. Van Den Ende. Computers and industrial organization: early sources of ‘just in time’ production in the Dutch steel industry. IEEE Annals of the History of Computing, 17(2):22–32, Summer 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/8635/00380269.pdf.

VanDormael:2009:BHF

[Van09] Armand Van Dormael. Biographies: Herbert F. Mataré. IEEE Annals of the History of Computing, 31(3):68–73, July/ September 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

VanDormael:2010:HW

[Van10a] Armand Van Dormael. Heinrich Welker. IEEE Annals of the History of Computing, 32(2):72–79, April/June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

VanLaer:2010:DEC

[Van10b] Arthe Van Laer. Developing an EC computer policy, 1965– 1974. IEEE Annals of the History of Computing, 32(1):16– 17, January/March 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

VanVleck:2012:EMT

[Van12] Tom Van Vleck. Electronic mail and text messaging in CTSS, 1965–1973. IEEE Annals of the History of Computing, 34(1):4–6, January/March 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

vanEmden:2019:RBL

[van19] M. van Emden. Reflecting back on the Lighthill Affair. IEEE Annals of the History of Computing, 41(4):119–123, October/December 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Vardalas:1994:DFT

[Var94] John Vardalas. From DATAR to the FP-6000: Technological change in a Canadian industrial context. IEEE Annals of the History of Computing, 16(2): 20–30, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a2020.pdf;   http://www.
computer.org/annals/an1994/
a2020abs.htm.

vandenEnde:1992:TCN

[vdE92] Jan van den Ende. Tidal calculations in The Netherlands. IEEE Annals of the History of Computing, 14(3):23–33, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3023.pdf;   http://www.
computer.org/annals/an1992/
a3023abs.htm.

vandenEnde:1994:NFP

[vdE94]  Jan van den Ende.  The

number factory: Punchedcard machines at the Dutch Central Bureau of Statistics. IEEE Annals of the History of Computing, 16 (3):15, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a3015.pdf;   http://www.
computer.org/annals/an1994/
a3015abs.htm.

Vehvilainen:1999:GCR

[Veh99] M. Vehvilainen. Gender and computing in retrospect: the case of Finland. IEEE Annals of the History of Computing, 21(2):44–51, April/June 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/16485/00761794.pdf.

Veraart:2011:LMC

[Ver11] Frank Veraart. Losing meanings: Computer games in Dutch domestic use, 1975– 2000. IEEE Annals of the History of Computing, 33(1):52– 65, January/March 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

vonHerrmann:2023:LAI

[vH23] Hans-Christian von Herrmann. Literature and artificial intelligence. IEEE Annals of the History of Computing, 45(3):11– 18, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Viales-Hurtado:2015:BMI

[VHCSCC15] Ronny J. Viales-Hurtado, Ana Lucia Calderon-Saravia, and David Chavarria-Camacho. Between Matilde and the Internet: Computerizing the University of Costa Rica (1968– 1993). IEEE Annals of the History of Computing, 37(4): 29–39, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Vidan:2020:CBP

[Vid20] G. Vidan. Checks and balances: Publics, interests, and the development of electronic fund transfers in 1970s US. IEEE Annals of the History of Computing, 42(3):11–25, July/ September 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Vidan:2023:RPC

[Vid23] Gili Vidan. Review of Proxies: The Cultural Work of Standing In, by Dylan Mulvin. IEEE Annals of the History of Computing, 45(3):103–104, July/September 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Varma:2018:RRP

[VM18] Roli Varma and Owen Marshall. Reviews of recent publications by Aspray and Nelson [2 books reviewed]. IEEE Annals of the History of Computing, 40(1):88–91, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2018/01/
man2018010088.pdf.

Versprille:2024:NRB

[VM24] Kenneth J. Versprille and Alexander B. Magoun. Nonuniform Rational B-Splines, Computervision, and me. IEEE Annals of the History of Computing, 46(4):63–69, October/ December 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

vonNeumann:1993:FDR

[vN93] John von Neumann. First draft of a report on the EDVAC. IEEE Annals of the History of Computing, 15(4):28–75, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a4027.pdf;   http://www.
computer.org/annals/an1993/

a4027abs.htm. Edited and corrected by Michael D. Godfrey.

Valorinta:2011:IEU

[VN11] Mikko Valorinta and Tomi Nokelainen. Introduction and early use of computers in the Finnish retail industry. IEEE Annals of the History of Computing, 33(4):45–55, October/ December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Vogel:2017:QSI

[Vog17] William F. Vogel. “The spitting image of a woman programmer”;: Changing portrayals of women in the American computing industry, 1958-1985. IEEE Annals of the History of Computing, 39(2): 49–64, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

                                        URL http://ieeexplore.
                                        ieee.org/document/7974717/

.

Vinsel:2022:INS

[VR22] Lee Vinsel and Andrew L. Russell. Introductory note to seeking high IMP reliability in maintenance of the 1970s ARPAnet. IEEE Annals of the History of Computing, 44 (2):6–7, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Vardouli:2021:WIW

[VT21] Theodora Vardouli and David Theodore. Walking instead of working: Space allocation, automatic architecture, and the abstraction of hospital labor. IEEE Annals of the History of Computing, 43(2):6–17, April/ June 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wade:2012:CSS

[Wad12] Bradford W. Wade. Compiling SQL into System/370 machine language. IEEE Annals of the History of Computing, 34 (4):49–50, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wallace:2000:IIB

[Wal00]  H. D. Wallace.  Inventing

the Internet [book review]. IEEE Annals of the History of Computing, 22(4):66– 67, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887994.pdf.

Walden:2011:IWH

[Wal11]  David Walden.  Interview:

Willis H. Ware. IEEE Annals of the History of Computing, 33(3):67–73, July/September 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2012:PJD

[Wal12] David Walden. Peter J. Denning. IEEE Annals of the History of Computing, 34(4):72– 77, October/December 2012.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2018:ICB

[Wal18] David Walden. Interview with Charles Bigelow. IEEE Annals of the History of Computing, 40(3):95–103, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8509521/.

Walden:2019:AAM

[Wal19] D. Walden. ARPANET 50th anniversary marked at 2019 AAAS Annual Meeting. IEEE Annals of the History of Computing, 41(2):61–63, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2020:ILB

[Wal20] David Walden. Interview with Larry Bohn. IEEE Annals of the History of Computing, 42(1):51–62, January/March 2020. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2021:NE

[Wal21] D. Walden. Notes from the Editor. IEEE Annals of the History of Computing, 43(1): 94, January/March 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Warner:2015:NEC

[War15] Michael Warner. Notes on the evolution of computer security policy in the US Government, 1965–2003. IEEE Annals of the History of Computing, 37(2):8– 18, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://www.computer.
org/csdl/mags/an/2015/02/
man2015020008-abs.html.

Warnock:2018:OP

[War18] John E. Warnock. The origins of PostScript. IEEE Annals of the History of Computing, 40(3):68–76, July/September 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8509542/.

Wasiak:2013:CDD

[Was13] Patryk Wasiak. Computer dealer demos: Selling home computers with bouncing balls and animated logos. IEEE Annals of the History of Computing, 35(4):56–68, October/ December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wade:2012:IRD

[WC12] Bradford W. Wade and Donald D. Chamberlin. IBM relational database systems: The early years. IEEE Annals of the History of Computing, 34 (4):38–48, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wolcott:1998:IPM

[WD98] P. Wolcott and M. N. Dorojevets. The Institute of Precision Mechanics and Computer Technology and the El’brus family of high-speed computers. IEEE Annals of the History of Computing, 20(1): 4–14, January/March 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/14098/00646204.pdf.

Walden:2014:ED

[WE14] David Walden and Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 36(3):2– 3, July/September 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Weiss:1992:H

[Wei92a] E. W. Weiss. Happenings. IEEE Annals of the History of Computing, 14(1):55–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a1055.pdf.

Weiss:1992:Ba

[Wei92b] Eric A. Weiss. Biographies. IEEE Annals of the History of Computing, 14(1):61–??, January/March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a1061.pdf.

Weiss:1992:Bb

[Wei92c] Eric A. Weiss. Biographies.

IEEE Annals of the History of Computing, 14(2):56–59, April/June 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a2056.pdf.

Weiss:1992:Bd

[Wei92d] Eric A. Weiss. Biographies.

IEEE Annals of the History of Computing, 14(4):67–??, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4067.pdf.

Weiss:1992:BEA

[Wei92e]  Eric A. Weiss.  Biogra-

phies: Eloge: Arthur Lee Samuel (1901–90). IEEE Annals of the History of Computing, 14(3):55–69, July/ September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3055.pdf.

Weiss:1993:Bb

[Wei93a] Eric A. Weiss. Biographies. IEEE Annals of the History of Computing, 15(2):55–??, April/June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a2055.pdf.

Weiss:1993:Bc

[Wei93b] Eric A. Weiss. Biographies. IEEE Annals of the History of Computing, 15(3):70–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a3070.pdf.

Weiss:1993:Bd

[Wei93c] Eric A. Weiss. Biographies. IEEE Annals of the History of Computing, 15(4):84–??, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a4084.pdf.

Weiss:1993:BEW

[Wei93d] Eric A. Weiss. Biographies: Eloge: An Wang, 1920–1990. IEEE Annals of the History of Computing, 15(1):60–69, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.

org/an/books/an1993/pdf/
a1060.pdf.

Weierud:1994:RCI

[Wei94a]  Frode Weierud.  Reviews:

Codebreakers: The Inside Story of Bletchley Park — a book review. IEEE Annals of the History of Computing, 16(1): 78–79, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1076.pdf.

Weiss:1994:Ba

[Wei94b] Eric A. Weiss. Biographies.

IEEE Annals of the History of Computing, 16(1):71– ??, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a1071.pdf.

Weiss:1994:Bb

[Wei94c] Eric A. Weiss. Biographies.

IEEE Annals of the History of Computing, 16(2):55– 64, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a2055.pdf.

Weiss:1994:Bc

[Wei94d]  Eric A. Weiss.  Biogra-

phies. IEEE Annals of the History of Computing, 16(3): 61–??, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a3061.pdf.

Weiss:1994:Bd

[Wei94e] Eric A. Weiss. Biographies. IEEE Annals of the History of Computing, 16(4):76– ??, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1994/pdf/
                                        a4076.pdf.

Weiss:1995:Ba

[Wei95a] Eric Weiss. Biographies. IEEE Annals of the History of Computing, 17(1):64–73, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Weiss:1995:Bb

[Wei95b] Eric Weiss. Biographies. IEEE Annals of the History of Computing, 17(3):79, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/8988/00397065.pdf.

Weiss:1995:BCM

[Wei95c]  Eric Weiss.  Biographies:

Calvin Mooers (1919–1994). IEEE Annals of the History of Computing, 17(3):79–81, Fall 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Weiss:1996:BRL

[Wei96a] Eric Weiss. Biographies: Roger Lee Sisson. IEEE Annals of the History of Computing, 18(2):67–70, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489725.pdf.

Weiss:1996:GCT

[Wei96b] Eric Weiss. Grosch claims title of “Oldest Living Computer”. IEEE Annals of the History of Computing, 18(1):70, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476567.pdf.

Weiss:1996:KZO

[Wei96c]  Eric Weiss.  Konrad Zuse

obituary. IEEE Annals of the History of Computing, 18(2):3, Summer 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10525/00489747.pdf.

Weiss:1996:NWM

[Wei96d] Eric Weiss. Neil Wiseman Memorial Fund. IEEE Annals of the History of Computing, 18(1):70, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476566.pdf.

Weiss:1996:O

[Wei96e]  Eric Weiss.  Obituaries.

IEEE Annals of the History of Computing, 18(4): 61, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/11673/00539918.pdf.

Weiss:1996:OGS

[Wei96f] Eric Weiss. Obituaries: Gerard Salton. IEEE Annals of the History of Computing, 18(1):67, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476564.pdf.

Weiss:1998:HH

[Wei98] E. A. Weiss. Helmut Hoelzer: 1912–1996. IEEE Annals of the History of Computing, 20 (2):62–63, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/14609/00667310.pdf.

Weiss:2000:MRF

[Wei00]  E. A. Weiss.  Magnetic

recording, the first 100 years. IEEE Annals of the History of Computing, 22(1):90, January/March 2000. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/17665/00815472.pdf.

Weischedel:2006:NLU

[Wei06] Ralph Weischedel. Naturallanguage understanding at BBN. IEEE Annals of the History of Computing, 28(1):46– 55, January/March 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Weizenbaum:2008:SPI

[Wei08] Joseph Weizenbaum. Social and political impact of the long-term history of computing. IEEE Annals of the History of Computing, 30(3):40– 42, July/September 2008. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Westergard:2017:DCF

[Wes17] Björn Westergard. Do computers follow rules once followed by workers? IEEE Annals of the History of Computing, 39(4):5–8, ???? 2017. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Weiss:1992:CQD

[WF92] Eric A. Weiss and Mervin Frank. Comments, queries, and debate: a mechanical delay line? IEEE Annals of the History of Computing, 14(3):6, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a3006.pdf.

Weiss:2007:R

[WFE07] Eric Weiss, Gerald Friedland, and Margarita Esponda. Reviews. IEEE Annals of the History of Computing, 29(3):72–75, July/ September 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2007/03/
man2007030072.pdf.

Warnock:2019:FGA

[WG19] John E. Warnock and Charles Geschke. Founding and growing Adobe Systems, Inc. IEEE Annals of the History of Computing, 41(3):24–34, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2017:DPP

[WH17] David Walden and Hiromichi Hashizume. Desktop publishing pioneer meeting at Computer History Museum. IEEE Annals of the History of Computing, 39(3):65–67, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL https://www.computer.
org/csdl/mags/an/2017/03/
man2017030065-abs.html.

Wheeler:1992:EPS

[Whe92a] David J. Wheeler. The ED-SAC programming systems.

IEEE Annals of the History of Computing, 14(4):34–40, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4034.pdf;   http://www.
computer.org/annals/an1992/
a4034abs.htm.

Wheeler:1992:AE

[Whe92b] Joyce M. Wheeler. Applications of the EDSAC. IEEE Annals of the History of Computing, 14(4):27–33, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/
a4027.pdf;   http://www.
computer.org/annals/an1992/
a4027abs.htm.

Walden:2014:AIP

[WI14] David Walden and “IMP Software Guys”. The Arpanet IMP program: Retrospective and resurrection. IEEE Annals of the History of Computing, 36(2):28–39, April/June 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wiehle:2010:ECC

[Wie10] Hans Rudiger Wiehle. External characteristics of computer operations: Toward large conversational time-sharing systems. IEEE Annals of the History of Computing, 32(3):4– 19, July/September 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wilkes:1992:E

[Wil92a] Maurice V. Wilkes. EDSAC 2. IEEE Annals of the History of Computing, 14(4):49–56, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/
                                        a4049.pdf;   http://www.
                                        computer.org/annals/an1992/
                                        a4049abs.htm.

Williams:1992:BJC

[Wil92b] M. R. Williams. Biographies — Joseph Clement: the first computer engineer. IEEE Annals of the History of Computing, 14(3):69–76, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). Joseph Clement was Charles Babbage’s superb machinist who worked for years on building Babbage’s Difference Engine.

Williams:1993:AI

[Wil93a] Michael R. Williams. About this issue. IEEE Annals of the History of Computing, 15(3):3–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1993/pdf/
                                        a3003.pdf.

Williams:1993:OUF

[Wil93b] Michael R. Williams. The origins, uses, and fate of the EDVAC. IEEE Annals of the History of Computing, 15(1):22–38, January/March 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/
a1022.pdf;   http://www.
computer.org/annals/an1993/
a1022abs.htm.

Williams:1994:AI

[Wil94a] Michael R. Williams. About this issue. IEEE Annals of the History of Computing, 16(4): 3–??, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4003.pdf.

Williams:1994:UFU

[Wil94b] Michael R. Williams. UTEC and Ferut: The University of Toronto’s Computation Centre. IEEE Annals of the History of Computing, 16(2):4– 24, Summer 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a2004.pdf;   http://www.
computer.org/annals/an1994/
a2004abs.htm.

Williams:1996:AI

[Wil96a]  M. R. Williams.  About

this issue. IEEE Annals of the History of Computing, 18(1):4, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476555.pdf.

Williams:1996:NE

[Wil96b] M. R. Williams. New EIC. IEEE Annals of the History of Computing, 18(1):3, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/10202/00476554.pdf.

Wilkes:1997:AE

[Wil97a] M. V. Wilkes. Arithmetic on the EDSAC. IEEE Annals of the History of Computing, 19(1):13–15, January/March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/12228/00560726.pdf.

Williams:1997:AI

[Wil97b] M. R. Williams. About this issue. IEEE Annals of the History of Computing, 19(4): 3–4, October/December 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/13659/00627894.pdf.

Williams:1998:ASH

[Wil98a] M. R. Williams. Alternate sources of history. IEEE Annals of the History of Computing, 20(4):79–80, October/December 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728233.pdf.

Williams:1998:LWC

[Wil98b] M. R. (Michael R.) Williams.

The “Last word” on Charles Babbage. IEEE Annals of the History of Computing, 20 (4):10–14, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728225.pdf.

Wilkes:2000:CBC

[Wil00a]  M. V. Wilkes.  The cog-

wheel brain: Charles Babbage and the quest to build the first computer [book review]. IEEE Annals of the History of Computing, 22(4):62– 63, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887990.pdf.

Wilkes:2000:CQD

[Wil00b] Maurice Wilkes. Comments, queries, and debates: Was Babbage caught in the act?

IEEE Annals of the History of Computing, 22(4):69–70, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4069.pdf.

Wilkes:2000:IBA

[Wil00c] Maurice V. Wilkes. Introduction to “Babbage’s Analytical Engine Plans 28 and 28a — The Programmer’s Interface” by Allan Bromley. IEEE Annals of the History of Computing, 22(4):4–5, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4004.pdf;   http://ieeexplore.
ieee.org/iel5/85/19196/00887985.
pdf;   http://www.computer.
org/annals/an2000/a4004abs.
htm.

Williams:2000:AB

[Wil00d] Michael R. Williams. Allan Bromley. IEEE Annals of the History of Computing, 22(4):3–??, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/
a4003.pdf.

Williams:2001:BWC

[Wil01a] Michael R. Williams. Building a world-class book collection: The Tomash Library.

IEEE Annals of the History of Computing, 23(4):39–43, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a4039.pdf;   http://www.
computer.org/annals/an2001/
a4039abs.htm.

Williams:2001:CMN

[Wil01b] Michael R. Williams. Calvin Mooers, the NOL computer project, and John Vincent Atanasoff: An introduction. IEEE Annals of the History of Computing, 23(2):50–51, April/June 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a2050.pdf.

Williams:2002:UPC

[Wil02] Robert V. Williams. The use of punched cards in US libraries and documentation centers, 1936–1965. IEEE Annals of the History of Computing, 24(2):16–33, April/June 2002. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2002/pdf/
a2016.pdf;   http://www.
computer.org/annals/an2002/
a2016abs.htm;   http://www.
computer.org/annals/punchedcards.
htm.

Williams:2003:SAC

[Wil03] Elizabeth Phillips Williams. A systems analyst’s computer watch: 1943–2003. IEEE Annals of the History of Computing, 25(4):27–39, October/ December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL a4027abs.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/04/a4027.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/04/a4027.pdf.

Wilkes:2006:WRE

[Wil06] Maurice Wilkes. What I remember of the ENIAC. IEEE Annals of the History of Computing, 28(2):30–31, April/ June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Williams:2018:FPD

[Wil18] Michael Williams. The first public discussion of the secret Colossus Project. IEEE Annals of the History of Computing, 40(1):84–87, January/March 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/an/2018/01/
                                        man2018010084.pdf.

Winegrad:1996:CBM

[Win96] D. Winegrad. Celebrating the birth of modern computing: the fiftieth anniversary of a discovery at the Moore School of Engineering of the University of Pennsylvania. IEEE Annals of the History of Computing, 18(1):5–9, Spring 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10202/00476556.pdf.

Winkler:2003:BRK

[Win03] Jonathan Reed Winkler. Book review: Ken Beauchamp, A History of Telegraphy, IEE History of Technology Series no. 26, IEE Press, 2001, ISBN 085-296-7926, xxiv + 413 pp., $95.00. IEEE Annals of the History of Computing, 25 (4):93–94, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wirth:2008:BHS

[Wir08] Niklaus Wirth. A brief history of software engineering. IEEE Annals of the History of Computing, 30(3):32–39, July/ September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Williams:2005:RR

[WJ05] Bruce O. B. Williams and Roger G. Johnson. Ready reckoners. IEEE Annals of the History of Computing, 27(4):64– 80, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2022:SHI

[WMB22] David C. Walden, Alexander A. McKenzie, and W. Ben Barker. Seeking high IMP reliability in maintenance of the 1970s ARPAnet. IEEE Annals of the History of Computing, 44(2): 8–19, April/June 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walden:2005:CBB

[WN05] David C. Walden and Raymond S. Nickerson. Computing at Bolt Beranek and Newman: The first 40 years, part 1. IEEE Annals of the History of Computing, 27(2):4–5, April/ June 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/an/2005/02/a2004.
                                        htm;   http://csdl2.computer.
                                        org/comp/mags/an/2005/02/
                                        a2004.pdf.

Walden:2006:CBB

[WN06] Dave Walden and Ray Nickerson. Computing at Bolt Beranek and Newman: The first 40 years, part 2. IEEE Annals of the History of Computing, 28 (1):4–5, January/March 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wohl:2006:HWP

[Woh06] Amy Wohl. How we process words: The marketing of WP software. IEEE Annals of the History of Computing, 28 (4):88–91, October/December 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Woitkowski:2025:SYH

[Woi25] David Woitkowski. A special year at the Heinz Nixdorf MuseumsForum. IEEE Annals of the History of Computing, 47(1):76–80, January/March 2025. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wood:1995:CSC

[Woo95] H. M. Wood. Computer Society celebrates 50 years. IEEE Annals of the History of Computing, 17(4):6, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/10129/00477432.pdf.

Worthy:1995:CDC

[Wor95] James C. Worthy. Control Data Corporation: The Norris era. IEEE Annals of the History of Computing, 17(1): 47–53, Spring 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Whiting:1994:HDF

[WP94] Paul G. Whiting and Robert S. V. Pascoe. A history of data-flow languages. IEEE Annals of the History of Computing, 16(4):38– 59, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1994/pdf/
a4038.pdf;   http://www.
computer.org/annals/an1994/
a4038abs.htm.

Wright:1998:HTC

[Wri98]  T. Wright.  History and

technology of computer fonts. IEEE Annals of the History of Computing, 20(2):30– 34, April/June 1998. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667294.pdf.

Walden:2017:ES

[WS17] David Walden and Dag Spicer. Events and sightings. IEEE Annals of the History of Computing, 39(4):93–96, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Williams:2003:SHS

[WT03] Michael R. Williams and Erwin Tomash. The sector: Its history, scales, and uses. IEEE Annals of the History of Computing, 25(1):34–47, January/March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/01/a1034.
htm;   http://csdl.computer.

org/dl/mags/an/2003/01/a1034.
pdf;   http://csdl.computer.
org/dl/mags/an/2003/01/a1034abs.
htm.

Wu:2015:HIU

[Wu15] Angela Xiao Wu. Historicizing Internet use in China and the problem of the user figure. IEEE Annals of the History of Computing, 37(4):2–4, ???? 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wu:2021:CHC

[Wu21] Di Wu. Cripping the history of computing. IEEE Annals of the History of Computing, 43(3): 68–72, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wyvill:1998:ECG

[Wyv98]  B. Wyvill.  Experiences

with computer graphics in the United Kingdom in the 1970’s. IEEE Annals of the History of Computing, 20 (2):52–54, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/14609/00667302.pdf.

Yamada:2018:ICI

[Yam18] Akihiko Yamada. IPSJ certifies 100 IP technology heritage artifacts in 10 years. IEEE Annals of the History of Computing, 40(3):105–112, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yamada:2019:NCC

[Yam19a] A. Yamada. 2018 NEC C C Prize Ceremony. IEEE Annals of the History of Computing, 41(2):64–66, April 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yamada:2019:PRC

[Yam19b] Akihiko Yamada. Parametron and Relay Computers Exhibition held at Tokyo University of Science in November, 2018. IEEE Annals of the History of Computing, 41(1):61–63, January/March 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

https://ieeexplore.ieee.
org/document/8667952/.

Yates:1997:EIB

[Yat97] J. Yates. Early interactions between the life insurance and computer industries: the Prudential’s Edmund C. Berkeley. IEEE Annals of the History of Computing, 19(3): 60–73, July/September 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/13172/00601736.pdf.

Yood:2005:HCC

[Yoo05] Charlie Yood. The history of computing at the Consumption Junction. IEEE Annals of the History of Computing, 27 (1):88, 86–87, January/March 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2005/01/a1088.
htm;   http://csdl.computer.
org/dl/mags/an/2005/01/a1088.
pdf.

Yost:2001:CTF

[Yos01] Jeffrey R. Yost. CBI/Tomash Fellowship: Sponsoring a generation of scholars in the history of information processing. IEEE Annals of the History of Computing, 23(4):24–28, October/December 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2001/pdf/
a4024.pdf;   http://www.
computer.org/annals/an2001/
a4024abs.htm.

Yost:2008:EDa

[Yos08a]  Jeffrey R. Yost.  From

the Editor’s desk. IEEE Annals of the History of Computing, 30(1):2–3, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://csdl.computer.
org/comp/mags/an/2008/01/
man2008010002.pdf.

Yost:2008:EDb

[Yos08b] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 30(2): 2, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/an/2008/02/
                                        man2008020002.pdf.

Yost:2008:EDc

[Yos08c] Jeffrey R. Yost. From the editor’s desk. IEEE Annals of the History of Computing, 30 (3):2–3, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2008:EDd

[Yos08d] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 30(4): 2–3, October/December 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2009:EDa

[Yos09a] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 31(1): 2, January/March 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2009:EDb

[Yos09b] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 31 (2):2, April/June 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2009:EDc

[Yos09c] Jeffrey R. Yost. From the editor’s desk. IEEE Annals of the History of Computing, 31(3): 2, July/September 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2009:EDd

[Yos09d] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 31 (4):2, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2010:EDa

[Yos10a] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 32 (1):2–3, January/March 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2010:EDb

[Yos10b] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 32 (2):2–3, April/June 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2010:EDc

[Yos10c] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 32 (3):2–3, July/September 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2010:EDd

[Yos10d] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 32(4): 2–4, October/December 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2011:EDa

[Yos11a] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 33 (1):2–3, January/March 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2011:EDb

[Yos11b] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 33 (2):2–3, April/June 2011. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2011:EDc

[Yos11c] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 33 (3):2–3, July/September 2011.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2011:EDd

[Yos11d] Jeffrey R. Yost. From the Editor’s desk. IEEE Annals of the History of Computing, 33(4): 2–4, October/December 2011. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2013:AIB

[Yos13a] Jeffrey R. Yost. Appropriation and independence: BTM, Burroughs, and IBM at the advent of the computer industry. IEEE Annals of the History of Computing, 35(4):5–17, October/December 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2013:CIP

[Yos13b] Jeffrey R. Yost. Computer industry pioneer: Erwin Tomash (1921–2012). IEEE Annals of the History of Computing, 35 (2):4–7, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2015:CSG

[Yos15a] Jeffrey R. Yost. Computer security [Guest Editors’ introduction]. IEEE Annals of the History of Computing, 37(2):6– 7, April/June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/an/2015/02/
                                        man2015020006-abs.html.

Yost:2015:OEH

[Yos15b] Jeffrey R. Yost. The origin and early history of the computer security software products industry. IEEE Annals of the History of Computing, 37(2):46–58, April/ June 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/an/2015/02/man2015020046-
                                        abs.html.

Yost:2016:CSP

[Yos16a] Jeffrey R. Yost. Computer security, part 2. IEEE Annals of the History of Computing, 38(4):10–11, ???? 2016. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2016:MIE

[Yos16b] Jeffrey R. Yost. The march of IDES: Early history of intrusion-detection expert systems. IEEE Annals of the History of Computing, 38(4): 42–54, ???? 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2021:CBI

[Yos21] J. R. Yost. Charles Babbage Institute for Computing, Information, and Culture: a major interdisciplinary turn. IEEE Annals of the History of Computing, 43(1):92–93, January/March 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2022:NSN

[Yos22] Jeffrey R. Yost. New symposia, NSF Project, book series, and more at CBI. IEEE Annals of the History of Computing, 44(2):78–79, April/June 2022. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2023:CAD

[Yos23] Jeffrey Yost. CBI’s automation by design, new fellows, interfaces, JHUP series, NAE, NSF, and more. IEEE Annals of the History of Computing, 45(2): 72–76, April/June 2023. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yost:2024:IJP

[Yos24] Jeffrey Yost. Interview of J. P.

(Jack) London. IEEE Annals of the History of Computing, 46(2):72–87, April/June 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Yamada:2012:ES

[YS12] Akihiko Yamada and Dag Spicer. Events and sightings. IEEE Annals of the History of Computing, 34(3):70– 73, July/September 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Zoppke:2006:VLE

[ZR06] Till Zoppke and Raúl Rojas. The virtual life of ENIAC: Simulating the operation of the first electronic computer. IEEE Annals of the History of Computing, 28(2): 18–25, April/June 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Zhao:2023:TDH

[ZS23] Yue Zhao and Rebecca Slayton. Tool, doctor, and hooligan: History of antivirus software in China, 1989–2010. IEEE Annals of the History of Computing, 45(1):43–54, January/March 2023. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Zwerman:1999:POH

[Zwe99]  W. L. Zwerman.  Profes-

sion/occupation without a history. IEEE Annals of the History of Computing, 21(1): 66–70, January/March 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        ieeexplore.ieee.org/iel4/
                                        85/16436/00759371.pdf.